A Complete Bibliography of Publications in Bernoulli

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

14 August 2024 Version 1.11

Title word cross-reference

(1 ≤ α ≤ 2) [MMR06]. 1/Γ(1/α) [Har17]. 2 [Cun15, IKP14, ZZ15]. α [ALT15, BYY17, BM16, CLD08, CG99b, Har17, HP14, HG97, NZ05, OS15, SXX22, WXX17, WYZ24, ZZ23]. AR(1) [CLLLR17]. AR(p) [EH02]. C_{p} [BCFH22]. L^{p} [GMZ21]. COGARCH(1, 1) [Ste10]. Δ [CS13, Jac02, GLS07]. e [Wan24]. ℓ_{1} [TV06]. F [Zho14a]. G [AM07]. H > 1/2 [DN08, FR06]. H_{α} [Har17]. ∞ [RS01a]. K [CY21, BL20, DGS08]. L [BW10]. L(X) = L[B(X + C)] [Duf96]. L^{2} [GMD05]. L^{α} [MMR06]. L^{p} [Wan16]. L^{r} [PS12]. L_{1} [DHKV15]. L_{2} [DGU05, GS16, PLH09]. L_{p} [SW19a]. M [AM17, CL14, DL19, HLP12, SGRV18]. L_{p} [Lep16]. R [JLL04]. R^{d} [AM07]. Z^{d} [BL22]. AR(1) [Erh14]. μ [FK21]. n [CP12, HW19a, Tru19]. Gamma(k) [BB17]. p [AM16, Cha03, CP12, Gra99, Man07, Wan24]. p∗ [Wan24, BNW98]. p > N [AH17a]. p ≫ n [VV04]. Φ [DL17]. Q [LXY13]. R [JKNP16]. r∗ [BNW98]. R^{d} [Dor97]. R^{d+1} [LS95]. ρ [KP22a]. S [Ben03]. \sqrt{n} [GK12]. t [BJSZ07, CK11a, Jon11, Zho14a]. U [BZ12, BZ14, BG00, DM08b, DDL23, FPW17, Han22, Leu12, LS24, MW20, PT18a, Tho23, VW17, ZHZL19]. V [BZ12, BZ14, Leu12]. ε [BPR12, PJR16]. φ [LMZ24]. x log x [Olo09]. y [BRS07]. Z [AP00, DM24, BRS07].

-approach [DHKV15]. -coalescents [IMM09]. -consistent [GK12]. -convergence [GMD05, MMR06]. -D [ZZ15]. -densities [GMZ21]. -dependence [Kre21b]. -distributions [BB17]. -Dvoretzky [FK21]. -estimator [CL14, DL19]. -estimators [AM17, HLP12, JKNP16, SGRV18]. -exponential [ADH21]. -leaf [HW19a]. -matrix [LXY13]. -mean [DM24]. -means [AM16, CY21]. -mixing [DL17]. -norm [KP23]. -norms [Lep16]. -optimal [Gra99, PS12]. -optimality [Jac02]. -points-based [BL20]. -processes [Han22]. -quantile [VW17]. -quantiles [KP22a]. -records [GLS07]. -smooth [Cun15]. -Sobolev [LMZ24]. -stable [ALT15, BYY17, BM16, HP14, NZ05, OS15, SXX22, WXX17, WYZ24, ZZ23]. -statistic [BJSZ07, Jon11, ZHZL19]. -statistics [DGS08, BZ12, BZ14, BG00, BW10, DM08b, DDL23, FPW17, Leu12, LS24, MW20, PT18a, Tho23]. -Strong [BPR12, PJR16]. -sun [HG97]. -tests [BJSZ07, CK11a]. -theorems [AP00]. -transform [Ben03]. -trimming [CLD08]. -type [IKP14, Küh22]. -valued [BL22]. -values [Cha03, Wan24]. -variation [Man07, GS16]. -Wasserstein [Wan16].

1 [HP06]. 1-Wasserstein [STC^{+}24]. 1056 [DLPZ12]. 15 [DLPZ12].

2 [FFZ22]. 2-GREM [FFZ22]. 21st [Tay13].

A/B [HR22b]. Aalen [RB06]. abelian [Fri24, CR17]. Absolute [FP10, Neu11, FM02, KMS17, Nie13]. absolutely [Pon07]. Accelerated [FP17c, YG19, ZWY17]. acceleration [MCC^{+}21]. accessible [BBS16]. Accuracy [Ma12, Nov14, PY01]. Acknowledgement [Ano95a, Ano97a, Ano98a, Ano99a, Ano00a, Ano05a, Ano06a]. Acknowledgements [Ano07a, Ano08a, Ano09a, Ano10a, Ano11a, Ano12a, Ano13a]. across [Cho18]. activity [FLLO14, KMM24]. acyclic [GK18]. ad [DFK16]. ad-hoc [DFK16]. Adaptation [Arm21, BDKS20, Ran22, SZ03]. Adaptive [BCFK15, BBJ08, Bel21, BB22, Bul15, BC09, CL05, CKLN18, Cas14, Cha23, CL19, Che02, CCL23, Com01, DRT16, GG22, GN10, IMS01, JS13b, Nau22, PHO17, Rei05, Rig06, SG16, VG18, WZ16, AGV18, AZ15, AB11, AR05, AF10, AF12, Bia14, CCJM24, Cav01, CL14, DDJ12, DR23, DJ96, Efr98, Efr00, FB22b, FMPV14, GN08, GD17, Gol99, GP03, HST01, KPT96, Kuj16, MPS19a, MS18, Ore22, Reb15, CT19]. Adaptive-treed [Bul15]. Adaptively [HIW19]. Adaptiveness [FZ22]. Adaptivity [Cat11]. Addendum [GK01, PS13c]. Additive [HLP23, CZH14, DL13, HMP18a, HKM06, HYY11, JLMP22, LLTY07, Wal13, Xia09, YLC19]. additivity [DV01, PV07]. ADF [Kut15]. adjusted [LK08, ZGV20]. adjustment [BZ22]. admissibility [CP19]. admissible [MS23b]. admitting [BNMS06]. admixture [Ngu15]. AEP [AEP11]. Affine [KF14, Jab21, KS13b, MM20, Rei05]. affinely [DEGR14]. after [BC13]. against [EF04, HP02a, Jur99, WT21]. age [AAI11, BBD20, FHJK20]. age-dependent [AAI11]. aggregate [CT12, LM13]. aggregated [GLST19]. Aggregating [Yan04]. Aggregation [PS16, Ald99, ÁEdBCAM18, Bel17, JT22, Lec07, Lec13, LM16, Sau18, SW20]. AIC [BCFH22]. Akaike [Mat24]. algebra [Bou15]. Algebraic [MR17, AGG23, Dob23]. algorithm [AKT10, AEP11, AM16, BP03a, BPR^{+}13, BBW20, CCZ13, DR07, Din98, DM19, FB22b, GLO22, HST01, LM10, LE23, Nie00, OW17, OPPS16, Wan22b]. algorithms [AR05, AF10, AF12, BB22, CK19, CBL22, ELP24, GC17, HVM99, JLM14, ŁMN13, San16, SZ22, Wan22a, WLH16]. aligned [HM20a]. alignment [BP21, HMP18b, HU22]. alignments [Cha03, GY04, LMV14]. allocation [AZ15, PY19]. Almost [AW96, Gas23, AZ15, BHM22, BS23, CK11b, Jir23, Len11, Syn07]. almost-circular [BS23]. Almost-sure [Gas23]. alpha [IPP23]. alpha-stable [IPP23]. alphabet [BH10]. alphabets [Zha18, Zha22]. alternative [ČK00a, LS99, Mon15, ZL17]. alternatives [BL04, Jur99]. Amendments [Ano96a]. American [EI18]. among [HTM06, VW15]. analog [MCC^{+}21]. analogue [KVZ19]. Analysis [Fer01, JN16, AH17a, AW09, Bac18, BW22, BCG19, CLZ18, CT22, DNN10, DG13, DHKV15, DMR21, GGKL22, GC05, GD17, HL17, HXZ18, HPY18, HMWY10, HK11, Hös05, JCL20, JR17, KVDH16, KSX19, Las14, Lau13, LW04, Len11, LM15, ML20, Mar13, MZ14, NRC24, Nda23, NC24, RSGW03, RS10, RvS19, Ruk14, ST04, SGL^{+}24, SPP21, Tru20b, VW17, WAP17, Wan22a, ZP19, ZL16, dUÁV21, BRJN19]. analytic [Bel98, CGER07, GLT96]. analytical [NVV99]. analyticity [BNH08, MR19]. analyzing [CA19]. anchored [Pai22, PT21]. ancillary [BNW98, FFS10]. Anderson [BS19]. angle [ADM18]. angular [CJL^{+}23]. anisotropic [Ber04, Gra21, LX23, PS16]. annealed [GD17]. Annual [Ano95b, Ano97b, Ano98b, Ano99b, Ano00b, Ano05b, Ano06b]. anomaly [LS20]. ANOVA [SN13]. ANOVA-fit [SN13]. anti [GNSU19]. anti-concentration [GNSU19]. Anticipating [MPS98, Tud04]. any [DN08]. aperiodicity [CA19]. appear [KRX18]. applicable [BP19b]. Application [KKS02, ABF12, BGMSB14, BP21, BS22, BN17, BBD20, BL15b, BS14, BMS21, CY21, CGER07, DMP11, DR16, DVSS14, DVSS16, Dor97, DL09b, FdH14, FM02, FO24, GW17a, GMD06, GY04, GQ23, HXZ18, HMP18a, HKM^{+}16, JB21a, KM23, Kre21b, LN02, Mar06, Mas04a, MZZ20, OCDM08, PP09, Syn07, Tal13, WC14, ÁEdBCAM18, HLT17, RT15]. Applications [BVP22, BP19a, DMMW18, GL95, Sio18, APP11, AC01, AI99, Bar10, BS21, BIPZ22, BHBO17, BJSZ07, BPR06, BZ23, BZ19, BG10, BD13b, BX15, BC09, CCR20, CG22a, CY05, CK15, CJL^{+}23, CR07, Cun15, DSS10, DGU05, Din20, DRRS18, Dre03, ET23, FR15, FY20, GGT22, GKP12, GZ18, GCJL00, Gin96a, GST08, GMR19, HPY18, Han22, HZBH23, HR22b, KF14, KK17, KM95, LC22, LSW20, MVY19, MW20, Möh99, NS01, Pai22, PX10, Pen07, PTA07, PWZ17, SS20, Sch13, SZ22, SK23, Sie07, Tan06, TC03, VV04, Vet15, VW17, WXX17, Yan08, ZZW22, ZBY17, vDD21, CLD08, Del17]. Applying [Mon15]. approach [AG99, ALT15, ABF12, BBG19, Ben03, BZ14, CEL20, CTP21, CLLLR17, CJP21, CS07a, CDH21, CM14a, DDT14, DHKV15, ELP24, GP04, Gob01, HLL23, Hay22, HKMS24, KS18, KP22a, LLBF24, Lep18, LSW20, MRS20, MCSS99, Mon15, MPV22, NN22, ELP22, NVV99, PS13a, PR13, Røy11, RS01b, SBP24, SGL^{+}24, TC24, TY23, Zha06, ZHZL19, dCCGZ23]. approaches [CT02, PY03]. Approximate [BHL99, BP98, But24, GW17a, BS09, BBW23, FCH15, Lua11, Ped95]. Approximating [BP10, CR13, Dan96, DDT14, DR19, MPS19b]. Approximation [Alt21, BD16, DVSS14, DVSS16, GT16a, HMW21, LP00, MSS00, PP09, Röl05, AN12, AL05, AKT10, ACV09, AR17, AV14, ABH06, BX00, BX06, BGX15, BJ00, BP21, BN17, BG19, BG95, Bou06, BV10, Car06, ČK00a, CS07a, CX11, CRT21, CR07, DH13, DLR06, DP16, DGP09, DJL23, Fan14, FR15, Fri18, GR23b, GLL20, GS21, GP11, Hir97, JMW05, KG00, LMK19, LP24, LLTY07, MS23c, OU23, PRR13, RT15, Roo99, San10, Sch09, SBP24, Sko96, TY16, Tor17, Vil14, XZ09, Zac23, ZC18, Zha23]. Approximations [LD11, BM07, BL14, BM09, BP08, CHP11, ČV10, HW10, Han03, HJQZ04, Kee13, KL10, LMC23, NP08, RT96, Roo05, SvS17, UCV17, Zho14a]. Arbitrary [ZSC22, Bru20, PTL04]. ARCH [Bor01, DR07, FR11, KL00]. Archimedean [KFT21, Qiu19]. architectures [LV15]. arcsine [KVZ19, MPAiS12]. Area [LS20, ADM14, Kar24, NRT98]. areas [NS07]. arising [BC00, CP16, FS10, Möh99]. arm [HR22b]. ARMA [ADFS15, BKT19, FZ04, HP02a, YB06]. ARMA-GARCH [FZ04]. array [JK21, OKM^{+}98]. arrays [Has13, Mcc00, ST20]. arrival [AIZ16, GK03]. arrivals [AIS19]. Aspects [Rei13, FS13, Ruk14]. Assessing [HS07, VBLN20]. assessment [DGS20]. asset [TKT09]. associated [ALL09, CPP11, DK18, EGG03, HY13, HM15, KF14, Kra21, PDL19]. association [WC98, Wie22]. associative [LV15]. assumption [Lec07]. assumptions [Ber97, LM07, Seg12, ZH20]. asymmetric [EI18, FLS01]. Asymptotic [AI99, AT24, AGM06, Bac18, BC16, BAT19, Beg07, Bia14, Bor01, BW10, But17, CDG14, CMMA20, CV97, DN08, DH02, DLPZ09, DLPZ12, Fas10, FRW15, FM00, FFZ22, FGK16, Gho99, Gor96, GLS07, GJ13, GK99, GK01, HS10, HLS14, HZ04, HY08, IKP14, KVDH16, Klo06, Kuj16, Len11, Leu12, LS19c, NRC24, Nie13, Qia21a, TY16, Xia20, XL10, Xu21, ZL17, Bar02, BZ12, BR96, CHWW05, CG23, CG99b, CSY99, Dic16, DY14, DMR21, DDFD04, DvdAW09, DL13, FK12, Fri24, FO24, GL18, GK13, GL06, Gob01, Gra99, HR22a, HM23, Jan07, KN99, LMMR16, LGS08, NR12, Nie00, NIR21, Ogi15, OU23, Ped95, Tho23, WB19, Yan24, Zha96, ZZW22, ZSC22, dUÁV21]. Asymptotically [BPB^{+}22, Ber04, FT19, GLT96, KVD04, Oud98, SSH14, TCP22, BS12, BK23, Van05b]. Asymptotics [AD07, ACG19, AD11, BCFH22, BHS11, BBC18, BH10, Cap24, CM13b, DGU05, DFK10, DM01, DN21, FLP12, GS21, Gra21, IMM17a, IMM17b, Mas04a, McG09, MS23d, PS12, Roo99, Seg12, ZWL14, ADM14, BDK17, BNC96, BMSZ21, BZ14, BDZ15, BN05, CD17, CHI17, FFP23, FLT14, Gas23, HP14, Hör08, HWXZ24, LLZ17, LS23, PR05, PB15, SU03, ZX17]. asynchronous [MV18a]. attached [DMY00]. attachment [BEMR17, HM20b, MM20]. attainment [BLO19]. attraction [HG97, KZ21a, Zha18]. attractors [MZ17]. Augmentation [CK19]. Augmented [Hör08, ABH06]. Author [Ano07b, Ano08b, Ano09b, Ano10b, Ano11b, Ano12b, Ano13b, Ano14a, Ano16a]. autocorrelation [BSS05, Bor01]. autocovariance [BBS14, Bor01, Fas10, ICG16, WAP17]. autocovariances [HK08]. Automorphism [DZ17]. autoregression [EH02, FSTD20]. autoregressions [GQ23, ILY09]. autoregressive [ACLZ17, Ber01, DNV99, JWY22, KS97, LLS11, PS16, Tru19, Tsa09]. average [BPW19, KL05, LLT12, LTL07, Mar06, SW04, SN13, Tsa09]. averaged [CCZ13]. averages [KY22, PT03, Sur04]. Averaging [BYY17, CHR24, DT12]. Azadkia [SDH24]. Azéma [CC00, CH07, PSV00].

b [IMM09, HR22b]. backfitting [HMP18a, Xia09]. Backward [MR15, NS01, Ama13, BFT15, BVM07, BRS07, ELP24, GT16a, Ham03, LGW06, LS02a, Qiu18]. Bahadur [Kre24, Kul07]. balanced [New16]. Ball [EI10, CFKR16, GNSU19]. ballistic [BRS16]. ballisticity [IRVZ22, IRVZ24]. balls [BCG19, PFEP23, ZX24]. Banach [BZ19, Cun15, Min15, Yar19, ZGD24]. Banach-valued [BZ19]. bandits [Bul15, Yu17]. bands [Arm21, CDJ13, Düm03, FL11, PFEP23, PBD15]. bandwidth [BGL22, DM08b, DMT24]. Barak [MR19]. Bartlett [Myk01]. base [Ngu16]. based [ADH21, ACG20, BBJ08, BBG19, BS15a, BD14, BDL^{+}23, BW22, BL20, BS18b, CHP11, CS07b, CMMA20, CSvS07, DGU05, DV01, DF19, DL17, EM03, EH02, EM13, FL11, GP15, GGKL22, GPNECA23, HP02a, HPV13, HLL23, HWW11, IKP13, JJW24, KS99a, KLW18, KP23, KW13, KKP21, Kuj16, Lau13, Lua11, Lud00, MCC^{+}21, Mar13, MSS14, New16, OCBG19, OW17, PV15, PT18a, PTA07, PVD13, QT05, QLC21, RCBR00, RSS12, RS01b, Sch10, Sev04, TC24, VAD22, ÝSJ08, ZL03, ZHZL19, ZS21, dBM17]. bases [Dob03, HAT10, TA05]. Basic [DR09]. bathtub [JW09]. bathtub-shaped [JW09]. Baum [LS03]. Baxter [IKP18, MJK17]. Bayes [BL04, CS20, DF98, DRRS18, GMR19, Hof23, KS13a, MMW17, Ngu16, Riv06, WM16]. Bayes-optimal [Hof23]. Bayesian [Abr19, AH17a, AW24, AKT10, BES07, BGSS22, BDPR21, BG17, BSG22, CG22a, CW19, Cha23, DJL10, DR19, DZ17, DM19, FLP12, FGW12, FCH15, Fus05, GL95, GHN21, GP12, HP99, KSJ19, Kuj16, LLD19, LNP14, MRS20, Nau22, NJG20, NC24, NPM21, OL23, PR13, RSBG22, SG16, SRR24, YK20, YG19, Yu17, vdMvZ13]. be [Efr00]. Beckner [GZ21b]. Behavior [JKNP16, BLLS12, BRS16, BDT21, CFKR16, FFZ22, FGK16, GT16b, IKP14, JLM14, LM20, Yan24]. behaviors [Bru19]. behaviour [AI99, Bor01, BHVD17, CK11b, FM00, Fra06, Fra07, dBCAM22]. Benamou [HT19]. Benamou-Brenier [HT19]. Berg [AGH18]. Bernoulli [DLPZ12, Ano95c, Ano95d, Ano95e, Ano95g, Ano95h, Ano95i, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano97d, Ano97e, Ano97g, Ano97h, Ano97i, Ano97j, Ano98c, Ano98d, Ano98e, Ano98f, Ano98h, Ano98i, Ano98j, Ano98k, Ano99c, Ano99d, Ano99e, Ano99f, Ano99g, Ano99h, Ano99i, Ano99j, Ano99k, Ano99l, Ano99m, Ano99n, Ano00b, Ano00c, Ano00d, Ano00e, Ano00f, Ano00g, Ano00h, Ano00i, Ano00j, Ano00k, Ano00l, Ano00m, Ano00n, Ano01a, Ano01b, Ano01c, Ano01d, Ano01e, Ano01f, Ano01g, Ano01h, Ano01i, Ano01j, Ano01k, Ano01l, Ano02a, Ano02b, Ano02c, Ano02d, Ano02e, Ano02f, Ano02g, Ano02h, Ano02i, Ano02j, Ano02k, Ano02l, Ano03a, Ano03b, Ano03c, Ano03d, Ano03e, Ano03f, Ano03g, Ano03h]. Bernoulli [Ano03i, Ano04a, Ano04b, Ano04c, Ano04d, Ano04e, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano04l, Ano05c, Ano05d, Ano05e, Ano05f, Ano05g, Ano05h, Ano05i, Ano05j, Ano05k, Ano05l, Ano05m, Ano06c, Ano06d, Ano06e, Ano06f, Ano06g, Ano06h, Ano06i, Ano06j, Ano06k, Ano06l, Ano06m, Ano06n, Ano07c, Ano07j, BGH01, DDW13, Gne04, JK23, Kar24, Ken95, MMR23, NMR20, NLS23]. Bernoullicity [HJL02]. Bernstein [Bar10, BC19, KM11, Mau19]. Bernstein-type [Bar10, BC19, Mau19]. Berry [BGT97, BGL24, CH08, HLP12, Jir23, LS24, Rai19, SZ22, Tur04]. Besicovitch [Man07]. Besov [BDKS20, GT09, Mor99, Riv06]. Bessel [BL23, DH17, DD15, GJY03, GM19, dV13, HS97, JW13, YZ01]. best [HR22b, Ose20]. best-arm [HR22b]. Beta [BHPZ15, BS15c, BW08, HR16, IMM09, Jia13]. beta-Jacobi [Jia13]. Between [GJS10, NRT98, Wan24, ACJ18, Alm03, AIS19, BP11, CLZ07, CG99b, DPSW10, Kur16]. Beyond [PB15, JW13]. Bias [HMW01, FLT14, Gau17]. biased [BGI18, Lin19, Olo09, PT15, RA17, SSY21, Van95]. biasing [BZ22]. BIC [BCFH22]. bifractional [TX07]. bifurcating [PHO17, PEBG17]. bilinear [HLPŠ15]. bin [MR19]. Binary [You05, BWW00, KMS^{+}99, MS18, SV13, ZS12]. binned [BK02]. binomial [BGX15, ČV10, HR16, MH10, PV07, Roo99, XZ09]. bio [VV04]. bio-medical [VV04]. birth [ALS09, CJ13, FM00, Hub12, OP10, OP11, YY24]. birth-and-death [YY24]. bisexual [MMR06]. Bivariate [DDFD04, AFGS08, GKNY11, GV19, GP12, HT00a, KFT21, KW97, ZX17]. BJ [Ano95b, Ano95a, Ano95f, Ano97b, Ano97a, Ano97f, Ano98b, Ano98a, Ano98g, Ano99b, Ano99a, Ano05b, Ano05a, Ano06b, Ano06a]. block [AH17a, BC06, BS18b, CK23, CDH^{+}98, Dom15, DF19, GP18, HLY20, MM15, NL14, Sch20b, Syn07, TGL23]. Blocking [FLS01]. blockmodels [TCP22]. blocks [DN21]. blockwise [Rig06]. Blumenthal [Sch13]. blur [IS98]. Board [Ano09c, Ano09d, Ano09e, Ano09f, Ano10c, Ano10d, Ano10e, Ano10f, Ano11c, Ano11d, Ano11e, Ano11f, Ano12c, Ano12d, Ano12e, Ano12f, Ano13c, Ano13d, Ano13e, Ano13f, Ano13g, Ano13h, Ano14b, Ano14c, Ano14d, Ano14e, Ano15a, Ano15b, Ano15c, Ano15d, Ano16b, Ano16c, Ano16d, Ano16e, Ano17a, Ano17b, Ano23a, Ano23b, Ano24a, Ano24b, Ano24c, Ano24d, Ano95c, Ano95d, Ano95e, Ano96b, Ano96c, Ano97d, Ano97e, Ano98c, Ano98d, Ano98e, Ano98f, Ano99c, Ano99d, Ano99e, Ano99f, Ano99g, Ano99h, Ano00c, Ano00d, Ano00e, Ano00f, Ano00g, Ano00h, Ano01a, Ano01b, Ano01c, Ano01d, Ano01e, Ano01f, Ano02a, Ano02b, Ano02c, Ano02d, Ano02e, Ano02f, Ano03a, Ano03b, Ano03c, Ano04a, Ano04b, Ano04c, Ano04d, Ano04e, Ano04f]. board [Ano05c, Ano05d, Ano05e, Ano05f, Ano05g, Ano06c, Ano06d, Ano06e, Ano06f, Ano06g, Ano06h, Ano07c]. Bochner [US24]. bodies [Bru20]. bold [ERS19]. Boltzmann [FM02]. book [EF04]. boosting [PLH09]. Bootstrap [FKM02, Fus05, GP23, KKP21, SW12, TGL23, Woo00, YDN23, Ber97, BC06, BB99, BD13b, BK16, Büh97, CDH^{+}98, DL19, FKA05, GP18, HMW01, HLPP09, HB23, Leu12, MJK17, MP23, NP08, NL14, PGV17, Syn07, WP22]. bootstrapped [TW14]. Bootstrapping [JW19, JK21, LEM23]. border [JT22]. Borel [GL06]. Boris [Ano97c]. Borrowing [Ngu16]. both [CP12, HPY18, PTL04]. bound [BD15, BGL24, CF15, CRT21, GL95, HMP18b, Jir23, Lua11]. boundaries [EK11, GJZ23, YS22]. Boundary [CCJM24, AAL23, BHM18, BVM07, Dan96, Dan00, DLR06, DNS19, FN97, FGS22, KM23, Kee13, RS17, RW19, XY23]. Bounded [BGI18, AHS15, AGM06, BZ23, GLT23, Mau19, PT03, Sur04]. Bounding [DSW21]. Bounds [AR17, GR23b, PTA07, APP11, AP14, Arm21, BRR19, BGMZ18, Bel17, BGT97, BM98, Cas14, CD21, CKHM23, Cha03, CL19, CBL22, CH08, CJL^{+}23, CDG14, ER23, GL22a, GW17b, GS20c, Han22, HNW14, Kim14, KL17, Kre21b, KL13, ŁMN13, LM10, LW17, LS24, MS10, MFWB22, Nov14, PJT19, PRR13, Roo99, San10, Sau19, SZ22]. box [ABH08]. branch [LW23]. Branching [Jag95, Mal17, Ath00, AAI11, BGMSB14, BV17, BHM18, Bha23, BR18b, BGT07, BHM22, BGL24, BD22, CLR14, FB14, GL18, GMD05, GMD06, Gor96, JS08, KS01, MS23a, Mar19, MMR06, MO08, Olo09, Pak99, Pro24, Rév96, Tka20, Wak95]. break [WS23]. breakdown [BG95]. breaks [BHVD17]. Bregman [Ovc18]. Brenier [HT19]. Breuer [HNTX15]. Bridge [CR17, BP11, GS22]. Bridges [Bra14, BS14, BMS21, CKM20, Cor14, FLT14, dV13, SvdMvZ17]. broad [DPX23]. Brownian [Coe06, LP18a, AD18, ADM14, ACPM24, Ass23, BM15a, BP11, BP06, BJ00, BNCP11, Bar22, BŁ15a, Ben03, BP10, BR12, BPR12, BDP23, BDG07, BS18a, CI21, Che01, Coe05, CRS00, Dan00, Das05, DS06, DS05, DS19, DP11, DMY00, Dor98, DW19, ES99b, FR06, FR19, GS22, Gra21, GV19, HS97, HWXZ24, JLT08, JM09, Kat09, KJ13, Kee13, KL02, LS19a, LD11, LS02b, LS20, MWX08, NVV99, NPS20, Obl06, PR16, Pec03, Pey17, PT01, Pit97, Sau12, SW19b, TX07, VY13, VVV06, Van05a, WYZ21, Wer95, Zei15, ZX24]. BS [CS13]. BSDE [EI18, PX10]. BSDEs [BR18a, BM09, BEM18, BGGL21, CS13, DM24, GS16, GLL20, IR20]. buffer [RS01a]. buffered [BD20b]. builder [IRVZ22, IRVZ24]. Building [BM13, FSW17]. built [KZ21b]. Bump [EMW18, EMPW20]. bundled [ZWY17]. Burkholder [BS15b, SS18, Sio18]. Busemann [ARAS20]. bushes [AMV18]. bypassing [GK12].

c [EBR23]. c-optimal [EBR23]. Côtes [NS07]. Cahn [CW01]. calculations [DKM16]. calculus [BM08, BP10, FO24, Gob01, Tud04, Wal12]. calibrated [GT22]. Calibration [Tra14]. Campbell [DL09a]. Can [Efr00]. canonical [GS13, HMWY10, HR15, ML20, MY23]. Cantelli [DR19]. Capacity [LV15]. Carlo [ARV18, AJDD11, AR05, Atc16, BB22, BPR^{+}13, BBLG17, BRE23, DS03, DDJ12, Dia13, ELP24, GC17, GLO22, GRS11, GT01, HVM99, Kam18, LBBG19, NIR21, OCDM08, PJT19, SZ03, VFJ18, WLH16]. CARMA [FK17, SS12]. cascades [RSGW03, TV99]. cascading [FFZ22]. Case [CGMST95, Bac18, BCM^{+}21, BHM18, BR96, BRW00, CR22, DVSS16, GZ09, LS99, Ma10, MV18a, PR02b, TT20a]. case-control [BRW00, Ma10]. cases [AGM06, Ber01]. catalytic [Fri24]. categorial [GGT22]. categorical [DLT09, ML11, NC17, Tru20a]. Cauchy [AH13, KJ13, KM20b, LP18c]. causal [Yan24]. Cayley [JHD20]. CCA [HPY18]. celestial [CCF^{+}20]. cell [HKM^{+}16]. censored [BL10, JV01, KVD04, LT13, Pon07, VV02, VD13]. censoring [Akr00, EFVG08, LPV13, RA17]. censorship [Bel98, Zha96]. Central [BST12, CL12, EK11, JS05, JS06, KLLM13, LR16, LTX22, NP08, Nic18, PT08, PZ19, PT03, RT15, Tho23, YZZ23, AP19, Akr00, ALOV19, Beg07, BJSZ07, BB99, Bra99, CLD08, CF15, DJ97, EO07, FMPV14, GP11, GV19, LRMT17, MP10, Mou13, Per02, PTA07, SW08, Tur04, BLP19, PDL19]. century [Tay13]. certain [AW96, CS24, LS03, LS02a]. CEV [BO18]. CEV-like [BO18]. Chain [Atc16, ML11, AJDD11, AR05, DS03, Dia13, Drt09, GT01, HVM99, HMS04, HTL07, Kam18, KT15, KSW01, MS10, NC17, NC21, PT18b, RZ13, SZ03, VFJ18, WK21]. Chains [CCN^{+}21, AI99, Ath00, BC06, BP19a, BC00, CJL17, CA19, Dor97, DMS07, DDL23, FS18, FGW12, FMPV14, FP17b, GMM05, GW99, HM11, HLP12, HK11, HM23, Hub12, Jou07, KM05, MV18b, PHO17, PEBG17, RS18b, San13, Sch01a, Tru20b, WCM06]. challenges [FZV23]. Change [Ath00, KL00, NLL18, Bei97, BW22, CLLLR17, EK18, Fer01, GR00, HS10, IS98, KL09, LL00, Nku21, PR02a, RA17, VW17, WYR21]. Change-point [KL00, NLL18, Bei97, BW22, Fer01, IS98, Nku21, VW17]. change-points [CLLLR17, LL00]. changed [FLZ21, LS16a, Wan23b]. changes [AHHK08, CN15, Küh19, Sch24]. channel [DS03]. chaos [BT20, MCSS99, Pan22, Pec03, PZ14]. Chaotic [PSV00, BDT21]. Characteristic [ZGD24, AK23, Bah21, CPY13, Sch20a]. characteristics [IKP13, Sch01b]. Characterization [CCN^{+}21, GKT24, HVM99, HNR23, KP23, KT18, PV07, ZP11, AG07, BC00, Cha12, DLS14, FP17b, GS22, KM06, KS20, Obl06, Wan17].

Characterizations [AM07, BW08]. Characterizing [Fri24]. chart [New16]. Chatterjee [ADN24, SDH24]. checks [VD12]. Chen [Hir97]. Chernoff [AP14, BW14]. Chernoff-type [AP14]. Chervonenkis [AN12]. chi [GR23b]. chi-square [GR23b]. Chinese [RW22]. choice [BNW98, DJL10, ERS19]. choosing [HP01]. Chung [ADS13, LX23, NRT98]. Chung-type [ADS13, LX23]. circle [AM16, FK21, JMV20]. circular [AAG24, BS23, HKM^{+}16, JS13b, KK16, KJ13]. circulas [AAG24]. claim [BBS15]. class [AH13, BV17, Bat17, BBS15, BS02, BB17, BDPR21, BBD17, BH96, BW10, CHP11, CK19, CD23a, CDH21, CH07, DDT14, EO07, FGW12, GPNECA23, Gay20, Gay22, GP23, GMM05, JMM17, Jam10, KS21, Lau13, LX23, MPS98, MPAiS12, OCBG19, PS11, PS13c, RVW01, San13, Sch13, Sch09, Ste13, WC14, WAP17, WP22, ZP11, ZWY17]. Classes [RS96, AN12, BNMS06, Ber04, GS18a, JLP10, PT01, PR23]. classical [DJ96]. classification [CDH21, GGM20, KS16, Lec07, Mei16]. Classifiers [TV06, PV15]. classifying [MW05]. clinical [HZ04]. Cliques [BCvdH20]. clock [YZ01]. closed [ACV09, SSH14, Vil14, Vil23]. closed-form [SSH14]. Closeness [Roo10, LWCS22]. Closing [HKK19]. clouds [BEN12]. CLS [IKP14]. CLT [BWZ10, BC16, CP15, JB21a, KMB21, MPU19, ZBY17]. Cluster [ACG13, HJL02, RNS23, Wu09]. clusterable [BBW22]. clustered [CW22, CZH14, MSW13]. Clustering [RZ13, CY21, MR13, NMR20, RNS23, vDD21]. clusters [CS24, PT21]. coagulation [Ald99]. Coalescence [CH18, Ald99]. coalescent [LS19a, Lim19, Möh06, Rát19]. coalescents [IMM09]. coefficient [ADN24, BRS07, DKS01, Erh14, FH05, Hof99, HM20c, IMS01, LMP12, SDH24, WFM10, XW12, ZLL17]. coefficients [BKMP09, BJM21, CGCR07, GG22, GLT23, Ham03, IKP18, KHLN97, LTC14, MY23, PS20, SS15, XY23]. cointegrating [WC14]. collecting [Mla08]. collisions [BW23b, BS18a, IMM09, Möh06]. color [SFP23]. colors [BT17b, SFP23]. coloured [MPV07]. combinations [MR17, PXH13]. Combinatorial [NLS23, CF15, HR16]. Combinatoric [GKP12]. Common [HPV13, LMV14, LS23]. community [BCvdHV21]. commutation [CJ13]. commutative [DS19]. compact [Bor21, Cun15, GS22, LBDM19, OSZ18]. compactly [BGMZ18]. Comparing [DW23, Kim21]. Comparison [CS13, DCC14, FS00, WYZ24, DLL20, EM18, GNSU19, HXZ18, NJ14]. comparisons [GRS11]. compatibility [LR21]. competing [BS18a, IK22, LC22]. complete [CMr19, Nak15, Obl06, PT01, PS11, PS13c, SS12]. completely [BWJ18, LC22, LMC23, LNP14]. completion [CKLN18, Klo14, KLTZ19]. complex [Ass23, dBCAM22]. Complexity [MN22, BP19a, Lee08, TV06]. component [BD18, DR11, LMMR16, QJ22, ST04]. Components [HPV13, BBW23, DR16, HHH24, HK13, HMRB08, KW20, KSX19, OL23]. composite [BP19b, KF14, NJ14, RK10]. Compound [BP21, BP03b, Bou06, ČV10, Sch99, UDS23, BX00, BV10, But17, CD11, Roo05]. Computable [DMS07]. computation [AEP11, BP06, Jor13, LP02]. computations [Myk01]. Computing [Cut00]. concave [ABD14, BP11, BW14, BS21, BGI18, CG22a, CEG17, DRD20, DR09, MRS20]. concavity [BD18, HR22a, HJ17, Yu10]. Concentration [BM15b, BHBO17, BZ19, BG10, CJL^{+}23, DDL23, GT18, GS20c, HP02b, HMRB08, KL17, LMZ20, Mar18, Ver20, BGI18, BŁ15a, BH00, CS07a, DRRS18, GNSU19, GS20b, HT21, LvdG14, Lei20, MMW17, Mau12, Ngu16, PEBG17, Roo05, Sau19, SV13, SM23, WW14, YP23, Zha22]. concept [Möh99, MSS14]. concern [LW04]. concomitant [Che02]. concomitants [GGT22, JN95]. condition [BD09, FN21, GGM20, GLL20, Hei18, MS23a, Olo09, Riv05, Riv07, Sto00]. Conditional [BC19, BL10, FB22a, GKP12, LMV22, LV17, Loc12, dCCGZ23, ALV18, ALR23, BL14, CCJM24, CL02, CPY13, CP05, DM08b, FZ22, Hoo16, KFT21, KK17, KSS21, KL10, KLS98, LS19a, LS19b, Mal09, Pak99, SDH24, SL18, SC11, Tor17]. conditionally [BL15b, Röl05]. conditioned [BL14, Bra14, DW20, FV03a, FGK16, OH23]. Conditioning [DR11, Cut00, Kos02, LN02, Mal09]. Conditions [Erh14, Mal16, BRS16, BVM07, BW19, CG23, CK15, CA19, GS16, GZ21b, JT03, Jir23, PTL04, RS01b, XY23]. cone [VY13]. cones [DR11]. Confidence [CDJ13, CP01, LR95, PBD15, Arm21, Atc16, BN20, Bel21, BP19b, CCF22, CKLN18, DR23, Düm03, EHS05, FL11, FSW17, HS07, MH10, Qia21a, Ren08, SW12, TZ17]. Conformal [FZV23]. conjecture [HJ17]. conjugacy [BWJ18]. conjugate [Lau13]. Conjunctive [BES07]. connected [KW20]. connecting [CH07]. Connection [DFK16, ST13, LP24]. connections [CG99b, GS11]. consensus [ÁEdBCAM18]. consequences [KFT21]. conservation [JMW05]. conservative [HK11, RSGW03]. considered [Lau04]. Consistency [CLD08, CS07b, DF98, ES99a, Hoo16, HK13, KP23, KSJ19, MPS19a, Ped95, Syn07, WM16, AAGGR03, BMT17, CW19, DR11, DJL10, Dom15, DM08b, DR09, DL17, Gay20, Gay22, GQ23, HY08, Leu12, LLD19, LPW04, NPM21, TT06, VFJ18, VS05]. Consistent [BBD20, CK19, Leh19, RP23, Sch20b, Van05b, WH10, vdMvZ13, AH17b, BKK23, BD14, BP19b, GK12, PV15, Tru19]. constancy [Möh18, Pap09]. constant [Har17, Jor23, KN99, Ose20, PT18b, SW19a, SS18]. constants [BX00, DY14, Rai19]. Constrained [GZ17, FG19, KS13a, PG21, TCP22, VZ23]. constraint [BEM18, LWLL22, PX10]. constraints [BR12, CGS18, HNR23]. Constructing [CSSC95]. Construction [AKT10, HLS19, ST20, LMC23, PR23]. constructions [Haf13, MY02]. Contact [FGS21, AMPV10, CD21, FV03a, FV03b, Mac23]. contamination [LP99]. content [RS01a]. Contents [Ano07g, Ano07h, Ano07i, Ano08g, Ano08h, Ano08i, Ano08j, Ano09g, Ano09h, Ano09i, Ano09j, Ano10g, Ano10h, Ano10i, Ano10j, Ano11g, Ano11h, Ano11i, Ano11j, Ano12g, Ano12h, Ano12i, Ano12j, Ano13i, Ano13j, Ano13k, Ano13l, Ano13m, Ano13n, Ano14f, Ano14g, Ano14h, Ano14i, Ano15e, Ano15f, Ano15g, Ano15h, Ano16f, Ano16g, Ano16h, Ano16i, Ano17d, Ano17c, Ano20, Ano21a, Ano21b, Ano21c, Ano21d, Ano22a, Ano22b, Ano22c, Ano22d, Ano23c, Ano23d, Ano23e, Ano23f, Ano24e, Ano24f, Ano95g, Ano95h, Ano95i, Ano96d, Ano96e, Ano96f, Ano96g, Ano97g, Ano97h, Ano97i, Ano97j, Ano98h, Ano98i, Ano98j, Ano98k, Ano99i, Ano99j, Ano99k, Ano99l, Ano99m, Ano99n, Ano00i, Ano00j, Ano00k, Ano00l, Ano00m, Ano00n, Ano01g]. contents [Ano01h, Ano01i, Ano01j, Ano01k, Ano01l, Ano02g, Ano02h, Ano02i, Ano02j, Ano02k, Ano02l, Ano03d, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano04l, Ano05h, Ano05i, Ano05j, Ano05k, Ano05l, Ano05m, Ano06i, Ano06j, Ano06k, Ano06l, Ano06m, Ano06n, Ano07j]. Context [NC21, KLLM13]. Context-specific [NC21]. contiguity [FY20]. contingency [HAT10]. continued [LS95]. Continuity [FMSY12, GJQS20, BGMSB14, BS21, BL15b, DPX23, FM02, FP10, FFS10, Hie17, LX23, LWX15, Nie13]. Continuous [BZ14, JT22, BS02, Bei97, BPW19, BHVD17, Car06, CR11, Deh05, DS96, FP17a, FB14, GLT23, GLL20, GW98, GHL05, GZ17, Ham03, JLT08, Jou07, LPSS18, MMS08, Pon07, Røy11, San10, SS12, SK23, TKT09, TC03, Tsa09]. continuous-state [FB14]. Continuous-time [JT22, BPW19, Deh05, GW98, GHL05, GZ17, Jou07, Røy11, SS12, TC03, Tsa09]. continuously [DW20]. continuum [BDR14, GT19]. contour [LR95]. contours [LR95]. contraction [Abr19, ADH21, GHN21, KS18, Ngu15, YG19]. contrast [BM98]. control [BIPZ22, BDR14, BRW00, GZ09, GMD05, Hof23, Ma10, SC23]. controlled [FLT14, GMD05]. Convergence [AH17a, BY05, BHPZ15, Bar22, Bra22, CP12, CP05, Del17, Dio12, FHJK20, FPW17, GK03, GC17, IMM17b, JM17, Lei20, Mac23, MM15, NL14, OCBG19, OB20, QJ22, VVV06, WW14, Yan08, dBM17, AP19, AKT10, ALOV19, AZ15, BD09, BBV17, BCPR09, BM98, BK02, BCG14, BBC12, BO18, BJM21, Bou15, BGGL21, CMS10, CK11b, CHR24, CS13, CS99, CP23, DPTV15a, DR19, DF20, DMS07, EO07, Erh14, FR15, FPSS19, FFP23, FRW04, GT09, GMD05, GT04, HMP18b, HW19a, HHH24, HNTX15, IS19, Jir18, JS05, JS06, KFT21, KK23b, KMW14, KT18, KS01, LGW06, MS10, MS23a, MMR06, Mou13, RWZ17, RT96, SW04, SW13, SXX22, Van05a, Var19, WC14, Wan22a, WB19, ZH20]. converging [CP12, KM05]. convex [AS19, BDK17, BG09, BEM18, Bru20, BKY21, CS07b, CK01, EGG03, FG19, FMNW18, JW09, JM22, KM98, Lec13, LV17, LP18c, Mar18, RK10, Zha13]. convex-constrained [FG19]. convexity [MFWB22]. convoluted [BM08]. convolution [BM05, BC09, FMNW18, GST08, JW21, LW17, SW13, Tan06, Ued21]. convolutions [BGMZ18, BL23, BMMS12, BB17, BOJGMR15, BLM19, Roo10]. copula [BBV17, BK16, FRW04, GCP17, GNR14, GP12, Haf13, KW97, RWZ17, Seg12, WZ16, ZGV20]. Copulas [Lag10, BD13b, CEL20, DHKV15, FRW15, GKNY11, HNW14, KFT21, WZ16]. corrected [Mat24]. Correcting [NS07]. Correction [DLPZ12, Fra07, OKM^{+}98, HMW01, WXZ19]. corrections [Ano96a]. correctness [DL19]. Correlated [LPSS18, BDT21, HLS14, Len11, LRMT17, MH22]. Correlation [Kre21b, ADN24, BMP08, DMMW18, DKS01, DEGR14, HL17, HR15, LWY24, ML20, MY23, WZ16, YZZ23, ZHZL19]. correlations [ACBL15, FV03a, GS13, MY23]. correlogram [DM09]. correlograms [BP98]. Corrigendum [BRJN19, BMS21]. count [FSTD20, GNR14, HLZZ23, LSY^{+}22, Neu11, PV07, SGL^{+}24, YKK21]. countable [AI99, Zha18]. counterexample [HL06]. Counting [PW04, ER23, GLS07, RNS23]. counts [BHBO17, DGP18, HMTvB24, LP24]. coupled [HLS19]. Coupling [HJL02, Ken15, SSW12, Tru20a, Wan11, BDV21, Bou06, FLS01, Hir97, KMM24, LSW20]. couplings [BGI18, Gau17, LV17]. Covariance [LP18b, WCM06, APP11, Bac18, BWZ10, BLP19, Bat17, BTT19, BD14, BGH01, BK23, BPD23, BX15, CM13a, CP12, CP15, DMM^{+}20, Dia23, DRW20, ERS20, HY05, HM19, HP02b, JMRX18, JB21a, JJLW22, JR17, KTU18, KL17, KR24, KLP21, LW18, LW22, LGS08, LWYZ22, LEM23, Lou14, MW20, MGS03, OR24, PTA07, PY01, SW19a, Sch10, SvS17, SPP21, TRB24, VBLN20, WHP18, WYR21, Yin22, ZSC22]. covariances [EGG03, ZP11]. covariates [Kur16, LPV13, SGL^{+}24, dUÁV21]. covariation [FPS95, GV19, Koi16]. Coverage [CCF22, CG22b, Hof23, YK20]. covering [FK21]. Cox [CHWW05, GW98, Kre21b, MPV22, NC24, PB15]. Cox-processes [Kre21b]. crackle [OB20]. Cramér [BL22, BGL24, FGLS19, FLS20, FLS23, GJZ23, GL95, HL06, JWY22, Len05, LSW13, LTX22, Lua11, Riv05, Riv07, SZ16, Zha23]. Cramér-type [FLS20, Zha23, FLS23, JWY22, LTX22]. Crandall [CR22]. Crandall-Lions [CR22]. credible [CS20, YK20]. Criteria [DSS10, CdLOR20, FK17, NJ14]. criterion [BS15a, FM02, MYT06, Mat24]. Critical [Nic18, BWW00, CK11a, CTP21, CD21, CS18b, Gor96, IKP14, Mal17, Rév96, RZ23, TK03, ZX24]. criticism [GLF16]. Cross [HP01, MYT06, PFEP23, ZDY23]. cross-validated [MYT06]. Cross-validation [HP01, ZDY23]. crossing [Dan96, Dan00, HWW99, Kee13, Pey17]. crossing-time [Dan96, Dan00]. Crump [BGMSB14]. cube [KPT21]. cubic [Stu23]. cumulants [BO20, ZS12]. cumulative [AAG24, BBC12]. cure [MPV22, PGV17]. current [Cha23, DPTV15a, PR06]. current-status [PR06]. Curvature [FS18, GZ21b, Jou09, SCK23, YP23]. curvature-dimension [GZ21b]. curve [CHT98, HP01, HWW99, HY08, HYY11]. curve-crossing [HWW99]. curved [Dan96, Jou07]. curves [Düm03, LP18c, LR95]. CUSUM [BHS11]. cut [ACPM24, HW19a]. cut-off [ACPM24]. cut-trees [HW19a]. cuts [CK00b]. cycles [BMSZ21]. Cylindrical [CGMST95].

d [Tho99, ZZ15]. DAG [ER19]. Dagum [BMP08]. damping [BW23b]. Dantzig [CCL23, Kol09]. Data [CK19, Abr19, ABF12, BCM^{+}21, BCPR09, BZ12, BK02, CLZ18, CDJ13, CDH^{+}98, Cha23, CL24, CY21, CZH14, CX24, CC19, CK00b, CGER07, DS03, DT23a, Dre03, ET23, GGT22, GP15, GNR14, GC05, GT22, HTW23, HHH24, HRY13, HK13, HLZZ23, JV01, KLW18, KW13, KK23b, KM24, LN14, Lee08, LT13, LMNV11, LSY^{+}22, MSW13, MMS08, ML11, Mar13, Mei16, NC17, PR06, PV07, SG16, SGRV18, SGL^{+}24, VV04, Van95, VD13, WS23, Xia20, YKK21, ZS15, ZWL14, ZWS21, dUÁV21]. data-based [KW13]. data-generating [Lee08]. Dating [WS23]. Davis [BS15b, SS18, Sio18]. Dawson [ALR23]. Dawson-Watanabe [ALR23]. DDSDE [CHR24]. De-biasing [BZ22]. death [ALS09, CJ13, Hub12, OP11, YY24]. decay [BKT19, CJL17, HW19b]. Decaying [FV03a]. decision [GZ17, Ruk14]. decision-theoretic [Ruk14]. decomposability [BNT02]. decomposable [Dob03, ER23, HAT10, Tra14]. decomposition [FY20, Hös05, KJ20, ZZW22]. decompositions [Yar19]. decompounding [vEGS07]. deconvolution [BTT19, BM05, DRT16, Gol99, JS13b, MNN^{+}22]. Deep [FFP23, KK21, KK23b]. Defective [KM22]. defined [BGMSB14, Hir16, WC98]. definite [Gne13]. Definition [BBS15, PS09]. deformable [AKT10]. deformed [Din20]. Degeneracy [Muk20]. Degenerate [IK22, Leu12, Bra22, HLT17]. degree [BC20, Lei19]. Degrees [PP22, BZ22, Jan07, Muk20]. degrees-of-freedom [BZ22]. Delaunay [BC20, CH18]. delay [BR12, DNR20, FR06, GK99, GK01, KS13b, Kut21, MMP00, Rei05]. delayed [GS08a, HZ04]. deletion [Rát19]. delta [CGER07]. denoising [Din20]. dense [FR15, YS22]. Densities [HLT17, ACV09, BEN12, BHL99, CG22b, DM21, FL11, GMZ21, HL12a, JS13a, JP15, KM05, LWLL22, LPW04, MS23d, SW04, Vil14]. Density [DM08a, HLT01, HNTX15, Sar23, VV02, AF19, BW14, Bel98, Bel17, BK02, BL10, Bra02, BN05, BM05, BDKS20, CD17, CW01, CCJM24, CKHM23, CCFH00, CC04, CGK^{+}20, CD11, CM14b, DRD20, Dan96, Dan00, DR09, DSW21, EH02, Gay01, GCP17, GM04, GN08, GN10, GL22a, GL22b, Hei18, JLL04, KPT96, KN99, KMS17, Lau97, Lei06, LW17, MPAiS12, MRS20, MCSS99, MYT06, MR16, MSS04, Mor99, Nau22, OU23, PRvS13, Qia21a, Reb15, Rig06, RV09, SW13, SG16, Tin98, Tru19, VSV03, VS05, Vil23, XL10, YKK21, Zha96, vEGS07]. Dependence [Hör08, APGMV04, BGMZ18, BS12, Buc18, CS24, CAA22, DLPZ09, DLPZ12, DDFD04, EKS08, HT00a, HP02a, HNZ18, JD16, Jir23, KN06, KKP07, Kre21b, Lag10, LS19b, Leu12, PQY13, PS16, SY24, SDH24, SDS24, VW17, ZC18]. dependence-function [HT00a]. dependency [EMPW20]. dependent [AW12, AEP11, AAI11, BW23b, BCM^{+}21, BGT97, BCPR09, BW22, BZ12, BZ19, Bou06, BV10, BK16, BST12, CDH^{+}98, CX11, CR13, CSvS07, CR22, CM99, DNR20, DDT14, Dre03, Gan21, GWY13, HTW23, HW24, JMM17, KY22, KK23b, Kul07, LR16, LW04, LNP14, Lud00, MPU19, MR21, MMR06, MZ17, PBD15, RZ21, RZ23, Sev04, Yan01, ZS15, ZWL14]. depending [LPV13]. depth [CC19, DGC11, Gao20, GC05, HLPŠ15, HWW11, Mas04a, PV15]. depth-based [PV15]. depths [NHV^{+}21]. Derivation [AG99]. derivative [MS23a, QT05]. derivatives [HN10, Lau97]. derived [KM20b, OH23, Pit97]. Deriving [BZ12]. descent [SZ22, SK23]. description [SPP21]. Design [EBR23, BBG19, Bir04, CHT98, CM99, HYY11, KP04, Kur22, MS18]. Designs [ST18, DLPZ09, DPW09, DPSW10, DLPZ12, HLS19]. Detecting [ACBL15, ACBLV18, BBW23, BCvdHV21, Sch24, BHVD17, GR00, JN15]. Detection [BI13, ACG13, ADN24, Bar02, Bei97, Cas14, CT19, CCF^{+}20, Dit19, EMW18, EMPW20, LMMR16, LMZ24, YS22]. determinant [BPZ15]. Determinantal [BTUA23, MNPR18, BAT19, BL16, BCG19, Qiu19]. determinants [DRW20]. determined [Lin09]. Determining [Mal09]. Deterministic [Ald99, BB22, CA19, KS21, PSS24, PT01]. development [BC16]. Deviation [AK18, Bru20, JWY22, BGL24, BW19, DL13, EX23, FLS20, FLS23, Fis14, GO00, GM04, HP02b, Jou07, Jou09, LTX22, Mel00, NZ05, PS98, SZ16, Sko96, Zha23]. deviations [AS03, ALL09, BNW98, Ber01, BD20b, Bha23, BDM10b, CG99a, DDK22, EZ03, FGLS19, GKP12, GNR19, GZ18, GZ21a, GJZ23, GY04, GT18, HM20b, HM15, HNZ18, JWY22, Kra21, KMB21, KZ21b, Lei06, LSW13, MR21, NR00, Pan22, Sid18, TW14, XZ18, YZ01, ZZ15]. device [BL22, HL06]. Diaconis [Din98]. diagonally [KM20a]. diagonals [BT17a]. diagrams [Blo04, BH10, HM15, OB20]. Dickman [BG19]. diffeomorphisms [BDM10b]. difference [Alm03, BP11, BGL22, GMD06]. differences [GT16b, vDD22]. differencing [IMS01]. different [CL05, Kur16, MYT06]. differentiability [CCR20, FMSY12, Hir16, WSX20]. differential [Ama13, AKH17, BM15a, BYY17, BGSS22, BE16, BG17, BSG22, BVM07, BRS07, BDKS20, DNR20, DDBM21, FR06, FK11, FN21, Fou01, FFS10, GT16a, GK12, GK99, GK01, Ham03, HDS22, JM09, KK23a, KHLN97, KS13b, LWLL22, Lau04, LGW06, LS02a, MMP00, Mar03, MSS04, Mor99, NS01, PS20, Pet97, Rei05, RTT20, Sar23, Sau12, SU03, SXX22, Tka20, Wan16, XZ18, Yar19]. differentially [Osȩ09, YRZ24]. differentiation [HKM^{+}16, TC03]. Diffusion [BSS05, EHHS23, AG01, BR07, BPR06, BC21, BS95, BS14, BMS21, BJ22, BN05, CGCR07, DV18, DJ97, DH02, FM00, FLLO14, Fri18, Gob01, HY05, Hof99, JW21, KS99a, KLW18, KMW14, KL13, Kut97, KY07, LP02, LS21, LLBF24, McG09, Ogi15, Ogi18, OU23, OPPS16, Ped95, PY99, RS10, Sau14, SvdMvZ17, Sør02, SU03]. Diffusion-type [BSS05]. Diffusions [IKP13, Abr19, AHS15, ALR23, BR97, BŁ15a, BS14, BMS21, BPD23, CK11b, CCTY19, Cho18, DRD20, Deb99, FP17a, FO24, GP15, Han03, IS19, Jac06, JS17, KM05, KSJ19, KT18, LPSS18, MFWB22, PR05, PY03, PJR16, RSBG22, Sch13, TT20b, TY23, Wan17, vdMvZ13]. diffusive [ELL22, MR16]. diffusivity [GR23a]. digital [JT22]. Dimension [ML20, BGMZ18, BH00, CP15, Cut00, DH17, DVSS14, DVSS16, Dic16, FB22a, GZ21b, Giu18, Gor96, JLT08, Kam18, Man07, QSSS04, SW19a, Sau19, SXXZ18, WW14, WM16, YLC19, ZWS21, ZS21]. dimension-free [BH00, Giu18]. dimensional [ARV18, AP19, AMPV10, ACPM24, BLP19, BCFH22, BC13, BTT19, BS16, BBW23, BJ22, BRS07, Büh13, BI13, CM13a, CGZJ22, CP19, DR16, Din20, DM19, EBR23, FP10, GS23, Gho99, GT22, Gra21, GR04, GS11, GQ23, GKT24, HTW23, HM19, HWZZ23, JCL20, JB21a, JMW05, JJLW22, JR17, KKP21, KMB21, LYG19, LW18, LAP20, LWYZ22, LWY24, LLL22, Lou14, MY23, Mac23, MMW17, NK21, New16, PV16, PR05, PY99, PY03, RTT20, SvdMvZ17, SG16, SL18, SvS17, TT20a, TRB24, VV04, VG18, WS23, Wan23b, WH10, Yin22, YZZ23, ZC18, ZSC22, Zha22, ZBY17]. dimensionality [RS01b]. dimensions [BH17, BGT07, CK11a, Car06, Erh14, GS20a, HP02b, KP23, LS19c, LEM23, MS23c, MSS00, NPS20, SX10, Tho99, WAP17, WYR21, YK20]. direct [Kur16]. Directed [CSY03, Cor16, GK18]. Directional [CCR20, JRSK21, LN14]. directions [PRV20]. Dirichlet [BR18a, BDL^{+}23, CS99, DRRS18, GO00, GS13, GT01, Han09, Hir97, Jam10, Lau04, LP18c, MR16, Ngu16, Pec08, Sie07]. disconnection [Wer95]. discontinuity [Tod20b]. discontinuous [Gra21]. discounted [GHL05]. Discovery [BDR14, SC23]. Discrete [DGP09, Drt09, DMS21, AK23, BDK17, BP03a, Bar18, BP10, CM18, DMZ11, DDBM21, ER19, Fan21, FS18, FK12, GP15, GKP12, GLS07, Han03, HM23, JS08, KS13b, LWCS22, LT16, LT15, PP05, PZ14, RT96, RS01b, SV13, TA05, TZ17, UCV17, Var19, WK21, Zha22]. discrete-time [BP03a, CM18, Han03, HM23, KS13b, LT15, PP05, Var19]. Discretely [IS98, BS95, BN05, GCJL98, GCJL99, Jac06, KS99a, KSJ19, Mar03, Ped95, Sør02, SU03, vdMvZ13]. discretization [MFWB22]. discretizations [GT09, RSBG22]. Discretized [Fan14, DMM^{+}20, Dio12]. discriminant [BL04, GC05, JCL20, LM15]. discriminating [DPSW10]. Discussion [Wak95, Wel96]. disorder [BvdHH13, CSY03, RASY17, UDS23]. disparity [Hoo16]. displacement [Pro24]. displacements [Bha23]. dissipation [TY23]. dissipative [Wan23a]. Distance [DMM^{+}20, LWY24, TA05, ACJ18, BC16, BD09, BX06, BGX15, BvdHH13, BL20, DMMW18, DEGR14, GJS22, HW99, JS05, JS06, Lei20, RS18b, STC^{+}24, Wan16, WB19, ZS21]. Distance-reducing [TA05]. Distances [HM20b, DGU05, DR23, Kre21b, MS10, NWR22, Wie22]. distortion [LP07]. distributed [CL24, EKS15, HK13, LS03, NZ05, PT15, UDS23]. Distribution [GNT17, HT00a, Rév96, AAG24, AEP11, AD11, AGM06, AG07, BHPZ15, BP11, BBS14, BZ12, BSS05, BG10, BBC12, BS18b, BDDF18, CEL20, CS18b, DDW13, DRT16, Deh05, DR16, DFK08, DLL20, Dor97, DR09, DLT09, FR19, FZ22, GL18, GR14, GC05, GN10, GW17b, GR00, GJ13, HW99, HW03, HL17, HLS14, HW24, JRSK21, JLP10, JS13b, JN95, KM20a, Kam18, Kar24, Kat09, Khm16, KL09, Klo14, KKP07, Kut97, LP02, LL00, LS95, LP18c, LGS08, LYT21, LH01, MM15, MR10, MRST07, NZ05, Nak15, PP09, PQY13, PS09, RS01a, RZ21, Röl05, Roo99, SY00, VT13, VW15, Wan17, WXZ19, YY24, ZSC22, BLL15]. distribution-free [CEL20, GC05]. distribution-freeness [HW03]. Distributional [BFM16, CRW24, DSW21, GS00b, HNR23, HKMS24, KP22b]. Distributions [BMMS12, ADH21, Alm03, AM07, ALR23, ABH00, AI99, BD18, BNMS06, BGI18, BBS15, BL23, BB17, BJSZ07, BL22, Ber03, Ber06a, BCPR09, BW10, BS15c, BP98, But17, CL12, CG99b, CEG17, Cut00, DCC14, DLS14, Duf96, EK95, FP17c, Gho99, GK13, Gru00, HT00a, HLPP09, HL12a, HL14, HG97, HMW21, IS19, JJLW22, KJ13, Kim21, KP06, LWCS22, Len11, MPAiS12, Mla08, Möh06, Nie13, PS22a, PJT19, Pon07, PV07, RT96, RR01, RT06, RW02, Sie07, TT06, TZ17, VVV06, Wer11, Woo00, XZ09, YY24, ZZZ17]. Divergence [Tal13, GHM^{+}17, JZ19, Ovc18]. divergences [Dit19, KF14]. divergent [Dia23, HW11]. diverging [Tho23]. divisibility [BL22, BS15c, SW05]. divisible [BNMS06, Ber03, Ber06a, HMRB08, JM97, KS16, MPAiS12]. DNA [Cha03]. do [KRX18, KK21]. Does [EMPW20, Har17]. domain [BM09, CHI17, HG97, KJ20, LS02b, LS23, MP23, VBLN20]. Domains [Zha18, DS06, Tod20a]. Dominance [KS13a, BFM16]. dominant [SFP23]. dominated [Osȩ07]. domination [KP23]. Donsker [BGGL21, BZ23]. Donsker-type [BGGL21]. Doob [LPW04]. dose [DPW09, DPSW10]. dose-finding [DPW09]. dose-response [DPSW10]. double [But24, HL14, PT08]. double-transform [But24]. Doubly [GT22, Ama13, BVM07, CP18, Eis00, FR15, HW95, IKP14, dUÁV21]. down [RW22]. drawdowns [LLZ17]. Drift [KKR24, Ass23, Bar22, Bei97, BC21, CHR24, DDBM21, Gra21, GY04, KP22b, WYZ21]. drifts [RZ21, RZ23]. driven [BM15a, BR18a, BYY17, BNBV13, BN17, BGSS22, BR12, BSG22, FR06, FK11, Fou01, GNNS21, GS16, HP14, HMW21, HW24, JM09, JVV05, KN08, KKR24, LSW20, Mai14, Mas04b, MH22, OR24, PS22b, RBM19, Sau12, SSX20, SXX22, Wan16, WXX17, WYZ24, XZ18, ZZ15, ZZ23]. drivers [CS13, IR20]. Dual [BLO19, Gat06, PR14]. duality [CK01, Möh99, RK10]. duration [LLZ17]. Dutch [EF04]. Dvoretzky [FK21]. Dynamic [JCL20, MR16]. dynamical [Kut15]. dynamically [CT19]. Dynamics [AGV18, Ald13, AS03, BW23b, BCM^{+}21, Cor16, DGLO19, FFZ22, Jag95, PR13, Var19, vDD22]. Dynkin [Mon15].

early [HKM^{+}16]. earthquake [OKM^{+}98, Wu09]. ease [EMPW20]. Eaton [HTL07]. economy [Che02]. edge [CGZJ22]. Edgeworth [BG00, CHP11, HS07, KM05]. Edgeworth-type [KM05]. Editorial [Ano95c, Ano95d, Ano95e, Ano96b, Ano96c, Ano97d, Ano97e, Ano98c, Ano98d, Ano98e, Ano98f, Ano99c, Ano99d, Ano99e, Ano99f, Ano99g, Ano99h, Ano00c, Ano00d, Ano00e, Ano00f, Ano00g, Ano00h, Ano01a, Ano01b, Ano01c, Ano01d, Ano01e, Ano01f, Ano02a, Ano02b, Ano02c, Ano02d, Ano02e, Ano02f, Ano03a, Ano03b, Ano03c, Ano04a, Ano04b, Ano04c, Ano04d, Ano04e, Ano04f, Ano05c, Ano05d, Ano05e, Ano05f, Ano05g, Ano06c, Ano06d, Ano06e, Ano06f, Ano06g, Ano06h, Ano07c, Ano07d, Ano07e, Ano07f, Ano08c, Ano08d, Ano08e, Ano08f, Ano09c, Ano09d, Ano09e, Ano09f, Ano10c, Ano10d, Ano10e, Ano10f, Ano11c, Ano11d, Ano11e, Ano11f, Ano12c, Ano12d, Ano12e, Ano12f, Ano13c, Ano13d, Ano13e, Ano13f, Ano13g, Ano13h].

Editorial [Ano14b, Ano14c, Ano14d, Ano14e, Ano15a, Ano15b, Ano15c, Ano15d, Ano16b, Ano16c, Ano16d, Ano16e, Ano17a, Ano17b, Ano23a, Ano23b, Ano24a, Ano24b, Ano24c, Ano24d, DM13]. effect [BDKS20, PY01]. effective [BX15, Che02, GW17a]. effects [Kur16]. Efficiency [GZ09, Kut97, MS17, PGV17, YMP11, BRW00, HW03, Hoo16, PS98]. Efficient [BKK23, Bel98, BG17, CCZ13, CZH14, GP15, JS17, Kam18, KW97, KS97, KP20, Mai14, OW17, RS17, WC18, WSZ12, Zha06, Cas07, FH07, FT19, GLT96, Ma10, QLC21, SSH14, TCP22]. Efron [dUÁV21]. Eigen [Bat17]. eigenfunctions [KS99a, Tod20a]. eigenproblem [CKM20]. Eigenstructure [SPP21, Dia23, HM19]. eigenvalue [CP12, EX23, GS00a, HPY18, JJLW22, ZBY17]. eigenvalues [BJ22, HLPP09, HZBH23, JMRX18, JB21a, KLP21, LWYZ22, Ma12]. Elbow [BDKS20]. electrophysiology [RT15]. Element [ZC23]. Element-wise [ZC23]. elementary [BL23, NVV99]. elicitable [CLZ24]. ellipsoids [BDKS20]. elliptic [AK22, BR07, Gob01, HP02a, Tin98]. elliptical [AFGS08, BGHL21, Has13, JJLW22, KKP07, WZ16]. elliptically [SR09]. ellipticity [BRS16]. embedding [AC17]. embeddings [AHS15, CH07]. emission [HN10]. emission-related [HN10]. emphasis [GLF16]. Empirical [BY05, BIPZ22, Ber06b, DP09, DJL23, EM03, EM13, GGT22, Gin96a, JJW24, Lec13, MMW17, PS13a, PR22, QT05, RST17, Wu03, XW12, ALT15, ACG19, ACG20, Bah21, BHPZ15, BRJN19, BBV17, BZ12, BZ23, BG10, BBC12, BK16, Buc18, CS20, CZ13, CJL^{+}23, CSY99, DDT14, DGU05, DR16, Dob23, DRRS18, EK11, FRW04, Fis14, FZ22, GNR14, GMR19, GW99, HKMS24, JN16, Khm16, KK21, KS13a, Kut97, LM10, LM16, LvdG14, LS19b, Lei20, LGW06, MZZ20, NL14, PT18a, PVD13, RWZ17, Ren08, RSS12, Sau18, Seg12, TW14, WB19, ZL03]. empty [ZX24]. encoded [CLR20]. encoding [Lim19]. end [HW99, HV09]. end-point [HW99]. end-points [HV09]. energy [BNBV13, MRST07, PT21]. Engen [Hir97]. enhancement [CK23]. enlargement [BH17]. ensembles [GNTT18, Jia13, Ran22]. Entrance [BC02]. entries [BDT21]. entropic [CR20]. entropies [Zha22]. Entropy [CS18a, GL06, ALL09, BK02, BDDF18, Gra99, HJ17, HW24, MMR23, RST17, TY23]. environment [BV17, BRS16, BD22, CTP21, CSY03, DPTV15b, FS10, GL18, KM22, LØ01, MPV07]. environments [AP19]. epidemic [BGMSB14]. Equal [BG21, Har17]. Equal-tailed [BG21]. equality [DT23b, JP15, ZZZ17, ZHZL19]. equation [BG17, BKT19, CW01, CRW24, CR22, DVSS14, DVSS16, DDBM21, DL09a, DP18, Duf96, Fou01, FM02, GK99, GK01, dV13, HNZ18, HW19b, JW21, KN08, Kra21, KS01, MNN^{+}22, Mar03, MV18b, MSS00, Mor99, PS20, QSSS04, SSX20, WXX17]. equations [AG01, Ama13, AKH17, AP00, AW09, Bad00, BM15a, BYY17, BGSS22, BE16, BR12, BSG22, BVM07, BRS07, BK06, DSS10, DNR20, FR06, FS00, FK11, FK12, FN21, GT16a, GMD06, GK12, Ham03, HDS22, Jab21, JM09, KS99a, KY22, KK23a, KHLN97, KS13b, Lau04, LGW06, LS02a, MMP00, Mel00, MN00, MSS04, NR00, NS01, PT18a, Pet97, Rei05, RTT20, Sau12, SU03, SXX22, Tin98, Tka20, Wan16, WYZ24, XZ18, ZZ15].

Equidistribution [Bor21]. Equilibrium [AMV18, DL09a, Var19]. equirepartition [GL06]. Equivalence [BK23, GS18a, GT19, Kur16, BAT19, CT02, DH02, NN22, Tan06]. equivalences [DFK08]. equivalent [AG99, BPB^{+}22]. erased [LW23]. Erdos [LMZ20, MR19]. ERGM [MX23]. ERGMs [Muk20]. Ergodic [CK11b, DMS07, DDL23, GP15, HLP12, Jac06, Kut97, KY07, Nob06, OU23, Tal13, WK21]. Ergodicity [KL13, LTL07, ZZ23, ALV18, BW23b, CH15, Han03, HW24, JT03, KW13, LS21, LBBG19, Neu11, Wan23a, Wan23b]. Erratum [Coe06, Gay22, IRVZ24]. Error [BRR19, LEM23, PJT19, BF01, BV10, CCF22, CHI17, CF15, CRT21, DRT16, DM08a, DLPZ09, DLPZ12, GT09, JS13b, Klo06, KMS17, ŁMN13, ML10, MCJ23, PRR13, Ros09, San10, ZC23]. errors [BGL22, Bor01, BW19, BTV17, CM99, DJ97, DSW21, HV09, HMP18a, JKNP16, LM15, Woo00, Yan01, YK20]. errors-in-variables [HMP18a]. Esseen [BGT97, BGL24, CH08, HLP12, Jir23, LS24, Rai19, SZ22, Tur04]. essential [TKT09]. estimate [BLL15, KK23b]. estimates [AL05, BJ22, BV10, Ček04, DF98, DLM14, EX23, GM15, Giu18, KKS02, MN22, Mus07, SSW12, SW19b, Van05b]. Estimating [BM07, BX00, Che24, DdH15, DGLO19, DLT09, GS00a, HW99, JN20, KS99a, KKP07, KW20, LYT21, LS23, LTC14, LS14, RBM19, Rob08, SC11, AK97, Bad00, BN05, CLLLR17, Cha97, CDG14, HP95, HLPP09, Jac02, PT18a, TMW13]. Estimation [AFGS08, APGMV04, BDG05, BKY21, CCC^{+}22, CGL12, CV19, DNS19, EH02, FRS07, FH07, GR23a, HRY13, HYY11, HMRR20, ICG16, JLT08, Lau97, MOP20, MH10, NWR22, PV09, Pon07, Sør02, TV99, Vet15, AAL23, AD18, AB15, ARV18, ACP23, AAG24, AF19, AH17b, Bac18, Bah21, BDJ19, BS21, BVT96, Bel98, BPW19, BTT19, BGSS22, BGL22, BF01, BS09, BS12, BHK03, Ber04, BG17, BC21, BS95, BK02, BMT17, BL10, BHM22, BS18b, BDKS19, BM05, BC09, BDKS20, CL05, CLZ07, CD17, CW22, CCZ13, CGLZ24, Cas07, Cas14, Cav01, CK19, CGS18, CL02, CPY13, CHT98, CCFH00, CC04, CZH14, CAA22, CL14, CCL23, Cho18, CM14a, CL17, CCT20, Com01, CGCR07, CSvS07, CM14b, CM18, DCG97, DNV99, DR07, DFS10, DRT16]. estimation [DJL10, Deh05, DM08a, DL09a, Dic16, DJ96, DL09b, DHKR15, DDFD04, Dre03, DR09, DL17, DM21, Efr98, EK18, ELP24, FN23, FV03b, FZ04, GG22, GP15, GL05, Gay20, Gay22, GCP17, GCJL99, GT22, GN10, GP03, GL08, GL22a, GL22b, GLT96, GT01, GK12, GP12, Haf13, HT00a, HLT01, HL17, HY05, Hay22, HHH24, HP99, Hil15, Hir16, HNR20, Hof99, HM20c, HKM06, HR22b, HF99, IMS01, JS17, JW09, JLL04, KFT21, KD95, KPT96, KK18, Kim14, KLW18, KW97, KLTZ19, Koi16, KL00, KN99, KLLM13, KS97, KS21, KMS17, KP20, Kuj16, KY07, Kut21, ŁMN13, LW18, LMP12, Leh19, LP18a, LP18b, LT16, LEM23, Lou14, Lud00, MSW13, Mai14, MM23, MRS20, MS23b, MR16, MCJ23, MSS04, Min15, MW20, MPV22]. estimation [Nau22, NR09, NZM23, Nie00, OKM^{+}98, OL23, OCDM08, OD19, Oud98, PW04, PHO17, PQY13, PTA07, PP22, PY01, Qia21b, QLC21, Reb15, Rei05, RS17, RW19, Rig06, Riv06, SSH14, Sar16, Sar23, Sch20b, SK23, Sri16, SRR24, Tal13, Tan15, TRB24, TMW13, Tru19, US24, VSV03, VV02, VG18, WXZ15, WC18, WZ16, WK21, WSZ12, Xia09, XL10, YB06, YG19, YLC19, YRZ24, Zha96, Zha00, Zha06, ZC23, ZWY17]. estimator [AR17, BBJ08, BBC18, BX15, Cat11, CHWW05, CL14, DL19, DDBM21, DNP06, DMR21, DM01, DMS21, EKS08, FLP12, FB22a, GJZ23, GM04, GN08, GZ09, HS07, HY08, JK21, Lei06, LT13, Lep18, Ma10, NR12, Ped95, TT06, TCP22, Van95, Van05a, vEGS07]. estimators [AM17, BD13a, BP06, Bel17, BBD20, Bia14, BM98, CDJ13, CL12, CCJM24, CG23, CGK^{+}20, CV97, Dom15, DF19, DN21, Efr00, FG19, Fer01, FT19, GGKL22, Gir08, GP23, GRS11, GMR19, GW99, HP01, HS10, HLP12, IKP14, JKNP16, KK16, KS13a, LM07, LGS08, NRC24, NLL18, OCBG19, PG21, PR06, RB06, RV09, SW04, SW13, SW12, SZ22, SGRV18, SC23, Tal13, TCP22, VK97, VFJ18, VS05, VW15, dBM17]. eternal [Lim19]. Euclidean [CY21, CH18, GJS22]. Euler [BJM21, Fri18, LTX22]. Euler-type [Fri18]. Evaluation [DL09b, Lee08, HMS04, VT13]. event [BBD20, HLZZ23]. events [ABH00, CR13, DM09, DN21, MZ14]. Evolution [ACJ18, Mel00]. evolving [CT19]. Ewens [Hir97]. Exact [ADN24, BN05, EO07, LWX15, LM20, OSZ18, TZ17, Wan22a, Ber04, BPR06, FSW17, GS20a, HVM99, LX23, LBDM19, PJR16]. example [FN97, PS20]. examples [JS13a]. exceedances [BP03b, RZ13]. exceedences [HW10]. exceptional [SS15]. Excess [BVT96, AFGS08]. exchange [MZ17, Wan22b]. exchangeability [CSY^{+}23]. Exchangeable [MSS16, dCQM09, BFT15, BCRT24, Mcc00, Zha23]. exciting [CP18, ET23]. exclusion [BFMP01, FLS01, MMVW10]. exclusion-voter [MMVW10]. Excursion [CX16, BST12, PY03, Yan08]. excursions [HS97, Kar24, Lim19, PY99]. Existence [CD11, DPX23, Dom15, CW01, Jab21, LS02a, Tin98]. Exit [AIZ16, GM15, LS02b, TT20b, VY13]. exogenous [MSS16]. expanding [AV14]. Expansion [MVY19, MCSS99, RW02, Beg07, BG00, BGL24, PT21]. expansions [AKH17, But17, DN08, FLLO14, GL18, HS07, HWW99, KM05, Nda23, Pan22]. expectancies [KM23]. expectation [BM16, KSS21]. expectations [AK18, CP05, FPSS19, KK17, KLS98]. Expected [CS18b, AW09, CD21, EHS05, Lin09]. expectile [DGS20, ZH20]. expectiles [PS22a]. experiments [BBG19, BM12, DPW09, EBR23, KMW14]. Explicit [BP06, BGMZ18, CPP11, CD21, Pec03, AH13, GCK04, MY02, PSS24, PR23, Rai19, Sko96]. exploratory [DG13]. Exploring [LLTY07]. explosion [Wan17]. Explosive [KM95, Ber01, JWY22]. exponent [TKT09]. Exponential [BW23b, CH15, DPTV15a, DV18, Fil00, GW17b, RT96, Sie07, TT20b, Wan23a, Wan23b, YP23, ADH21, AIM07, BKR99, BL23, BMMS12, Bel17, BC19, BM98, BJM21, BHL99, BWJ18, BKT19, Che24, CS24, DLR06, HMS04, HW24, KN06, LR21, LM13, MS23d, PR02a, PSS24, PRvS13, Pat11, PR99, SST16, Sch20b, SM23, VW15, Woo00, Xu21]. exponential-family [Sch20b]. Exponential-polynomial [Fil00]. exponentially [EGG03, TT06]. exponentials [TC03]. exponents [BS19, Hie17, TV99, Wer95]. exposure [Lin09]. expression [FR19]. expressions [PSS24, SSH14]. Extended [PWZ17, Wal12, APP11, HT00b, KJ13]. Extending [LPW04]. extension [FPS95, GS00b, JMV20, KT18, LN02, PFEP23]. extensions [GR16, GS11, KMB21, Riv05, Riv07]. Extinction [DPTV15b, BHM22, CD21, FB14]. extracted [BS18b]. extrapolated [Ber97]. extrapolation [LP17]. Extrema [QP18]. Extremal [KLP21, RNS23, CS24, DGG13, JD16, LS16a, MR13, Rob08, SS15]. Extreme [Dre03, DZ24, HS97, KL05, BVT96, BDG05, BDV21, DFS10, DPS24, DR11, DM09, Dom15, FFZ22, GKNY11, HT00a, Ma12, MZ14, PS22a, PQY13, Ued21, ZHZL19, KFT21]. extreme-value [BVT96, BDG05, GKNY11, HT00a, PQY13, ZHZL19]. extremes [Bha23, BC00, BJ24, DG13, DY14, EFVG08, Fas10, ZX17]. extremogram [DM09].

F [HPY18]. facets [HSS07]. factor [BBW22, KLW18, WM16]. factor-based [KLW18]. factories [NLS23]. factoring [GW98]. factorization [Pat11]. factors [BX06, BGX15]. failure [BBD17, DdH15]. fair [HN10]. False [BDR14, SC23]. families [ÁEdBCAM18, AIM07, BKR99, BWJ18, Che24, Fil00, KM06, KMS17, Nob06, VW15]. family [APP11, BMP08, GZ21b, KJ13, KM20b, Sch20b, SXXZ18]. FARIMA [BF01]. FARIMA-GARCH [BF01]. fast [AEP11, BBW20, CCZ13, FLS23, LM15, OSZ18, PTL04, SXX22]. favorable [HL12b]. favors [ERS19]. favourable [KW97]. Favourite [CRS00]. Feature [HJM14, LMZ24, Zha13]. features [BBW22, HJM14, RP23]. feedback [LLT12]. Fefferman [Ose20]. Feller [Rát19]. Ferguson [LV12]. Feynman [GD17, GST08, IR20, NS01]. field [Ald99, BvdHH13, DJL23, EZ03, ELL22, ERS19]. fields [ACBLV18, BP21, BCI04, BL15b, BK23, BSW24, BST12, Cap24, CX16, CS18b, DK18, GS08a, HNR23, JK23, JRSK21, KM20a, Kur22, Lee22, LX23, LO14, LRMT17, Lep16, LT16, LWX15, LBDM19, LS23, MP10, MJK17, PZ19, PS22b, PFEP23, PS16, QP18, Qiu19, RNS23, SW20, SX10, VAD22, WSX20, ZX17]. filter [CM14b, GCK04, HB23, Yan24]. Filtered [DS05, DNR20, LMNV11]. filtering [BHL99, PP05, PR14, Wu09]. filters [CH08, CM18, Kün13, OD19, RS01b]. finance [Dre03, EK95, FS13, NS01, PP16]. financial [ET23, KLW18, LT15]. finding [AM16, DPW09]. Finetti [BCRT24, DF20]. Finite [AHS15, CAA22, DGP18, DMR21, PT21, AP11, BG00, BBC12, BBC18, CD21, CP16, CLZ06, FGS21, GR16, HB04, HNR20, HT00b, HWZZ23, IKP18, KMM24, LMC23, MY23, Ngu15, PT15, RS01b, SS20, Sur04, TT06, WB19, Yan24, Zha22]. Finite-energy [PT21]. finite-population [BG00]. finite-sample [HT00b, WB19]. finite-state [HB04]. finiteness [KS20]. First [CD11, ERS20, But17, Dan96, Dan00, GJS22, HP14, LS02b, McG09, PB15, PT18b, WXZ19, YY24]. First-order [ERS20, PB15]. first-passage [But17, GJS22, McG09]. Fisher [BL04, BCG14, HL12a, HL14, HZBH23, KL13, LO14, SSH14, ZBY17, dSJMT23]. fit [CEL20, CV09, DGU05, GKNY11, GLF16, HLS14, Kut15, LN14, LWLL22, Loc12, NP08, PS13a, PT18a, SN13, Vet15]. fitting [ZL03]. fixed [BAT19, BDZ15, BKT19, GM15, JS17, LS23, MNN^{+}22, Tod20b]. fixed-domain [LS23]. fixed-size [BAT19]. flat [BTUA23]. Fleming [ALR23]. Fleming-Viot [ALR23]. Flexible [DMT24, NHV^{+}21, JD16, SBP24]. Flexible-bandwidth [DMT24]. flows [BT20, BDM10b, TY23]. fluctuation [AC01, DKM16]. Fluctuations [Zei15, BGT07, Gor96, LX21, PZ14, RT15]. fluid [RS01a]. Flux [Kra21]. Fokker [HW19b, MNN^{+}22]. force [AG01]. forced [FK12]. forcing [TKT09]. forecasting [AW12]. forecasts [BG21, ILY09]. forest [DLWZ17, RP23]. forests [Lei19]. form [Fis14, SSH14]. forms [BT17a, BP08, FLS23, JN20, PR06, Zho19]. formula [BR07, BDM10a, FPS95, Hir97, JW21, NVV99, NS01, PSV00, TK03]. formulae [ALW11, Pec03]. formulas [DD15, HM23, PT18b, RASY17]. formulation [HT19, TKT09]. Fortune [ERS19]. Forward [GLT23, JN16, BRJN19, MSS04]. forward-reverse [MSS04]. foundation [Ken95]. foundations [BBLG17, BM13]. four [JB21a, JB21b]. Fourier [BP16, Bar18, CdLOR20, Efr00, MZ14, Nic18, PZ19]. fourth [Sch20a, KT18]. fPCA [BX15]. Fractal [CGMST95, Cut00]. fraction [LS95]. Fractional [BK06, FK11, Mar06, OP10, Pey17, SSX20, AD18, BM15a, BP06, BŁ15a, BCI02, BCI04, Ben03, BP10, BR12, BK23, Che01, CD23a, CNW06, DN08, DS19, DP11, DW19, FR06, GJQS20, IMS01, JM09, JMW05, KL02, LPSS18, LD11, MOP20, Nku21, NVV99, NPS20, OP11, PR16, PS22b, PT00, PT01, Sau12, Sau14, Zei15]. fractional-order [BK23]. fractionally [DNR20, LRS23, Tsa09]. fragmentation [AD07, Ber17, DHKR15, HK11, MPW08, OH23]. fragments [AD07]. framework [CMMA20, DGS08, Dom15, LTC14, Mar13, Nda23, VAD22, Wu09]. Fréchet [BS18b, Elt22, ZP19]. free [AH13, BNT02, BH00, Bou15, CEL20, DN14, DM21, GC05, Giu18, Khm16, MNN^{+}22, PP16, Ued21, WP22]. freedom [BZ22, PP22]. freeness [HW03]. frequencies [Cho18]. Frequency [KJ20, Abr19, BPW19, CI21, DM21, FT19, JS17, KLW18, MP23, NR09, RvS19, VBLN20]. frequentist [Hof23, YK20]. Friedman [GR23b]. Frontier [DFS10]. full [Mal09]. fully [CV19, SS15]. Function [RSS12, AF19, AH17b, Bah21, BHPZ15, BM07, BSS05, BL04, BBC12, Bor01, Buc18, CLZ07, CMS10, CL12, Che24, Dai17, Deh05, DR16, DNP06, DJ96, DR09, EH02, Fis14, GL05, GN10, GL08, GMD05, GS20c, HT00a, HP99, Hir16, JW09, KKP07, LS14, MM21, OPPS16, Oud98, PR06, PQY13, PY01, QT05, RA17, Rob08, Sch24, TMW13]. Function-indexed [RSS12, Buc18, CMS10]. Functional [BWZ10, BGMZ18, CG22a, CP21, HMWY10, HR15, IPP23, Kab11, Mar13, MZZ20, MPU19, PR16, RW19, SW04, BB99, CDJ13, CL24, CC19, Cor14, CM13b, CGER07, DDK22, GGM20, GP11, GQ23, HMP18a, HMP18b, HHH24, HK13, HR13, KJ20, KS01, Lei20, LRS23, LTC14, Mei16, MCJ23, NHV^{+}21, SGRV18, WFM10, Xia20, Yan08, YDN23, ZS15, vDD21, vDD22]. functionals [Alt21, BIZ02, BC09, CL05, CCR20, CLZ24, CCT20, DNN10, DGU05, Dio12, DL13, FMSY12, GNNS21, GP03, HSS07, HF99, Kre24, Lau97, MS23d, PSS24, PRvS13, Pec08, PT03, PV09, RS17, Wal13, Xu21, Zäh15]. functions [ARAS20, AK23, AK97, AAG24, AM17, BBCM11, BGMSB14, BVT96, BS12, BMP08, BZ23, BS95, BH96, BG10, Bra02, CPY13, CGER07, DDT14, DJ97, DD15, Dor98, EGG03, Fas10, Gne13, GLT96, GP16, HP95, HLPP09, HMP18a, Hös05, Jac02, JRSK21, JM97, KSX19, Küh22, Lep16, Mar18, Mau19, MGS03, Muk20, NN22, Obl06, PTA07, RWZ17, Roo05, Sev04, SW19b, Ste13, Sur04]. Fundamental [GS20a]. Further [JS13a]. Fused [ZC23]. fuzzy [HKK19, JK17].

G [HP06, RS01a]. gains [BEM18, YMP11]. gains-process [BEM18]. Galton [Alt13, BWW00, CTP21, HW19a, IKP14, KM22, KZ21a, Lin19, OH23, Sid18, dBCAM22]. Gambler [Mus07]. gamblers [CS06]. game [Hei18]. games [GHL05, Hei18, TKT09]. gamma [Alm03, BMMS12, BB17, BGSS22, Ber06a, BLM19, Duf96, LRMT17, Sie07, DP18]. gamma-correlated [LRMT17]. gamma-driven [BGSS22]. gap [AIS19, HKK19]. Gaps [PY19]. GARCH [ABH06, BF01, BHK03, BM12, FZ04, Hil15, Hör08, MMS08]. Gauss [AG07]. Gaussian [AD11, ALW11, AGG23, Bac18, Bar10, BBJ08, BGMZ18, BBG19, Beg07, BD15, Ber01, BG95, BL15b, Bir04, BGH01, BK23, BSW24, BDM10a, Car06, CP16, CX16, CS18b, CKM20, CR07, Com01, DR16, DY14, DZ17, DLWZ17, DRW20, DLM14, EKS15, EMW18, FT19, GLT96, GNSU19, HXZ18, HNW14, HNTX15, JZ19, KM20a, Kar23, KM06, KLLM13, Lee22, LX23, Lep16, LS95, LWX15, LMK19, LS23, Lud00, LP07, MR17, MS23c, NN22, Nda23, Nor15, OR24, OL23, Pai22, PZ14, PR02b, QP18, SSH14, SW20, SSX20, SN13, Vii16, WSX20, YB06, YRZ24, ZC18, ZGV20, ZX17]. gene [CHWW05]. genealogical [DHKR15]. genealogy [Cor16]. General [JS08, PR23, AG99, BDZ15, CL12, CGK^{+}20, DRW19, Dit19, DM01, FHJK20, HM23, JMM17, KPT21, Klo14, KS18, KLP21, LAP20, Mas04b, MP23, Nda23, OW17, PT18a, VBLN20, VAD22, WP22, ZWY17, ZBY17, ZH20]. generalised [BLM19, IR20, PWZ17]. Generalization [Sch13]. generalizations [DGS08, GJY03]. generalize [KK21]. Generalized [BVM07, CK01, JB21a, Wan97, BMMS12, BB17, BS22, BDG05, BSG22, BW10, BOJGMR15, CZH14, Cor14, Dem07, Fas10, FdH14, FGK16, GHM^{+}17, HL17, HTW23, HY08, ICL19, JW21, JB21b, KM06, LGW06, LX21, Nku21, PR16, RT06, RVW01, SW05, WC18, ZC23]. generated [Kat09]. generating [Lee08]. generation [GGKL22, Klo06]. genetics [Möh99]. geodesics [CH18]. Geometric [GMD06, Han03, Min15, ALV18, BP10, BBLG17, BL20, CCZ13, DL16, DMS07, EHHS23, Gan21, JT03, Kab11, LS21, LBBG19, PRR13, SBP24, YP23]. geometric-median-of-means [YP23]. Geometrical [Elt22]. geometrically [HLP12]. Geometry [ST13, AK97, Hay22, Kre21a, MV04, Pen07, Tay13, ZS12]. geostatistics [BPB^{+}22]. Getoor [Sch13]. GGC [JS13a]. Gibbs [AH17a, ALV18, BFT15, BD13a, Bra22, CS15, CL17, DL09a, EZ03, FP17c, HP99, HKK19, JK17, JM17, LN02, MSWW11, MPW08, MK20, SM23, WW14]. Gibbs-non-Gibbs [JK17]. Gibbs-type [BFT15]. Gibbsianness [JK23]. Ginzburg [WXX17]. Girsanov [NVV99, PS20]. given [BSS05, BDDF18, SL18]. glass [AS03]. Glivenko [DR19]. Glivenko-Cantelli [DR19]. Global [GGKL22, CLZ07, CK23, Dor97, Efr98, GY04, LX21, MV04, RT15]. Gnedenko [Ano97c]. Gneiting [ZP11]. Goldenshluger [PG21]. good [ST18]. Goodness [CEL20, GLF16, LN14, NP08, CV09, GKNY11, HLS14, Kut15, LWLL22, Loc12, PS13a, PT18a, Vet15]. Goodness-of-fit [CEL20, LN14, NP08, CV09, GKNY11, HLS14, Kut15, LWLL22, Loc12, PT18a, Vet15]. governed [BJM21]. Gradient [TY23, BCM^{+}21, CCZ13, CBL22, MCC^{+}21, SSW12, SZ22, SK23]. gradient-based [MCC^{+}21]. gradients [MWZ10]. Gram [Giu18, ZZW22]. Gram-type [ZZW22]. graph [ADM14, BCvdHV21, BC20, CDL24, CMr19, Drt09, DGLO19, GST08, GS18b, KW20, LV15, ML11, NC17, NPM21, Sch20b, SV13, Stu23, Wer11]. Graphical [MH22, Dob03, DRW20, KT15, Kos02, NC21]. graphs [BCvdH20, BSW24, CD21, DT23b, Fan21, FR15, FGS21, Gan21, GK18, HMT00, Jan07, LR23, LMZ20, MR19, Man07, Sad13, SL14, TAS^{+}17, Tur04, WC98]. grass [AMV18]. Greedy [San16, ELP22]. Greenwood [Möh18]. GREM [FFZ22]. Grenander [GJZ23]. grid [Cha23, PP22]. Griffiths [Hir97]. Gromov [HW19a]. Group [MP10, CR17, ES99a, HIW19, NR12, NJG20, WH10]. group-invariant [ES99a]. group-lasso [NR12]. groups [Bor21, CD01, Dob23, DZ17, Fri24, MM15]. grow [BH10]. growing [Dic16, OL23]. grows [WM16]. growth [Ber17, BJM21, Bra99, DHKR15, GMM05, GMD06, HY08, HYY11, PTL04, ÝSJ08]. growth-fragmentation [Ber17, DHKR15]. guaranteed [MH10]. guarantees [BBW20, BRE23]. Guided [SvdMvZ17]. Gundy [BS15b, SS18, Sio18].

Hadamard [ST18]. Hahn [GS11]. half [BR97, DGC11, Yan08]. half-line [Yan08]. half-space [BR97, DGC11]. halfspaces [Tho23]. Hamilton [Kra21]. Hamiltonian [BW23b, BBLG17, BRE23, LBBG19]. Hanson [CY21, Zho19]. hard [CFKR16]. Harmonic [Lin19, Nda23, JN15, LP00, MPS19b, SU08, WXZ19]. harmonizable [BCI02]. Harnack [CCTY19]. Hartman [Dem07]. harvesting [LØ01]. Hastings [BP19a, JLM14, MV18b, Wan22a]. Hausdorff [HW19a, Bru19, DF20, Lee22, Man07, SX10]. Hawkes [ELL22, GZ18, GZ21a, MM23, RvS19, SRR24]. haystack [BN20, CT19]. hazard [BM07, But17, GJ13, JW09, RA17, VV02]. hazards [ZWY17]. HCM [JS13a]. heat [CR22, FK12, GS20c, HNZ18, KY22, SW19b, WYZ24]. Heavy [BD20b, ABH00, DFK10, FGS21, HM19, Hil15, HMW21, JMRX18, Jon11, Jur99, Kam18, LT00, Nov14, PS22a, PT00, PTL04, Xu21, ZZ23]. Heavy-tailed [BD20b, ABH00, FGS21, HMW21, JMRX18, Jur99, Kam18, LT00, PS22a, PT00, PTL04, Xu21, ZZ23]. heavy-tailedness [Jon11]. height [CS18b]. Hellinger [Sar23]. help [HLP23, SC11]. hereditary [DL09a]. Hermite [AT24, GNNS21]. Hermite-driven [GNNS21]. heterogeneities [BG95]. heterogeneous [EMW18, KKR24]. heteroscedastic [BJ24, CL14, DM08a, Sch24]. heteroscedasticity [Tan15]. heuristic [LT16, MO08]. hidden [ACBLV18, BD20b, BR96, DS03, DM01, DL17, FH07, GR14, GR16, GCJL00, Hay22, KK18, Leh19, LK08, LS20, OW17, SV13, Van05b]. hierarchical [JM17, Ngu16]. hierarchies [RR01]. hierarchy [BG95]. High [Din20, DM19, HTW23, KMB21, LAP20, Lou14, LP07, NK21, VV04, Abr19, ACPM24, BCFH22, BH17, BP03b, BC13, BTT19, BBW23, Büh13, BI13, CM13a, CK11a, CI21, DR16, DM21, EBR23, FT19, GS20a, Gho99, GT22, GR04, GQ23, GKT24, HM19, HWZZ23, JS17, JCL20, JB21a, JJLW22, JR17, Kam18, KLW18, KP23, KKP21, LYG19, LS19c, LWYZ22, LWY24, LLL22, LEM23, MY23, MP10, MMW17, MS23c, PV16, SG16, SL18, SvS17, TRB24, VG18, WW14, WAP17, WYR21, WS23, WH10, YK20, Yin22, YZZ23, YLC19, ZC18, ZSC22, Zha22, ZS21]. high-dimension [Kam18]. High-dimensional [DM19, KMB21, LAP20, Lou14, NK21, VV04, BCFH22, BC13, BTT19, BBW23, Büh13, BI13, DR16, EBR23, Gho99, GT22, GR04, GQ23, GKT24, HM19, HWZZ23, JB21a, JJLW22, JR17, KKP21, LYG19, LWYZ22, LWY24, LLL22, MY23, MMW17, PV16, SG16, SL18, SvS17, TRB24, VG18, WS23, WH10, YZZ23, ZSC22, Zha22]. high-frequency [Abr19, CI21, DM21, KLW18]. high-level [BP03b]. High-resolution [LP07, MP10]. higher [CG23, Erh14, GLF16, US24]. higher-order [US24]. Hilbert [Cap24, CCZ13, CP19, CY21, GP03, Jir18, KSS21, ZGD24]. Hilbert-valued [Cap24]. Hilbertian [JLMP22]. Hill [HS07, JK21]. Hilliard [CW01]. Hipp [Roo05]. history [Sen13]. hit [DW20, HMTvB24]. hit-and-run [HMTvB24]. Hitting [Möh18, PY03, BL23, DSS10, Das05, DH17, Gra21, dV13, YY24]. HMC [OPPS16]. hoc [DFK16]. Hodgkin [HLT17]. Hölder [Ber04, DVSS14, DVSS16, GLT23, GLL20, Mor99]. homogeneous [DN14, GMM05, NZM23, Sch13]. Homogenization [HDS22]. Hopf [KK23a]. horizon [PSS24]. Hörmander [Qiu18]. Hörmander-type [Qiu18]. Horvitz [CDJ13]. hulls [KM98]. Hurst [AH17b, DN08, FR06, GJQS20]. Huxley [HLT17]. HWBI [TY23]. Hybrid [CP19, CD95, BPR^{+}13, MMVW10, RT15]. Hydrodynamics [FGS22]. hydrostatics [FGS22]. Hyperbolic [EK95, BS19, BE16, GM15, Gru00, JW13]. hypercube [BBS16, HT00b]. hypergeometric [GW17b]. hyperplanes [ST13]. hypersphere [GPNECA23]. hypocoercivity [HW19b]. hypoelliptic [FO24]. hypotheses [BP19b, BDR14, Dob23, DNS19, DX16, RK10]. Hypothesis [DT23b, Nob06, CM13a, EM03, Fer01, Gay01, HW10, LS99, LAP20, Nak15, Reb15, RW19, TAS^{+}17, TZ17, ZL17].

i.i.d [BG10, LS00, Zho14a]. identically [EKS15, LS03, NZ05, PT15]. Identifiability [CGLZ24, CP21, DL17]. identifiable [ER19]. Identification [BCI02, Coe05, SV13, HR22b, NLL18, Wu09, Coe06]. identifying [GP04]. identities [AIZ16, CPP11, Jam10, Myk01, PSS24]. identity [APP11, AC01, NK21, Van95]. II [AF12, BvdHH13, Bor21, GL22b, HG97, IMM17b, NC17, PXH13, Riv07, YB06]. III [HG97]. illustrated [Wu09]. images [LRD02, SX10]. immersions [Ken15]. immigration [CLR14, FB14, IMM17a, IMM17b, IKP14, PRR19, Sid18]. impact [BT17a, BD20a]. implicitly [Hir16]. Importance [AHK15, BP06, Cha03, FH07, MPS19a, Ore22, RP23]. improper [BD16, HMS04]. improve [Yan04]. Improved [MFWB22, Tan15]. improvement [GMR19]. Improvements [KMB21, VZ23]. Improving [Kee13, KK16]. impulse [DMR21]. imsets [KT15]. Inadmissibility [Mat24]. INAR [JW19]. including [HVM99, HPY18]. inclusion [FGS22, Yu12]. incomplete [PX10, SCK23]. inconsistency [ES99a]. increasing [BG09, BC02, GS21, HM11]. increment [BS11, BMMS12, MR10]. increments [DJ97, DGP09, KM20a, Mus07, OS15, WSX20]. Independence [KV12, BD14, BW08, BO20, DDFD04, LN18, LS19c, LWY24, Mal09, Reb15, Tho23, WC98]. independencies [NC21]. independent [BMMS12, DLT09, EKS15, GY04, HU22, HMRB08, Kab11, LS03, MP06, NZ05, PXH13, PT15, RP23, Röl05, ST04, Wan22a]. Index [Ano97f, Ano07b, Ano08b, Ano09b, Ano10b, Ano11b, Ano12b, Ano13b, Ano14a, Ano16a, Ano95b, Ano95f, Ano97b, Ano98b, Ano98g, Ano99b, Ano00b, Ano05b, Ano06b, BDJ19, BVT96, BDG05, BL10, CP01, DPS24, DN08, Dom15, GJQS20, HS07, KL09, KP20, LMV22, LLD19, LPV13, NK21, Rob08, Sch13, WXZ15, WC18, XW12, ZLL17]. indexed [BZ23, Buc18, CMS10, DDT14, Eis00, FR15, FPW17, RWZ17, RSS12]. indices [BCI04, CV97]. indirect [JN20, Kur16, MRST07]. Inequalities [Osȩ07, AS19, ASC02, BM15b, BGMZ18, BHBO17, BC19, Bru20, CG22a, CC00, CCTY19, CEG17, DL13, ERS20, FS18, GZ21b, GS23, GS20b, HRBR15, HP02b, HT21, JWY22, Jou07, Kol09, KL17, LvdG14, LYG19, LT16, Lev18, LMZ24, MSWW11, MMR23, Mar18, MPRZ10, MR15, Osȩ09, Osȩ11, PXH13, PEBG17, RR15, SS20, SW19a, Sau19, Sau12, Sio18, TY23, Ver20, WW14, Yu10, Yu11, Zho19]. inequality [AS19, AGH18, Bar10, BS15b, Bou06, CS07a, CY21, CR20, DDL23, FMNW18, GL95, HXZ18, HR22a, IKP18, Jou09, JM22, KN06, KM11, Mau19, MJK17, Ose20, SS18, Zha22, ZK20]. inert [Bar22]. infection [CD21]. Inference [BD18, BM18, BR96, BBW22, BPD23, CT12, LR21, Nku21, RA17, Rob12, ZL13, ZGV20, ACG20, BDPR21, BS14, BMS21, BL20, BEMR17, Cha23, CP18, CSSC95, DPS24, DM19, GHM^{+}17, GT22, GP23, GK99, GK01, HB04, HP06, Hil15, Jac06, JVV05, KVDH16, KL10, KSJ19, KKP21, KS13b, LNP14, LSY^{+}22, Mar03, MRST07, MS18, NIR21, Nov14, Ogi18, PS22a, PS98, QT05, Rei13, RCBR00, RSBG22, Sau14, Tod20b, WZ18, YDN23, YRZ24, ZS15, vdMvZ13]. inferences [DR19, ES99a, FH05]. inferred [HKM^{+}16]. infimum [FMNW18]. Infinite [New16, RR01, SFP23, SW05, ARAS20, AGH18, BCRT24, BHBO17, BS16, BL22, BGT07, BS15c, CMS10, CP19, CL17, FRS07, FLLO14, GS11, HP02b, HS08, MR19, PT21, PT03, PTA07, RSS12, Stu23]. infinite-bin [MR19]. Infinite-color [SFP23]. Infinite-dimensional [New16, ARV18, GS11]. infinite-source [FRS07]. infinite-variance [BGT07, PT03]. infinitely [BT17b, BNMS06, Ber03, Ber06a, HMRB08, JM97, KS16, MPAiS12]. infinity [BH10, CP12, FLZ21]. Influence [AM17, MM20, CL12]. influential [HJM14]. Information [AK97, FK17, Hay22, HNW14, BCG14, EI18, HL12a, HL14, KSW01, Kuj16, Lua11, MV04, Mat24, MV99, NJ14, RCBR00, SSH14, SY24]. information-based [Kuj16]. informative [BBC12, BBC18, HLZZ23, PR02b]. informed [CT22]. inhomogeneity [AW24, BCvdH20]. Inhomogeneous [JN00, BCvdHV21, DV18, DZ24, GP18, HB04, HLT17, LS99, Mus07, Vil23]. initial [GZ18, YY24]. innovation [EH02]. innovations [KL05, SS12]. input [RBM19, RS01a]. instability [KTU18]. instantaneous [BM07, CH08]. instrument [ZWS21]. insurance [LT15]. integer [DH17, DvdAW09]. integer-valued [DvdAW09]. Integrability [BŁ15a, CK15, VY13, TT20b]. integrable [AHS15]. Integral [DP16, DP11, FR19, Vii16, AB15, AM07, BNMS06, BK06, CNW06, DZ03, GNNS21, KS16, KG00, Lau97, MN00, Pec08, Tud04]. integrals [BMMS12, BOJGMR15, CK15, FPW17, GT09, KS20, Küh19, MPS98, Nor15, NS07, OR24, PT08, PP16, Qia21b, US24, dUÁV21]. integrands [PT01]. Integrated [Ros09, CMS10, Fas10, GM04, GN08, HP14, KMS17, LRS23, NHV^{+}21, Tsa09, Vet15]. Integration [BS16, DS19, Ben03, BDP23, GLT23, GS23, GK12, Wal13]. integro [KK23a, Tka20]. integro-differential [KK23a, Tka20]. Intensity [LW04, Bar22, CCL23, CM14a, GZ18, NZM23, PW04, RB06, Tor17]. Intensity-dependent [LW04]. inter [LT00, LYT21]. inter-occurrence [LYT21]. inter-renewal [LT00]. Interacting [Ald13, ACG20, Ban24, CK11b, CJ22, ELL22, Fan21, FM00, Fis14, FMPV14, Kra21, LØ01]. interaction [CL17, DJL23, DGLO19, HKK19, JK17, Mau19, WLH16, ZL16]. interdirections [HP02a]. interest [Efr00, Fil00]. interface [AMPV10, AMV18]. interfaces [CD95, SSY21]. interleaving [PY19]. intermediate [Kur16]. Interplay [LT15]. Interpoint [ZS21]. interpretability [GHN21]. interpretation [TT20a]. Intertwining [CJ13]. interval [BG21, CLZ06, DW20, JS17, MR15, PT01]. intervals [Atc16, CCF22, CP01, FL11, HS07, MGS03, Ren08, SW12, TZ17]. intractable [ARV18]. intractable/infinite [ARV18]. intractable/infinite-dimensional [ARV18]. intriguing [DGC11]. intrinsic [Ste13]. Invariance [DMP11, DN14, KZ21a, MSS14, MZ17, OH23, Bad00, Bar18, BWW00, DW23, GL06, HW03, KJ20, Kre21a, SSY21]. invariant [BS15a, Deb99, Deh05, DJL23, DEGR14, ES99a, HAT10, KF14, KT18, LP02, LJ03, WXZ19]. Inverse [BKT19, CL24, KR24, LRD02, AW24, Bat17, BS15c, GP03, HMTvB24, ICG16, KS18, LS95, MH10, PY03, PBD15]. inversion [But24]. invertibility [LTL07]. inviscid [JMW05]. involving [PV07]. ion [DS03]. Irreducibility [WXX17, CA19]. irreducible [HTL07]. Irregular [Per02, PS20]. irregularly [Koi16, MMS08]. Ising [BM18, HKK19]. isolated [JK23]. isomorphism [Mon15]. isoperimetric [BH96, SW19a]. isoperimetry [AK18]. isotonic [DMP11, SW12]. isotropic [BMP08, BDKS19, Cap24, CS18b]. Itô [BR07, BS15a, FPS95, JLT08, JW21, Qiu18, Wal12]. iterated [ALV18, CD01, ES99b, GM04, KKW17, Kre21a, LX23]. iteration [ST13]. iterative [MM21].

Jack [CRT21]. Jackknife [PT18a]. Jacobi [GS11, Jia13, Kra21]. Jagers [BGMSB14]. Jeffreys [GP12]. jellium [CGZJ22]. Joint [CKHM23, JN95, PS22a, DPX23, DLT09, Kar24, LGS08, SK23]. Jong [DP18]. Journal [Ano95c, Ano95d, Ano95e, Ano95g, Ano95h, Ano95i, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano97d, Ano97e, Ano97g, Ano97h, Ano97i, Ano97j, Ano98c, Ano98d, Ano98e, Ano98f, Ano98h, Ano98i, Ano98j, Ano98k, Ano99c, Ano99d, Ano99e, Ano99f, Ano99g, Ano99h, Ano99i, Ano99j, Ano99k, Ano99l, Ano99m, Ano99n, Ano00b, Ano00c, Ano00d, Ano00e, Ano00f, Ano00g, Ano00h, Ano00i, Ano00j, Ano00k, Ano00l, Ano00m, Ano00n, Ano01a, Ano01b, Ano01c, Ano01d, Ano01e, Ano01f, Ano01g, Ano01h, Ano01i, Ano01j, Ano01k, Ano01l, Ano02a, Ano02b, Ano02c, Ano02d, Ano02e, Ano02f, Ano02g, Ano02h, Ano02i, Ano02j, Ano02k, Ano02l, Ano03a, Ano03b, Ano03c, Ano03d, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i]. Journal [Ano04a, Ano04b, Ano04c, Ano04d, Ano04e, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano04l, Ano05c, Ano05d, Ano05e, Ano05f, Ano05g, Ano05h, Ano05i, Ano05j, Ano05k, Ano05l, Ano05m, Ano06c, Ano06d, Ano06e, Ano06f, Ano06g, Ano06h, Ano06i, Ano06j, Ano06k, Ano06l, Ano06m, Ano06n, Ano07c, Ano07j]. jump [Bar22, BC21, BHVD17, BFM16, CI21, CR21, CCTY19, CRS00, DS03, ET23, FLLO14, FM02, Jou09, KKW17, KSJ19, Kra21, KS21, MR10, OU23, PJR16, SW19b]. jump-diffusion [BC21, FLLO14, OU23]. jumps [CDG14, Fou01, IPP23, Koi16, ŁS13, MV18a, Oud98, PS13b, PV09, Sch13, Wan11].

Kabanov [PSV00]. Kac [GD17, HKK19, IR20, JK17, NS01]. Kac-type [HKK19, JK17]. Katz [LS03]. Keller [TT20a]. Kemeny [PT18b]. Kendall [BD14]. Kernel [BDL^{+}23, CX24, CGK^{+}20, KD95, AG99, BOPG22, BP19a, BK02, BMT17, CS07b, CM14b, DGG13, DP16, DV01, FKM02, GCP17, GM04, GS20c, KC19, Lei06, LLD19, PLH09, SW19b, WK21, Zha96, vEGS07]. kernel-based [CS07b, DV01]. kernels [AF10, AF12, GS13, Sau19, ZGD24]. Kesten [AGH18, HL06]. Keyword [Ano95f, Ano97f, Ano98g]. Kiefer [Kul07]. killing [DV18]. kinetic [DRD20, HW19b]. Kingman [LS19a]. Klass [LV12]. Knight [RW22]. Kolmogorov [ER23, RR15]. Krein [Sto00]. kriging [PY01]. Kronecker [LP18b]. Kshirsagar [BW08].

L [ZWL14]. L-1 [ZWL14]. L1 [Lei06]. lag [HRY13]. Laguerre [Dem07, GS11]. Lamperti [CPR13]. Landau [RR15, WXX17]. Langevin [AS03, BCM^{+}21, DRD20, DM19, LE23, MFWB22, OPPS16, RT15, RT96, RSBG22]. Laplace [AW24, DZ03, Jac96, PY99, RS01b, dBM17]. Laplace-transform [dBM17]. LAQ [EM18]. LARCH [BS09]. LARCH-processes [BS09]. Large [AS03, Bha23, BDM10b, CG99a, FFP23, GZ18, GNSU19, GY04, HNZ18, KLW18, KZ21b, Lei06, NR00, PR05, Sen13, SvS17, TW14, XZ18, YZ01, ZZ15, BLP19, BNW98, BBS15, Ber01, BD20b, BBD17, BGT07, BDT21, BEMR17, BW19, DFK16, Fis14, GKP12, GNR19, GO00, HXZ18, HM20b, HM15, Kim21, KLS98, Kra21, KMB21, LW18, LW22, Mel00, MR21, NZ05, Pan22, Pen07, PS98, Sko96, SN13, Wan97, ZZW22, ZBY17, ZHZL19]. large-average [SN13]. large-deviation [GO00, Mel00, NZ05, PS98, Sko96]. large-dimensional [LW18, ZBY17]. Large-noise [PR05]. large-order [Wan97]. Large-sample [SvS17]. Large-width [FFP23]. larger [CP15]. largest [CP12, EX23, GS00a, HPY18, JJLW22, MRS22]. Laslett [Van04]. Lasso [CHS19, DHL17, HRBR15, HIW19, ZC23, BZ22, CCL23, NR12]. Latent [RS10, Bia14, BBW22, BBW23, DLL20, DLWZ17, DT23b, ER19, HL17, HMTvB24, LR23, MM15]. Latin [HT00b]. lattice [DH13, JK23]. lattice-valued [DH13]. Law [CD01, Sen13, AK22, BPZ15, ČV10, CD11, Dem07, GM04, GJQS20, GT04, GNTT18, GT01, Hir97, HLY20, KVZ19, KSW01, KLS98, Kre21a, LX23, Pat11, PS13b, QSSS04, Rob12, WHP18]. Laws [GS08a, KKW17, Pen07, AK23, AH13, BKR99, BD09, BCG14, ČK00a, CP16, Jam10, JS05, JS06, JMW05, LS03, Loc12, McG09, Wan97]. layered [HK07]. lead [HRY13]. lead-lag [HRY13]. Leading [ERS19]. leaf [HW19a]. learned [Dia13]. Learning [DLL20, MN22, MWZ10, AB11, BZ19, CJL^{+}23, GJS22, LXY13, QLC21]. Least [BDJ19, BC13, ABD14, ADM18, Cat11, CJP21, CV97, DMR21, FG19, Gir08, GT16a, HL12b, HY08, KW97, PG21, Tay13, VZ23]. least-angle [ADM18]. least-squares [Cat11, CV97, Gir08, GT16a, PG21]. Lebesgue [FY20]. lemma [CK01, DZ03]. lemmas [KK23a]. Length [CCN^{+}21, BH10, NL14]. lengths [HS97, Tod20a]. LePage [LV12]. Lépingle [ZK20]. Lepski [PG21]. Level [KN08, BP03b, BM05, Gor96, HP01, HW10, Lee22, LR95, Qia21b, RV09]. levels [GLF16]. Lévy [AIZ16, AC01, ADS13, BNT02, BNH08, BNBV13, BOHP18, BŁ15a, BN17, BPW19, BCI02, BM08, BOJGMR15, Bra14, BFM16, BLM19, CPP11, CI21, CY05, CR07, CD11, DKM16, DM21, FL11, FLT14, FK11, FLZ21, FPW17, GS16, GST08, GS00b, GS20c, HP14, HDS22, JVV05, Kab11, KMM24, KN08, KS20, KP22b, Küh22, KKR24, KP06, LLZ17, LSW20, Lim19, LCP13, Mai14, MS23a, Mal16, Man07, Mar06, Mas04b, NRT98, NR09, NS07, NS01, PSS24, PRvS13, PS22b, Pro24, RBM19, RW02, SSW12, SU08, Tra14, Wan16, XZ18, Xu21, ÝSJ08, ZZ15]. Lévy-based [ÝSJ08]. Lévy-driven [BNBV13, BN17, FK11, JVV05, Mai14, RBM19]. Lévy-type [Lim19]. life [KM23]. Lift [KM98]. light [KL05]. light-tailed [KL05]. like [BO18, Fra06, Fra07, MPU19]. Likelihood [LP99, Mar03, Myk01, RCBR00, Sev04, ALT15, AR17, AGM06, BS09, Ber06b, BBD17, Bia14, BS14, BMS21, BBC18, BS18b, Cha99b, CHWW05, CT22, DL09a, Dom15, DF19, DM01, DLWZ17, DMS21, DLS14, DR09, DL17, EM03, EM13, FH05, FCH15, FZ04, GZ09, GS18b, HS10, Hir16, JZ19, JJW24, Mai14, MYT06, McC08, NJ14, Ped95, PS13a, PT18a, QT05, Rei13, Ren08, Ruk14, TT06, Van95, Van05a, VW15, XW12, YB06, ZL03, ZWY17]. Likelihood-based [RCBR00, EM13]. likelihood-informed [CT22]. likelihoods [GW98]. LIL [ADS13, Cun15, GJS10]. Limit [AF19, AF10, AF12, BBCM11, BLP19, BT20, BKR99, BM12, BD22, DNN10, DL16, FPSS19, GNNS21, GCJL98, HSS07, Jia13, KY22, Mla08, PDL19, AP19, Akr00, ALOV19, BT17a, BEN12, BRR19, BP16, BTUA23, BOHP18, BM16, Beg07, BP03b, BB99, BL14, Bra99, BST12, CTP21, CL12, CF15, DS05, DJ97, EK11, EO07, FMPV14, Fra06, Fra07, GW17a, GP11, GV19, HLY20, IPP23, JS05, JS06, JJLW22, Kab11, KLLM13, KS99b, LR16, LS16b, LS19b, Lei19, LRMT17, Loc12, LS20, LTX22, MP10, MZZ20, MR10, Mou13, NZ05, NP08, OPPS16, Pak99, PR16, PR06, PT08, PZ19, Per02, PT00, PT03, PTA07, RT15, RR15, SW08, Tho23, Tur04, Ued21, VY13, Yan08, YZZ23]. limited [GG22]. Limiting [BBS14, BJSZ07, BDT21, BW19]. Limits [Wak95, AP00, AAI11, BD20a, DG13, DFK08, ELL22, Fan21, GS20a, GKT24, HM11, HKMS24, IMM17a, JMW05, KMS03, KMW14, Ma12, Nic18, PS22b, Sur04]. Lindeberg [BD09]. line [Van04, Yan08]. Linear [DPX23, MCJ23, AH17a, APGMV04, BLP19, BL04, BK23, Büh13, BC09, CL05, CV21, CP15, CHT98, CH08, CZH14, CJ22, CM13b, CNOH17, Dan00, DL19, DMP11, EBR23, FH05, GG22, GR23a, Gho99, GK18, GP03, GNT17, GR04, GK99, GK01, HLT01, HLY09, HTW23, HMTvB24, HMWY10, HNR23, JCL20, JN20, KSS21, Kul07, LR16, LR21, LM16, LAP20, LX21, MMW17, MOP20, NR12, NJG20, OP11, PXH13, PT00, Rát19, Sau18, TRB24, VZ23, WC14, WAP17, WC18, WH10, WFM10, WSZ12, YDN23, YZZ23, YLC19, ZGV20, AH17b]. linearity [KM23]. lines [HTM06]. linkage [Hös05]. Lions [CR22]. Lipschitz [BH96, BRS07, CS13, Cra15a, XY23]. LlogL [Ath00]. load [EMPW20]. Local [AK22, AIM07, BS21, BF01, BR96, BDM10a, BDKS20, CSY^{+}23, FB14, GP18, Gob01, HLY09, HLPŠ15, HLY20, KM24, LWCS22, MWX08, Ogi15, OU23, PS22b, RR15, ZL03, ABD14, ACP23, AGM06, BW23b, BRR19, BS02, BDP23, BDG07, CLZ07, CCF22, CHP11, CCJM24, Cav01, Ček04, CP21, CC00, CHT98, CCF^{+}20, CK00b, CSY99, DPX23, EK11, ES99b, Eis00, FLS20, FLLO14, FO24, GW17a, GLF16, GNTT18, JW21, KM20a, Ken15, LWLL22, LS16b, Lee22, LP00, MZZ20, NZ05, NIR21, Obl06, PY03, RS18a, Sar23, SCK23, Tho23, Vil14]. localization [CSY03, WYR21]. Locally [JLMP22, BZ23, Bou06, BV10, CCFH00, DP09, DRW19, Kur22, MM23, PR22, QP18, RvS19, XZ18]. Locating [HTM06]. location [ÁEdBCAM18, BGHL21, CD17, EHS05, FZ22, Gan21, KMS17, Mat00, OKM^{+}98, SGRV18, TT06]. location-scale [CD17, Mat00, TT06]. location-scatter [ÁEdBCAM18]. log [AGM06, BW14, BD18, BS21, BGI18, CG22a, CEG17, DRD20, DR09, GT16b, HR22a, KLLM13, Lud00, MRS20, MYT06, NR12, Yu10]. log-concave [BW14, BS21, CG22a, CEG17, DRD20, DR09, MRS20]. log-concavity [BD18, HR22a, Yu10]. log-likelihood [AGM06]. log-linear [NR12]. log-normal [GT16b]. log-periodogram [Lud00]. log-regression [KLLM13]. logarithm [CD01, GM04, GS08a, KKW17, Kre21a, LX23]. Logarithmic [SS20, BPZ15, GS20b, WHP18]. logarithmically [LBDM19]. logconcave [BCM^{+}21]. logistic [BRW00, ZSC22]. logit [DJL10]. Long [CFKR16, GWY13, SW19b, APGMV04, BT17a, BT20, BS12, BW22, Buc18, CT12, CS24, CM99, CNOH17, DNR20, DLPZ09, DLPZ12, HK08, ICG16, IKP18, LR16, LM20, Lud00, Mar06, OS15, PR02b, PS16, Sur04, Wu03, WZ18]. long-memory [APGMV04, BT20, CT12, HK08, IKP18, PR02b, Sur04, Wu03, WZ18]. Long-range [GWY13, BS12, BW22, Buc18, CM99, DNR20, DLPZ09, DLPZ12, Lud00, PS16]. Long-time [SW19b]. longest [AIS19, BMSZ21, LMV14]. longitudinal [CLZ18, CZH14, MSW13]. longitudinal/clustered [CZH14, MSW13]. longtime [JLM14]. look [KPT21]. Looking [BFT15]. Looking-backward [BFT15]. Loop [LW23, CDL24]. Loop-erased [LW23]. loss [ABF12, GN10, KN99, KMS17, LW18, Sar23, SC11]. losses [JJW24, SM23]. Low [BPW19, FFZ22, Klo14, Lec07, LM13, Mal09, NR09]. Low-frequency [BPW19, NR09]. low-order [Mal09]. low-rank [Klo14]. low-temperature [FFZ22]. Lower [LW17, Nov14, Sam98, APP11, Cas14, CBL22, CDG14, DFK08, LM10, SL18, dCQM09, GL22a]. lower-dimensional [SL18]. Lp [Hof99, KPT96]. LSE [BDK17]. LSL [GJS10]. Lyapunov [BS19, Hie17]. Lyons [RVW01]. Lyons-Zheng [RVW01].

M [FMSY12, SZ22, Kre21b, HP06, RS01a]. M-estimators [SZ22]. M/G/ [RS01a]. M/G/1 [HP06]. machine [QLC21, TV06]. machines [WY11]. Macroscopic [BCG19, BMSZ21]. magnitude [LLZ17]. Mahalanobis [HP02a]. Major [HNTX15]. majorant [BP11]. majorants [ABD14]. Making [MY02]. Malliavin [FO24, Gob01, GT16a]. Mallows [BD09, ICL19, JS05, JS06]. management [KK17]. manifold [CX24, GS22, PS13b, PR99]. manifold-valued [PS13b]. manifolds [BHL99, CGK^{+}20, LLBF24, MWZ10, New16, QP18]. many [BT17b, BL04, HJM14]. mapping [BZ14]. Marchenko [GT04]. Margin [AB11]. Margin-adaptive [AB11]. Marginal [DLWZ17, McC08, BBCM11, BSS05, BN05, CG99a, GLO22, KSW01, KL09, Røy11, Tru19].

Marginalizing [Kos02]. marginally [HLS19]. marginals [ACJ18, BDDF18, GMZ21, MY02]. margins [KW97]. mark [HLS14]. Marked [CZ13, GP18, Sch01b]. market [Ogi18, PX10, PR13]. Markov [ACJ18, AJDD11, ACBLV18, AI99, AR05, Atc16, Ath00, AAI11, BS22, BC06, BP19a, BC02, BR96, BC00, ČV10, CW19, CT02, CPR13, CPY13, CJL17, CA19, CH15, CM18, DS03, Dia13, Dob03, DMY00, Dor97, DM01, DMS07, DDL23, FS18, FGW12, FMPV14, FP17b, FH07, GS00a, GR14, GR16, GCJL00, GL06, GMM05, GW99, GT01, GHL05, GZ17, HM11, HVM99, HB04, HAT10, Hay22, HLP12, HMS04, HTL07, HM23, IR20, JN00, Jou07, Jou09, Kam18, KSW01, KM05, KMW14, KS21, Leh19, Lei06, LK08, LS20, MY02, MS10, MR17, Möh99, Mon15, OW17, PP09, PR02a, PHO17, PEBG17, PS13b, PT18b, RZ13, Riv05, Riv07, RS18b, SL14, SZ03, San13, Sch01a, SXXZ18, TA05, Tal13, Tka20, Tru20b, Van05b, VFJ18, Wal12, WK21, XZ09, Yan08]. Markov-modulated [BS22]. Markovian [AV14, Ber17, FN97, GZ18, KMM24, Lag10, MMP00, NP08]. marks [HLS14]. Martingale [BS95, JM22, KS01, Tud04, Yar19, BLO19, BGL22, DS96, GP11, Gra99, HT19, Jac02, LV17, MS23a, MPU19, Mou13, Ose20, PY03, Røy11]. martingale-like [MPU19]. Martingale-type [Tud04]. martingales [BS02, BRJN19, Cun15, EO07, FGLS19, JN16, KS99b, MPS98, MY02, Obl06, Osȩ07, Osȩ09, Osȩ11, PSV00, RS18a, WC14]. Maruyama [LTX22]. mass [BHBO17]. Matched [CDH^{+}98]. Matched-block [CDH^{+}98]. Matching [CLR20, ACZ23]. matchings [ICL19]. Matérn [BSW24]. mathematical [Cha99b, VV04]. mating [JMM17, MMR06]. Matrices [BY05, BWZ10, BHPZ15, BLP19, BBS14, Bat17, Blo04, BW08, BDM09, BDT21, BPD23, BX15, BS23, CM13a, CP12, CP15, Dia23, Din20, DZ24, EX23, GNT17, HXZ18, HPY18, HM19, HYY11, HMRR20, ICG16, JHD20, JB21a, JJLW22, JR17, KLP21, LW22, LX21, LEM23, Ma12, MP06, ST18, SvS17, WAP17, WHP18, YZZ23, ZZW22, ZBY17, ZHZL19]. Matrix [LLL22, Ass23, BTT19, BX15, BI13, CKLN18, CGS18, CJP21, Din20, FP17c, GNR19, GS00a, HL17, HPY18, HZBH23, HT21, JMRX18, KLW18, Klo14, KLTZ19, KG00, LW18, LWYZ22, LXY13, Lou14, MYT06, RCBR00, SN13, TC03, WZ16, Yin22]. Matsumoto [KV12, MW04]. Matsumoto-Yor [MW04]. Max [GK18, Ued21, BDKS19, DK18, KS16, KZ21b, LBDM19, OSZ18, OS15, SS15]. Max-convolution [Ued21]. max-increments [OS15]. max-infinitely [KS16]. Max-linear [GK18]. max-stability [KZ21b]. max-stable [BDKS19, DK18, LBDM19, OSZ18, SS15]. Maxima [EKS15, OS15, BS18b, CCF^{+}20, Dom15, DF19, GY04, Has13, JN95]. Maximal [Pro24, BP11, BC20, BFM16, FP17a, KM23, KM11, Osȩ11, PS12, SN13]. Maximum [BS18b, BDDF18, CCTY19, DF19, DLS14, DR09, FZ04, VW15, AR17, BS09, Bia14, BBC18, CHWW05, DKS01, Dom15, DM01, DMS21, FGK16, GZ09, GS18b, HS10, LSW13, Mai14, MR10, Obl06, Ped95, TT06, Van95, Van05a, YB06, YG19, ZWY17]. maxisets [KP02]. McCloskey [Hir97]. MCD [CL12]. McKean [DJL23, TT20a, TY23, Wan23a]. MCMC [AB15, AF10, AF12, BCFK15, ŁMN13, MCC^{+}21, QJ22]. Mean [DJL23, ELL22, TC24, Ald99, Alm03, ACV09, Bah21, BvdHH13, CLLLR17, CGCR07, DGP18, DM24, EZ03, GR00, GT01, HJQZ04, HK13, Jam10, KMS03, LSY^{+}22, MS23b, Mas04a, Nie13, PR99, PT18b, Tan15, TMW13, Vil14, Vil23, YRZ24]. Mean-field [DJL23, Ald99, BvdHH13, EZ03]. means [ACG19, ACG20, AM16, CY21, CCT20, CNOH17, DH13, Elt22, EHHS23, JLP10, LMMR16, LLL22, LM19, YP23, ZP19]. measurable [LV17]. measure [AJLS18, BR18a, BGI18, Blo04, BH00, CJL^{+}23, DS96, DMY00, EHS05, FGW12, GS22, Gra99, HLL23, Hei18, LS16a, Lee22, Lin19, NZM23, Ngu16, Pai22, PR02a, TT20b, WXZ19, dSJMT23, vDD21]. measure-valued [DMY00, FGW12]. measurement [CHI17, JKNP16, ML10, MCJ23, MRST07]. measurements [ACP23, BKY21]. measures [Ath00, BGMZ18, BNH08, BS15a, BH96, BK23, Bra22, BWJ18, CL05, CHHB19, CG22a, CCN^{+}21, CRT21, DMZ11, DJL23, EZ03, FLS01, Fis14, GNR19, GS13, HW95, HNR20, JM22, KPT21, Kol08, KT18, Lau13, LC22, LMC23, Lei20, LO14, LNP14, LS14, MSWW11, Olo09, Roo10, Sri16, SDS24, TW14, WB19, ZGD24]. Measuring [BS11, Wie22]. median [Alm03, CCZ13, Cha99a, LM19, Min15, YP23]. median-of-means [LM19]. medians [Pai22]. medical [VV04]. meet [MY02]. meets [Nda23]. Meixner [GS11]. Mellin [HY13]. membership [CK23]. Memory [CCN^{+}21, APGMV04, BT17a, BT20, BBJ08, CT12, CNOH17, FRS07, HK08, ICG16, IKP18, KLLM13, LV15, Mar06, OS15, PR02b, Sur04, Var19, Wu03, WZ18]. mesoscopic [LX21]. meta [BEN12, Ruk14]. meta-analysis [Ruk14]. method [AHK15, BIZ02, BRR19, BOPG22, BBD17, BDV21, BV10, Ček04, Cha97, CN15, CGER07, Dai17, DF19, ER23, EKS08, ERS20, Fan14, Gau17, GR23b, Hir97, KVD04, KKR24, LLBF24, LGW06, LRD02, MPS19b, OCBG19, PG21, Rig06, Riv06, Sch09, VBLN20]. Methods [KK16, BSG22, BZ19, CCFH00, CT22, DDJ12, DV01, HP95, HW99, HMW01, HB04, HLPP09, Hoo16, KC19, KK17, ŁS13, PP05, PW04]. metric [BSW24, Bor21, Bru19, CDL24, CGK^{+}20, TT20b]. metrics [BBV17, BLL15, Sch09]. Metropolis [BP19a, FB22b, HST01, JLM14, MV18b, SR09, Wan22a]. microarray [LW04]. Microscopic [HM20a]. microstructure [Ogi18, PV09, VD12]. mid [Wan24]. mighty [GGKL22]. Milstein [BO18]. Mimicking [FHK15]. Minima [Has13]. Minimal [GMZ21, HAT10, Mat00, Gra99]. minimal-entropy [Gra99]. Minimax [AAL23, BM05, CDH21, EHS05, Gay01, GL22a, GL22b, KP02, Kim14, LWLL22, LS99, LM15, YKK21, YLC19, ADN24, Bar02, Ber04, BBW20, Dic16, DJ96, JLL04, KN99, LWCS22, ML20, MS12, Oud98, RST17, Riv06, SC23, Tan15, XL10]. minimization [BIPZ22, CS07b, KKS02, LM10, Lec13, LM16, Sau18]. minimize [VW15]. minimizing [KK21]. Minimum [BM98, Gan21, SY24, BHM18, Bra22, HW99]. minimum-distance [HW99]. Mirror [DT12]. Mises [Gru00]. missing [BHBO17, JR17, Lou14, MS17, Van95]. misspecification [BS21, CAA22, FG19, NRC24, NPM21]. misspecified [Bac18, Fus05]. Mixed [CHI17, Che01, Gut01, AH17a, CK23, DJL10, DT23a, FO24, Gob01, Mar13, Ogi15, Sad13, SL14, WSZ12]. mixed-model [Mar13]. Mixing [BRE23, DLN22, FR11, Gir08, PDL19, BB99, Bra99, BK16, DV18, DL17, HMTvB24, HNR20, Kre21b, KL13, RS96, Zäh15]. mixture [AIM07, BD18, BFMP01, BTV17, DCG97, GHN21, HNR20, LLD19, LMK19, MPV22, NMR20, SST16]. mixtures [AL05, AGM06, BL23, BB17, CD17, CR17, Dit19, DRRS18, FP17b, JZ19, Kim14, LMMR16, MR16, OL23, Sch10, TT06, Van05b]. MLE [TCP22, ZSC22]. MMD [CAA22]. mobile [DFK16]. Möbius [KM20b]. modal [BG21]. Mode [BGMSB14, YG19]. Model [AW12, Bir04, DR11, DS03, NJ14, VD12, WP22, APGMV04, AB11, BCFH22, BW23a, BDJ19, Bar10, BS21, BKK23, BFMP01, BC13, BvdHH13, BL10, BEMR17, BM05, BC09, CK23, CV21, CW19, CHWW05, CHI17, CPY13, Che24, CC04, Cor16, CD95, DCG97, DNR20, DHKR15, DLWZ17, DLT09, EM18, GK03, GG22, GW98, HPY18, Hay22, HKK19, HM20b, HY08, HYY11, ICL19, IS98, JK17, JT22, JD16, JMRX18, JJW24, JS13b, JV01, KVD04, KLW18, KW97, LR21, LW17, LT15, LMK19, LP24, LV15, MMVW10, Mar13, MN22, Möh18, MPV22, NC24, NPM21, PP16, PY19, PGV17, Sar16, SSY21, TT20a, TGL23, WM16, Wu09, Zha00, ZLL17, ZGV20, NWR22]. Model-free [WP22]. modeling [ACLZ17, GHN21, HMP18a, LM07, MSW13, MH22, NZM23, PR13, SY24, YMP11]. Modelling [BNBV13, BGMSB14, MMS08, TC03]. models [AFGS08, AH17a, ARV18, Ald99, AKT10, AK97, AIM07, AJLS18, BD13a, BS19, BNS03, BGI18, BC13, Ber06b, BDPR21, BBD17, BBD20, BM18, BG17, Bia14, BSS05, BR96, BBW20, BG10, BDM09, Büh13, BKT19, BTV17, CS18a, CGLZ24, Cha97, CV09, CZH14, CA19, CK00b, CV19, Cra15b, CM14b, CM18, DJL10, DS03, DT23a, DLR06, DPSW10, DLL20, DMR21, Dob03, DM01, DNS19, DvdAW09, Drt09, DLWZ17, DRW20, DMS21, DL13, EBR23, EM18, ERS19, ER19, FH05, FLLO14, FZ22, FH07, Fus05, GR14, GR16, GCJL98, GCJL99, GCJL00, Gho99, GD17, GK18, GLS07, Gut01, HJL02, HLL23, HTW23, HAT10, HS10, HM19, HL12b, HNR20, HM20c, HIW19, HKM06, HZ04, HLY20, JW19, KM23, KT15, KL00, Kos02, KS97, KKP21, KP20, LLZ17, LMV22]. models [LC22, LMP12, Lee08, Leh19, LK08, LP99, LLT12, LLD19, LTL07, LLS11, LSY^{+}22, LPV13, ML10, MSS16, MR19, ML11, MM15, MV04, MMW17, MYT06, Mat00, MR13, Möh99, MNPR18, MS17, MZ17, MV99, NK21, NMR20, NP08, Ngu15, NC17, NC21, OCDM08, OW17, Pon07, PBD15, RS17, RW19, RT15, RS10, Røy11, SST16, Sch10, SS12, Sch20b, SV13, TRB24, Tra14, Tru19, Tru20b, Van95, VVV06, VV02, Van05a, Van05b, WXZ15, WC18, WC98, WSZ12, Xia09, XW12, YB06, YDN23, ÝSJ08, YLC19, ZWL14, ZL16, ZWY17, ZSC22, ZS12]. Moderate [EZ03, BGL24, CS24, FGLS19, FLS20, FLS23, GKP12, GJZ23, GT18, JWY22, LSW13, LTX22, SZ16, Zha23]. moderately [WAP17, YK20]. modifications [MW20]. modified [AT24, PR06]. modulated [BNBV13, BS22, BS16, BE16, ZL13]. moduli [LX23, LWX15, WSX20]. Modulus [BL15b]. molecular [GP04]. Moment [HWXZ24, KT18, KY07, DF20, Gut02, JB21a, JB21b, KL17, Sto00]. Moments [BL23, PSS24, BP06, DL09b, EKS08, Kol08, MMVW10, MVY19, Osȩ11, SL18, WSZ12]. Monge [HMRR20]. monitoring [ADFS15]. monotone [AF19, AH13, BDJ19, Cat11, CG22b, DNP06, XZ18]. monotonic [BRS07]. Monotonicity [Alm03, Jan07, ABD14, BGMSB14, CGS18, GJ13, KP23, Yu17]. Monte [BBLG17, BRE23, LBBG19, ARV18, AJDD11, AR05, Atc16, BB22, BPR^{+}13, DS03, DDJ12, Dia13, ELP24, GC17, GLO22, GRS11, GT01, HVM99, Kam18, NIR21, OCDM08, PJT19, SZ03, VFJ18, WLH16]. mortal [Pak99]. MOSUM [EK18]. motion [AD18, Ass23, BM15a, BP06, Bar22, Ben03, BP10, BR12, BDG07, Che01, Coe05, CRS00, Dan00, Das05, DS06, DS19, DMY00, ES99b, FR06, FR19, Gra21, Kat09, KJ13, LD11, LS02b, MOP20, NPS20, Obl06, Pey17, PT00, PT01, Pit97, Sau12, SW19b, TX07, VY13, WYZ21, Zei15, ZX24, AH17b, LP18a, Coe06]. motions [ACPM24, BŁ15a, BCI02, DS05, DP11, HWXZ24, JM09, LV12, NVV99]. moving [BPW19, CT19, KL05, LLT12, LTL07, Mar06, PT03, SW04, Sur04, Syn07, Tsa09]. moving-average [LLT12, LTL07]. much [CP15]. Multi [Zha13, Alt13, CDH21, Gay20, Gay22, Kut21, KS01, Olo09, SvdMvZ17, TRB24, Zha06]. multi-class [CDH21, Gay20, Gay22]. multi-dimensional [SvdMvZ17]. multi-scale [Zha06]. Multi-stage [Zha13]. multi-step [Kut21]. multi-task [TRB24]. multi-type [Alt13, KS01, Olo09]. Multicolor [BDM09]. Multidimensional [Ham03, KP22b, BP03a, CdLOR20, KVZ19, Mas04b, Nic18]. Multifractal [ALS09, DL16, GL05, LS14]. Multifractional [AH17b, BDG07, BDM10a, Coe05, Coe06, DPX23, LV12, MWX08]. MultiGraphex [NRC24]. Multilevel [BN17, HB23, LP17]. multilinear [GNR14]. multimodal [PJT19]. multinomial [Ber03, DJL10]. multinomials [LWLL22]. Multiple [BS18a, CCF^{+}20, Pec08, BW23a, Che02, CN15, DL19, DR16, EK18, FB22b, HW10, HLPŠ15, HMP18a, HS10, LL00, Möh06, SCK23, WXZ15]. multiple-index [WXZ15]. multiple-output [HLPŠ15]. multiple-stable [BW23a]. multiple-try [FB22b]. multiplicative [Lim19, MCJ23, Rát19, RB06, Ued21, ZZ15]. Multiplier [BD13b, Han22, BK16]. Multipower [BNCP11]. multiresponse [CV09]. multisample [HL12b]. multiscale [NC24]. multistable [LV12]. multistep [ILY09]. multitype [GMM05]. Multivariate [CG06, CSvS07, DDW13, DT23a, ET23, FSTD20, GS11, KL10, KP22a, RT06, SS12, Ste10, ACBL15, ABF12, BPB^{+}22, BNMS06, Ber06a, BL14, BLM19, CD17, CL12, Cha99a, CR07, Cut00, DCC14, DGS08, FR15, FK17, GL08, HP02a, HPY18, Han03, HRBR15, IKP18, JMRX18, JK12, Kim21, KKS02, KR24, KS99b, MPAiS12, ML11, Mas04a, MP23, NJG20, PS22a, Pai22, PBD15, Rai19, Reb15, Rob08, SZ22, Tan15, VFJ18, WT13]. mutation [BR18b]. mutivariate [Gao20]. myopic [Kuj16].

naive [BL04, TCP22]. Navier [ZZ15]. Near [FN23, MFWB22, EK11, GJZ23, Mal16, Möh18, WHP18]. Near-optimal [FN23, MFWB22]. nearest [CLZ06]. Nearly [Bah21, DvdAW09]. necessarily [Zho14a]. Necessary [CG23, JT03, MS23a]. Needles [BN20, CT19]. needlet [BKMP09, KLP12]. needlets [DMT24]. negative [BGX15, Ber03, CY05, DvdAW09, GY04, HR16, LCP13, Osȩ11, Pro24, San10]. negativity [BR12]. neighborhood [LS19b]. Neighbour [MZ17]. Neighbour-dependent [MZ17]. Neo [DJ96]. Neo-classical [DJ96]. Nested [CM18, DRW20, GP04, GZ09, JJW24]. Nesterov [MCC^{+}21]. nets [FSW17]. network [Cra15b, KRX18, KK23b, Kre21b]. Networks [ACG19, ACG13, BES07, DFK16, DZ17, FFP23, KK21, MN22]. Neumann [BVM07]. neural [DGLO19, FFP23, KK21, KK23b]. neutral [BR18b, Möh99]. Nevanlinna [PS13c, PS11]. Newton [NS07]. Neyman [CK01]. nilpotent [CD01]. no [Ano95c, Ano95d, Ano95e, Ano95g, Ano95h, Ano95i, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano97d, Ano97e, Ano97g, Ano97h, Ano97i, Ano97j, Ano98c, Ano98d, Ano98e, Ano98f, Ano98h, Ano98i, Ano98j, Ano98k, Ano99c, Ano99d, Ano99e, Ano99f, Ano99g, Ano99h, Ano99i, Ano99j, Ano99k, Ano99l, Ano99m, Ano99n, Ano00c, Ano00d, Ano00e, Ano00f, Ano00g, Ano00h, Ano00i, Ano00j, Ano00k, Ano00l, Ano00m, Ano00n, Ano01a, Ano01b, Ano01c, Ano01d, Ano01e, Ano01f, Ano01g, Ano01h, Ano01i, Ano01j, Ano01k, Ano01l, Ano02a, Ano02b, Ano02c, Ano02d, Ano02e, Ano02f, Ano02g, Ano02h, Ano02i, Ano02j, Ano02k, Ano02l, Ano03a, Ano03b, Ano03c, Ano03d, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano04a]. no [Ano04b, Ano04c, Ano04d, Ano04e, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano04l, Ano05c, Ano05d, Ano05e, Ano05f, Ano05g, Ano05h, Ano05i, Ano05j, Ano05k, Ano05l, Ano05m, Ano06c, Ano06d, Ano06e, Ano06f, Ano06g, Ano06h, Ano06i, Ano06j, Ano06k, Ano06l, Ano06m, Ano06n, Ano07c, Ano07j, BL13]. Nodal [Tod20a]. nodes [AD07]. Noise [TRB24, ACBLV18, BM05, DNN10, FP17a, Fou01, GM15, GJQS20, GLT96, HP02a, HP14, HDS22, HW24, IS98, KN08, KM95, KMS03, Koi16, KP22b, KKR24, Lec07, LSW20, MH22, Ogi18, Pan22, PR05, PV09, SSX20, VG18, VD12, XZ18]. noises [BYY17, FT19, WXX17, WYZ24, ZZ15]. Noisy [Klo14, BKY21, BI13, CDJ13, LN14, Zha06]. Non [Bar02, CP16, DL13, LMMR16, LRMT17, NZM23, PR02b, ZZW22, ALOV19, BW23b, Bar10, BP03b, BJSZ07, BR12, BG10, BRS07, CZ13, CGL12, CY21, CH08, CS13, CJ22, DH17, Deb99, DVSS16, DL09a, DS19, DMR21, DLN22, DvdAW09, EKS15, FN97, FFZ22, GS16, GM15, GP23, Gne13, HM11, HY05, HKK19, HRY13, ILY09, Jac06, JK17, JK23, JJW24, KMB21, LS02a, LAP20, LLS11, MZZ20, Osȩ11, Qiu19, San10, SS18, Seg12, Syn07, VZ23, WC14, WSX20, Wan23a, WZ18, XY23, XL10, YP23, Zho14a, Zho14b, vDD21]. non-Archimedean [Qiu19]. Non-asymptotic [Bar02, DL13, LMMR16, ZZW22, DMR21]. non-cascading [FFZ22]. Non-central [LRMT17, ALOV19, BJSZ07]. non-commutative [DS19]. non-differentiability [WSX20]. non-dissipative [Wan23a]. non-ergodic [Jac06]. non-Euclidean [CY21]. non-existence [LS02a]. Non-Gaussian [CP16, Bar10]. non-Gibbs [HKK19]. non-Gibbsianness [JK23]. non-hereditary [DL09a]. Non-homogeneous [NZM23]. non-hyperbolic [GM15]. non-i.i.d [BG10]. non-identically [EKS15]. non-increasing [HM11]. Non-informative [PR02b]. non-instantaneous [CH08]. non-integer [DH17]. non-linear [CJ22, LAP20, VZ23, WC14]. non-Lipschitz [BRS07, CS13, XY23]. non-local [BW23b]. non-Markovian [FN97]. non-negative [DvdAW09, Osȩ11, San10]. non-negativity [BR12]. non-nested [JJW24]. non-normal [Zho14a]. non-normality [BJSZ07]. non-parametric [XL10]. non-polar [Deb99]. non-positive [YP23]. non-regular [GP23]. non-restrictive [Seg12]. non-smooth [GS16, SS18]. non-stationary [BP03b, CZ13, CGL12, DVSS16, DLN22, ILY09, LLS11, MZZ20, Syn07, WZ18, Zho14b, vDD21]. non-strictly [Gne13]. non-synchronous [HRY13]. non-synchronously [HY05]. non-uniform [KMB21]. Nonasymptotic [GD17, ŁMN13]. noncentral [DP18, HZBH23]. Noncommutative [FY20]. nondeterminism [Lee22]. Nonhomogeneous [GJS22, XY23]. nonignorable [ZWS21]. Nonlinear [BK02, DNV99, Riv06, AJDD11, BVM07, BHL99, CS18a, CS07a, DRW19, DNN10, DL13, FK12, FKM02, GCK04, HP01, HT21, KY22, KS97, OCDM08, PP05, SZ22, SRR24]. nonlinearity [LLTY07]. nonlocal [HDS22]. Nonmonotonicity [HR22a]. nonnegative [IKP13]. nonnested [WM16]. Nonparametric [Abr19, AAG24, BG09, BGSS22, BHVD17, CL02, Cho18, CC19, CM11, DL17, FKA05, GR16, GMR19, HB04, HV09, HP06, HF99, JW09, JVV05, JK12, KS21, Kur22, LMNV11, LSY^{+}22, MM23, MRST07, NR09, Qia21b, ST04, Sau14, VSV03, VD13, Yan01, Zho14b, BGL22, CLZ07, CD17, CLZ18, Car06, CW19, CHWW05, CV19, CGCR07, DJL10, DNP06, DF98, FLP12, GP15, GCP17, Gol99, HM20c, JZ19, KFT21, KS18, KN99, KSJ19, Leh19, LS99, LPW04, MVY19, MRS20, Mei16, Pon07, PR13, QT05, RS17, RW19, SW13, SCK23, TAS^{+}17, Van95, VS05, YMP11, Yu17, ZWL14, vEGS07, vdMvZ13]. Nonparametrically [PV15]. nonparametrics [FGW12]. nonresponse [ZWS21]. nonsmooth [CCT20]. Nonstationary [LRS23, ACLZ17, KR24, MPU19, MS23c, SPP21]. nonsynchronously [Ogi15, Ogi18]. Nonuniform [LS24]. norm [Ber04, DVSS14, DVSS16, GN10, KP23, KN99, LEM23, LMZ20, LP07, Nau22, Yin22]. Normal [CS07a, DH13, LP24, AR17, AG07, BS23, ČK00a, CRT21, CSSC95, CCT20, DF98, Fan14, FR15, FP17c, Gau17, GT16b, HL14, Kim14, KW97, Sch10, SBP24, Tan15, Van05b, Zha23, Zho14a]. normality [AT24, BJSZ07, BR96, CG23, CV97, FO24, Gho99, Gob01, GLS07, HZ04, HY08, Jan07, KC19, NIR21, Ogi15, OU23, Ped95]. normalization [LW04]. normalizations [NRT98]. normalized [CP12, CP15, DJ97, FGLS19, LNP14, LSW13, LTX22, OSZ18, SZ16]. norms [GL22a, GL22b, HMRB08, Lep16]. note [BD09, BJ22, CT02, DZ03, FMNW18, GS23, Gru00, HU22, KM11, KS99b, LS16b, LW04, TC03, Xia09, ZX24]. notion [BB99]. novel [GGKL22, GP04, LLL22]. nugget [LS23]. nuisance [VK97]. null [BDR14, DH02, Dob23, Fer01, SY00]. number [AW09, BBS16, BBW20, CS18b, CP23, GR14, IMM09, KVZ19, Kar24, Kim21, KW20, OL23]. Numbers [Sen13, KLS98, Pen07]. numerical [Ama13, GK12, VT13]. Numerically [OD19].

objective [EF04]. observation [Cho18, MRS22]. observational [Che02]. observations [BL04, CI21, CSvS07, DDBM21, FT19, JN20, JR17, Lou14, MV18a, NR09, PP05, WT21, Zha06]. Observed [MV99, AIZ16, BS95, BPD23, BN05, DHKR15, FP17a, GCJL98, GCJL99, HY05, Jac06, KS99a, Koi16, KSJ19, Mar03, Ogi15, Ogi18, Ped95, Sør02, vdMvZ13]. observing [IS98]. obtaining [BV10]. occupancy [BGI18, BHBO17, DGP18]. Occupation [BGT07, Alt21, DMY00, PY03]. occurrence [LYT21]. occurrences [Wu09]. O’Connell [ARAS20]. off [ACPM24, DJ97, PP22, Ros09]. off-the-grid [PP22]. offs [CLZ07]. Olkin [HJ17]. One [BRS07, AP19, AMPV10, BJ22, But24, CGZJ22, FP10, GS23, JMW05, KC19, Mac23, PR05, PY99, PY03, RTT20, SW19a, Sau19, Sko96, SV13, TT20a, Wan97, Wan23b]. One-dimensional [BRS07, AP19, AMPV10, BJ22, CGZJ22, FP10, GS23, JMW05, Mac23, PR05, PY99, PY03, RTT20, TT20a, Wan23b]. one-parameter [Sko96]. one-sample [KC19]. Online [BC21, BCFK15, CM18, DR07, GLO22, GS21, OW17, OD19, SK23]. operational [ABF12]. Operator [NN22, SW20, AD18, AGH18, Del17, DP11, Giu18, HG97, KM20a, LWX15, LEM23, Mar13, dCCGZ23]. operator-based [Mar13]. Operator-scaling [SW20, LWX15]. operator-self-similar [KM20a]. operator-theoretical [dCCGZ23]. operators [AB15, ALL09, AGG23, BK23, BJ22, CGER07, KG00, KL17, KR24, UCV17]. Optimal [BGHL21, Bel17, BPR^{+}13, CM13a, CV21, CN15, CHT98, DPS24, DPW09, DPSW10, Düm03, FMR19, GGM20, HP02a, HPV13, HKM06, JLM14, KMM24, KL02, Lec07, LW18, LN18, LMV14, LØ01, Mei16, MS18, OL23, PP05, PR14, RV09, SR09, SC23, STC^{+}24, TMW13, WYR21, ASC02, Bah21, BP03a, BM07, BIPZ22, BS12, BBW20, BK23, BDKS20, CCF22, CHP11, DS96, DLPZ09, DLPZ12, EBR23, FN23, Gay20, Gay22, GL22b, Gra99, GY04, HMP18b, Hof23, HM20c, HT19, HKMS24, ILY09, KVD04, KW13, LWLL22, Lec13, LBDM19, MFWB22, NV17, PS12, SK23, Sri16, Yan24, YLC19, Zha22]. Optimality [Cho18, Jac02, ADN24, Cat11, GW99, Kuj16, LM13, Sau18, SW04]. optimally [HM20a]. optimization [Dor97, Lin09]. Optimizing [Hin03]. optimum [MH10]. option [PP09, WY11]. options [EI18]. Oracle [CBL22, LYG19, Kol09, LT16, Lev18]. order [BS19, BP06, BG09, Ber97, BK23, BDDF18, Cas07, CG23, CSSC95, DCG97, DL19, DDL23, ERS20, FFS10, GS20b, HNR23, JN95, JM22, Leh19, LP18c, Mal09, OPPS16, PB15, PS09, RSBG22, Sch01b, Tal13, US24, Wan97, WXZ19, ZS15, dBCAM22, vDD21, vDD22]. order-restricted [CSSC95]. ordered [BO20, Cha12]. ordering [EM13, WT21]. orders [Fri24, GL18, LV17]. ordinal [AC17]. ordinary [BG17, GK12]. origin [Jac96]. Orlicz [SST16]. Ornstein [BPD23, BNBO11, BS22, Fas10, FK11, GLST19, HP14, Jac96, JVV05, LP00, Mai14, Mas04b, Nku21, PS09, Wan11]. orthogeodesic [CK00b]. Orthogonal [GS13, HYY11, JHD20, PW04, ST20]. orthogonality [PR99]. Oscillating [LP18a]. oscillation [AW96, Möh18]. oscillations [ES99b, JN15]. oscillator [MPS19b]. oscillators [LP00]. Osgood [FN21]. other [NVV99]. out-of-sample [Lee08]. outcomes [KVD04]. output [HLPŠ15]. outputs [CM13b]. Over-parametrized [KK21]. overparametrization [GR04]. overview [Gin96a].

p [ADH21, KP23]. P. [GS00b]. p.m.f [BDK17]. packet [GK03]. packing [SX10]. pair [Haf13, HP99, Kat09, Yu10]. pair-copula [Haf13]. paired [DLL20, WT21]. pairs [Mla08, Zha23]. pairwise [CL17]. Palm [HLS14, Tho99]. Panel [CK00b, HLZZ23, LSY^{+}22, SGL^{+}24]. Papangelou [Tor17]. parabolic [BS19, CD23a, Fou01, LS02b, MPRZ10, Mor99, TT20a]. parabolic-parabolic [TT20a]. parallel [McC08]. Parameter [ACP23, BHM22, CPY13, FV03b, GCJL99, Haf13, Bac18, BBJ08, CJL17, CCFH00, CM18, DL19, FRS07, FR06, GK12, Han09, Hay22, HMS04, HRY13, KK18, KLLM13, LR21, MS12, Mat00, OCDM08, SSH14, Sev04, SK23, Sko96, TC03, Van05b]. parameter-dependent [Sev04]. parameterized [SS15]. parameters [BD13a, BP19b, Cha97, DR07, EHS05, GHN21, HKM^{+}16, KK16, LS23, PR99, Sør02, TT06, TZ17, Tru20b, VK97, ZWY17]. Parametric [CL17, Jac06, Lud00, Ogi18, AK97, AAGGR03, BBD17, BRW00, Cas07, CV09, CGL12, CCFH00, CAA22, FZ22, HW03, HL12b, Hir16, HLP23, JV01, KLLM13, MV04, MV99, Pon07, Sar16, WXZ15, XL10, YMP11, Yu17, Zha00]. parametrix [AKH17]. Parametrization [LJ03]. Parametrization-invariant [LJ03]. Parametrized [AJLS18, KK21]. Pareto [FdH14, KL09, RT06]. Pareto-type [KL09]. PaRIS [OW17]. Parisian [LCP13]. Parking [CG21]. Part [AF12, YMP11]. Partial [Cor14, JB21b, APGMV04, BYY17, BE16, BSG22, FN21, Fou01, GZ09, HR15, HDS22, Lau04, Mar03, MK20, Mor99, PT03, XZ18]. partially [BPD23, CZH14, FH05, WC18, YLC19]. Particle [CM14b, Kün13, Ald13, ALV18, CLZ06, CS15, CM18, GD17, GM19, HB23, IK22, OW17, OD19]. particle-based [OW17]. Particle-kernel [CM14b]. particles [BS18a, GL18, Mac23, Rév96]. particular [CK11b]. Partition [Pit97, Cra15a]. partitions [BFT15, BO20, DW19, You05]. partly [GGT22]. parts [CI21]. Passage [CY05, DKM16, MMVW10, But17, CD11, GJS22, HP14, McG09, PT18b].

Passage-time [MMVW10]. passively [JV01]. pasting [SS18]. Pastur [GT04]. path [BN17, BPR12, BDG07, CSY03, CR22, GS22, HM20a, LMV14, LS20, TX07]. path-dependent [CR22]. paths [BS11, BŁ15a, BBS16, BPR06, BDP23, Küh22, OR24]. Pathwise [BS15b, Pan22, PP16, CM11, Sio18]. Paving [AB22]. payoff [GHL05]. PDE [HNR23]. PDEs [CD23a, CR22, Fan21]. PDMP [RT15]. peaks [CCF^{+}20, Kar24]. Pearson [APP11, CK01, LPSS18]. Penalisation [RTT20]. Penalization [ŁS13, Pet97]. Penalized [CGCR07, HHH24, RB06, AM17, Gay20, Gay22, Xia20, ZL16]. pendant [BL13]. Percolation [CDL24, ACG13, GT19, GJS22, HU22, TGL23]. Perfect [MK20]. Performance [LM16, CL05, Cas14, CT22, DHL17, LM10, Yan04]. perils [BM13]. perimeter [Vil23]. Perimeters [BH17]. period [Cas07]. periodic [BG95, Fra06, Fra07, Syn07]. periodically [Len11]. periodogram [BP16, CMS10, Lud00, PZ19]. periodograms [GK13]. permutation [FR15, PT15, RP23]. permutations [BMSZ21]. Perpetual [Küh19, KS20]. perpetuities [BG19, Erh14, HW11]. perpetuity [HY13]. Persistence [GR04, HKM^{+}16, OB20]. persistent [CdLOR20]. perspective [ML20]. Perturbation [Las14, RS18b, DL19, GCK04, Tru20a, Tru20b]. perturbed [CG22a, Dan00, IPP23, Kut15]. Petersburg [CS06, FGK16]. Petrosian [dUÁV21]. phantom [JRSK21]. Phase [CJ22, CK23, HU22, JLM14, LR23]. phenomenon [ACPM24, Elt22, LLL22]. Phi [Dit19]. Phi-divergences [Dit19]. photon [MRST07]. physical [ZC18]. Pick [PS13c, PS11]. Pickands [GP16, Har17]. picture [PS11, PS13c]. Piecewise [ACLZ17, BB22, KS21]. piecing [ABF12]. piecing-together [ABF12]. pinball [SC11]. pinned [GS22]. Pivotal [vDD22, Cha12]. pivots [CNOH17]. placement [SK23]. planar [Stu23, VY13]. Planck [HW19b, MNN^{+}22]. plane [BJ00, CG21, Dor98]. plans [Yu12]. planted [BCvdHV21]. plants [Hin03]. plots [BDG05, DG13, DCC14]. plug [RV09, SW04]. plug-in [RV09, SW04]. Poincaré [AS19, CEG17, GS23, HT21, Jor23, SW19a, Sau19]. Point [Tho99, BD13a, BM15a, BAT19, BTUA23, Bei97, BW22, BL16, BKT19, CP16, CX11, CCL23, CM14a, CL17, DMZ11, DL09a, Fer01, FN97, GP18, HVM99, HW99, Han09, HRBR15, HP99, HLS14, IS98, JN00, KL00, LS19a, MK20, MNPR18, MZ17, NLL18, Nie13, Nku21, PY19, PDL19, RA17, Sch01b, Tor17, VW17, WYR21, Yan08]. points [BL20, BDZ15, CLLLR17, CS18b, Deb99, DLR06, EK18, GM15, HTM06, HV09, HS10, LL00, Nic18]. Pointwise [Reb15, VS05, CL14, Cut00, Efr98, GP15, Gol99, GL08]. Poisson [AL05, AIZ16, AH13, BM15a, BIZ02, BX00, BX06, BJ00, BP21, BP03b, BC19, BC20, Bou15, Bou06, BV10, ČK00a, ČV10, CX11, CH18, CD11, DS05, FRS07, GP18, GT18, Han09, Hir97, Jou07, Jou09, KRX18, KM95, KMS03, KM06, Las14, MSWW11, MV18b, Neu11, NZM23, PT08, PZ14, RBM19, Röl05, Roo99, Roo05, RSS12, ST13, Sch09, TY16, UDS23, Zac23]. Poisson-binomial [Roo99]. Poisson-Delaunay [BC20]. Poisson-type [Jou07, Jou09]. Poissonian [Fou01]. polar [Deb99]. Pólya [BT17b, PRR19, Ran22]. polygons [LW23]. polyhedra [GT18]. polykays [DGS08]. polymers [ARAS20, CSY03, CMr19, Cor16]. Polynomial [GP16, MGS03, AW09, AHHK08, BT17a, BF01, CCF22, CCJM24, Fil00, GS13, JT03, JLMP22, KM24, ZL03]. polynomials [AB22, AK18, AGG23, CP23, FP17c, GS11, SU08]. polytope [Ngu15]. polytopes [Bru19, Bru20, GT18]. Pooling [CS06, DPS24]. population [BG00, BBC12, BBC18, BX15, Cor16, FHJK20, HWZZ23, JMM17, Jag95, KLP21, Mar19, Möh99, MMR06, Ngu15, Olo09, Yin22]. population-dependent [JMM17]. population-size [MMR06]. populations [Ban24, CMMA20, FN23, GMD06, LØ01, Wak95]. Portfolio [Lin09, Hin03]. posedness [Kra21, RZ21]. positioned [HV09]. positions [DT23b]. positive [ACBL15, Bra02, Bra14, FLZ21, Gne13, IPP23, Jou09, KVZ19, Mal16, PS09, Sie07, YP23]. possible [ČK00a, MR13]. possibly [BN20]. post [Bei97]. post-change [Bei97]. Posterior [ARV18, AAGGR03, CD17, DRRS18, HR22a, Ngu15, Ngu16, Abr19, ADH21, CG22b, GR14, Gho99, GHN21, JLP10, KS18, Lau13, LLD19, MM15, MMW17, SM23, VVV06]. Posteriors [BWJ18, GO00]. potential [BM13, HP99, Hie17]. potentials [RASY17]. Potts [HJL02, HKK19, JK17]. Power [ADM18, CK23, CD23a, CNW06, AIS19, AGM06, BNS03, BJSZ07, Hin03, LS19c, Nor15, Rob12, Zei15]. power-producing [Hin03]. power-rate [AIS19]. powers [Del17]. practice [LN18]. Precise [BMSZ21, BDZ15, GZ21a, MR21]. Prediction [BNC96, Gay20, Gay22, BPB^{+}22, BK23, CG06, CM13b, DHL17, FZV23, Fus05, FKA05, Hof23, LYG19, Lee08, San16]. Predictive [FP17b, BCPR09, BDPR21, CG99b, ES99a, JJW24, KMS17, XL10, YKK21]. predictor [GR04, HMP18a, IKP18]. predictors [PY01, VZ23]. preferential [HM20b, MM20]. preprocessing [BM13]. Prequential [DV99]. prescribed [GMZ21, Lei19]. presence [DJ97, DFK10, EMPW20, IS98, LPV13, Ogi18, PV09]. preserving [CSvS07]. presmoothing [MPV22]. previsions [dCQM09]. price [BD20a]. prices [BNBV13]. pricing [HN10, PP09]. Principal [HPV13, KSX19, LMZ24, HHH24, HK13, PRV20]. principle [CHR24, GO00, KJ20, LN02, Mel00, OH23, SSY21, WYZ24]. principled [CMMA20]. principles [BYY17, Bar18, CCTY19, DV99, DMP11, DN14, GL06, Kre21a]. prior [GP12, RR01]. priority [Mac23]. priors [ADH21, AW24, BD16, DT12, EF04, HMS04, PR02b]. privacy [BDKS20, LWLL22, Sar23]. private [YRZ24]. Probabilistic [AP11, DD15, FS13, JMW05, ASC02, CTP21, GS11, HRBR15, LR23, Sto00]. probabilists [Ald99]. probabilities [AFGS08, BFT15, CLR14, DSS10, DGP18, DdH15, GNSU19, HR22a, HWW99, HS08, LT16, Möh18, MO08, SY00, WT13, Yu12]. Probability [BNH08, Tor17, Wer11, AK23, BV17, BNT02, CLD08, CCN^{+}21, CX16, DV99, GL22a, GL22b, GT04, Gra21, HW99, JM22, KM20a, Kam18, KL10, Lau13, LCP13, Mal16, MH10, MN00, Roo10, Sri16, TK03, Yu12]. Probit [GCP17]. problem [BLO19, BH96, Cav01, DF20, FN97, Gut02, KK23b, ŁS13, LL00, Lec13, Mla08, Nku21, SW12, TAS^{+}17, UDS23, Van04]. problems [AW24, AAGGR03, BP03a, BGL22, CG99a, CG22a, DJ96, HMTvB24, Kee13, KS18, MW20, Sto00]. procedure [DL17, EK18]. procedures [AZ15, BKK23, DRRS18, GL22b, HP02a, Yan04]. Process [BY05, AMPV10, AMV18, AD11, AT24, BD13a, BM15a, BJ00, Bar18, BBG19, BFMP01, BF01, Ber01, BBV17, BZ23, BR18b, Bor01, BEM18, BDM10a, Bou06, BK16, BHVD17, BFM16, Buc18, CD21, CI21, CKHM23, CLLLR17, CT02, CHI17, CX11, CR11, CY05, CP18, CM14a, CSY99, DGS20, DGU05, DJ97, Dem07, DGP09, DRRS18, Eis00, EH02, FRS07, FdH14, FV03a, FV03b, FP17a, FGS21, FGS22, FS10, GGT22, GNR14, GCK04, GLS07, GV19, GR00, GT01, Han09, Jac96, JLT08, JWY22, KS99a, KK18, LS16a, LS19a, Lee08, Mac23, MS23a, Mal16, Mas04a, Mas04b, MZZ20, MMR06, MNPR18, Mon15, NRT98, NZM23, Nie13, Nor15, OP11, PP09, Pak99, PR14, PW04, PR22, PT00, RBM19, Rév96, Sch09, Sid18, SU08, SS15, SXX22]. processes [ALT15, AIZ16, ACG19, ACG20, ACJ18, AC01, AB22, ALS09, AP11, Ath00, AAI11, ALL09, ADS13, AW96, Bac18, BW23a, BGMSB14, BR18a, BIZ02, BV17, Bar10, BBJ08, BR07, BNT02, BNCP11, BNBO11, BNBV13, Bar22, BAT19, BTUA23, BOHP18, BŁ15a, BL23, Beg07, BMMS12, BS15a, BS22, BP03b, BPW19, BM08, BP10, BS16, BE16, BS09, BHK03, BC02, Ber17, BZ12, BC21, BS95, BL16, BO20, BOJGMR15, BHM22, Bra14, BFM16, BLM19, BDKS19, BD22, BN05, CPP11, CMS10, CTP21, CJ13, CZ13, CP16, CPR13, CX11, CPY13, CS24, CH08, CCL23, CLR14, CH15, CL17, CR07, CGCR07, CS99, CNW06, CD11, Cra15a, CNOH17, DNV99, DP09, DRW19, DNN10, DNR20, DMZ11, DH17, DS05, DMP11, Deh05, DDT14, DMM^{+}20, DV18]. processes [DS96, DL16, DL09a, DDK22, DY14, Dob23, DMY00, DKM16, DW20, DJL23, EK11, Erh14, ELL22, ET23, FHK15, FHJK20, Fas10, FK17, FRW04, FM00, FLS01, FK11, FB14, FLZ21, FM02, FP10, FZ04, Fra06, FR11, GK03, GZ18, GZ21a, GNNS21, GP18, GLT23, GT09, Gin96a, GS03, GL06, GJY03, GMD05, GLST19, GW98, GY04, GS20c, GZ17, HVM99, HB04, HLT01, Han22, HRBR15, Has13, HY05, HR16, HP99, HLS14, HP14, HK07, IMM17a, IMM17b, ICG16, IKP18, IKP14, JMM17, Jag95, JS08, JW13, JN00, Jir18, JN16, JVV05, Jou09, Kab11, KS16, Kar23, KMM24, KM22, Khm16, KKW17, KVDH16, KM95, KL05, KS20, Kra21, Kre21b, KS21, Küh22, Kul07, KY07, KP06, LR16, Las14, Lau04, LS21, LV12, LvdG14, LS19b]. processes [Lei06, LT00, LYT21, Lim19, LM20, LCP13, Lud00, LP07, Mai14, MM23, Man07, Mar06, Mas04b, MP23, MH22, Möh99, Möh06, MK20, MSS14, MW05, NRC24, Nda23, NR09, Neu11, Nku21, Nob06, NS01, OSZ18, Ogi15, Ogi18, OU23, OR24, OH23, OP10, OS15, PR02a, PSS24, PRvS13, Pec08, Ped95, PR22, PR02b, PS13b, PS09, PTL04, PY19, PDL19, PVD13, Pro24, Qiu18, Qiu19, RWZ17, RS18a, Riv05, Riv07, RS96, RSS12, RvS19, RW02, RVW01, Sam98, SW04, SSW12, Sch01b, SS12, Seg12, SZ16, SW19b, Sør02, Ste10, SRR24, SXXZ18, SPP21, Tal13, Tka20, Tor17, Tsa09, Tud04, Vii16, VW17, Wal12, Wan11, Wan16, Wan23b, Wu03, Xu21, YY24, Yan08, YB06, ZZ23, ZL13, ZH20, BRJN19, Fra07]. Procrustes [ZP19]. producing [Hin03]. Product [PR06, BH96, DD15, GNT17, KPT21, LP07, PFEP23]. Product-limit [PR06]. production [CS18a]. products [ALS09, CJP21, DL16, Gau17]. Profile [FH05, BBD17, Hir16]. programs [LE23]. Projection [LMP12, BS21, BHL99, CG22b, GPNECA23, HWW11, RB06]. projection-based [GPNECA23]. projection-posterior [CG22b]. Projection-type [LMP12]. projections [AV14, GN10, KPT21, SL18]. Prokhorov [HW19a]. promotion [PGV17]. proof [DD15, HJ17]. proofs [CEG17]. Proper [Ovc18, KF14]. Properties [CCTY19, VT13, AF19, BGMSB14, BBS15, BCI02, Ber97, Bia14, BDP23, BDG07, CAA22, CM11, CG99b, CSY99, DV99, DGP18, DV18, DP11, DLN22, DR09, DGC11, FR11, GCK04, GL06, GT19, GLST19, HS10, HLT17, KM20b, KV12, KZ21a, KS13a, LMV14, MSS14, NR12, PDL19, SL14, TX07, VY13, Wan22b, Xia20, Yu17, ZL17, ZZW22, vDD21]. property [AB22, BW08, BC00, Duf96, Gob01, MW04, Mon15, NLL18, Ogi15, Rát19, San13, SSW12, ZS15]. Prophet [ASC02]. proposals [SvdMvZ17]. protein [Cha03]. pseudo [BS09, CMMA20, DL09a, DL17, GLO22, HP02a]. pseudo-likelihood [DL09a, DL17]. pseudo-Mahalanobis [HP02a]. pseudo-marginal [GLO22]. pseudo-maximum [BS09]. pseudo-populations [CMMA20]. Pulses [CGMST95]. Pure [CR21, FZ04, OP10]. Pure-jump [CR21]. purely [BW19].

Quadrangulations [BL13]. quadrant [IKP13]. Quadratic [CD23b, FPS95, Koi16, LW22, AT24, Beg07, BP08, FLS23, HR13, HF99, JN20, Wal13, Zho19]. Qualitative [Zäh15]. quantification [BG17]. Quantifying [BL16, DDK22]. Quantile [KVDH16, Zha00, ZLL17, ACLZ17, BDG05, CL02, CC19, DGG13, DRT16, DGU05, DCC14, Dre03, EH02, FCH15, HLY09, HLPŠ15, HIW19, LT13, LTC14, VW17, VD13, WZ18, ZH20]. quantile/depth [HLPŠ15]. quantiles [BBCM11, CL02, DPS24, Das05, DHKV15, DW23, EH02, HR22b, KP22a, Kre24, Kul07, MVY19, Ren08, SC11, dCCGZ23]. Quantitative [MS10]. quantization [BP03a, Cor14, Lev18, LP07, ELP22, PP05]. quantizers [PS12]. quantum [FY20, Lua11, MPS19b]. Quasi [BNBO11, KP06, GC17]. quasi-Monte [GC17]. Quasi-stationary [KP06]. Quasilinear [Tin98]. Quenched [AP19, BP16, Bar18]. queue [RBM19]. queues [BD20b, RS01a].

Rademacher [BD15, BP19a]. radius [PS12, TK03]. Radon [KLP12]. rainfall [Cha97]. Ramanujan [Alm03]. Random [AW09, BS23, CMr19, CM99, DK18, GLL20, GCK04, JHD20, KG00, MPV07, AP19, ACV09, AEP11, ACBLV18, ALL09, AW96, BFT15, BR18a, BV17, BPZ15, Bar10, BS11, BHM18, Bel98, BD15, BD20b, BMSZ21, Bha23, BL15b, BBW23, Bir04, BL14, Blo04, BCvdH20, BCvdHV21, BK23, Bor21, BRS16, Bou06, BV10, BH10, BCG19, BWJ18, Bru19, Bru20, BGL24, BST12, BD22, CHHB19, Cap24, CD01, CD21, CdLOR20, CJP21, CG21, CX16, CS18b, CH15, CSY03, CR17, CP23, CRS00, CGER07, DMZ11, DPTV15b, DH13, DN14, Dio12, Dor97, DT23b, DZ24, DW19, ER23, EK18, EFVG08, Erh14, FK21, FM00, FS10, FPW17, GG22, GKP12, Gan21, GL18, Gas23, Gat06, Gau17, GW17a, GMD05, GNT17, GT16b, GS08a, HMT00, HJL02, HXZ18, HM20a, HSS07]. random [HNR23, Hie17, HM20c, HM15, HS08, IRVZ22, IRVZ24, IMM17a, IMM17b, IPP23, JMM17, JS08, JRSK21, JLP10, Jan07, JT03, KVZ19, KPT21, KMS^{+}99, KP04, KHLN97, KM98, Kre21b, Küh19, Kur22, KS01, LS16a, LS16b, LS00, LS03, Lau13, LC22, Lee22, LX23, LMC23, Lei19, LO14, LRMT17, LPSS18, Lep16, LT16, LS95, LWX15, LNP14, Lin19, LBDM19, LW23, LP24, LS23, LV15, LS14, LMZ20, MY23, Mal17, MP10, MM21, MR15, Mel00, MJK17, MR10, MSS14, MS18, MS17, MZ17, Mus07, NZ05, NN22, Nic18, PXH13, PRR19, PZ19, PS22b, PTA07, PFEP23, PS16, RASY17, RNS23, San10, Sch20b, SW20, SR09, SX10, Ste13, SL18, SW08, SN13, TAS^{+}17, Tod20a, TV99, VAD22, Ver20, Vii16, Vil14, Vil23, WXZ19, WSX20, You05, Zha96]. random [Zho14a, ZX17, RP23]. random-cluster [HJL02]. random-connection [LP24]. Random-design [CM99]. random-walk-type [JT03]. randomization [JP15]. Randomized [CNOH17, Dob23, GŠ08b]. randomly [AGV18, DFK08, DFK10, FK12, SFP23]. range [AS03, BS12, BW22, Buc18, CS24, CL17, CM99, DNR20, DLPZ09, DLPZ12, GWY13, LR16, Lee22, Lud00, PS16]. Rank [BW22, HLL23, BCvdH20, BX15, CP21, FP17a, GGKL22, HPV13, IKP13, Jur99, JK12, Klo14, MY23, TCP22, Tur04, dBCAM22]. rank-1 [BCvdH20]. Rank-based [BW22, HLL23, HPV13, IKP13]. rank-constrained [TCP22]. rank-order [dBCAM22]. rank-scores [Jur99]. ranked [Che02]. ranked-set [Che02]. ranks [ACZ23, DHKV15, HMT00, HP02a, ZGV20]. Rao [GL95, Len05, LS19c, Lua11]. rapidly [HMTvB24]. Rare [ABH00, BR18b, CR13, DN21]. Rasch [LMK19]. Rate [BCPR09, BDR14, BGGL21, GT04, HM20c, JZ19, Jir18, LGW06, Var19, ALOV19, AIS19, BJM21, DPTV15a, Del17, FPSS19, FLS23, Fis14, GW17a, HMP18b, HMS04, KK23b, Mal17, Mou13, SW19b, Van05a, WW14, GL22b]. Rate-optimal [HM20c, GL22b]. Rates [ADH21, CG22b, CP23, DF20, FR15, HHH24, Abr19, AP19, BM07, Bar02, BM98, Bra22, But17, CD21, DR19, DRRS18, DMS07, EO07, FFP23, Fil00, FMR19, GHL05, HR22a, JM17, Klo06, KS21, LWCS22, Lec07, LM15, ML20, MFWB22, NL14, OCBG19, QJ22, RV09, SXX22, SM23, Tal13, VVV06, WC14, WB19, YG19, dBM17]. ratio [AGM06, CP12, CC04, DL09b, JZ19, JJW24, LP99, Ren08]. rational [DMS21]. ratios [BS11, BP08, DSW21, LS14, VG18]. Ray [RW22]. Rayleigh [AB22]. Rd [KM24, Lei06]. Reaction [ADFS15, AG01, GR23a]. reaction-diffusion [AG01]. real [AB22, BCI02, Bou06, CPR13, WXX17]. real-valued [Bou06, CPR13]. realised [GV19]. Realized [BNS03]. Rearranged [SDS24]. rearrangement [AF19]. reasonable [DV01]. recall [ASC02]. recombination [CS18a]. records [GLS07]. Recovering [CI21, LS19a]. recovery [GS20a, LM19, MV04]. rectangular [Din20, DZ24]. Recurrence [BR97, CdLOR20, IRVZ22, IRVZ24, San13]. Recurrent [Riv05, Riv07, BBD20, DH02, KK23b]. Recursive [LP02, BIZ02, CRW24, DR07]. recycling [MPS19a]. reduced [BX15]. reducible [BDM09]. reducing [TA05]. reduction [BIPZ22, FB22a, HT00b, KKR24, ML20, ZWS21]. refereeing [Ano95a, Ano97a, Ano98a, Ano99a, Ano00a, Ano05a, Ano06a]. refined [FLS20, Pat11]. refitting [CHS19]. Reflected [PX10, WYZ21, DM24, EI18, FR19, RTT20]. reflecting [JW21, ŁS13, Pet97]. reflection [Tin98]. Regenerative [BC06, BW23a, SZ03]. regime [BS23, DM21, ELL22, LM20]. regimes [RASY17]. regions [BW19, CLD08, Qia21a]. registered [JV01]. registration [CP21]. Regression [Jur99, KP04, AAGGR03, APGMV04, AF19, AHHK08, ADM18, BCFH22, Bar10, Bel21, BSG22, BBW22, Bir04, BL10, BRW00, CCF22, CLZ18, CV21, Car06, Cha99a, CP19, CL24, CL19, CL02, CV09, CGL12, CN15, Che24, CL14, CC19, CS07b, CM13b, CM99, DGG13, DL19, DMP11, DNP06, DW23, DF98, Dic16, DL09b, DNS19, DL17, EMW18, FMSY12, FCH15, Gao20, GT16a, GW98, HV09, HLY09, HLPŠ15, HMP18a, HMWY10, HM20c, HIW19, HLP23, HKM06, HR13, HLZZ23, HF99, JLMP22, KM23, KKS02, KK23b, KLLM13, Kur22, KM24, LMV22, LMP12, LGW06, LT13, LMNV11, LPV13, LTC14, MVY19, ML11, Mei16, MS18, MS17, NC21, NJG20, PLH09, PB15, PBD15, QT05, SW13, SG16, SCK23, SGL^{+}24, VV02, VD13, WC14, WH10, WFM10, XL10, Yan01, Yan04, YDN23, YMP11, ZWL14]. regression [ZLL17, ZGV20, Zho14b]. regressors [CGL12, GG22]. regret [RST17]. Regular [KMS03, DR11, GP23, HL06, LT15]. regularisation [CP19]. regularising [BDP23]. Regularities [HW24]. Regularity [BEM18, CKHM23, Car15, Hei18, Mor99, Neu11, SS12]. Regularization [LM19, TV06]. regularized [GT22, Lev18]. regularly [BP21, Bha23, BL14, FN23, HS08, JK21, KSX19, MR10]. Reimer [AGH18]. reinforced [AGV18, ACG19, ACG20, SFP23]. reject [WY11]. rejection [Dai17, MK20]. Rejoinder [Gin96b]. relabeling [BCFK15]. related [Alm03, AH13, BCRT24, BMMS12, BP10, BD14, BW08, Cor16, DCC14, HW10, HN10, HDS22, KJ13, LV12, LBDM19, MPAiS12, NRC24, PY99, Qiu18, Sau19, Tka20]. relation [Gra99]. relations [CJ13, Möh18]. relationship [CG99b]. Relative [Yu10, BW19, KD95, Lee08, SU08, TY23, Woo00]. relatively [LS00]. relaxation [Zha13]. relevant [vDD22]. Reliable [NMR20, LR95]. remainder [HY13]. Remarks [DKS01, AB15]. removal [JK23]. Renewal [LT00, AP11, GK03, GS21, LYT21, PT00, PTL04]. renewal-reward [PTL04]. renewals [FGS21, PTL04]. Renorming [HW11]. Rényi [LMZ20, MMR23, TV99]. Reparametrization [HL12b]. repeated [DDBM21, dSJMT23]. replacement [BM15b, Ber97, BDM09, HWZZ23, Yu12]. replicate [BBW23]. replication [BD20a]. representation [CPR13, DF20, GS03, GST08, Kre24, KL02, LV12, OSZ18, Pec08, RW22, Ruk14, Vii16, WT13]. Representations [CJL17, AK23, AM07, BNMS06, BGH01, BFM16, DP11, HP02b, KS16, LMC23, MP10, MR17, MSS04, You05]. Reproduction [YY24]. repulsiveness [BL16]. Resampling [Mcc00, CMMA20, DDJ12]. rescaled [NIR21, QP18]. reserves [KM95]. reservoirs [DPTV15a]. residual [PY19, ZGV20]. residuals [FZ22]. resolution [HP01, LP07, MP10]. respect [Ben03, BBV17, BDP23, DS19, GLT23, GJQS20, LM07, Nor15, Vii16, Wal13]. respecting [HLPP09]. response [DPSW10, DMR21, HZ04, SW13, YDN23]. responses [MS17, NJG20]. restaurants [RW22]. Restricted [Ruk14, Bel21, BP19b, CSSC95, DRW20, Düm03, EHS05, MS12]. restrictive [Seg12]. result [IR20, MS12]. results [BCRT24, BRJN19, BZ23, CP05, Fas10, HVM99, JN16, Nie00, NVV99, SST16, ST20, SXXZ18, Xu21, Zha96]. Retrospective [BPR06]. returns [Lin09]. reverse [MSS04]. reversed [LSY^{+}22]. Reversible [GW99, CLZ06, DS03, Lei06]. review [Ald99, FZV23]. revisited [JB21b, Len05, PS20]. Revisiting [Pak99]. reward [LT00, PT00, PTL04]. rewards [LT00, PT00, PTL04]. Rice [ALW11]. Richardson [LP17]. Ridge [Dic16, CP19, DDBM21]. ridges [Qia21a]. Riemann [BOPG22, LH01, Nak15]. Riemannian [BPD23, Gas23, GS22, LLBF24, LE23, TT20b]. right [KVD04, RA17]. Rigid [Qiu19]. risk [BS22, CL19, CY05, CS07b, DGS20, FG19, KD95, KK17, KM95, KKS02, KK21, LM10, Lec13, LM16, Lin09, RST17, Sau18, XL10]. risks [CLZ07, LC22, LT15]. RKHSs [PG21]. robin [DLL20]. Robust [BN20, Gao20, Giu18, Hil15, HNR20, MW20, Ore22, Sar16, Bah21, BL20, CGLZ24, CLLLR17, CL14, GT22, KLLM13, LM07, LTC14, Min15, SGL^{+}24, YRZ24]. Robustness [CAA22, CS13, CS07b, Hoo16, Zäh15]. role [BCvdH20, WLH16]. Romberg [AHK15, LP17]. Root [Tru19]. Root- [Tru19]. rooted [CG21]. roots [AW09]. Rosenblatt [VT13]. rotationally [WXZ19]. Rough [OR24, HNZ18, LS20, RTT20, SSX20]. roughness [BS11, BCI04, CD23b]. round [DJ97, DLL20, Ros09]. round-off [DJ97, Ros09]. round-robin [DLL20]. rows [MP06]. Rudin [PFEP23]. ruin [BBS15, BL14, LCP13, Mus07]. rule [GL08]. rules [ASC02, BCFH22, Din98, GNR19, KF14, Ovc18]. run [HMTvB24].

Saddlepoint [DLR06, HJQZ04, KK17, BP08, BW19, KL10]. Sample [BDG07, HK08, MY23, Sch01a, TX07, ACBL15, BWZ10, BLP19, BEN12, BG09, BBS14, BŁ15a, BPR06, BBC18, Bor01, BX15, CP12, CP15, CC04, CX24, CAA22, DGP18, Fas10, HM19, HT00b, JMRX18, JR17, Kar23, KC19, KL17, Kre24, Küh22, Kul07, KLP21, Lee08, LGS08, MRS22, Pon07, SvS17, TA05, TAS^{+}17, WAP17, WHP18, WB19, Yin22, YZZ23, Zha00, ZS15, ZZZ17, ZS21]. sampled [CDJ13, JS17, SS12, SU03]. sampler [JM17, LN18]. samplers [AH17a, ALV18, HMTvB24, PJT19, QJ22]. samples [ÁEdBCAM12, BGT97, KS13b, dSJMT23]. Sampling [GR00, HWZZ23, AHK15, AP11, BP06, Ban24, BM15b, BC19, BRW00, Cha03, CBL22, CP16, Cha12, Che02, CS15, Dai17, DRD20, ELP24, FP17c, FH07, GZ09, GRS11, HMW01, Hir97, HT00b, Klo14, KW20, Kur22, LS19a, Mar19, MPS19a, MH10, Möh06, MK20, Ore22, PY19, RA17, RBM19, TA05, Van95, WW14, Yu12, ZZ23]. Sanov [LN02]. satisfies [Pey17]. scalability [Jor13]. Scalable [NIR21]. scalar [Abr19, Cho18, DN08, YDN23, vdMvZ13]. Scale [Deb99, BYY17, BB17, CD17, FZ22, GCK04, MS23b, Mat00, SSH14, Sch10, TT06, Zha06]. Scale-invariant [Deb99]. scaled [GK03]. scales [BG95, FFZ22, GP03]. Scaling [BD20a, IMM17a, Lei19, ALS09, AAI11, Fan21, Fra06, Fra07, HM11, JLM14, LN18, LWX15, LS20, LS14, PS22b, SW20, SR09]. scan [ZL17]. scanning [SY00]. scatter [ÁEdBCAM18]. scattered [HTM06]. scenery [FS10, GW17a]. scheduling [DLL20]. scheme [Ama13, BHBO17, BO18, BJM21, Fri18, HMW01, LTX22, NV17]. schemes [ADFS15, BT17b, BN17, BC19, BB22, FHJK20, FLT14, MPS19a, Pet97]. Schoenberg [PS13c, PS11]. Schrödinger [dV13]. Schwarz [EM18]. score [Hös05, Kre24, LS19c]. scores [Jur99]. Scoring [BG21, Gay20, Gay22, KF14, Ovc18]. SDE [KP22b, Kut21]. SDEs [BN17, BO18, BJM21, DN08, HLT17, HMW21, HW24, KKR24, ŁS13, LSW20, RZ21, RZ23, Wan23a, ZZ23]. Search [JN16, BRJN19, JT22, NV17, ZWS21]. seasonal [CT12]. Second [BS19, Ber97, DL19, FFS10, GL18, GS20b, BP06, Cas07, GS00a, HMP18b, HNR23, OPPS16, PS09, Sch01b, ZS15, vDD21, vDD22]. Second-order [Ber97, BP06, Cas07, Sch01b, ZS15]. section [Yan24]. Sectional [GKT24]. sectorial [Lee22]. seed [BEMR17, MM20]. seeds [BEMR17]. Segel [TT20a]. segment [Van04]. seismic [OKM^{+}98]. Selecting [MYT06]. Selection [ABH08, AW12, AB11, BCFH22, Bar10, BKK23, BC13, BGL22, Bir04, BBC12, BBC18, Che24, Cor16, DS03, DLWZ17, DLT09, GS21, GL08, GR04, HJM14, JJW24, JS13b, KVD04, Lee08, Lep18, Lev18, ML10, MSW13, Mal17, MN22, NPM21, RP23, WXZ15, WM16, WH10, Zha13]. selector [CCL23, Kol09]. selectors [NL14]. Self [BNT02, BP19b, Cha99b, FGLS19, HM11, LSW13, SZ03, Arm21, BWW00, CK11b, CPR13, CP18, DS06, DGP09, DJL23, ET23, FHK15, FS10, FT19, HK11, KM20a, LXY13, LTX22, OS15, Riv05, Riv07, Sam98, SZ16, Tra14]. Self-consistent [BP19b]. Self-decomposability [BNT02]. self-decomposable [Tra14]. self-exciting [CP18, ET23]. self-interacting [CK11b]. self-interaction [DJL23]. self-learning [LXY13]. Self-normalized [FGLS19, LSW13, LTX22, SZ16]. Self-regenerative [SZ03]. Self-similar [HM11, BWW00, CPR13, DS06, DGP09, FHK15, FS10, FT19, HK11, OS15, Riv05, Riv07, Sam98]. Self-similarity [Cha99b, Arm21]. Semi [Cas07, HW03, YMP11, ARAS20, AK97, AAGGR03, BOHP18, BC02, BBD17, BRW00, CP16, CV09, CGL12, CS24, CDH21, FZ22, GR23a, GL06, HL12b, Hir16, JV01, KLLM13, MV99, Pon07, WXZ15, Zha00]. semi-exponential [CS24]. semi-infinite [ARAS20]. semi-linear [GR23a]. semi-Markov [GL06]. Semi-parametric [Cas07, HW03, YMP11, AK97, AAGGR03, BBD17, BRW00, CV09, CGL12, FZ22, HL12b, Hir16, JV01, KLLM13, MV99, Pon07, WXZ15, Zha00]. semi-stable [BC02, CP16]. semi-stationary [BOHP18]. semi-supervised [CDH21]. semicircular [GNTT18]. semidefinite [LE23, PS09]. Semigroup [GS03]. semigroups [DMZ11, GST08, Sie07, Ued21]. semilinear [ACP23]. semimartingale [DNR20, Koi16, Sch20a, VVV06, Van05a]. semimartingales [CR21, CM11, Dio12, Tod20b]. Semiparametric [BDKS19, BTV17, CC04, HLZZ23, Ber06b, BBD20, BM05, CZH14, DLM14, FH05, GT22, HLL23, HIW19, LTC14, Ma10, MSW13, MPV22, Tru19, WZ16, ZGV20]. semistationary [BNCP11, GV19]. sense [BOJGMR15, DPX23]. sensing [Cas14, CT19]. Sensitivity [BEN12, GGKL22]. sensor [SK23]. separability [DDK22]. separable [BLP19]. separately [Mar18]. separating [GC05]. separation [ACPM24, DRW20, GGM20]. sequence [Alm03, BP21, Cha03, Com01, Gat06, GY04, HMP18b, PS12, PT15]. sequences [ABH06, BCRT24, BN20, BDL^{+}23, Bra99, CLR20, DDT14, HM20a, Hör08, HNTX15, Lei19, Wu03]. Sequential [Bei97, CLZ24, HR22b, MS23c, OCDM08, BBG19, BK16, Buc18, DDJ12, ELP24, GC17, GLO22, PJT19, WLH16, You05]. serial [BP98, HMT00, Tur04]. seriation [FMR19]. series [AW12, ADFS15, ACLZ17, BKK23, BW22, BB99, BS18b, BJ24, Büh97, CdLOR20, CT12, CZ13, Cut00, DP09, DMMW18, DT23a, Efr00, FKM02, GWY13, GMR19, GQ23, HK08, HS08, JK21, KL09, KJ20, KS97, KKP21, KR24, KL02, LV12, LMC23, Len11, LRS23, MMS08, MZZ20, McC08, MS23c, Nda23, NP08, Nic18, Pap09, SSH14, Sch24, SvS17, Syn07, TC24, Tru19, Tru20a, WAP17, WP22, WZ18, Zho14b, vDD21, vDD22]. set [BST12, Che02, CDH21, LBDM19, OSZ18]. set-valued [CDH21]. sets [ACV09, Bel21, BP19b, CKLN18, CS20, DR23, EK11, EHS05, HLPP09, KN08, Kur16, Lee22, LR95, Per02, Qia21b, RV09, Vil14, Vil23, YK20, ZGD24]. setting [Zha22]. sex [JMM17]. shape [BEN12, Bel21, CS24, CSvS07, Düm03, Gat06, HM15, HKM^{+}16]. shape-preserving [CSvS07]. shape-restricted [Düm03]. shaped [JW09]. share [PR13]. Sharp [AL05, BRS16, GM15, Han22, JK17, LT16, Osȩ09, Osȩ11, WB19, YS22, BV10, Jir23, Roo99, SS18]. Sharper [LM10]. sheet [BJ00, BDP23, FPW17, PR16]. sheets [DPX23, DW19, KL02, MWX08]. Shepp [HJ17]. shift [DW23]. shifted [Ran22]. shifts [MZ17]. shock [Gut01, MSS16]. short [AS03, CNOH17]. short-range [AS03]. shortest [BG21]. shot [KM95, KMS03]. Shrinkage [US24, GMR19, LW22, LAP20, LLL22]. shrinking [Tod20a]. sided [MR19]. Sieve [Büh97, FL11, ZWY17, BB99, Gne04, MJK17, QLC21]. Sieve-based [FL11, QLC21]. sieves [BM98]. Sieving [MM21]. Sign [PRV20, BD14, LWYZ22, PV16]. Signal [Dit19, Bar02, Car15, Cas07, Cas14, TC24, VG18]. signal-to-noise [VG18]. signals [CW22, CT19, KRX18]. Signature [BO20]. Signed [ČK00a, AL05, ČV10]. significance [Büh13, HR13]. similar [BWW00, CPR13, DS06, DGP09, FHK15, FS10, FT19, HM11, HK11, KM20a, OS15, Riv05, Riv07, Sam98]. Similarity [ÁEdBCAM12, Arm21, Cha99b, vDD21]. Simple [BS14, BMS21, DNP06, GN08, Pak99]. simplex [SCK23]. simplices [Pai22]. simplified [CJP21]. simulating [SvdMvZ17]. Simulation [BE16, DH17, AKH17, ABH00, BN17, BPR06, BPR12, BG19, BS14, BMS21, CR07, FdH14, FLT14, HVM99, LBDM19, OSZ18, PJR16]. Simultaneous [CK11a, CLZ18, HM15, MSW13, WZ18, MV18a, Möh06, PV09]. simultaneously [BH10]. sin [HL14]. Single [LPV13, BDJ19, BL10, GS08a, KP20, LMV22, WC18, XW12, ZLL17]. single-index [BL10, LMV22, WC18, XW12, ZLL17]. singular [AG01, Ass23, CJP21, CHR24, DX16, DZ24, GNT17, KKR24, MP06, OR24, RZ21, RCBR00, WYZ21]. singularity [WHP18]. sites [CRS00, JK23]. Size [Olo09, PT15, BAT19, BGI18, BBS15, BH10, CP15, Lee08, MR10, MMR06, SN13]. Size-biased [Olo09, PT15]. SiZer [HKM^{+}16]. sizes [CRS00]. sketching [LEM23]. Skew [HL12a, HL14, Fri18, Lua11]. skew-normal [HL14]. Skew-symmetric [HL12a, HL14]. Skewing [CCFH00]. Skorokhod [CH07, ŁS13]. slab [CS20]. sliding [DN21]. slope [LT16]. Slow [PTL04, FGS22, SXX22]. slow-fast [SXX22]. Small [ALL09, ADS13, CLR14, EX23, FLT14, FLLO14, HP14, Kar23, MO08, Sid18, SU03, AD07, BLLS12, GR23a, Gra21, HMP18a, Har17, Jac02, Lee08, OKM^{+}98, RW02, TT06]. Small-diffusion [SU03]. small-seismic-array [OKM^{+}98]. Small-time [FLT14, FLLO14, BLLS12]. SMC [ALV18]. smeariness [Elt22]. Smooth [ER19, HMP18a, AW09, Car15, CS18b, Cun15, GS16, HP95, NC17, SS18, ZZZ17]. Smoothed [AAG24, Ren08, BZ23, DL13]. Smoothing [ALR23, Ran22, AG99, BP19a, BDZ15, CHT98, DGG13, DP16, FKM02, GC17, GLO22, GK12, HP99, KKS02, KP20, OCDM08, OW17]. Smoothness [PS13b, DH13, QSSS04, Seg12]. Snedecor [KL13, LO14]. Sobolev [CD23a, GS20b, JMV20, LMZ24, SS20]. social [Ald13]. Society [Ken95]. Sojourn [LO14, NPS20]. solution [CW01, DDBM21, Fou01, Mor99]. Solutions [Lau04, CR22, LS02a, MPRZ10, RZ23, Tka20]. solvable [Cor16]. solving [BP03a, LGW06, LE23]. Some [AB15, AC17, BNMS06, BL04, CK11b, CSY99, Dia13, DMY00, DGC11, HMTvB24, KM20b, PXH13, Sch10, Yu11, Yu17, Zha96, AC01, AI99, AF10, AF12, BP03b, BRJN19, Ber06b, BL15b, BO18, CG22a, CC00, CNW06, CD11, Dio12, EF04, FP10, GJY03, GRS11, GT19, HVM99, JN16, LSW20, LM20, LS23, Man07, Sch13, Yu12]. soup [CDL24]. source [FRS07, RSS12]. space [ÁEdBCAM18, BR97, BK06, BDKS19, CR11, CK15, CM14b, CM18, DDK22, DL13, DGC11, FM00, GJQS20, GS22, HNZ18, JCL20, Jir18, KSS21, LS16b, LGS08, LR23, MSS00, OCDM08, OPPS16, Pec03, PR99, SS12, SU08, SSX20, Ste13, TA05, WK21, ZP11, ZP19]. space-time [BDKS19, FM00]. spaced [MMS08]. spaces [AGH18, CCZ13, CP19, CY21, CD23a, CGK^{+}20, Cun15, EHHS23, FS18, GT09, GL06, Kar23, Lei20, MS12, Min15, PFEP23, Riv06, SST16, TT20b, Yar19, YP23, ZGD24]. spacing [ADM18]. spanning [Gan21, LW23]. Sparse [BTT19, Lev18, Zho19, BN20, BC13, BI13, CLZ18, CS20, Cas14, CCT20, DZ24, Gay20, Gay22, HHH24, HLY20, KKP21, LP18b, LM19, MMW17, Muk20, PP22, YKK21]. sparsity [CV21, DT12, GQ23, Kol09, NJG20]. Spatial [Hub12, YB06, AW24, BG95, CCL23, CM14a, HVM99, HMW01, HLT01, HLY09, HM20b, KP22a, KM24, LR16, LWYZ22, LLTY07, LTC14, Rob12, RNS23, VBLN20, Wak95, ZL17]. spatial-sign [LWYZ22]. spatially [CL24, HK13, Mus07]. Spatio [RT15]. Spatio-temporal [RT15]. SPDE [VAD22]. SPDE-based [VAD22]. SPDEs [ACP23, BVM07, GR23a, GJQS20, Qiu18, XY23]. Special [BR18a, GLF16, Nda23]. species [FLP12]. specific [KS21, NC21]. specification [GWY13, LT16, Zho14b]. Specified [BGHL21]. specifying [CT02]. spectra [AB15, DHKV15, KG00]. Spectral [AK23, BY05, DM21, Hös05, JR17, NN22, WAP17, Yin22, BHPZ15, BLP19, BBS14, Bra02, Cha97, CP15, DP09, DHKV15, GNR19, GS03, HXZ18, JP15, KVDH16, LL00, Len11, LAP20, LMZ20, MYT06, OSZ18, Pap09, TCP22, VFJ18, YZZ23, ZZW22]. Spectral-free [DM21]. spectrally [CY05, FLZ21, LCP13, Pro24]. spectrum [CK19, Com01]. sphere [BKMP09, CX16, CCF^{+}20, DLM14, Gra21, JMV20, KM20b, MNPR18]. spheres [ACPM24, Dic16, EI10, Gne13]. spherical [Cap24, Has13, MP10]. Spike [CS20]. Spiked [HZBH23, JB21a, NWR22]. spikes [Dia23]. spin [AS03, SS20]. Spine [Ban24]. spline [AG99, GN10, HHH24, KKS02]. splines [KP20, Xia20]. Splitting [MRS22, Sch01a]. splittings [FP17c]. spot [BNBV13]. Spread [VK97]. square [CGCR07, GR23b, GN08]. squared [GM04, GM19]. squares [ABH06, BDJ19, BC13, Cat11, CV97, DMR21, FG19, Gir08, GT16a, HY08, PG21, Tay13, VZ23]. St [CS06]. St. [FGK16]. Stability [DMZ11, HMS04, Jor23, Yu13, KZ21b, Tka20]. Stable [AH17b, CGMST95, DW20, LS19b, Sad13, Sur04, ALT15, AB22, AH13, ABH08, ALL09, BW23a, BYY17, BD09, BM16, Ber01, BC02, BL15b, BCG14, BDKS19, CPP11, CKHM23, CP16, CRW24, CR07, DPX23, DGP09, DK18, Eis00, FFP23, Fra06, Fra07, HP14, HK07, HMW21, IPP23, JS05, JS06, KMS03, LBDM19, MOP20, NZ05, OSZ18, OH23, OD19, OS15, PT00, PTA07, Pit97, RS96, Sam98, ST13, SS15, SXX22, WXX17, Wan23b, WYZ24, ZZ23]. stable-driven [HMW21]. stable-like [Fra06, Fra07]. stable-regenerative [BW23a]. Staff [Ano07d, Ano07e, Ano07f, Ano08c, Ano08d, Ano08e, Ano08f]. stage [HY08, Zha13]. staged [GS18a]. Standard [KT15]. star [MX23]. started [FLZ21]. state [BW23b, BK06, CR11, CM14b, CM18, DL13, FB14, GL06, HB04, OCDM08, RS01a, SS12, WK21]. state-dependent [BW23b]. state-space [CM14b, CM18, DL13]. states [GR14]. stationarity [IMM17b, KW13, Nda23, Nie13, PVD13, Tho99, TC24]. Stationary [Dor97, EGG03, IS19, SPP21, AI99, BBJ08, Bar18, BOHP18, BP03b, BR96, CCN^{+}21, CW19, CZ13, CGL12, CS24, Com01, DP09, DRW19, Deh05, DVSS16, DL16, DGP09, DLN22, FR19, FT19, GS03, GR00, HSS07, HLS14, HNTX15, ILY09, IKP18, JRSK21, KM20a, KL09, Kur22, KP06, LLS11, MM23, MZZ20, MP23, OS15, PP09, PR22, PS09, QP18, Qiu19, RvS19, SW04, Syn07, Tal13, VAD22, WSX20, WP22, WZ18, ZL13, Zho14b, vDD21]. statistic [BJSZ07, GR23b, Jon11, ZHZL19, dBCAM22]. Statistical [ACG20, Büh13, CCR20, CP18, DHKR15, HL17, HK11, KMW14, KS13b, LP18a, MNN^{+}22, WK21, AHK15, AK97, AB11, BZ19, CJL^{+}23, DMS21, FMR19, GCJL00, GHM^{+}17, NRC24, New16, Pen07, PR99, PS98, Tal13, Zäh15]. Statistics [BJ24, EFVG08, MX23, BLP19, BGT97, BHS11, BZ12, BZ14, BG00, BW10, BDDF18, CS07a, CP15, DGS08, DM08b, DDL23, FR15, FLS20, FPW17, FY20, GGT22, GGKL22, GST08, GNT17, Jor13, JN95, Leu12, LS24, LX21, Mat00, MW20, Muk20, PT18a, SZ22, SBP24, SY00, TY16, Tho23, Tur04, VV04, Wan97, Yin22, YZZ23, ZL17, ZBY17, Zho14a]. status [Cha23, PR06]. stay [Bra14, Mal16]. Steady [RS01a]. Steady-state [RS01a]. Stein [APP11, AGG23, BX00, BX06, BGX15, BRR19, BOPG22, Ček04, DP18, ER23, ERS20, Fan14, Gau17, GR23b, GS23, Hir97, LLBF24, LW18, MPS19b, NK21, OCBG19, Rig06, Sau19, Sch09, UCV17]. Stein-Tikhomirov [ER23]. Stein-type [APP11, GS23]. stem [HKM^{+}16]. step [BM07, BSG22, Kut21, WXZ19]. step-function [BM07]. stereological [SW12]. Sticky [RS18a]. Stieltjes [AH13]. Stochastic [BM15a, BGMSB14, BM08, BR12, BDP23, BKT19, DNR20, Fan21, FR06, GCJL00, GRS11, JM09, KS16, KHLN97, MMP00, MN00, MR13, RSBG22, Wal13, AG01, Ald99, Ald13, ACG19, ACG20, AKT10, Ama13, AKH17, AV14, AP00, AM07, AM16, BYY17, BCM^{+}21, BNS03, BNMS06, Bar22, BGSS22, BE16, BvdHH13, BO20, BOJGMR15, BVM07, BRS07, BDM10b, CK23, CW01, CCZ13, CFKR16, CHI17, CBL22, CK15, CD23a, CP18, DSS10, DMM^{+}20, DVSS14, DVSS16, Del17, DDBM21, DGLO19, Duf96, EM13, FN97, FS00, FK11, FN21, Fou01, FPW17, Fri24, GT09, GCJL98, GCJL99, GT16a, GMD06, GK99, GK01, Ham03, HM19, HW95, HNZ18, HDS22, HLY20, Jab21, JD16, JMRX18, KN08, KY22, Kre21a, KS13b, Kur22, Lau04, LGW06, LS02a, LM20, LØ01]. stochastic [MM15, Mar03, MR15, MSS04, MH22, MSS14, Mor99, Nie00, NR00, NS01, PS20, Pen07, PP16, Pet97, QSSS04, Rei05, Sau12, SZ22, SK23, SSX20, SU03, SXX22, TT20a, TCP22, Tin98, TGL23, Tud04, Var19, Wan16, WXX17, WT21, WYZ24, XZ18, Yu11, ZZ15, Zha06]. stock [Lin09]. Stokes [ZZ15]. stop [NV17]. stopped [DFK08, DFK10, FLT14]. stopping [ASC02, BP03a, GŠ08b, KW13]. storage [ADM14]. Strassen [NRT98]. strategies [CHS19, CS06, DDJ12, Kuj16, QLC21]. strategy [Kam18]. Stratified [SCK23, GRS11, NC21]. Stratonovich [OR24, PS20]. straw [BN20]. stream [RSS12]. streams [BCM^{+}21]. strengh [Ngu16]. strength [ST20]. Strengthened [AP14]. Strictly [Gne13, DNP06]. strings [CLR20]. Strong [ABH06, BPR12, BO18, BM09, CHR24, HW10, LMK19, SXX22, TT06, VFJ18, BK16, CSY03, ES99a, KLS98, Kre21a, PJR16, ST20, Sri16, Wan23b, Zäh15]. Strongly [HLT17, AB22, AH17b, Bra99, LT15]. Stroock [JW21].

Stroock-Williams [JW21]. structural [CN15, KKS02, Røy11]. structure [Bat17, BHK03, BBW23, DLPZ09, DLPZ12, DvdAW09, FHJK20, Fil00, HM20a, HR16, Hör08, JD16, LW17, Pap09, PS09, Sch20b, Syn07, Wer11]. Structured [KLTZ19, CGLZ24, Mar19]. structures [KTU18, LP18b, MYT06, MPU19, Pit97, VBLN20]. Student [Jon11, LO14, Zho14a]. studentized [LS24, VW17]. studies [DPW09, DPSW10, Ma10]. study [AKT10, HMT00]. Sturmfels [Din98]. sub [BL15b, DMS07, LS21, SM23, TT20b, Woo00]. sub-exponential [SM23, Woo00]. sub-Gaussian [BL15b]. sub-geometric [DMS07, LS21]. sub-Riemannian [TT20b]. subclass [NC17]. subclasses [AM07]. subcritical [BV17, GT19]. Subexponential [HW19b, DKM16, MR21]. subexponentiality [Sch99, SW05]. subgaussian [Zho19]. subgeometric [AF10, AF12]. subgraph [ER23, KW20, LP24]. subhypergraph [YS22]. subject [HNR23]. sublinear [BM16, FPSS19]. submartingales [Osȩ11]. submatrices [SN13]. submatrix [BI13]. submodel [HL12b]. Suboptimality [VZ23]. subordinated [LRMT17, MP10, Osȩ09]. subordination [BLM19, Yar19]. subordinator [HY13]. subordinators [BLLS12, Jam10, MS23d, Pit97]. Subsampling [BKMP09, Len11, PV07]. subsequence [GS21]. subsequences [LMV14]. subsets [JN95]. subsolutions [KK23a]. Subspace [ML20, CT22, HLS19, SPP21]. subtypes [GP04]. succession [Din98]. sufficiency [BC00]. Sufficient [RS01b, ZWS21, CG23, FB22a, MS23a, Mat00, Muk20]. Sum [GNR19, AEP11, FGK16, GHL05, KS16]. sum- [KS16]. summary [Wer11]. summations [BG19]. Sums [CGMST95, ABH06, BD15, BZ19, BV10, DH13, DFK08, DFK10, DN14, FLS20, GKP12, GK13, GT16b, GS08a, JK21, Kab11, LS00, LS03, LSW13, MMR23, NZ05, PT03, Röl05, Sch99, Sur04, Yu11]. sun [HG97]. sup [Ber04, GN10, KN99, LS16a, LP07, Nau22]. sup-norm [Ber04, GN10, KN99, LP07, Nau22]. super [BD20a, HWXZ24, ZX24]. super-Brownian [HWXZ24, ZX24]. super-replication [BD20a]. superadditive [GY04]. Supercritical [AAI11, BGL24, BD22, CLR14, FV03a, FV03b, Lin19, ZZ23]. superlinear [BJM21]. supermartingale [BBG19]. supermartingales [CLZ24]. supermodels [ER19]. superposed [LYT21]. superposition [CX11]. superpositions [GLST19]. supervised [CDH21, GGM20]. Support [Fou01, WY11, Cas14, DVSS14, DVSS16, GS20a, GJZ23, Gay01, HLT17, MSS00, TV06]. supported [BGMZ18]. Supports [HW95, BKY21]. suprema [Bar10, LvdG14]. supremum [AD11, CCR20, CKHM23]. supremum-type [CCR20]. sure [AZ15, AW96, BHM22, CK11b, Gas23]. surface [Qia21b]. surfaces [GC05]. survey [BC19, GJY03]. survival [BV17, FV03a, PR06, RS10]. swap [Hub12]. swept [ADM14]. switching [CCTY19, CH15, LM20, MMP00]. symbol [BS15a, SSW12]. Symmetric [FGS22, Kol08, ALT15, BD18, BTV17, CPP11, Eis00, HL12a, HL14, IKP14, KN08, KKW17, Mus07, OR24, OS15, SR09, SW19b, Wal12, Wan23b]. symmetric-Stratonovich [OR24]. Symmetrization [KMS^{+}99]. symmetrized [BO18, GW99]. symmetry [BGHL21, CHP11]. synchronous [HRY13]. synchronously [HY05]. system [BGT07, CLZ06, Gor96, OKM^{+}98]. systematic [AG99]. systems [Ald13, AW09, BS18a, CFKR16, DSS10, Del17, Fan21, Fis14, GM19, GK12, HLT17, IK22, Klo06, Kut15, MM21, MPRZ10, SS20].

Table [Ano95g, Ano95h, Ano95i, Ano96d, Ano96e, Ano96f, Ano96g, Ano97g, Ano97h, Ano97i, Ano97j, Ano98h, Ano98i, Ano98j, Ano98k, Ano99i, Ano99j, Ano99k, Ano99l, Ano99m, Ano99n, Ano00i, Ano00j, Ano00k, Ano00l, Ano00m, Ano00n, Ano01g, Ano01h, Ano01i, Ano01j, Ano01k, Ano01l, Ano02g, Ano02h, Ano02i, Ano02j, Ano02k, Ano02l, Ano03d, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano04l, Ano05h, Ano05i, Ano05j, Ano05k, Ano05l, Ano05m, Ano06i, Ano06j, Ano06k, Ano06l, Ano06m, Ano06n, Ano07j, Ano07g, Ano07h, Ano07i, Ano08g, Ano08h, Ano08i, Ano08j, Ano09g, Ano09h, Ano09i, Ano09j, Ano10g, Ano10h, Ano10i, Ano10j, Ano11g, Ano11h, Ano11i, Ano11j, Ano12g, Ano12h]. Table [Ano12i, Ano12j, Ano13i, Ano13j, Ano13k, Ano13l, Ano13m, Ano13n, Ano14f, Ano14g, Ano14h, Ano14i, Ano15e, Ano15f, Ano15g, Ano15h, Ano16f, Ano16g, Ano16h, Ano16i, Ano17d, Ano17c, Ano20, Ano21a, Ano21b, Ano21c, Ano21d, Ano22a, Ano22b, Ano22c, Ano22d, Ano23c, Ano23d, Ano23e, Ano23f, Ano24e, Ano24f]. tables [HAT10]. Tail [BW23a, DGS20, GT16b, HS08, SY00, TK03, Zho14a, ZX17, ADM14, BD13b, BDZ15, CP01, CV97, DPS24, DDFD04, EKS08, FN23, HS07, HMS04, JK21, KL09, KKP07, KL10, LLD19, Sau19, WT13]. tailed [ABH00, BD20b, BG21, FGS21, Hil15, HMW21, JMRX18, Jur99, Kam18, KL05, LT00, PS22a, PT00, PTL04, Xu21, ZZ23]. tailedness [Jon11]. tails [BL23, DFK08, DFK10, HM19, LS00, Nov14, Sam98]. Talagrand [CR20]. tale [HL12a]. Tan [BW08]. Tanaka [BDM10a]. target [Kam18]. Targeted [ZDY23]. targets [SR09]. task [TRB24]. tau [BD14]. Tauber [JM97]. Tauberian [BW19]. Taylor [HWW99]. technique [PR02a]. techniques [FO24, GRS11, JP15, RTT20, Tru20a]. telegraph [KK18]. temperature [FFZ22]. temperatures [LM13]. tempered [CR07]. Template [ACZ23, BLL15]. temporal [Fou01, HTW23, Pap09, RT15]. temporally [ZS15]. tend [CP12]. Tensor [HWW11, MN22]. Tensor-based [HWW11]. tensors [Ver20]. term [AG01, Fil00, GR23a]. terminal [GS16, GLL20, HLZZ23]. terms [GLF16, KVZ19, PT15]. tessellations [DK18, GKT24, HSS07, ST13]. Test [KL09, AGM06, ADM18, BD14, CV09, DV01, GKNY11, HR13, JZ19, KC19, LN14, LS19c, LWY24, PVD13, TC24]. Testing [ABD14, AHHK08, BDR14, CW22, Car15, DV01, FP17a, GJ13, JP15, KM23, KTU18, MV18a, Pap09, RK10, Tod20b, WT21, Wan24, APGMV04, Bar02, CM13a, CK23, CEL20, CV21, CLZ24, CPY13, CCF^{+}20, DR16, DNS19, DT23b, EM03, Gay01, HW10, HP02a, HLL23, HR22b, Khm16, LWCS22, LWLL22, LS99, MVY19, Nob06, PS13a, RW19, Sar16, SPP21, TAS^{+}17, Vet15, YS22, ZHZL19, vDD21]. tests [BGHL21, BG09, BJSZ07, BP19b, BHVD17, CK11a, CHP11, CX24, CSSC95, CM11, DGU05, DCC14, Dob23, DX16, EM13, FRW15, GPNECA23, GLF16, HPV13, HLS14, JMV20, JJW24, Jur99, JK12, KP23, Kut15, LJ03, LP99, LAP20, Loc12, NP08, PV16, PRV20, PT18a, Sko96, SCK23, TZ17, ZZZ17, ZS21, vDD22]. their [AK18, AAI11, BJSZ07, CP12, Das05, DGS08, DR19, FO24, GS11, JK12, KF14, LMV14, RT96, Tal13, Xu21]. theorem [AP19, Akr00, Ath00, BCRT24, BM16, Beg07, Bra99, BGGL21, CL12, CF15, DJ97, DVSS14, DVSS16, DF20, EO07, FMPV14, Fou01, GP11, GV19, HLP12, HNTX15, JB21a, JB21b, JS05, JS06, KLLM13, LS16b, LN02, LPW04, LTX22, MZZ20, MSS00, Mou13, PZ19, Qiu18, Rai19, RT15, SW08, YZZ23, BLP19, KT18, PDL19]. theorems [ALOV19, AP00, AF10, AF12, BBCM11, BT17a, BT20, BRR19, BP16, BP03b, BB99, BL14, BST12, BD22, CG99a, DNN10, DL16, DP18, EK11, FPSS19, GNNS21, GCJL98, GW17a, HSS07, IPP23, Jia13, Kab11, KY22, KS99b, LR16, LS19b, LRMT17, MP10, NZ05, Pak99, PR16, PT08, Per02, PT03, PTA07, PFEP23, RR15, SZ16, Tho23, Tur04, Ued21, VY13, Yan08, ZP11]. theoretic [Ruk14]. theoretical [Wan22b, dCCGZ23]. Theory [CK15, LN18, LXY13, Ald99, AC17, BVP22, BOHP18, BBS15, BS22, BS16, BL04, BDV21, CHWW05, CY05, CMMA20, DRW19, DFS10, DF98, DY14, FZV23, FK12, GJ13, HLS19, JM97, KL05, KJ20, Kul07, Leu12, NP08, PR22, RS18b, TY16, Tho99]. there [BL04, CT19, HJM14, MCC^{+}21]. Thermodynamics [Mau12]. thin [LS00]. things [Dia13]. Thinned [LC22]. third [GL18, Klo06]. third-generation [Klo06]. Thompson [CDJ13]. three [DVSS14, DVSS16, IK22, ST20]. three-particle [IK22]. threshold [CDL24, GS18b, LTL07, LLS11]. thresholding [DJ96, KLP12]. thresholds [ADN24, JK17]. throwing [EI10]. Thurstonian [ERS19]. Tie [HLPP09]. Tie-respecting [HLPP09]. tight [BD15]. Tightness [AMPV10]. Tikhomirov [ER23]. tilting [Lau13]. Time [Bad00, BG95, Cra15b, FLZ21, LS16a, RvS19, SU08, AW12, Alt21, AHS15, AD11, ADFS15, ACLZ17, ADS13, BP03a, BYY17, BLLS12, BKK23, BNH08, Bei97, BPW19, BW22, BB99, BDP23, BGT07, BRE23, BDG07, BDM10a, BHVD17, BS18b, BJ24, BDKS19, Büh97, But24, CFKR16, CT12, CZ13, CHI17, CK15, CD11, CM18, CSY99, Cut00, DNV99, DR07, DP09, Dan96, Dan00, DN08, DMMW18, DPTV15b, DT23a, Deh05, DV18, DDK22, DW23, DKM16, ES99b, FM00, FLT14, FLLO14, FFZ22, FKM02, FR11, GWY13, GL18, GMR19, GW98, GHL05, GZ17, GQ23, GK99, GK01, Han03, HLT17, HK08, HM23, JS08, JS17, JW21, JT22, JK21, Jou07, KL09, KY22, KJ20, KS97, KKP21, KR24, Kre21b, KS13b, Küh19, Len11, LP00, LPSS18, LGS08, LT15, LYT21]. time [LRS23, LS02b, LM20, MMVW10, MNN^{+}22, MMS08, MZZ20, MS23c, NP08, NPS20, PP05, Pap09, Pec03, PFEP23, PGV17, PBD15, RZ23, Røy11, RS01b, RW02, SSH14, SS12, Sch24, SK23, SW19b, Ste13, SvS17, SGL^{+}24, Syn07, TKT09, TC24, Tru19, Tru20a, Tru20b, TC03, Tsa09, UDS23, VY13, Var19, WAP17, Wan17, WP22, Wan23b, WZ18, YY24, ZP11, Zei15, Zho14b, vDD21, vDD22]. Time-changed [FLZ21, LS16a, Wan23b]. time-continuous [BHVD17]. time-dependent [KY22]. Time-frequency [RvS19]. Time-invariance [Bad00]. Time-scales [BG95]. time-series [KS97]. time-space [Pec03].

Time-varying [Cra15b, DNV99, DR07, FR11, Kre21b, SGL^{+}24, Tru20b]. times [AIZ16, ADFS15, BL23, BBD20, CD21, CC00, CY05, Das05, DH17, DPX23, DMY00, DFK16, Eis00, GM15, dV13, HP14, KM20a, Ken15, LT00, MWX08, PRR19, PY03, PT18b, TT20b, Tod20b]. timescale [SK23]. together [ABF12]. tomography [Cav01]. tools [QLC21]. topic [BBW20]. topics [BBW20]. topographical [BTV17]. Topological [BMT17, Fri24, LW23, OB20, SBP24]. topologies [Sri16]. Total [BL14, PRR13, BC16, FRW15, GZ17, KT18, MS10]. tournaments [LM19]. toxicity [DPW09]. toxicology [DPSW10]. trace [CK19]. Tracy [HLY20, JJLW22, Ma12]. Trade [CLZ07]. Trade-offs [CLZ07]. trading [TKT09]. training [LK08]. trajectorial [TY23]. transelliptical [HL17]. transfer [MS17]. transform [Ben03, BDZ15, But24, JHD20, PZ19, RS01b, dBM17]. transform-based [RS01b]. transformation [CV19, GCP17, JN00]. transformations [KM20b, Khm16, PR99]. transforms [AH13, BP16, Bar18, BK06, HY13, PY99]. transience [BR97, Deb99, IRVZ22, IRVZ24, San13]. transient [BD20a, JLM14]. transit [GM15]. Transition [MSS04, GS00a, GHL05, HU22, KM05, OU23, RW02]. transitions [CK23, CJ22, HKK19, JK17, LR23]. translated [Röl05]. translation [GR16]. transmission [RSS12]. Transport [CEG17, NWR22, BVP22, BLO19, FS18, HT19, HKMS24, NZM23]. Transportation [Blo04, MSWW11, PEBG17, Sau12, AS19, HLL23]. traveling [Tka20]. Tree [IRVZ22, PT18b, ZS12, CGLZ24, CRW24, CG21, DHKR15, DLWZ17, JT22, Lin19, LW23, MN22, MO08, Ran22, IRVZ24]. treed [Bul15]. trees [Alt13, AMV18, BEMR17, BWW00, CL02, Gan21, GS18a, HW19a, JS08, KZ21a, MM20, MW04, MPW08, MPV07, OH23, GL95]. trek [DRW20]. trend [BGL22, BS12, CHI17, KM24, Sch24]. trends [Rob12]. trials [HZ04]. triangle [Yu10]. triangles [Gat06]. triangulation [CH18]. Tricentenary [Sen13]. trigonometric [CP23]. trimmed [BHS11, HJQZ04, Mas04a]. trimming [ÁEdBCAM12, CLD08]. true [NLL18, RS18a]. Truncated [CHHB19, BŁ15a, KK16, dUÁV21]. truncation [SW08]. try [FB22b]. tube [TK03]. tug [Hei18]. tug-of-war [Hei18]. Tukey [DGC11, Mas04a]. tuning [BPR^{+}13, CP21]. Turán [HR22a]. Turing [CG23]. turning [PFEP23]. twice [VD13]. Two [BYY17, BSG22, MR19, SK23, ZS15, ZZZ17, ACJ18, BD18, BG09, Car06, CT02, CCFH00, CC04, CX24, DDL23, FN97, Gor96, Gra21, HL12a, Han09, HY08, JMM17, LMMR16, Mac23, MSS00, MX23, PR06, Pey17, Pon07, QJ22, TAS^{+}17, Zha00, ZS21]. two-component [BD18, LMMR16, QJ22]. two-level [Gor96]. two-parameter [CCFH00, Han09]. two-point [FN97]. Two-sample [ZZZ17, BG09, CC04, CX24, Pon07, TAS^{+}17, Zha00]. two-sex [JMM17]. Two-sided [MR19]. two-stage [HY08]. Two-step [BSG22]. Two-time-scale [BYY17]. Two-timescale [SK23]. two-way [Pey17]. type [APP11, AP14, Alt13, AM07, ADS13, BFT15, Bar10, BC19, BSS05, BDM10a, BGGL21, CCR20, EM18, FGLS19, FLS23, Fri18, GJZ23, GS23, GLST19, HKK19, HG97, HLT17, IKP14, JK17, JT03, JWY22, Jou07, Jou09, KL09, KK23a, KM05, KV12, Küh22, KS01, LMP12, LX23, Lim19, LSW13, LTX22, ML11, Mau19, Olo09, Osȩ09, PS09, Qiu18, SZ16, SW19b, SM23, TV06, Tud04, ZZW22, vEGS07, NC17, FLS20, Zha23]. types [CJ22, Mac23].

Uhlenbeck [BNBO11, BS22, BPD23, Fas10, FK11, GLST19, HP14, Jac96, JVV05, LP00, Mai14, Mas04b, Nku21, PS09, Wan11]. UMD [Yar19]. unadjusted [BRE23, DM19]. Unbiased [ARV18, AKH17, RP23, VW15]. unbiasedness [JK12]. unbounded [AGM06, GHL05, Lei20, Yin22, Zha23]. uncensored [MRS22]. uncertainty [BG17]. underlying [CA19]. understand [Mon15]. undirected [KT15]. undiscounted [GZ17]. unequal [Yu12]. unified [CMMA20, CT22, FZV23, HPY18, LMC23, LSW20, MS12]. Uniform [AN12, ALV18, BBC12, Bru19, BP08, DR19, DM08b, ES99b, Hei18, Mar19, San10, SW13, SXXZ18, WC14, BH17, BEMR17, CSY99, DR09, HLP12, KMB21, LW23, Stu23, WT21, Zha22]. uniformity [GPNECA23, JMV20]. Uniformly [AH17b, DDL23, Ham03, UDS23, VW15]. unifying [CH07, DGS08, HKMS24]. unimodal [CL19, SR09]. Unimodularity [Alt13]. uniqueness [BS02, Jab21]. unit [DLM14, Kat09]. Unitary [Khm16]. Universal [GL08, NV17, PZ14, QLC21, KLS98, VS05]. Universality [CJP21]. unknown [Bei97, BBW20, DRT16, GR14, Gir08, JS13b, LMMR16, MS23b]. unlimited [GMM05]. unseen [FN23]. unspecified [BGHL21]. unstable [Ber01, DvdAW09, GM15]. unsupervised [GP04]. unusual [GLST19]. up-down [RW22]. Upper [Cha03, Dor98, Küh22, Lep16, CKHM23, HMP18b, SW19b]. urn [AGV18, BT17b, BHBO17, BDM09, HZ04]. urns [CJ22, PRR19, SFP23]. use [DR19]. Using [LR21, AKH17, ACG13, BRR19, BE16, BRJN19, BZ12, CDJ13, CG22b, CL02, CS07a, Dai17, DRD20, DCC14, FP17c, GC05, GT22, GT16a, GP12, HW99, JP15, JN16, LLTY07, LS14, MK20, NZM23, OSZ18, QLC21, RP23, Rig06, Zha06, ZGV20].

vaccination [BGMSB14]. validated [MYT06]. validation [HP01, ZDY23]. Vallois [CH07]. value [BVT96, BDG05, BDV21, CLR14, DFS10, Dom15, FN97, GKNY11, HT00a, KL05, MO08, PQY13, ZHZL19, KFT21]. valued [BL22, BZ19, Bou06, Cap24, CPR13, CDH21, DH13, DMY00, DvdAW09, FGW12, Jir18, LS16b, PS13b]. values [Ass23, CK11a, Cha03, CJP21, CRS00, DZ24, GNT17, IKP13, MP06, Ued21, Wan24]. Vapnik [AN12]. VAR [HLL23]. Varadhan [DZ03]. variability [KM98, dSJMT23]. Variable [CCN^{+}21, ML10, WXZ15, Bia14, Lev18, MSW13, RP23, SV13]. variables [AEP11, BD15, BL04, Bou06, BV10, But24, Che02, DH13, DN14, DLL20, DL17, DLT09, ER23, ER19, GKP12, Gau17, GT16b, HMP18a, HY13, KMS^{+}99, Kur16, LS16b, LS00, LS03, LM15, MR21, PXH13, PTA07, Pon07, Röl05, San10, Vii16, Zho14a]. Variance [BH96, ELP24, KJ20, APP11, AP14, BIPZ22, BGT07, BLM19, CMS10, DS96, FB22a, Gir08, HT00b, HM23, JM97, KM06, OD19, PT03, PTA07, SvS17, Sur04, TC24, TMW13, VFJ18, VW15]. variance-based [TC24]. variance-covariance [SvS17]. variance-optimal [DS96]. variances [Bra99, PWZ17]. variation [AT24, BC16, BNS03, BNCP11, BŁ15a, BZ23, BL14, CD23b, CNW06, DR11, FRW15, GG22, GLT23, GS16, HL06, KMS03, KT18, LT15, MS10, Man07, Nor15, PRR13, TKT09, Wal13, Zei15]. Variational [BD13a, CM14a, HM23, RASY17, CG99a, MPRZ10, MR15]. variations [Beg07, CP21, CD23a, PR16]. variety [FLP12]. variogram [PS11, PS13c]. various [BP98, GZ21b]. Varying [WFM10, BAT19, BP21, Bha23, BL14, CTP21, Cra15b, DNV99, DR07, DW23, FH05, FN23, FR11, HS08, JK21, KM22, KSX19, Kre21b, LMP12, MR10, SGL^{+}24, Tru20b, XW12]. Varying-coefficient [WFM10, FH05, XW12]. varying-size [BAT19]. vector [BBW23, DR16, MS23b, ELP22, TV06, WY11, ZC18]. vectors [EKS15, HMRB08, HS08, Kat09, KM98, MY23, MSS14, NZ05, SL18]. Verifiable [CA19]. version [Ose20]. versions [BS15b]. vertex [HW19a, Jan07]. vertices [BL13]. very [CK11a]. via [ABD14, AKT10, Bah21, BP10, BMT17, Bou06, CLZ24, CPY13, CK01, DNN10, Dit19, DSW21, DM19, DLT09, ERS20, FLS01, FH07, Gao20, GP11, GS20b, GK12, HMT00, HMWY10, HP02b, HMW21, IR20, JN20, KJ13, KW20, Kol08, Kra21, Küh19, LP18b, Lev18, LAP20, LMZ24, Lim19, MSS04, MO08, NK21, OU23, RR15, Roo05, RK10, RS18b, Sar16, SSW12, SW20, TK03, Tur04, VZ23]. view [FdH14, FFS10, LR23]. Viot [ALR23]. virtual [BBD20]. virtue [GR04]. Viscosity [MPRZ10, Wan17, CR22]. Viterbi [CR11, LK08]. Vladimirovitch [Ano97c]. Vlasov [DJL23, TT20a, TY23, Wan23a]. vol [Ano95b, Ano95a, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano95i, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano97b, Ano97a, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano97i, Ano97j, Ano98b, Ano98a, Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano98h, Ano98i, Ano98j, Ano98k, Ano99b, Ano99a, Ano99c, Ano99d, Ano99e, Ano99f, Ano99g, Ano99h, Ano99i, Ano99j, Ano99k, Ano99l, Ano99m, Ano99n, Ano00b, Ano00c, Ano00d, Ano00e, Ano00f, Ano00g, Ano00h, Ano00i, Ano00j, Ano00k, Ano00l, Ano00m, Ano00n, Ano01a, Ano01b, Ano01c, Ano01d, Ano01e, Ano01f, Ano01g, Ano01h, Ano01i, Ano01j, Ano01k, Ano01l, Ano02a, Ano02b, Ano02c, Ano02d, Ano02e, Ano02f, Ano02g, Ano02h, Ano02i, Ano02j]. vol [Ano02k, Ano02l, Ano03a, Ano03b, Ano03c, Ano03d, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano04a, Ano04b, Ano04c, Ano04d, Ano04e, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano04l, Ano05b, Ano05a, Ano05c, Ano05d, Ano05e, Ano05f, Ano05g, Ano05h, Ano05i, Ano05j, Ano05k, Ano05l, Ano05m, Ano06b, Ano06a, Ano06c, Ano06d, Ano06e, Ano06f, Ano06g, Ano06h, Ano06i, Ano06j, Ano06k, Ano06l, Ano06m, Ano06n, Ano07c, Ano07j]. volatility [BNS03, BNBV13, BGSS22, BS16, BE16, Cho18, FP17a, GCJL98, GCJL99, GCJL00, HM19, JD16, JMRX18, KLW18, LM07, MR13, NK21, PV09, Ros09, VSV03, VD12, Vet15, Zha06]. Volterra [BNBV13, BS16, BE16, BDM10a, DNN10, FS00, Jab21, NR00].

Volterra-type [BDM10a]. volumes [BST12]. Voronoi [GKT24, TY16]. voter [BFMP01, CD95, MMVW10, SSY21]. vs [DJL23].

Wald [DX16, LJ03]. walk [BHM18, Bha23, BGL24, CRS00, DPTV15b, FS10, FPW17, GL18, GLL20, Hie17, IRVZ22, IRVZ24, JT03, KVZ19, KS01, Lin19, LW23, Mal17, MR10, SR09]. walks [AP19, BD20b, BL14, Bor21, BRS16, CD01, CdLOR20, CR17, GW17a, IPP23, LPSS18, LS20, LS14, Mus07, RASY17, WXZ19]. Walsh [IS19]. war [Hei18]. warped [KP04]. Wasserstein [ACJ18, ÁEdBCAM18, BX06, BGX15, BLL15, Bor21, DGU05, DR23, Jou09, JM22, Lei20, MS10, NWR22, RS18b, Sch09, STC^{+}24, Wan16, WB19, Wie22, ZP19]. Watanabe [ALR23]. Watson [Alt13, BWW00, CTP21, Dem07, HW19a, IKP14, KM22, KZ21a, Lin19, OH23, Sid18]. wave [DSS10, DVSS14, DVSS16, KN08, MSS00, QSSS04, SSX20, Tka20]. Wavelet [AD18, GL05, RSGW03, BBJ08, BS12, CGK^{+}20, CSvS07, DNV99, GN10, HP95, HP01, KLLM13, PW04, PTA07]. wavelet-based [BBJ08, CSvS07, PTA07]. wavelets [Efr00, KP04]. waves [ALW11, Gas23]. way [Che02, Pey17]. Weak [BJ00, BBV17, BvdHH13, BH00, BLM19, CMS10, DG13, FRW04, GT09, GM15, HW19b, Jab21, RWZ17, RZ23, ZH20, AS19, BVP22, BR18a, BN17, Ber97, BJM21, CHR24, DS05, Dob23, FM02, Fri18, GLS07, KN06, KFT21, Leu12, MZZ20, Osȩ09, PRV20, Riv06, Sri16, SXX22, Yar19]. weak-type [Osȩ09]. weakly [AW12, BGT97, BZ19, BST12, Fis14, Kra21, Kul07]. wedge [FR19]. Weibullian [AD11]. Weighted [LS00, Nor15, PQY13, Sau19, ZK20, ACG20, BCRT24, BD15, BBV17, BZ12, CEG17, DPS24, DGU05, GK13, HIW19, LS03, Ren08, Yu11]. weighting [LR21]. weights [BDZ15, Gan21, GT16a, LM13]. Welch [Zho14a]. Well [RZ21, KK21, Kra21]. Well-posedness [RZ21, Kra21]. we’ve [Dia13]. WGANs [STC^{+}24]. Which [Ber03, Ber06a, Obl06]. white [ACBLV18, DNN10, Fou01, HP02a, IS98, Ma12, Pan22, WYZ24]. white-noise-driven [Fou01]. Whittle [BSW24]. whose [BRS07]. Wick [BP10]. Wide [ÁEdBCAM18]. Widom [HLY20, JJLW22, Ma12]. width [FFP23]. Wiener [BT20, LS16b, MCSS99, Rév96]. Wiener-chaos [MCSS99]. Wigner [BY05, EX23, GNTT18, LX21]. Wijs [Mon15]. Williams [JW21]. windings [Dor98]. Winsorization [Ore22]. wireless [KRX18, Klo06]. wise [ZC23]. Wishart [BDT21, Ma12]. within [Dom15]. without [BM15b, Ber97, BMSZ21, Dai17, HWZZ23, LR21, MN00, MFWB22, Myk01, PT21, TMW13, Yu12]. Wold [BL22, HL06]. word [BH10]. words [HM15]. work [Ano95a, Ano97a, Ano98a, Ano99a, Ano00a, Ano05a, Ano06a]. workload [HP06, RBM19]. wrapped [KJ13]. Wright [CY21, Zho19].

Yaglom [CTP21]. Yor [MW04, ARAS20, CH07, KV12]. Young [Blo04, BH10, HM15].

Zero [GHL05, CP12, Gau17, TK03, Wal13, Wan97]. zero-one [Wan97]. Zerosum [GHL05]. zeros [CP23]. zeroth [RSBG22]. zeroth-order [RSBG22]. zeta [LH01, Nak15]. Zheng [RVW01]. zones [CD95, MMVW10]. zonoid [MSS14]. zonoids [KM98].

References

Ameijeiras-Alonso:2024:SCN

[AAG24] Jose Ameijeiras-Alonso and Irène Gijbels. Smoothed circulas: Nonparametric estimation of circular cumulative distribution functions and circulas. Bernoulli, 30(4):2747–2769, November 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

Amewou-Atisso:2003:PCS

[AAGGR03] Messan Amewou-Atisso, Subhashis Ghosal, Jayanta K. Ghosh, and R. V. Ramamoorthi. Posterior consistency for semiparametric regression problems. Bernoulli, 9(2):291–312, April 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1068128979.

Athreya:2011:SAD

[AAI11] Krishna B. Athreya, Siva R. Athreya, and Srikanth K. Iyer. Supercritical age-dependent branching Markov processes and their scaling limits. Bernoulli, 17(1):138–154, February 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1297173836.

Aamari:2023:MBE

[AAL23] Eddie Aamari, Catherine Aaron, and Clément Levrard. Minimax boundary estimation and estimation with boundary. Bernoulli, 29(4):3334–3368, November 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-4/Minimax-boundary-estimation-and-estimation-
with-boundary/10.3150/23-BEJ1585.full.

Arlot:2011:MAM

[AB11] Sylvain Arlot and Peter L. Bartlett. Margin-adaptive model selection in statistical learning. Bernoulli, 17(2):687–713, May 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1302009243.

Adamczak:2015:SRM

[AB15] Radosław Adamczak and Witold Bednorz. Some remarks on MCMC estimation of spectra of integral operators. Bernoulli, 21(4):2073–2092, November 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1438777586.

Alishahi:2022:PPR

[AB22] Kasra Alishahi and Milad Barzegar. Paving property for real stable polynomials and strongly Rayleigh processes. Bernoulli, 28(2):1198–1223, May 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-2/Paving-property-for-real-stable-polynomials-
and-strongly-Rayleigh-processes/10.3150/21-BEJ1383.
full.

Akakpo:2014:TML

[ABD14] Nathalie Akakpo, Fadoua Balabdaoui, and Cécile Durot. Testing monotonicity via local least concave majorants. Bernoulli, 20(2):514–544, May 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1393593996.

Aulbach:2012:MPT

[ABF12] Stefan Aulbach, Verena Bayer, and Michael Falk. A multivariate piecing-together approach with an application to operational loss data. Bernoulli, 18(2):455–475, May 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1334580720.

Asmussen:2000:RES

[ABH00] Søren Asmussen, Klemens Binswanger, and Bjarne Højgaard. Rare events simulation for heavy-tailed distributions. Bernoulli, 6(2):303–322, April 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1081788030.

Aue:2006:SAS

[ABH06] Alexander Aue, István Berkes, and Lajos Horváth. Strong approximation for the sums of squares of augmented GARCH sequences. Bernoulli, 12(4):583–608, August 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1155735928.

Aue:2008:SSB

[ABH08] Alexander Aue, István Berkes, and Lajos Horváth. Selection from a stable box. Bernoulli, 14(1):125–139, February 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1202492787.

Abraham:2019:NBP

[Abr19] Kweku Abraham. Nonparametric Bayesian posterior contraction rates for scalar diffusions with high-frequency data. Bernoulli, 25(4A):2696–2728, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1568362040.

Alili:2001:NFI

[AC01] Larbi Alili and Loïc Chaumont. A new fluctuation identity for Lévy processes and some applications. Bernoulli, 7(3):557– 569, June 2001. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1080004766.

Arias-Castro:2017:STO

[AC17] Ery Arias-Castro. Some theory for ordinal embedding. Bernoulli, 23(3):1663–1693, August 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1489737621.

Arias-Castro:2015:DPC

[ACBL15] Ery Arias-Castro, Sébastien Bubeck, and Gábor Lugosi. Detecting positive correlations in a multivariate sample. Bernoulli, 21(1):209–241, February 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1426597068.

Arias-Castro:2018:DMR

[ACBLV18] Ery Arias-Castro, Sébastien Bubeck, Gábor Lugosi, and Nicolas Verzelen. Detecting Markov random fields hidden in white noise. Bernoulli, 24(4B):3628–3656, November 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1524038765.

Arias-Castro:2013:CDN

[ACG13] Ery Arias-Castro and Geoffrey R. Grimmett. Cluster detection in networks using percolation. Bernoulli, 19(2):676–719, May 2013. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1363192043.

Aletti:2019:NRS

[ACG19] Giacomo Aletti, Irene Crimaldi, and Andrea Ghiglietti. Networks of reinforced stochastic processes: Asymptotics for the empirical means. Bernoulli, 25(4B):3339–3378, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL https://projecteuclid.org/euclid.bj/1569398768.

Aletti:2020:IRS

[ACG20] Giacomo Aletti, Irene Crimaldi, and Andrea Ghiglietti. Interacting reinforced stochastic processes: Statistical inference based on the weighted empirical means. Bernoulli, 26(2):1098– 1138, May 2020. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1580461574.

Alfonsi:2018:EWD

[ACJ18] Aurélien Alfonsi, Jacopo Corbetta, and Benjamin Jourdain. Evolution of the Wasserstein distance between the marginals of two Markov processes. Bernoulli, 24(4A):2461–2498, November 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1522051215.

Aue:2017:PQA

[ACLZ17] Alexander Aue, Rex C. Y. Cheung, Thomas C. M. Lee, and Ming Zhong. Piecewise quantile autoregressive modeling for nonstationary time series. Bernoulli, 23(1):1–22, February 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1475001346.

Altmeyer:2023:PES

[ACP23] Randolf Altmeyer, Igor Cialenco, and Gregor Pasemann. Parameter estimation for semilinear SPDEs from local measurements. Bernoulli, 29(3):2035–2061, August 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-3/Parameter-estimation-for-semilinear-SPDEs-
from-local-measurements/10.3150/22-BEJ1531.full.

Arnaudon:2024:SCP

[ACPM24] Marc Arnaudon, Koléhè Coulibaly-Pasquier, and Laurent Miclo. On the separation cut-off phenomenon for Brownian motions on high dimensional spheres. Bernoulli, 30(2):1007–1028, May 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-30/issue-2/On-the-separation-cut-off-
phenomenon-for-Brownian-motions-on/10.3150/23-BEJ1622.
full.

Ambrosio:2009:AMD

[ACV09] Luigi Ambrosio, Vincenzo Capasso, and Elena Villa. On the approximation of mean densities of random closed sets. Bernoulli, 15(4):1222–1242, November 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1262962233.

Arias-Castro:2023:TMR

[ACZ23] Ery Arias-Castro and Lin Zheng. Template matching with ranks. Bernoulli, 29(1):52–74, February 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-1/Template-matching-with-ranks/10.3150/21-
BEJ1449.full.

Abraham:2007:ASF

[AD07] Romain Abraham and Jean-François Delmas. Asymptotics for the small fragments of the fragmentation at nodes. Bernoulli, 13(1):211–228, February 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1175287730.

Arendarczyk:2011:ASD

[AD11] Marek Arendarczyk and Krzysztof Dȩbicki. Asymptotics of supremum distribution of a Gaussian process over a Weibullian time. Bernoulli, 17(1):194–210, February 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1297173839.

Abry:2018:WEO

[AD18] Patrice Abry and Gustavo Didier. Wavelet estimation for operator fractional Brownian motion. Bernoulli, 24(2):895–928, May 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1505980882.

Aue:2015:RTM

[ADFS15] Alexander Aue, Christopher Dienes, Stefan Fremdt, and Josef Steinebach. Reaction times of monitoring schemes for ARMA time series. Bernoulli, 21(2):1238–1259, May 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1429624977.

Agapiou:2021:RCP

[ADH21] Sergios Agapiou, Masoumeh Dashti, and Tapio Helin. Rates of contraction of posterior distributions based on p - exponential priors. Bernoulli, 27(3):1616–1642, August 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-27/issue-3/Rates-of-contraction-of-
posterior-distributions-based-on-p-exponential/10.3150/
20-BEJ1285.full.

Arendarczyk:2014:TAA

[ADM14] Marek Arendarczyk, Krzysztof Dȩbicki, and Michel Mandjes. On the tail asymptotics of the area swept under the Brownian storage graph. Bernoulli, 20(2):395–415, May 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1393593992.

Azais:2018:PST

[ADM18] Jean-Marc Azaïs, Yohann De Castro, and Stéphane Mourareau. Power of the spacing test for least-angle regression. Bernoulli, 24(1):465–492, February 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1501142452.

Auddy:2024:EDT

[ADN24] Arnab Auddy, Nabarun Deb, and Sagnik Nandy. Exact detection thresholds and minimax optimality of Chatterjee’s correlation coefficient. Bernoulli, 30(2):1640–1668, May 2024. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
30/issue-2/Exact-detection-thresholds-and-minimax-optimality-
of-Chatterjees-correlation-coefficient/10.3150/23-BEJ1648.
full.

Aurzada:2013:STC

[ADS13] Frank Aurzada, Leif Döring, and Mladen Savov. Small time Chung-type LIL for Lévy processes. Bernoulli, 19(1):115–136, February 2013. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1358531743.

Alvarez-Esteban:2012:SST

[ÁEdBCAM12] Pedro C. Álvarez-Esteban, Eustasio del Barrio, Juan A. Cuesta-Albertos, and Carlos Matrán. Similarity of samples and trimming. Bernoulli, 18(2):606–634, May 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1334580726.

Alvarez-Esteban:2018:WCA

[ÁEdBCAM18] Pedro C. Álvarez-Esteban, Eustasio del Barrio, Juan A. Cuesta-Albertos, and Carlos Matrán. Wide consensus aggregation in the Wasserstein space. Application to location-scatter families. Bernoulli, 24(4A):3147–3179, November 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1522051236.

Arbenz:2011:AAF

[AEP11] Philipp Arbenz, Paul Embrechts, and Giovanni Puccetti. The AEP algorithm for the fast computation of the distribution of the sum of dependent random variables. Bernoulli, 17(2):562– 591, May 2011. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1302009237.

Atchade:2010:LTS

[AF10] Yves Atchadé and Gersende Fort. Limit theorems for some adaptive MCMC algorithms with subgeometric kernels. Bernoulli, 16(1):116–154, February 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1265984706.

Atchade:2012:LTS

[AF12] Yves F. Atchadé and Gersende Fort. Limit theorems for some adaptive MCMC algorithms with subgeometric kernels: Part II. Bernoulli, 18(3):975–1001, August 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1340887010.

Anevski:2019:LPM

[AF19] Dragi Anevski and Anne-Laure Fougères. Limit properties of the monotone rearrangement for density and regression function estimation. Bernoulli, 25(1):549–583, February 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1544605256.

Abdous:2008:EBE

[AFGS08] Belkacem Abdous, Anne-Laure Fougères, Kilani Ghoudi, and Philippe Soulier. Estimation of bivariate excess probabilities for elliptical models. Bernoulli, 14(4):1065–1088, November 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1225980571.

Abramovich:1999:DEK

[AG99] Felix Abramovich and Vadim Grinshtein. Derivation of equivalent kernel for general spline smoothing: a systematic approach. Bernoulli, 5(2):359–379, April 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1173147911.

Alabert:2001:SRD

[AG01] Aureli Alabert and István Gyöngy. On stochastic reactiondiffusion equations with singular force term. Bernoulli, 7(1): 145–164, February 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1080572343.

Azzalini:2007:GCN

[AG07] Adelchi Azzalini and Marc G. Genton. On Gauss’s characterization of the normal distribution. Bernoulli, 13(1):169–174, February 2007. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1175287727.

Azmoodeh:2023:ASO

[AGG23] Ehsan Azmoodeh, Dario Gasbarra, and Robert E. Gaunt. On algebraic Stein operators for Gaussian polynomials. Bernoulli, 29(1):350–376, February 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-29/issue-
1/On-algebraic-Stein-operators-for-Gaussian-polynomials/
10.3150/22-BEJ1460.full.

Arratia:2018:VBK

[AGH18] Richard Arratia, Skip Garibaldi, and Alfred W. Hales. The van den Berg–Kesten–Reimer operator and inequality for infinite spaces. Bernoulli, 24(1):433–448, February 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1501142450.

Azais:2006:ADL

[AGM06] Jean-Marc Azaïs, Élisabeth Gassiat, and Cécile Mercadier. Asymptotic distribution and local power of the log-likelihood ratio test for mixtures: bounded and unbounded cases. Bernoulli, 12(5):775–799, October 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1161614946.

Aletti:2018:DAR

[AGV18] Giacomo Aletti, Andrea Ghiglietti, and Anand N. Vidyashankar. Dynamics of an adaptive randomly reinforced urn. Bernoulli, 24(3):2204–2255, August 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1517540473.

Arizmendi:2013:CEC

[AH13] Octavio Arizmendi and Takahiro Hasebe. On a class of explicit Cauchy–Stieltjes transforms related to monotone stable and free Poisson laws. Bernoulli, 19(5B):2750–2767, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1386078619.

Abrahamsen:2017:CAB

[AH17a] Tavis Abrahamsen and James P. Hobert. Convergence analysis of block Gibbs samplers for Bayesian linear mixed models with p > N. Bernoulli, 23(1):459–478, February 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1475001361.

Ayache:2017:USC

[AH17b] Antoine Ayache and Julien Hamonier. Uniformly and strongly consistent estimation for the Hurst function of a Linear Multifractional Stable Motion. Bernoulli, 23(2):1365–1407, May 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1486177402.

Aue:2008:TCP

[AHHK08] Alexander Aue, Lajos Horváth, Marie Hušková, and Piotr Kokoszka. Testing for changes in polynomial regression. Bernoulli, 14(3):637–660, August 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1219669624.

Alaya:2015:ISS

[AHK15] Mohamed Ben Alaya, Kaouther Hajji, and Ahmed Kebaier. Importance sampling and statistical Romberg method. Bernoulli, 21(4):1947–1983, November 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1438777583.

Ankirchner:2015:FIB

[AHS15] Stefan Ankirchner, David Hobson, and Philipp Strack. Finite, integrable and bounded time embeddings for diffusions. Bernoulli, 21(2):1067–1088, May 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1429624971.

Aspandiiarov:1999:ABS

[AI99] Sanjar Aspandiiarov and Roudolf Iasnogorodski. Asymptotic behaviour of stationary distributions for countable Markov chains, with some applications. Bernoulli, 5(3):535–569, June 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1172617202.

Anaya-Izquierdo:2007:LMM

[AIM07] Karim Anaya-Izquierdo and Paul Marriott. Local mixture models of exponential families. Bernoulli, 13(3):623–640, August 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1186503479.

Asmussen:2019:LGB

[AIS19] Søren Asmussen, Jevgenijs Ivanovs, and Johan Segers. On the longest gap between power-rate arrivals. Bernoulli, 25(1):375– 394, February 2019. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1544605250.

Albrecher:2016:EIL

[AIZ16] Hansjörg Albrecher, Jevgenijs Ivanovs, and Xiaowen Zhou. Exit identities for Lévy processes observed at Poisson arrival times. Bernoulli, 22(3):1364–1382, August 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1458132985.

Andrieu:2011:NMC

[AJDD11] Christophe Andrieu, Ajay Jasra, Arnaud Doucet, and Pierre Del Moral. On nonlinear Markov chain Monte Carlo. Bernoulli, 17(3):987–1014, August 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1310042853.

Ay:2018:PMM

[AJLS18] Nihat Ay, Jürgen Jost, Hông Vân Lê, and Lorenz Schwachhöfer. Parametrized measure models. Bernoulli, 24(3):1692–1725, August 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1517540458.

Amari:1997:IGE

[AK97] Shun-Ichi Amari and Motoaki Kawanabe. Information geometry of estimating functions in semi-parametric statistical models. Bernoulli, 3(1):29–54, February 1997. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1178291931.

Arutyunyan:2018:DPT

[AK18] Lavrentin M. Arutyunyan and Egor D. Kosov. Deviation of polynomials from their expectations and isoperimetry. Bernoulli, 24(3):2043–2063, August 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1517540467.

Alt:2022:LEL

[AK22] Johannes Alt and Torben Krüger. Local elliptic law. Bernoulli, 28(2):886–909, May 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-28/issue-
2/Local-elliptic-law/10.3150/21-BEJ1370.full.

Alexeev:2023:SRC

[AK23] Ivan Alexeev and Alexey Khartov. Spectral representations of characteristic functions of discrete probability laws. Bernoulli, 29(2):1392–1409, May 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-2/Spectral-representations-of-characteristic-
functions-of-discrete-probability-laws/10.3150/22-BEJ1503.
full.

Andersson:2017:USS

[AKH17] Patrik Andersson and Arturo Kohatsu-Higa. Unbiased simulation of stochastic differential equations using parametrix expansions. Bernoulli, 23(3):2028–2057, August 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1489737632.

Akritas:2000:CLT

[Akr00] Michael G. Akritas. The central limit theorem under censoring. Bernoulli, 6(6):1109–1120, December 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1081194162.

Allassonniere:2010:CBD

[AKT10] Stéphanie Allassonnière, Estelle Kuhn, and Alain Trouvé. Construction of Bayesian deformable models via a stochastic approximation algorithm: A convergence study. Bernoulli, 16 (3):641–678, August 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1281099879.

Adell:2005:SES

[AL05] José Antonio Adell and Alberto Lekuona. Sharp estimates in signed Poisson approximation of Poisson mixtures. Bernoulli, 11(1):47–65, January 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1110228242.

Aldous:1999:DSM

[Ald99] David J. Aldous. Deterministic and stochastic models for coalescence (aggregation and coagulation): a review of the meanfield theory for probabilists. Bernoulli, 5(1):3–48, February 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1173707093.

Aldous:2013:IPS

[Ald13] David Aldous. Interacting particle systems as stochastic social dynamics. Bernoulli, 19(4):1122–1149, September 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1377612846.

Aurzada:2009:SDS

[ALL09] Frank Aurzada, Mikhail Lifshits, and Werner Linde. Small deviations of stable processes and entropy of the associated random operators. Bernoulli, 15(4):1305–1334, November 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1262962237.

Alm:2003:MDB

[Alm03] Sven Erick Alm. Monotonicity of the difference between median and mean of gamma distributions and of a related Ramanujan sequence. Bernoulli, 9(2):351–371, April 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1068128981.

Anh:2019:RCN

[ALOV19] Vo Anh, Nikolai Leonenko, Andriy Olenko, and Volodymyr Vaskovych. On rate of convergence in non-central limit theorems. Bernoulli, 25(4A):2920–2948, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1568362047.

Ascolani:2023:SDC

[ALR23] Filippo Ascolani, Antonio Lijoi, and Matteo Ruggiero. Smoothing distributions for conditional Fleming-Viot and Dawson-Watanabe diffusions. Bernoulli, 29(2):1410–1434, May 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-29/issue-2/Smoothing-distributions-
for-conditional-FlemingViot-and-DawsonWatanabe-diffusions/
10.3150/22-BEJ1504.full.

Anh:2009:MSP

[ALS09] Vo V. Anh, Nikolai N. Leonenko, and Narn-Rueih Shieh. Multifractal scaling of products of birth–death processes. Bernoulli, 15(2):508–531, May 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1241444900.

Altok:2013:UMT

[Alt13] Serdar Altok. Unimodularity for multi-type Galton–Watson trees. Bernoulli, 19(3):780–802, August 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1372251143.

Akashi:2015:ELA

[ALT15] Fumiya Akashi, Yan Liu, and Masanobu Taniguchi. An empirical likelihood approach for symmetric α-stable processes. Bernoulli, 21(4):2093–2119, November 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1438777587.

Altmeyer:2021:AOT

[Alt21] Randolf Altmeyer. Approximation of occupation time functionals. Bernoulli, 27(4):2714–2739, November 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-4/Approximation-of-occupation-time-functionals/
10.3150/21-BEJ1328.full.

Andrieu:2018:UEI

[ALV18] Christophe Andrieu, Anthony Lee, and Matti Vihola. Uniform ergodicity of the iterated conditional SMC and geometric ergodicity of particle Gibbs samplers. Bernoulli, 24(2):842–872, May 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1505980880.

Azais:2011:RFG

[ALW11] Jean-Marc Azaïs, José R. León, and Mario Wschebor. Rice formulae and Gaussian waves. Bernoulli, 17(1):170–193, February 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1297173838.

Aoyama:2007:CST

[AM07] Takahiro Aoyama and Makoto Maejima. Characterizations of subclasses of type G distributions on R^{d} by stochastic integral representations. Bernoulli, 13(1):148–160, February 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1175287725.

Arnaudon:2016:SAF

[AM16] Marc Arnaudon and Laurent Miclo. A stochastic algorithm finding p-means on the circle. Bernoulli, 22(4):2237–2300, November 2016. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1462297681.

Avella-Medina:2017:IFP

[AM17] Marco Avella-Medina. Influence functions for penalized Mestimators. Bernoulli, 23(4B):3178–3196, November 2017. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1495505089.

Aman:2013:NSB

[Ama13] Auguste Aman. A numerical scheme for backward doubly stochastic differential equations. Bernoulli, 19(1):93–114, February 2013. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1358531742.

Andjel:2010:TIO

[AMPV10] Enrique Andjel, Thomas Mountford, Leandro P. R. Pimentel, and Daniel Valesin. Tightness for the interface of the onedimensional contact process. Bernoulli, 16(4):909–925, November 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1290092889.

Andjel:2018:EIG

[AMV18] Enrique Andjel, Thomas Mountford, and Daniel Valesin. Equilibrium of the interface of the grass–bushes–trees process. Bernoulli, 24(3):2256–2277, August 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1517540474.

Adams:2012:UAV

[AN12] Terrence M. Adams and Andrew B. Nobel. Uniform approximation of Vapnik–Chervonenkis classes. Bernoulli, 18(4): 1310–1319, November 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1352727812.

Anonymous:1995:ARW

[Ano95a] Anonymous. Acknowledgement of refereeing work BJ vol. 01. Bernoulli, 1(4):??, December 1995. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1193758714.

Anonymous:1995:AIB

[Ano95b] Anonymous. Annual index BJ vol. 01. Bernoulli, 1(4):??, December 1995. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1193758716.

Anonymous:1995:EBBa

[Ano95c] Anonymous. Editorial board, Bernoulli journal, vol. 1, no. 1–2 (1995). Bernoulli, 1(1–2):??, March 1995. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1186078357.

Anonymous:1995:EBBb

[Ano95d] Anonymous. Editorial board, Bernoulli journal, vol. 1, no. 3 (1995). Bernoulli, 1(3):??, September 1995. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1193667814.

Anonymous:1995:EBBc

[Ano95e] Anonymous. Editorial board, Bernoulli journal, vol. 1, no. 4 (1995). Bernoulli, 1(4):??, December 1995. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1193758707.

Anonymous:1995:KIB

[Ano95f] Anonymous. Keyword index BJ vol. 01. Bernoulli, 1(4):??, December 1995. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1193758715.

Anonymous:1995:TCBa

[Ano95g] Anonymous. Table of contents, Bernoulli journal, vol. 1, no. 1–2 (1995). Bernoulli, 1(1–2):??, March 1995. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1186078356.

Anonymous:1995:TCBb

[Ano95h] Anonymous. Table of contents, Bernoulli journal, vol. 1, no. 3 (1995). Bernoulli, 1(3):??, September 1995. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1193667813.

Anonymous:1995:TCBc

[Ano95i] Anonymous. Table of contents, Bernoulli journal, vol. 1, no. 4 (1995). Bernoulli, 1(4):??, December 1995. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1193758706.

Anonymous:1996:AC

[Ano96a] Anonymous. Amendments and corrections. Bernoulli, 2(1): ??, March 1996. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1193758792.

Anonymous:1996:EBBa

[Ano96b] Anonymous. Editorial board, Bernoulli journal, vol. 2, no. 1 (1996). Bernoulli, 2(1):??, March 1996. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1193758785.

Anonymous:1996:EBBb

[Ano96c] Anonymous. Editorial board, Bernoulli journal, vol. 2, no. 2 (1996). Bernoulli, 2(2):??, June 1996. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1193839218.

Anonymous:1996:TCBa

[Ano96d] Anonymous. Table of contents, Bernoulli journal, vol. 2, no. 1 (1996). Bernoulli, 2(1):??, March 1996. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1193758784.

Anonymous:1996:TCBb

[Ano96e] Anonymous. Table of contents, Bernoulli journal, vol. 2, no. 2 (1996). Bernoulli, 2(2):??, June 1996. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1193839217.

Anonymous:1996:TCBc

[Ano96f] Anonymous. Table of contents, Bernoulli journal, vol. 2, no. 3 (1996). Bernoulli, 2(3):??, September 1996. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1178291718.

Anonymous:1996:TCBd

[Ano96g] Anonymous. Table of contents, Bernoulli journal, vol. 2, no. 4 (1996). Bernoulli, 2(4):??, December 1996. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1178291832.

Anonymous:1997:ARW

[Ano97a] Anonymous. Acknowledgement of refereeing work BJ vol. 03. Bernoulli, 3(4):??, December 1997. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1175882221.

Anonymous:1997:AIB

[Ano97b] Anonymous. Annual index BJ vol. 03. Bernoulli, 3(4):??, December 1997. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1175882222.

Anonymous:1997:BVG

[Ano97c] Anonymous. Boris Vladimirovitch Gnedenko (1912–1995). Bernoulli, 3(1):??, February 1997. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1178291935.

Anonymous:1997:EBBa

[Ano97d] Anonymous. Editorial board, Bernoulli journal, vol. 3, no. 3 (1997). Bernoulli, 3(3):??, September 1997. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1177334454.

Anonymous:1997:EBBb

[Ano97e] Anonymous. Editorial board, Bernoulli journal, vol. 3, no. 4 (1997). Bernoulli, 3(4):??, December 1997. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1175882212.

Anonymous:1997:KIB

[Ano97f] Anonymous. Keyword index BJ vol. 03. Bernoulli, 3(4):??, December 1997. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1175882223.

Anonymous:1997:TCBa

[Ano97g] Anonymous. Table of contents, Bernoulli journal, vol. 3, no. 1 (1997). Bernoulli, 3(1):??, February 1997. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1178291929.

Anonymous:1997:TCBb

[Ano97h] Anonymous. Table of contents, Bernoulli journal, vol. 3, no. 2 (1997). Bernoulli, 3(2):??, June 1997. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1180728929.

Anonymous:1997:TCBc

[Ano97i] Anonymous. Table of contents, Bernoulli journal, vol. 3, no. 3 (1997). Bernoulli, 3(3):??, September 1997. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1177334453.

Anonymous:1997:TCBd

[Ano97j] Anonymous. Table of contents, Bernoulli journal, vol. 3, no. 4 (1997). Bernoulli, 3(4):??, December 1997. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1175882211.

Anonymous:1998:ARW

[Ano98a] Anonymous. Acknowledgement of refereeing work BJ vol. 04. Bernoulli, 4(4):??, December 1998. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1173883820.

Anonymous:1998:AIB

[Ano98b] Anonymous. Annual index BJ vol. 04. Bernoulli, 4(4):??, December 1998. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1173883821.

Anonymous:1998:EBBa

[Ano98c] Anonymous. Editorial board, Bernoulli journal, vol. 4, no. 1 (1998). Bernoulli, 4(1):??, March 1998. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1175865486.

Anonymous:1998:EBBb

[Ano98d] Anonymous. Editorial board, Bernoulli journal, vol. 4, no. 2 (1998). Bernoulli, 4(2):??, June 1998. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1174937291.

Anonymous:1998:EBBc

[Ano98e] Anonymous. Editorial board, Bernoulli journal, vol. 4, no. 3 (1998). Bernoulli, 4(3):??, September 1998. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1174324981.

Anonymous:1998:EBBd

[Ano98f] Anonymous. Editorial board, Bernoulli journal, vol. 4, no. 4 (1998). Bernoulli, 4(4):??, December 1998. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1173883813.

Anonymous:1998:KIB

[Ano98g] Anonymous. Keyword index BJ vol. 04. Bernoulli, 4(4):??, December 1998. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1173883822.

Anonymous:1998:TCBa

[Ano98h] Anonymous. Table of contents, Bernoulli journal, vol. 4, no. 1 (1998). Bernoulli, 4(1):??, March 1998. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1175865485.

Anonymous:1998:TCBb

[Ano98i] Anonymous. Table of contents, Bernoulli journal, vol. 4, no. 2 (1998). Bernoulli, 4(2):??, June 1998. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1174937290.

Anonymous:1998:TCBc

[Ano98j] Anonymous. Table of contents, Bernoulli journal, vol. 4, no. 3 (1998). Bernoulli, 4(3):??, September 1998. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1174324980.

Anonymous:1998:TCBd

[Ano98k] Anonymous. Table of contents, Bernoulli journal, vol. 4, no. 4 (1998). Bernoulli, 4(4):??, December 1998. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1173883812.

Anonymous:1999:ARW

[Ano99a] Anonymous. Acknowledgement of refereeing work BJ vol. 05. Bernoulli, 5(6):??, December 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1143122306.

Anonymous:1999:AIB

[Ano99b] Anonymous. Annual index BJ vol. 05. Bernoulli, 5(6):??, December 1999. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1143122307.

Anonymous:1999:EBBa

[Ano99c] Anonymous. Editorial board, Bernoulli journal, vol. 5, no. 1 (1999). Bernoulli, 5(1):??, February 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1173707092.

Anonymous:1999:EBBb

[Ano99d] Anonymous. Editorial board, Bernoulli journal, vol. 5, no. 2 (1999). Bernoulli, 5(2):??, April 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1173147901.

Anonymous:1999:EBBc

[Ano99e] Anonymous. Editorial board, Bernoulli journal, vol. 5, no. 3 (1999). Bernoulli, 5(3):??, June 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1172617196.

Anonymous:1999:EBBd

[Ano99f] Anonymous. Editorial board, Bernoulli journal, vol. 5, no. 4 (1999). Bernoulli, 5(4):??, August 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1171899317.

Anonymous:1999:EBBe

[Ano99g] Anonymous. Editorial board, Bernoulli journal, vol. 5, no. 5 (1999). Bernoulli, 5(5):??, October 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1171290397.

Anonymous:1999:EBBf

[Ano99h] Anonymous. Editorial board, Bernoulli journal, vol. 5, no. 6 (1999). Bernoulli, 5(6):??, December 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1143122296.

Anonymous:1999:TCBa

[Ano99i] Anonymous. Table of contents, Bernoulli journal, vol. 5, no. 1 (1999). Bernoulli, 5(1):??, February 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1173707091.

Anonymous:1999:TCBb

[Ano99j] Anonymous. Table of contents, Bernoulli journal, vol. 5, no. 2 (1999). Bernoulli, 5(2):??, April 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1173147900.

Anonymous:1999:TCBc

[Ano99k] Anonymous. Table of contents, Bernoulli journal, vol. 5, no. 3 (1999). Bernoulli, 5(3):??, June 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1172617195.

Anonymous:1999:TCBd

[Ano99l] Anonymous. Table of contents, Bernoulli journal, vol. 5, no. 4 (1999). Bernoulli, 5(4):??, August 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1171899316.

Anonymous:1999:TCBe

[Ano99m] Anonymous. Table of contents, Bernoulli journal, vol. 5, no. 5 (1999). Bernoulli, 5(5):??, October 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1171290396.

Anonymous:1999:TCBf

[Ano99n] Anonymous. Table of contents, Bernoulli journal, vol. 5, no. 6 (1999). Bernoulli, 5(6):??, December 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1143122295.

Anonymous:2000:ARW

[Ano00a] Anonymous. Acknowledgement of refereeing work. Bernoulli, 6(6):??, December 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1081194165.

Anonymous:2000:AIB

[Ano00b] Anonymous. Annual index Bernoulli journal, vol. 6, (2000). Bernoulli, 6(6):??, December 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1081194164.

Anonymous:2000:EBBa

[Ano00c] Anonymous. Editorial board, Bernoulli journal, vol. 6, no. 1 (2000). Bernoulli, 6(1):??, February 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1082665376.

Anonymous:2000:EBBb

[Ano00d] Anonymous. Editorial board, Bernoulli journal, vol. 6, no. 2 (2000). Bernoulli, 6(2):??, April 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1081788025.

Anonymous:2000:EBBc

[Ano00e] Anonymous. Editorial board, Bernoulli journal, vol. 6, no. 3 (2000). Bernoulli, 6(3):??, June 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1081616696.

Anonymous:2000:EBBd

[Ano00f] Anonymous. Editorial board, Bernoulli journal, vol. 6, no. 4 (2000). Bernoulli, 6(4):??, August 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1081449591.

Anonymous:2000:EBBe

[Ano00g] Anonymous. Editorial board, Bernoulli journal, vol. 6, no. 5 (2000). Bernoulli, 6(5):??, October 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1081282687.

Anonymous:2000:EBBf

[Ano00h] Anonymous. Editorial board, Bernoulli journal, vol. 6, no. 6 (2000). Bernoulli, 6(6):??, December 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1081194153.

Anonymous:2000:TCBa

[Ano00i] Anonymous. Table of contents, Bernoulli journal, vol. 6, no. 1 (2000). Bernoulli, 6(1):??, February 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1082665375.

Anonymous:2000:TCBb

[Ano00j] Anonymous. Table of contents, Bernoulli journal, vol. 6, no. 2 (2000). Bernoulli, 6(2):??, April 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1081788024.

Anonymous:2000:TCBc

[Ano00k] Anonymous. Table of contents, Bernoulli journal, vol. 6, no. 3 (2000). Bernoulli, 6(3):??, June 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1081616695.

Anonymous:2000:TCBd

[Ano00l] Anonymous. Table of contents, Bernoulli journal, vol. 6, no. 4 (2000). Bernoulli, 6(4):??, August 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1081449590.

Anonymous:2000:TCBe

[Ano00m] Anonymous. Table of contents, Bernoulli journal, vol. 6, no. 5 (2000). Bernoulli, 6(5):??, October 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1081282686.

Anonymous:2000:TCBf

[Ano00n] Anonymous. Table of contents, Bernoulli journal, vol. 6, no. 6 (2000). Bernoulli, 6(6):??, December 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1081194152.

Anonymous:2001:EBBa

[Ano01a] Anonymous. Editorial board, Bernoulli journal, vol. 7, no. 1 (2001). Bernoulli, 7(1):??, February 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1080572336.

Anonymous:2001:EBBb

[Ano01b] Anonymous. Editorial board, Bernoulli journal, vol. 7, no. 2 (2001). Bernoulli, 7(2):??, April 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1080222080.

Anonymous:2001:EBBc

[Ano01c] Anonymous. Editorial board, Bernoulli journal, vol. 7, no. 3 (2001). Bernoulli, 7(3):??, June 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1080004756.

Anonymous:2001:EBBd

[Ano01d] Anonymous. Editorial board, Bernoulli journal, vol. 7, no. 4 (2001). Bernoulli, 7(4):??, August 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1079559462.

Anonymous:2001:EBBe

[Ano01e] Anonymous. Editorial board, Bernoulli journal, vol. 7, no. 5 (2001). Bernoulli, 7(5):??, October 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1079399537.

Anonymous:2001:EBBf

[Ano01f] Anonymous. Editorial board, Bernoulli journal, vol. 7, no. 6 (2001). Bernoulli, 7(6):??, December 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1078951124.

Anonymous:2001:TCBa

[Ano01g] Anonymous. Table of contents, Bernoulli journal, vol. 7, no. 1 (2001). Bernoulli, 7(1):??, February 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1080572335.

Anonymous:2001:TCBb

[Ano01h] Anonymous. Table of contents, Bernoulli journal, vol. 7, no. 2 (2001). Bernoulli, 7(2):??, April 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1080222079.

Anonymous:2001:TCBc

[Ano01i] Anonymous. Table of contents, Bernoulli journal, vol. 7, no. 3 (2001). Bernoulli, 7(3):??, June 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1080004755.

Anonymous:2001:TCBd

[Ano01j] Anonymous. Table of contents, Bernoulli journal, vol. 7, no. 4 (2001). Bernoulli, 7(4):??, August 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1079559461.

Anonymous:2001:TCBe

[Ano01k] Anonymous. Table of contents, Bernoulli journal, vol. 7, no. 5 (2001). Bernoulli, 7(5):??, October 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1079399536.

Anonymous:2001:TCBf

[Ano01l] Anonymous. Table of contents, Bernoulli journal, vol. 7, no. 6 (2001). Bernoulli, 7(6):??, December 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1078951123.

Anonymous:2002:EBBa

[Ano02a] Anonymous. Editorial board, Bernoulli journal, vol. 8, no. 1 (2002). Bernoulli, 8(1):??, February 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1078951086.

Anonymous:2002:EBBb

[Ano02b] Anonymous. Editorial board, Bernoulli journal, vol. 8, no. 2 (2002). Bernoulli, 8(2):??, April 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1078866864.

Anonymous:2002:EBBc

[Ano02c] Anonymous. Editorial board, Bernoulli journal, vol. 8, no. 3 (2002). Bernoulli, 8(3):??, April 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1078779870.

Anonymous:2002:EBBd

[Ano02d] Anonymous. Editorial board, Bernoulli journal, vol. 8, no. 4 (2002). Bernoulli, 8(4):??, August 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1078681377.

Anonymous:2002:EBBe

[Ano02e] Anonymous. Editorial board, Bernoulli journal, vol. 8, no. 5 (2002). Bernoulli, 8(5):??, October 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1078435217.

Anonymous:2002:EBBf

[Ano02f] Anonymous. Editorial board, Bernoulli journal, vol. 8, no. 6 (2002). Bernoulli, 8(6):??, December 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1076364801.

Anonymous:2002:TCBa

[Ano02g] Anonymous. Table of contents, Bernoulli journal, vol. 8, no. 1 (2002). Bernoulli, 8(1):??, February 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1078951085.

Anonymous:2002:TCBb

[Ano02h] Anonymous. Table of contents, Bernoulli journal, vol. 8, no. 2 (2002). Bernoulli, 8(2):??, April 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1078866863.

Anonymous:2002:TCBc

[Ano02i] Anonymous. Table of contents, Bernoulli journal, vol. 8, no. 3 (2002). Bernoulli, 8(3):??, April 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1078779869.

Anonymous:2002:TCBd

[Ano02j] Anonymous. Table of contents, Bernoulli journal, vol. 8, no. 4 (2002). Bernoulli, 8(4):??, August 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1078681376.

Anonymous:2002:TCBe

[Ano02k] Anonymous. Table of contents, Bernoulli journal, vol. 8, no. 5 (2002). Bernoulli, 8(5):??, October 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1078435216.

Anonymous:2002:TCBf

[Ano02l] Anonymous. Table of contents, Bernoulli journal, vol. 8, no. 6 (2002). Bernoulli, 8(6):??, December 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1076364800.

Anonymous:2003:EBBa

[Ano03a] Anonymous. Editorial board, Bernoulli journal, vol. 9, no. 1 (2003). Bernoulli, 9(1):??, February 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1068129007.

Anonymous:2003:EBBb

[Ano03b] Anonymous. Editorial board, Bernoulli journal, vol. 9, no. 2 (2003). Bernoulli, 9(2):??, April 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1068128973.

Anonymous:2003:EBBc

[Ano03c] Anonymous. Editorial board, Bernoulli journal, vol. 9, no. 6 (2003). Bernoulli, 9(6):??, December 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1072215195.

Anonymous:2003:TCBa

[Ano03d] Anonymous. Table of contents, Bernoulli journal, vol. 9, no. 1 (2003). Bernoulli, 9(1):??, February 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1068129006.

Anonymous:2003:TCBb

[Ano03e] Anonymous. Table of contents, Bernoulli journal, vol. 9, no. 2 (2003). Bernoulli, 9(2):??, April 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1068128972.

Anonymous:2003:TCBc

[Ano03f] Anonymous. Table of contents, Bernoulli journal, vol. 9, no. 3 (2003). Bernoulli, 9(3):??, June 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1065444809.

Anonymous:2003:TCBd

[Ano03g] Anonymous. Table of contents, Bernoulli journal, vol. 9, no. 4 (2003). Bernoulli, 9(4):??, August 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1066223268.

Anonymous:2003:TCBe

[Ano03h] Anonymous. Table of contents, Bernoulli journal, vol. 9, no. 5 (2003). Bernoulli, 9(5):??, October 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1066418875.

Anonymous:2003:TCBf

[Ano03i] Anonymous. Table of contents, Bernoulli journal, vol. 9, no. 6 (2003). Bernoulli, 9(6):??, December 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1072215194.

Anonymous:2004:EBBa

[Ano04a] Anonymous. Editorial board, Bernoulli journal, vol. 10, no. 1 (2004). Bernoulli, 10(1):??, February 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1077544600.

Anonymous:2004:EBBb

[Ano04b] Anonymous. Editorial board, Bernoulli journal, vol. 10, no. 2 (2004). Bernoulli, 10(2):??, April 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1082380216.

Anonymous:2004:EBBc

[Ano04c] Anonymous. Editorial board, Bernoulli journal, vol. 10, no. 3 (2004). Bernoulli, 10(3):??, June 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1089206402.

Anonymous:2004:EBBd

[Ano04d] Anonymous. Editorial board, Bernoulli journal, vol. 10, no. 4 (2004). Bernoulli, 10(4):??, August 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1093265629.

Anonymous:2004:EBBe

[Ano04e] Anonymous. Editorial board, Bernoulli journal, vol. 10, no. 5 (2004). Bernoulli, 10(5):??, October 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1099579154.

Anonymous:2004:EBBf

[Ano04f] Anonymous. Editorial board, Bernoulli journal, vol. 10, no. 6 (2004). Bernoulli, 10(6):??, December 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1106314842.

Anonymous:2004:TCBa

[Ano04g] Anonymous. Table of contents, Bernoulli journal, vol. 10, no. 1 (2004). Bernoulli, 10(1):??, February 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1077544599.

Anonymous:2004:TCBb

[Ano04h] Anonymous. Table of contents, Bernoulli journal, vol. 10, no. 2 (2004). Bernoulli, 10(2):??, April 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1082380215.

Anonymous:2004:TCBc

[Ano04i] Anonymous. Table of contents, Bernoulli journal, vol. 10, no. 3 (2004). Bernoulli, 10(3):??, June 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1089206401.

Anonymous:2004:TCBd

[Ano04j] Anonymous. Table of contents, Bernoulli journal, vol. 10, no. 4 (2004). Bernoulli, 10(4):??, August 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1093265628.

Anonymous:2004:TCBe

[Ano04k] Anonymous. Table of contents, Bernoulli journal, vol. 10, no. 5 (2004). Bernoulli, 10(5):??, October 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1099579153.

Anonymous:2004:TCBf

[Ano04l] Anonymous. Table of contents, Bernoulli journal, vol. 10, no. 6 (2004). Bernoulli, 10(6):??, December 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1106314841.

Anonymous:2005:ARW

[Ano05a] Anonymous. Acknowledgement of refereeing work BJ vol. 11. Bernoulli, 11(6):??, December 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1137421643.

Anonymous:2005:AIB

[Ano05b] Anonymous. Annual index BJ vol. 11. Bernoulli, 11(6):??, December 2005. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1137421644.

Anonymous:2005:EBBa

[Ano05c] Anonymous. Editorial board, Bernoulli journal, vol. 11, no. 2 (2005). Bernoulli, 11(2):??, April 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1116340290.

Anonymous:2005:EBBb

[Ano05d] Anonymous. Editorial board, Bernoulli journal, vol. 11, no. 3 (2005). Bernoulli, 11(3):??, June 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1120591181.

Anonymous:2005:EBBc

[Ano05e] Anonymous. Editorial board, Bernoulli journal, vol. 11, no. 4 (2005). Bernoulli, 11(4):??, August 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1126126760.

Anonymous:2005:EBBd

[Ano05f] Anonymous. Editorial board, Bernoulli journal, vol. 11, no. 5 (2005). Bernoulli, 11(5):??, October 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1130077592.

Anonymous:2005:EBBe

[Ano05g] Anonymous. Editorial board, Bernoulli journal, vol. 11, no. 6 (2005). Bernoulli, 11(6):??, December 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1137421634.

Anonymous:2005:TCBa

[Ano05h] Anonymous. Table of contents, Bernoulli journal, vol. 11, no. 1 (2005). Bernoulli, 11(1):??, January 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1110228238.

Anonymous:2005:TCBb

[Ano05i] Anonymous. Table of contents, Bernoulli journal, vol. 11, no. 2 (2005). Bernoulli, 11(2):??, April 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1116340289.

Anonymous:2005:TCBc

[Ano05j] Anonymous. Table of contents, Bernoulli journal, vol. 11, no. 3 (2005). Bernoulli, 11(3):??, June 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1120591180.

Anonymous:2005:TCBd

[Ano05k] Anonymous. Table of contents, Bernoulli journal, vol. 11, no. 4 (2005). Bernoulli, 11(4):??, August 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1126126759.

Anonymous:2005:TCBe

[Ano05l] Anonymous. Table of contents, Bernoulli journal, vol. 11, no. 5 (2005). Bernoulli, 11(5):??, October 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1130077591.

Anonymous:2005:TCBf

[Ano05m] Anonymous. Table of contents, Bernoulli journal, vol. 11, no. 6 (2005). Bernoulli, 11(6):??, December 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1137421633.

Anonymous:2006:ARW

[Ano06a] Anonymous. Acknowledgement of refereeing work BJ vol. 12. Bernoulli, 12(6):??, December 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1165269153.

Anonymous:2006:AIB

[Ano06b] Anonymous. Annual index BJ vol. 12. Bernoulli, 12(6):??, December 2006. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1165269154.

Anonymous:2006:EBBa

[Ano06c] Anonymous. Editorial board, Bernoulli journal, vol. 12, no. 1 (2006). Bernoulli, 12(1):??, February 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1141136645.

Anonymous:2006:EBBb

[Ano06d] Anonymous. Editorial board, Bernoulli journal, vol. 12, no. 2 (2006). Bernoulli, 12(2):??, April 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1145993971.

Anonymous:2006:EBBc

[Ano06e] Anonymous. Editorial board, Bernoulli journal, vol. 12, no. 3 (2006). Bernoulli, 12(3):??, June 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1151525126.

Anonymous:2006:EBBd

[Ano06f] Anonymous. Editorial board, Bernoulli journal, vol. 12, no. 4 (2006). Bernoulli, 12(4):??, August 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1155735926.

Anonymous:2006:EBBe

[Ano06g] Anonymous. Editorial board, Bernoulli journal, vol. 12, no. 5 (2006). Bernoulli, 12(5):??, October 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1161614944.

Anonymous:2006:EBBf

[Ano06h] Anonymous. Editorial board, Bernoulli journal, vol. 12, no. 6 (2006). Bernoulli, 12(6):??, December 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1165269144.

Anonymous:2006:TCBa

[Ano06i] Anonymous. Table of contents, Bernoulli journal, vol. 12, no. 1 (2006). Bernoulli, 12(1):??, February 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1141136644.

Anonymous:2006:TCBb

[Ano06j] Anonymous. Table of contents, Bernoulli journal, vol. 12, no. 2 (2006). Bernoulli, 12(2):??, April 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1145993970.

Anonymous:2006:TCBc

[Ano06k] Anonymous. Table of contents, Bernoulli journal, vol. 12, no. 3 (2006). Bernoulli, 12(3):??, June 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1151525125.

Anonymous:2006:TCBd

[Ano06l] Anonymous. Table of contents, Bernoulli journal, vol. 12, no. 4 (2006). Bernoulli, 12(4):??, August 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1155735925.

Anonymous:2006:TCBe

[Ano06m] Anonymous. Table of contents, Bernoulli journal, vol. 12, no. 5 (2006). Bernoulli, 12(5):??, October 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1161614943.

Anonymous:2006:TCBf

[Ano06n] Anonymous. Table of contents, Bernoulli journal, vol. 12, no. 6 (2006). Bernoulli, 12(6):??, December 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1165269143.

Anonymous:2007:A

[Ano07a] Anonymous. Acknowledgements. Bernoulli, 13(4):1195–1197, November 2007. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1194625608.

Anonymous:2007:AI

[Ano07b] Anonymous. Author index. Bernoulli, 13(4):1199–1204, November 2007. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1194625609.

Anonymous:2007:EBB

[Ano07c] Anonymous. Editorial board, Bernoulli journal, vol. 13, no. 1 (2007). Bernoulli, 13(1):??, February 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1175287716.

Anonymous:2007:ESa

[Ano07d] Anonymous. Editorial staff. Bernoulli, 13(2):??, May 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1179498749.

Anonymous:2007:ESb

[Ano07e] Anonymous. Editorial staff. Bernoulli, 13(3):??, August 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1186503477.

Anonymous:2007:ESc

[Ano07f] Anonymous. Editorial staff. Bernoulli, 13(4):??, November 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1194625593.

Anonymous:2007:TCa

[Ano07g] Anonymous. Table of contents. Bernoulli, 13(2):??, May 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1179498748.

Anonymous:2007:TCc

(electronic). URL http://projecteuclid.org/euclid.bj/
1186503476.

Anonymous:2007:TCd

[Ano07i] Anonymous. Table of contents. Bernoulli, 13(4):??, November 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1194625592.

Anonymous:2007:TCB

[Ano07j] Anonymous. Table of contents, Bernoulli journal, vol. 13, no. 1 (2007). Bernoulli, 13(1):??, February 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1175287715.

Anonymous:2008:A

[Ano08a] Anonymous. Acknowledgements. Bernoulli, 14(4):1191–1192, November 2008. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1225980576.

Anonymous:2008:AI

[Ano08b] Anonymous. Author index. Bernoulli, 14(4):1193–1197, November 2008. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1225980577.

Anonymous:2008:ESa

[Ano08c] Anonymous. Editorial staff. Bernoulli, 14(1):??, February 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1202492781.

Anonymous:2008:ESb

[Ano08d] Anonymous. Editorial staff. Bernoulli, 14(2):??, May 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1208872105.

Anonymous:2008:ESc

[Ano08e] Anonymous. Editorial staff. Bernoulli, 14(3):??, August 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1219669620.

Anonymous:2008:ESd

[Ano08f] Anonymous. Editorial staff. Bernoulli, 14(4):??, November 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1225980563.

Anonymous:2008:TCa

[Ano08g] Anonymous. Table of contents. Bernoulli, 14(1):??, February 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1202492795.

Anonymous:2008:TCb

[Ano08h] Anonymous. Table of contents. Bernoulli, 14(2):??, May 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1208872120.

Anonymous:2008:TCc

[Ano08i] Anonymous. Table of contents. Bernoulli, 14(3):??, August 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1219669634.

Anonymous:2008:TCd

[Ano08j] Anonymous. Table of contents. Bernoulli, 14(4):??, November 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1225980578.

Anonymous:2009:A

[Ano09a] Anonymous. Acknowledgements. Bernoulli, 15(4):1397–1398, November 2009. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1262962241.

Anonymous:2009:AI

[Ano09b] Anonymous. Author index. Bernoulli, 15(4):1399–1403, November 2009. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1262962242.

Anonymous:2009:EBa

[Ano09c] Anonymous. Editorial board. Bernoulli, 15(1):??, February 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1233669880.

Anonymous:2009:EBb

[Ano09d] Anonymous. Editorial board. Bernoulli, 15(2):??, May 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1241444891.

Anonymous:2009:EBc

[Ano09e] Anonymous. Editorial board. Bernoulli, 15(3):??, August 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1251463272.

Anonymous:2009:EBd

[Ano09f] Anonymous. Editorial board. Bernoulli, 15(4):??, November 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1262962220.

Anonymous:2009:TCa

[Ano09g] Anonymous. Table of contents. Bernoulli, 15(1):??, February 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1233669893.

Anonymous:2009:TCb

[Ano09h] Anonymous. Table of contents. Bernoulli, 15(2):??, May 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1241444904.

Anonymous:2009:TCc

[Ano09i] Anonymous. Table of contents. Bernoulli, 15(3):??, August 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1251463288.

Anonymous:2009:TCd

[Ano09j] Anonymous. Table of contents. Bernoulli, 15(4):??, November 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1262962243.

Anonymous:2010:A

[Ano10a] Anonymous. Acknowledgements. Bernoulli, 16(4):1415–1416, November 2010. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1290092912.

Anonymous:2010:AI

[Ano10b] Anonymous. Author index. Bernoulli, 16(4):1417–1422, November 2010. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1290092913.

Anonymous:2010:EBa

[Ano10c] Anonymous. Editorial board. Bernoulli, 16(1):??, February 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1265984701.

Anonymous:2010:EBb

[Ano10d] Anonymous. Editorial board. Bernoulli, 16(2):??, May 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1274821071.

Anonymous:2010:EBc

[Ano10e] Anonymous. Editorial board. Bernoulli, 16(3):??, August 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1281099876.

Anonymous:2010:EBd

[Ano10f] Anonymous. Editorial board. Bernoulli, 16(4):??, November 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1290092888.

Anonymous:2010:TCa

[Ano10g] Anonymous. Table of contents. Bernoulli, 16(1):??, February 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1265984713.

Anonymous:2010:TCb

[Ano10h] Anonymous. Table of contents. Bernoulli, 16(2):??, May 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1274821086.

Anonymous:2010:TCc

[Ano10i] Anonymous. Table of contents. Bernoulli, 16(3):??, August 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1281099889.

Anonymous:2010:TCd

[Ano10j] Anonymous. Table of contents. Bernoulli, 16(4):??, November 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1290092914.

Anonymous:2011:A

[Ano11a] Anonymous. Acknowledgements. Bernoulli, 17(4):1435–1437, November 2011. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1320417512.

Anonymous:2011:AI

[Ano11b] Anonymous. Author index. Bernoulli, 17(4):1438–1442, November 2011. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1320417513.

Anonymous:2011:EBa

[Ano11c] Anonymous. Editorial board. Bernoulli, 17(1):??, February 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1297173830.

Anonymous:2011:EBb

[Ano11d] Anonymous. Editorial board. Bernoulli, 17(2):??, May 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1302009233.

Anonymous:2011:EBc

[Ano11e] Anonymous. Editorial board. Bernoulli, 17(3):??, August 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1310042845.

Anonymous:2011:EBd

[Ano11f] Anonymous. Editorial board. Bernoulli, 17(4):??, November 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1320417497.

Anonymous:2011:TCa

[Ano11g] Anonymous. Table of contents. Bernoulli, 17(1):??, February 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1297173853.

Anonymous:2011:TCb

[Ano11h] Anonymous. Table of contents. Bernoulli, 17(2):??, May 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1302009249.

Anonymous:2011:TCc

[Ano11i] Anonymous. Table of contents. Bernoulli, 17(3):??, August 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1310042859.

Anonymous:2011:TCd

[Ano11j] Anonymous. Table of contents. Bernoulli, 17(4):??, November 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1320417514.

Anonymous:2012:A

[Ano12a] Anonymous. Acknowledgements. Bernoulli, 18(4):1465–1466, November 2012. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1352727820.

Anonymous:2012:AI

[Ano12b] Anonymous. Author index. Bernoulli, 18(4):1467–1471, November 2012. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1352727821.

Anonymous:2012:EBa

[Ano12c] Anonymous. Editorial board. Bernoulli, 18(1):??, February 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1327068614.

Anonymous:2012:EBb

[Ano12d] Anonymous. Editorial board. Bernoulli, 18(2):??, May 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1334580717.

Anonymous:2012:EBc

[Ano12e] Anonymous. Editorial board. Bernoulli, 18(3):??, August 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1340886999.

Anonymous:2012:EBd

[Ano12f] Anonymous. Editorial board. Bernoulli, 18(4):??, November 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1352727802.

Anonymous:2012:TCa

[Ano12g] Anonymous. Table of contents. Bernoulli, 18(1):??, February 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1327068630.

Anonymous:2012:TCb

[Ano12h] Anonymous. Table of contents. Bernoulli, 18(2):??, May 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1334580733.

Anonymous:2012:TCc

[Ano12i] Anonymous. Table of contents. Bernoulli, 18(3):??, August 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1340887015.

Anonymous:2012:TCd

[Ano12j] Anonymous. Table of contents. Bernoulli, 18(4):??, November 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1352727822.

Anonymous:2013:A

[Ano13a] Anonymous. Acknowledgements. Bernoulli, 19(5B):2769–2771, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1386078621.

Anonymous:2013:AI

[Ano13b] Anonymous. Author index. Bernoulli, 19(5B):2772–2779, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1386078622.

Anonymous:2013:EBa

[Ano13c] Anonymous. Editorial board. Bernoulli, 19(1):??, February 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1358531737.

Anonymous:2013:EBb

[Ano13d] Anonymous. Editorial board. Bernoulli, 19(2):??, May 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1363192030.

Anonymous:2013:EBc

[Ano13e] Anonymous. Editorial board. Bernoulli, 19(3):??, August 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1372251140.

Anonymous:2013:EBd

[Ano13f] Anonymous. Editorial board. Bernoulli, 19(4):??, September 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1377612843.

Anonymous:2013:EBe

[Ano13g] Anonymous. Editorial board. Bernoulli, 19(5A):??, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1383661191.

Anonymous:2013:EBf

[Ano13h] Anonymous. Editorial board. Bernoulli, 19(5B):??, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1386078597.

Anonymous:2013:TCa

[Ano13i] Anonymous. Table of contents. Bernoulli, 19(1):??, February 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1358531754.

Anonymous:2013:TCb

[Ano13j] Anonymous. Table of contents. Bernoulli, 19(2):??, May 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1363192044.

Anonymous:2013:TCc

[Ano13k] Anonymous. Table of contents. Bernoulli, 19(3):??, August 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1372251154.

Anonymous:2013:TCd

[Ano13l] Anonymous. Table of contents. Bernoulli, 19(4):??, September 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1377612863.

Anonymous:2013:TCe

[Ano13m] Anonymous. Table of contents. Bernoulli, 19(5A):??, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1383661217.

Anonymous:2013:TCf

[Ano13n] Anonymous. Table of contents. Bernoulli, 19(5B):??, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1386078623.

Anonymous:2014:AI

[Ano14a] Anonymous. Author index. Bernoulli, 20(4):??, November 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1411134462.

Anonymous:2014:EBa

[Ano14b] Anonymous. Editorial board. Bernoulli, 20(1):??, February 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1390407278.

Anonymous:2014:EBb

[Ano14c] Anonymous. Editorial board. Bernoulli, 20(2):??, May 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1393593991.

Anonymous:2014:EBc

[Ano14d] Anonymous. Editorial board. Bernoulli, 20(3):??, August 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1402488932.

Anonymous:2014:EBd

[Ano14e] Anonymous. Editorial board. Bernoulli, 20(4):??, November 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1411134439.

Anonymous:2014:TCa

[Ano14f] Anonymous. Table of contents. Bernoulli, 20(1):??, February 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1390407294.

Anonymous:2014:TCb

[Ano14g] Anonymous. Table of contents. Bernoulli, 20(2):??, May 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1393594015.

Anonymous:2014:TCc

(electronic). URL http://projecteuclid.org/euclid.bj/
1402488954.

Anonymous:2014:TCd

[Ano14i] Anonymous. Table of contents. Bernoulli, 20(4):??, November 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1411134463.

Anonymous:2015:EBa

[Ano15a] Anonymous. Editorial board. Bernoulli, 21(1):??, February 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1426597061.

Anonymous:2015:EBb

[Ano15b] Anonymous. Editorial board. Bernoulli, 21(2):??, May 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1429624956.

Anonymous:2015:EBc

[Ano15c] Anonymous. Editorial board. Bernoulli, 21(3):??, August 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1432732019.

Anonymous:2015:EBd

[Ano15d] Anonymous. Editorial board. Bernoulli, 21(4):??, November 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1438777582.

Anonymous:2015:TCa

[Ano15e] Anonymous. Table of contents. Bernoulli, 21(1):??, February 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1426597085.

Anonymous:2015:TCb

[Ano15f] Anonymous. Table of contents. Bernoulli, 21(2):??, May 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1429624979.

Anonymous:2015:TCc

[Ano15g] Anonymous. Table of contents. Bernoulli, 21(3):??, August 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1432732043.

Anonymous:2015:TCd

[Ano15h] Anonymous. Table of contents. Bernoulli, 21(4):??, November 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1438777607.

Anonymous:2016:AI

[Ano16a] Anonymous. Author index. Bernoulli, 22(4):2609–2614, November 2016. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1462367824.

Anonymous:2016:EBa

[Ano16b] Anonymous. Editorial board. Bernoulli, 22(1):??, February 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1450802413.

Anonymous:2016:EBb

[Ano16c] Anonymous. Editorial board. Bernoulli, 22(2):??, May 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1454335222.

Anonymous:2016:EBc

[Ano16d] Anonymous. Editorial board. Bernoulli, 22(3):??, August 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1458132982.

Anonymous:2016:EBd

[Ano16e] Anonymous. Editorial board. Bernoulli, 22(4):??, November 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1462297670.

Anonymous:2016:TCa

[Ano16f] Anonymous. Table of contents. Bernoulli, 22(1):??, February 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1450802414.

Anonymous:2016:TCb

[Ano16g] Anonymous. Table of contents. Bernoulli, 22(2):??, May 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1454335223.

Anonymous:2016:TCc

[Ano16h] Anonymous. Table of contents. Bernoulli, 22(3):??, August 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1458133004.

Anonymous:2016:TCd

[Ano16i] Anonymous. Table of contents. Bernoulli, 22(4):??, November 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1462297691.

Anonymous:2017:EBa

[Ano17a] Anonymous. Editorial board. Bernoulli, 23(1):??, February 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1475001345.

Anonymous:2017:EBb

[Ano17b] Anonymous. Editorial board. Bernoulli, 23(2):??, May 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1486177383.

Anonymous:2017:TCa

[Ano17c] Anonymous. Table of contents. Bernoulli, 23(1):??, February 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1475001375.

Anonymous:2017:TC

[Ano17d] Anonymous. Table of contents. Bernoulli, 23(2):??, May 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1486177404.

Anonymous:2020:TC

[Ano20] Anonymous. Table of contents. Bernoulli, 26(4):??, November 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-26/issue-4/Table-of-Contents/bj/1598493648.
full.

Anonymous:2021:TCa

[Ano21a] Anonymous. Table of contents. Bernoulli, 27(1):??, February 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-27/issue-1/Table-of-Contents/bj/1605841262.
full.

Anonymous:2021:TCb

[Ano21b] Anonymous. Table of contents. Bernoulli, 27(2):??, May 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-27/issue-2/Table-of-Contents/bej272c4.
full.

Anonymous:2021:TCc

[Ano21c] Anonymous. Table of contents. Bernoulli, 27(3):??, August 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-27/issue-3/Table-of-Contents/bej273c4.
full.

Anonymous:2021:TCd

[Ano21d] Anonymous. Table of contents. Bernoulli, 27(4):??, November 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-27/issue-4/Table-of-Contents/10.3150/
bej274c4.full.

Anonymous:2022:TCa

[Ano22a] Anonymous. Table of contents. Bernoulli, 28(1):??, February 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-28/issue-1/Table-of-Contents/bej281c4.
full.

Anonymous:2022:TCb

[Ano22b] Anonymous. Table of contents. Bernoulli, 28(2):??, May 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-28/issue-2/Table-of-Contents/bej282c4.
full.

Anonymous:2022:TCc

[Ano22c] Anonymous. Table of contents. Bernoulli, 28(3):??, August 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-28/issue-3/Table-of-Contents/bej283c4.
full.

Anonymous:2022:TCd

[Ano22d] Anonymous. Table of contents. Bernoulli, 28(4):??, November 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-28/issue-4/Table-of-Contents/bej284c4.
full.

Anonymous:2023:EBa

[Ano23a] Anonymous. Editorial board. Bernoulli, 29(3):??, August 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-29/issue-3/Editorial-
Board/bej293p1.full.

Anonymous:2023:EBb

[Ano23b] Anonymous. Editorial board. Bernoulli, 29(4):??, November 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-29/issue-4/Editorial-
Board/bej294p1.full.

Anonymous:2023:TCa

[Ano23c] Anonymous. Table of contents. Bernoulli, 29(1):??, February 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-29/issue-1/Table-of-Contents/bej291c4.
full.

Anonymous:2023:TCb

[Ano23d] Anonymous. Table of contents. Bernoulli, 29(2):??, May 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-29/issue-2/Table-of-Contents/bej292c4.
full.

Anonymous:2023:TCc

[Ano23e] Anonymous. Table of contents. Bernoulli, 29(3):??, August 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-29/issue-3/Table-of-Contents/bej293c4.
full.

Anonymous:2023:TCd

[Ano23f] Anonymous. Table of contents. Bernoulli, 29(4):??, November 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-29/issue-4/Table-of-Contents/bej294c4.
full.

Anonymous:2024:EBa

[Ano24a] Anonymous. Editorial board. Bernoulli, 30(1):??, February 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-30/issue-1/Editorial-
Board/bej301p1.full.

Anonymous:2024:EBb

[Ano24b] Anonymous. Editorial board. Bernoulli, 30(2):??, May 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-30/issue-2/Editorial-Board/bej302p1.
full.

Anonymous:2024:EBc

[Ano24c] Anonymous. Editorial board. Bernoulli, 30(3):??, August 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-30/issue-3/Editorial-
Board/bej303p1.full.

Anonymous:2024:EBd

[Ano24d] Anonymous. Editorial board. Bernoulli, 30(4):??, November 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

Anonymous:2024:TCa

[Ano24e] Anonymous. Table of contents. Bernoulli, 30(2):??, May 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-30/issue-2/Table-of-Contents/bej302c4.
full.

Anonymous:2024:TCb

[Ano24f] Anonymous. Table of contents. Bernoulli, 30(4):??, November 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

Anisimov:2000:TLS

[AP00] Vladimir V. Anisimov and Georg Ch. Pflug. Z-theorems: limits of stochastic equations. Bernoulli, 6(5):917–938, October 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1081282695.

Antunes:2011:PSF

[AP11] Nelson Antunes and Vladas Pipiras. Probabilistic sampling of finite renewal processes. Bernoulli, 17(4):1285–1326, November 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1320417505.

Afendras:2014:SCT

[AP14] G. Afendras and N. Papadatos. Strengthened Chernoff-type variance bounds. Bernoulli, 20(1):245–264, February 2014. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1390407288.

Ahn:2019:QCL

[AP19] Sung Won Ahn and Jonathon Peterson. Quenched central limit theorem rates of convergence for one-dimensional random walks in random environments. Bernoulli, 25(2):1386–1411, May 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1551862854.

Aneiros-Perez:2004:ETP

[APGMV04] Germán Aneiros-Pérez, Wenceslao González-Manteiga, and Philippe Vieu. Estimation and testing in a partial linear regression model under long-memory dependence. Bernoulli, 10 (1):49–78, February 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1077544603.

Afendras:2011:EST

[APP11] G. Afendras, N. Papadatos, and V. Papathanasiou. An extended Stein-type covariance identity for the Pearson family with applications to lower variance bounds. Bernoulli, 17(2):507–529, May 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1302009234.

Atchade:2005:AMC

[AR05] Yves F. Atchadé and Jeffrey S. Rosenthal. On adaptive Markov chain Monte Carlo algorithms. Bernoulli, 11(5):815–828, October 2005. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1130077595.

Anastasiou:2017:BNA

[AR17] Andreas Anastasiou and Gesine Reinert. Bounds for the normal approximation of the maximum likelihood estimator. Bernoulli, 23(1):191–218, February 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1475001353.

Alberts:2020:BFS

[ARAS20] Tom Alberts, Firas Rassoul-Agha, and Mackenzie Simper. Busemann functions and semi-infinite O’Connell–Yor polymers. Bernoulli, 26(3):1927–1955, August 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1587974529.

Armstrong:2021:ABC

[Arm21] Timothy B. Armstrong. Adaptation bounds for confidence bands under self-similarity. Bernoulli, 27(2):1348– 1370, May 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-27/issue-2/Adaptation-
bounds-for-confidence-bands-under-self-similarity/10.
3150/20-BEJ1277.full.

Agapiou:2018:UMC

[ARV18] Sergios Agapiou, Gareth O. Roberts, and Sebastian J. Vollmer. Unbiased Monte Carlo: Posterior estimation for intractable/ infinite-dimensional models. Bernoulli, 24(3):1726–1786, August 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1517540459.

Arous:2003:LDL

[AS03] Gérard Ben Arous and Michel Sortais. Large deviations in the Langevin dynamics of a short-range spin glass. Bernoulli, 9 (6):921–954, December 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1072215196.

Adamczak:2019:CPI

[AS19] Radoslaw Adamczak and Michal Strzelecki. On the convex Poincaré inequality and weak transportation inequalities. Bernoulli, 25(1):341–374, February 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1544605249.

Assaf:2002:PIO

[ASC02] David Assaf and Ester Samuel-Cahn. Prophet inequalities for optimal stopping rules with probabilistic recall. Bernoulli, 8 (1):39–52, February 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1078951088.

Assiotis:2023:SVC

[Ass23] Theodoros Assiotis. On the singular values of complex matrix Brownian motion with a matrix drift. Bernoulli, 29(2): 1686–1713, May 2023. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-29/issue-2/On-the-singular-
values-of-complex-matrix-Brownian-motion-with/10.3150/
22-BEJ1517.full.

Ayache:2024:ANM

[AT24] Antoine Ayache and Ciprian A. Tudor. Asymptotic normality for a modified quadratic variation of the Hermite process. Bernoulli, 30(2):1154–1176, May 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
30/issue-2/Asymptotic-normality-for-a-modified-quadratic-
variation-of-the-Hermite/10.3150/23-BEJ1627.full.

Atchade:2016:MCM

[Atc16] Yves F. Atchadé. Markov chain Monte Carlo confidence intervals. Bernoulli, 22(3):1808–1838, August 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1458133000.

Athreya:2000:CMM

[Ath00] Krishna B. Athreya. Change of measures for Markov chains and the LlogL theorem for branching processes. Bernoulli, 6(2):323– 338, April 2000. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1081788031.

Andrieu:2014:MSA

[AV14] Christophe Andrieu and Matti Vihola. Markovian stochastic approximation with expanding projections. Bernoulli, 20 (2):545–585, May 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1393593997.

Azais:1996:ASO

[AW96] Jean-Marc Azaïs and Mario Wschebor. Almost sure oscillation of certain random processes. Bernoulli, 2(3):257–270, September 1996. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1178291722.

Armentano:2009:RSP

[AW09] Diego Armentano and Mario Wschebor. Random systems of polynomial equations. The expected number of roots under smooth analysis. Bernoulli, 15(1):249–266, February 2009. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1233669890.

Alquier:2012:MSW

[AW12] Pierre Alquier and Olivier Wintenberger. Model selection for weakly dependent time series forecasting. Bernoulli, 18(3):883– 913, August 2012. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1340887007.

Agapiou:2024:LPS

[AW24] Sergios Agapiou and Sven Wang. Laplace priors and spatial inhomogeneity in Bayesian inverse problems. Bernoulli, 30(2): 878–910, May 2024. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-30/issue-2/Laplace-priors-
and-spatial-inhomogeneity-in-Bayesian-inverse-problems/
10.3150/22-BEJ1563.full.

Antognini:2015:ASC

[AZ15] Alessandro Baldi Antognini and Maroussa Zagoraiou. On the almost sure convergence of adaptive allocation procedures. Bernoulli, 21(2):881–908, May 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1429624964.

Bachoc:2018:AAC

[Bac18] François Bachoc. Asymptotic analysis of covariance parameter estimation for Gaussian processes in the misspecified case. Bernoulli, 24(2):1531–1575, May 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1505980903.

Baddeley:2000:TIE

[Bad00] A. J. Baddeley. Time-invariance estimating equations. Bernoulli, 6(5):783–808, October 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1081282689.

Bahmani:2021:NOR

[Bah21] Sohail Bahmani. Nearly optimal robust mean estimation via empirical characteristic function. Bernoulli, 27(3):2139– 2158, August 2021. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-27/issue-3/Nearly-optimal-

robust-mean-estimation-via-empirical-characteristic-
function/10.3150/20-BEJ1304.full.

Bansaye:2024:SIP

[Ban24] Vincent Bansaye. Spine for interacting populations and sampling. Bernoulli, 30(2):1555–1585, May 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
30/issue-2/Spine-for-interacting-populations-and-sampling/
10.3150/23-BEJ1645.full.

Baraud:2002:NAM

[Bar02] Yannick Baraud. Non-asymptotic minimax rates of testing in signal detection. Bernoulli, 8(5):577–606, October 2002. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1078435219.

Baraud:2010:BTI

[Bar10] Yannick Baraud. A Bernstein-type inequality for suprema of random processes with applications to model selection in non-Gaussian regression. Bernoulli, 16(4):1064–1085, November 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1290092896.

Barrera:2018:QIP

[Bar18] David Barrera. Quenched invariance principles for the discrete Fourier transforms of a stationary process. Bernoulli, 24(2): 1307–1350, May 2018. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1505980897.

Barnes:2022:CJP

[Bar22] Clayton Barnes. Convergence of jump processes with stochastic intensity to Brownian motion with inert drift. Bernoulli, 28(2): 1491–1518, May 2022. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-28/issue-2/Convergence-of-
jump-processes-with-stochastic-intensity-to-Brownian-
motion/10.3150/21-BEJ1399.full.

Battey:2017:ESN

[Bat17] Heather Battey. Eigen structure of a new class of covariance and inverse covariance matrices. Bernoulli, 23(4B):3166– 3177, November 2017. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1495505088.

Barthelme:2019:AEF

[BAT19] Simon Barthelmé, Pierre-Olivier Amblard, and Nicolas Tremblay. Asymptotic equivalence of fixed-size and varying-size determinantal point processes. Bernoulli, 25(4B):3555–3589, November 2019. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1569398777.

Bickel:1999:NMN

[BB99] Peter J. Bickel and Peter Bühlmann. A new mixing notion and functional central limit theorems for a sieve bootstrap in time series. Bernoulli, 5(3):413–446, June 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1172617198.

Behme:2017:CSM

[BB17] Anita Behme and Lennart Bondesson. A class of scale mixtures of Gamma(k)-distributions that are generalized gamma convolutions. Bernoulli, 23(1):773–787, February 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1475001374.

Bertazzi:2022:ASP

[BB22] Andrea Bertazzi and Joris Bierkens. Adaptive schemes for piecewise deterministic Monte Carlo algorithms. Bernoulli, 28(4):2404–2430, November 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-28/issue-
4/Adaptive-schemes-for-piecewise-deterministic-Monte-
Carlo-algorithms/10.3150/21-BEJ1423.full.

Bonnery:2012:UCE

[BBC12] Daniel Bonnéry, F. Jay Breidt, and François Coquet. Uniform convergence of the empirical cumulative distribution function under informative selection from a finite population. Bernoulli, 18(4):1361–1385, November 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1352727815.

Bonnery:2018:AMS

[BBC18] Daniel Bonnéry, F. Jay Breidt, and François Coquet. Asymptotics for the maximum sample likelihood estimator under informative selection from a finite population. Bernoulli, 24(2): 929–955, May 2018. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1505980883.

Babu:2011:LTF

[BBCM11] G. Jogesh Babu, Zhidong Bai, Kwok Pui Choi, and Vasudevan Mangalam. Limit theorems for functions of marginal quantiles. Bernoulli, 17(2):671–686, May 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1302009242.

Beutner:2017:FPL

[BBD17] Eric Beutner, Laurent Bordes, and Laurent Doyen. The failure of the profile likelihood method for a large class of semiparametric models. Bernoulli, 23(4B):3650–3684, November 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1495505105.

Beutner:2020:CSE

[BBD20] Eric Beutner, Laurent Bordes, and Laurent Doyen. Consistent semiparametric estimators for recurrent event times models with application to virtual age models. Bernoulli, 26(1):557– 586, February 2020. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1574758838.

Bect:2019:SAG

[BBG19] Julien Bect, François Bachoc, and David Ginsbourger. A supermartingale approach to Gaussian process based sequential design of experiments. Bernoulli, 25(4A):2883–2919, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1568362046.

Bardet:2008:AWB

[BBJ08] Jean-Marc Bardet, Hatem Bibi, and Abdellatif Jouini. Adaptive wavelet-based estimator of the memory parameter for stationary Gaussian processes. Bernoulli, 14(3):691–724, August 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1219669626.

Betancourt:2017:GFH

[BBLG17] Michael Betancourt, Simon Byrne, Sam Livingstone, and Mark Girolami. The geometric foundations of Hamiltonian Monte Carlo. Bernoulli, 23(4A):2257–2298, November 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1494316818.

Basak:2014:LSD

[BBS14] Anirban Basak, Arup Bose, and Sanchayan Sen. Limiting spectral distribution of sample autocovariance matrices. Bernoulli, 20(3):1234–1259, August 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1402488939.

Beck:2015:NCL

[BBS15] Sergej Beck, Jochen Blath, and Michael Scheutzow. A new class of large claim size distributions: Definition, properties, and ruin theory. Bernoulli, 21(4):2457–2483, November 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1438777601.

Berestycki:2016:NAP

[BBS16] Julien Berestycki, Éric Brunet, and Zhan Shi. The number of accessible paths in the hypercube. Bernoulli, 22(2):653– 680, May 2016. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1447077757.

Berghaus:2017:WCE

[BBV17] Betina Berghaus, Axel Bücher, and Stanislav Volgushev. Weak convergence of the empirical copula process with respect to weighted metrics. Bernoulli, 23(1):743–772, February 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1475001373.

Bing:2020:FAM

[BBW20] Xin Bing, Florentina Bunea, and Marten Wegkamp. A fast algorithm with minimax optimal guarantees for topic models with an unknown number of topics. Bernoulli, 26(3):1765–1796, August 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1587974523.

Bing:2022:ILF

[BBW22] Xin Bing, Florentina Bunea, and Marten Wegkamp. Inference in latent factor regression with clusterable features. Bernoulli, 28(2):997–1020, May 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-2/Inference-in-latent-factor-regression-with-
clusterable-features/10.3150/21-BEJ1374.full.

Bing:2023:DAR

[BBW23] Xin Bing, Florentina Bunea, and Marten Wegkamp. Detecting approximate replicate components of a high-dimensional random vector with latent structure. Bernoulli, 29(2):1368– 1391, May 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-29/issue-2/Detecting-
approximate-replicate-components-of-a-high-dimensional-
random-vector/10.3150/22-BEJ1502.full.

Bortot:2000:SPA

[BC00] Paola Bortot and Stuart Coles. A sufficiency property arising from the characterization of extremes of Markov chains. Bernoulli, 6(1):183–190, February 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1082665385.

Bertoin:2002:EIS

[BC02] Jean Bertoin and Maria-Emilia Caballero. Entrance from 0+ for increasing semi-stable Markov processes. Bernoulli, 8(2):195–205, April 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1078866867.

Bertail:2006:RBB

[BC06] Patrice Bertail and Stéphan Clémençon. Regenerative block bootstrap for Markov chains. Bernoulli, 12(4):689–712, August 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1155735932.

Butucea:2009:AEL

[BC09] C. Butucea and F. Comte. Adaptive estimation of linear functionals in the convolution model and applications. Bernoulli, 15(1):69–98, February 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1233669883.

Belloni:2013:LSA

[BC13] Alexandre Belloni and Victor Chernozhukov. Least squares after model selection in high-dimensional sparse models. Bernoulli, 19(2):521–547, May 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1363192037.

Bally:2016:ADC

[BC16] Vlad Bally and Lucia Caramellino. Asymptotic development for the CLT in total variation distance. Bernoulli, 22(4):2442–2485, November 2016. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1462297686.

Bertail:2019:BTE

[BC19] Patrice Bertail and Stephan Clémençon. Bernstein-type exponential inequalities in survey sampling: Conditional Poisson sampling schemes. Bernoulli, 25(4B):3527–3554, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1569398776.

Bonnet:2020:MDP

[BC20] Gilles Bonnet and Nicolas Chenavier. The maximal degree in a Poisson-Delaunay graph. Bernoulli, 26(2):948–979, May 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL https://projecteuclid.org/euclid.bj/1580461569.

Bhudisaksang:2021:ODE

[BC21] Theerawat Bhudisaksang and Álvaro Cartea. Online drift estimation for jump-diffusion processes. Bernoulli, 27(4):2494– 2518, November 2021. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-27/issue-4/Online-drift-estimation-
for-jump-diffusion-processes/10.3150/20-BEJ1319.full.

Bai:2022:AAB

[BCFH22] Zhidong Bai, Kwok Pui Choi, Yasunori Fujikoshi, and Jiang Hu. Asymptotics of AIC, BIC and C_{p} model selection rules in high-dimensional regression. Bernoulli, 28(4):2375–2403, November 2022. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-28/issue-4/Asymptotics-of-
AIC-BIC-and-Cp-model-selection-rules-in/10.3150/21-
BEJ1422.full.

Bardenet:2015:AMO

[BCFK15] Rémi Bardenet, Olivier Cappé, Gersende Fort, and Balázs Kégl. Adaptive MCMC with online relabeling. Bernoulli, 21 (3):1304–1340, August 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1432732021.

Bobkov:2014:FIC

[BCG14] S. G. Bobkov, G. P. Chistyakov, and F. Götze. Fisher information and convergence to stable laws. Bernoulli, 20(3): 1620–1646, August 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1402488952.

Breton:2019:MAD

[BCG19] Jean-Christophe Breton, Adrien Clarenne, and Renan Gobard. Macroscopic analysis of determinantal random balls. Bernoulli, 25(2):1568–1601, May 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1551862860.

Benassi:2002:IPR

[BCI02] Albert Benassi, Serge Cohen, and Jacques Istas. Identification and properties of real harmonizable fractional Lévy motions. Bernoulli, 8(1):97–115, February 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1078951091.

Benassi:2004:RIF

[BCI04] Albert Benassi, Serge Cohen, and Jacques Istas. On roughness indices for fractional fields. Bernoulli, 10(2):357–373, April 2004. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1082380223.

Barkhagen:2021:SGL

[BCM^{+}21] M. Barkhagen, N. H. Chau, É. Moulines, M. Rásonyi, S. Sabanis, and Y. Zhang. On stochastic gradient Langevin dynamics with dependent data streams in the logconcave case. Bernoulli, 27(1):1–33, February 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-1/On-stochastic-gradient-Langevin-dynamics-
with-dependent-data-streams-in/10.3150/19-BEJ1187.full.

Berti:2009:RCP

[BCPR09] Patrizia Berti, Irene Crimaldi, Luca Pratelli, and Pietro Rigo. Rate of convergence of predictive distributions for dependent data. Bernoulli, 15(4):1351–1367, November 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1262962239.

Barber:2024:FTR

[BCRT24] Rina Foygel Barber, Emmanuel J. Candès, Aaditya Ramdas, and Ryan J. Tibshirani. De Finetti’s theorem and related results for infinite weighted exchangeable sequences. Bernoulli, 30 (4):3004–3028, November 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

Bogerd:2020:CRR

[BCvdH20] Kay Bogerd, Rui M. Castro, and Remco van der Hofstad. Cliques in rank-1 random graphs: The role of inhomogeneity. Bernoulli, 26(1):253–285, February 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1574758828.

Bogerd:2021:DPC

[BCvdHV21] Kay Bogerd, Rui M. Castro, Remco van der Hofstad, and Nicolas Verzelen. Detecting a planted community in an inhomogeneous random graph. Bernoulli, 27(2):1159– 1188, May 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-27/issue-2/Detecting-

a-planted-community-in-an-inhomogeneous-random-graph/
10.3150/20-BEJ1269.full.

Barbosa:2009:NLC

[BD09] Euro G. Barbosa and Chang C. Y. Dorea. A note on the Lindeberg condition for convergence to stable laws in Mallows distance. Bernoulli, 15(3):922–924, August 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1251463287.

Baddeley:2013:VEP

[BD13a] Adrian Baddeley and David Dereudre. Variational estimators for the parameters of Gibbs point process models. Bernoulli, 19(3):905–930, August 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1372251147.

Bucher:2013:MBT

[BD13b] Axel Bücher and Holger Dette. Multiplier bootstrap of tail copulas with applications. Bernoulli, 19(5A):1655–1687, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1383661198.

Bergsma:2014:CTI

[BD14] Wicher Bergsma and Angelos Dassios. A consistent test of independence based on a sign covariance related to Kendall’s tau. Bernoulli, 20(2):1006–1028, May 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1393594014.

Bentkus:2015:TGB

[BD15] Vidmantas Kastytis Bentkus and Dainius Dzindzalieta. A tight Gaussian bound for weighted sums of Rademacher random variables. Bernoulli, 21(2):1231–1237, May 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1429624976.

Bioche:2016:AIP

[BD16] Christele Bioche and Pierre Druilhet. Approximation of improper priors. Bernoulli, 22(3):1709–1728, August 2016. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1458132996.

Balabdaoui:2018:ITC

[BD18] Fadoua Balabdaoui and Charles R. Doss. Inference for a two-component mixture of symmetric distributions under logconcavity. Bernoulli, 24(2):1053–1071, May 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1505980889.

Bank:2020:SLS

[BD20a] Peter Bank and Yan Dolinsky. Scaling limits for superreplication with transient price impact. Bernoulli, 26(3):2176– 2201, August 2020. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1587974537.

Bernhard:2020:HTR

[BD20b] Harald Bernhard and Bikramjit Das. Heavy-tailed random walks, buffered queues and hidden large deviations. Bernoulli, 26(1):61–92, February 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1574758822.

Buraczewski:2022:LTS

[BD22] Dariusz Buraczewski and Ewa Damek. Limit theorems for supercritical branching processes in random environment. Bernoulli, 28(3):1602–1624, August 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-3/Limit-theorems-for-supercritical-branching-
processes-in-random-environment/10.3150/21-BEJ1349.
full.

Butucea:2018:MED

[BDDF18] Cristina Butucea, Jean-François Delmas, Anne Dutfoy, and Richard Fischer. Maximum entropy distribution of order statistics with given marginals. Bernoulli, 24(1):115–155, February 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1501142438.

Beirlant:2005:EEV

[BDG05] J. Beirlant, G. Dierckx, and A. Guillou. Estimation of the extreme-value index and generalized quantile plots. Bernoulli, 11(6):949–970, December 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1137421635.

Boufoussi:2007:SPP

[BDG07] Brahim Boufoussi, Marco Dozzi, and Raby Guerbaz. Sample path properties of the local time of multifractional Brownian motion. Bernoulli, 13(3):849–867, August 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1186503490.

Balabdaoui:2019:LSE

[BDJ19] Fadoua Balabdaoui, Cécile Durot, and Hanna Jankowski. Least squares estimation in the monotone single index model. Bernoulli, 25(4B):3276–3310, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1569398766.

Balabdaoui:2017:ADC

[BDK17] Fadoua Balabdaoui, Cécile Durot, and François Koladjo. On asymptotics of the discrete convex LSE of a p.m.f. Bernoulli, 23(3):1449–1480, August 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1489737614.

Buhl:2019:SEI

[BDKS19] Sven Buhl, Richard A. Davis, Claudia Klüppelberg, and Christina Steinkohl. Semiparametric estimation for isotropic max-stable space-time processes. Bernoulli, 25(4A):2508–2537, November 2019. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1568362034.

Butucea:2020:LDP

[BDKS20] Cristina Butucea, Amandine Dubois, Martin Kroll, and Adrien Saumard. Local differential privacy: Elbow effect in optimal density estimation and adaptation over Besov ellipsoids. Bernoulli, 26(3):1727–1764, August 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1587974522.

Berti:2023:KBD

[BDL^{+}23] Patrizia Berti, Emanuela Dreassi, Fabrizio Leisen, Luca Pratelli, and Pietro Rigo. Kernel based Dirichlet sequences. Bernoulli, 29(2):1321–1342, May 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-2/Kernel-based-Dirichlet-sequences/10.3150/
22-BEJ1500.full.

Bose:2009:MUM

[BDM09] Arup Bose, Amites Dasgupta, and Krishanu Maulik. Multicolor urn models with reducible replacement matrices. Bernoulli, 15 (1):279–295, February 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1233669892.

Boufoussi:2010:LTT

[BDM10a] Brahim Boufoussi, Marco Dozzi, and Renaud Marty. Local time and Tanaka formula for a Volterra-type multifractional Gaussian process. Bernoulli, 16(4):1294–1311, November 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1290092907.

Budhiraja:2010:LDS

[BDM10b] Amarjit Budhiraja, Paul Dupuis, and Vasileios Maroulas. Large deviations for stochastic flows of diffeomorphisms. Bernoulli, 16(1):234–257, February 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1265984710.

Bogso:2023:SIR

[BDP23] Antoine-Marie Bogso, Moustapha Dieye, and Olivier Menoukeu Pamen. Stochastic integration with respect to local time of the Brownian sheet and regularising properties of Brownian sheet paths. Bernoulli, 29(4):2627–2651, November 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-29/issue-4/Stochastic-integration-
with-respect-to-local-time-of-the-Brownian/10.3150/
22-BEJ1555.full.

Berti:2021:CMB

[BDPR21] Patrizia Berti, Emanuela Dreassi, Luca Pratelli, and Pietro Rigo. A class of models for Bayesian predictive inference. Bernoulli, 27(1):702–726, February 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-1/A-class-of-models-for-Bayesian-predictive-
inference/10.3150/20-BEJ1255.full.

Blanchard:2014:TCN

[BDR14] Gilles Blanchard, Sylvain Delattre, and Etienne Roquain. Testing over a continuum of null hypotheses with false discovery rate control. Bernoulli, 20(1):304–333, February 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1390407291.

Bourguin:2021:LBL

[BDT21] Solesne Bourguin, Charles-Philippe Diez, and Ciprian A. Tudor. Limiting behavior of large correlated Wishart matrices with chaotic entries. Bernoulli, 27(2):1077–1102, May 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-27/issue-2/Limiting-behavior-of-large-
correlated-Wishart-matrices-with-chaotic-entries/10.
3150/20-BEJ1266.full.

Bobbia:2021:CME

[BDV21] Benjamin Bobbia, Clément Dombry, and Davit Varron. The coupling method in extreme value theory. Bernoulli, 27(3):1824–1850, August 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-27/issue-
3/The-coupling-method-in-extreme-value-theory/10.3150/
20-BEJ1293.full.

Buraczewski:2015:PTA

[BDZ15] D. Buraczewski, E. Damek, and J. Zienkiewicz. Precise tail asymptotics of fixed points of the smoothing transform with general weights. Bernoulli, 21(1):489–504, February 2015. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1426597079.

Benth:2016:SVM

[BE16] Fred Espen Benth and Heidar Eyjolfsson. Simulation of volatility modulated Volterra processes using hyperbolic stochastic partial differential equations. Bernoulli, 22(2):774–793, May 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1447077761.

Begyn:2007:AEC

[Beg07] Arnaud Begyn. Asymptotic expansion and central limit theorem for quadratic variations of Gaussian processes. Bernoulli, 13(3):712–753, August 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1186503484.

Beibel:1997:SCP

[Bei97] M. Beibel. Sequential change-point detection in continuous time when the post-change drift is unknown. Bernoulli, 3 (4):457–478, December 1997. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1175882219.

Belitser:1998:EEA

[Bel98] Eduard Belitser. Efficient estimation of analytic density under random censorship. Bernoulli, 4(4):519–543, December 1998. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1173883819.

Bellec:2017:OEB

[Bel17] Pierre C. Bellec. Optimal exponential bounds for aggregation of density estimators. Bernoulli, 23(1):219–248, February 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1475001354.

Bellec:2021:ACS

[Bel21] Pierre C. Bellec. Adaptive confidence sets in shape restricted regression. Bernoulli, 27(1):66–92, February 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-1/Adaptive-confidence-sets-in-shape-restricted-
regression/10.3150/20-BEJ1223.full.

Bouchard:2018:RBC

[BEM18] Bruno Bouchard, Romuald Elie, and Ludovic Moreau. Regularity of BSDEs with a convex constraint on the gains-process. Bernoulli, 24(3):1613–1635, August 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1517540455.

Bubeck:2017:TSI

[BEMR17] Sébastien Bubeck, Ronen Eldan, Elchanan Mossel, and Miklós Z. Rácz. From trees to seeds: On the inference of the seed from large trees in the uniform attachment model. Bernoulli, 23(4A):2887–2916, November 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1494316836.

Bender:2003:TAI

[Ben03] Christian Bender. An S-transform approach to integration with respect to a fractional Brownian motion. Bernoulli, 9 (6):955–983, December 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1072215197.

Balkema:2012:SLS

[BEN12] Guus Balkema, Paul Embrechts, and Natalia Nolde. Sensitivity of the limit shape of sample clouds from meta densities. Bernoulli, 18(4):1386–1404, November 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1352727816.

Bertail:1997:SOP

[Ber97] Patrice Bertail. Second-order properties of an extrapolated bootstrap without replacement under weak assumptions. Bernoulli, 3(2):149–179, June 1997. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1177526727.

Bercu:2001:LDG

[Ber01] Bernard Bercu. On large deviations in the Gaussian autoregressive process: stable, unstable and explosive cases. Bernoulli, 7(2):299–316, April 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1080222089.

Bernardoff:2003:WNM

[Ber03] Philippe Bernardoff. Which negative multinomial distributions are infinitely divisible? Bernoulli, 9(5):877–893, October 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1066418882.

Bertin:2004:AEM

[Ber04] Karine Bertin. Asymptotically exact minimax estimation in sup-norm for anisotropic Hölder classes. Bernoulli, 10(5):873– 888, October 2004. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1099579160.

Bernardoff:2006:WMG

[Ber06a] Philippe Bernardoff. Which multivariate gamma distributions are infinitely divisible? Bernoulli, 12(1):169–189, February 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1141136656.

Bertail:2006:ELS

[Ber06b] Patrice Bertail. Empirical likelihood in some semiparametric models. Bernoulli, 12(2):299–331, April 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1145993976.

Bertoin:2017:MGF

[Ber17] Jean Bertoin. Markovian growth-fragmentation processes. Bernoulli, 23(2):1082–1101, May 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1486177393.

Beerenwinkel:2007:CBN

[BES07] Niko Beerenwinkel, Nicholas Eriksson, and Bernd Sturmfels. Conjunctive Bayesian networks. Bernoulli, 13(4):893–909, November 2007. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1194625594.

Beran:2001:LPE

[BF01] Jan Beran and Yuanhua Feng. Local polynomial estimation with a FARIMA-GARCH error process. Bernoulli, 7(5):733– 750, October 2001. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1079399539.

Buchmann:2016:DRD

[BFM16] Boris Buchmann, Yuguang Fan, and Ross A. Maller. Distributional representations and dominance of a Lévy process over its maximal jump processes. Bernoulli, 22(4):2325–2371, November 2016. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1462297683.

Belitsky:2001:MEP

[BFMP01] Vladimir Belitsky, Pablo A. Ferrari, Mikhail V. Menshikov, and Serguei Yu. Popov. A mixture of the exclusion process and the voter model. Bernoulli, 7(1):119–144, February 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1080572342.

Bacallado:2015:LBP

[BFT15] Sergio Bacallado, Stefano Favaro, and Lorenzo Trippa. Looking-backward probabilities for Gibbs-type exchangeable random partitions. Bernoulli, 21(1):1–37, February 2015. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1426597062.

Bhattacharya:1995:TSG

[BG95] Rabi N. Bhattacharya and Friedrich Götze. Time-scales for Gaussian approximation and its breakdown under a hierarchy of periodic spatial heterogeneities. Bernoulli, 1(1–2):81–123, March 1995. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1186078363.

Bloznelis:2000:EEF

[BG00] Mindaugas Bloznelis and Friedrich Götze. An Edgeworth expansion for finite-population U-statistics. Bernoulli, 6(4):729– 760, August 2000. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1081449604.

Baringhaus:2009:NTS

[BG09] Ludwig Baringhaus and Rudolf Grübel. Nonparametric twosample tests for increasing convex order. Bernoulli, 15(1): 99–123, February 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1233669884.

Bobkov:2010:CED

[BG10] S. G. Bobkov and F. Götze. Concentration of empirical distribution functions with applications to non-i.i.d. models. Bernoulli, 16(4):1385–1414, November 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1290092911.

Bhaumik:2017:EBE

[BG17] Prithwish Bhaumik and Subhashis Ghosal. Efficient Bayesian estimation and uncertainty quantification in ordinary differential equation models. Bernoulli, 23(4B):3537–3570, November 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1495505101.

Bhattacharjee:2019:DAS

[BG19] Chinmoy Bhattacharjee and Larry Goldstein. Dickman approximation in simulation, summations and perpetuities. Bernoulli, 25(4A):2758–2792, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1568362042.

Brehmer:2021:SIF

[BG21] Jonas R. Brehmer and Tilmann Gneiting. Scoring interval forecasts: Equal-tailed, shortest, and modal interval. Bernoulli, 27(3):1993–2010, August 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-3/Scoring-interval-forecasts-Equal-tailed-
shortest-and-modal-interval/10.3150/20-BEJ1298.full.

Briand:2021:DTT

[BGGL21] Philippe Briand, Christel Geiss, Stefan Geiss, and Céline Labart. Donsker-type theorem for BSDEs: Rate of convergence. Bernoulli, 27(2):899–929, May 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-2/Donsker-type-theorem-for-BSDEs-Rate-of-convergence/
10.3150/20-BEJ1259.full.

Bobkov:2001:GBC

[BGH01] Sergey G. Bobkov, Friedrich Götze, and Christian Houdré. On Gaussian and Bernoulli covariance representations. Bernoulli, 7(3):439–451, June 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1080004759.

Babic:2021:OTE

[BGHL21] Sladana Babić, Laetitia Gelbgras, Marc Hallin, and Christophe Ley. Optimal tests for elliptical symmetry: Specified and unspecified location. Bernoulli, 27(4):2189–2216, November 2021. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-27/issue-4/Optimal-tests-
for-elliptical-symmetry-Specified-and-unspecified-location/
10.3150/20-BEJ1305.full.

Bartroff:2018:BSB

[BGI18] Jay Bartroff, Larry Goldstein, and Ümit Islak. Bounded size biased couplings, log concave distributions and concentration of measure for occupancy models. Bernoulli, 24(4B):3283– 3317, November 2018. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1524038755.

Benhenni:2022:BSP

[BGL22] Karim Benhenni, Didier A. Girard, and Sana Louhichi. On bandwidth selection problems in nonparametric trend estimation under martingale difference errors. Bernoulli, 28(1):395– 423, February 2022. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-28/issue-1/On-bandwidth-selection-
problems-in-nonparametric-trend-estimation-under-martingale/
10.3150/21-BEJ1347.full.

Bui:2024:BEB

[BGL24] Thi Thuy Bui, Ion Grama, and Quansheng Liu. Berry–Esseen bound and Cramér moderate deviation expansion for a supercritical branching random walk. Bernoulli, 30(2):1401– 1415, May 2024. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-30/issue-2/Berry--Esseen-
bound-and-Cram%c3%a9r-moderate-deviation-expansion-
for-a/10.3150/23-BEJ1636.full.

Ball:2014:SMC

[BGMSB14] Frank Ball, Miguel González, Rodrigo Martínez, and Maroussia Slavtchova-Bojkova. Stochastic monotonicity and continuity properties of functions defined on Crump–Mode–Jagers branching processes, with application to vaccination in epidemic modelling. Bernoulli, 20(4):2076–2101, November 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1411134454.

Bardet:2018:FIG

[BGMZ18] Jean-Baptiste Bardet, Nathaël Gozlan, Florent Malrieu, and Pierre-André Zitt. Functional inequalities for Gaussian convolutions of compactly supported measures: Explicit bounds and dimension dependence. Bernoulli, 24(1):333–353, February 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1501142446.

Belomestny:2022:NBV

[BGSS22] Denis Belomestny, Shota Gugushvili, Moritz Schauer, and Peter Spreij. Nonparametric Bayesian volatility estimation for gamma-driven stochastic differential equations. Bernoulli, 28(4):2151–2180, November 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-28/issue-
4/Nonparametric-Bayesian-volatility-estimation-for-
gamma-driven-stochastic-differential-equations/10.3150/
21-BEJ1413.full.

Bentkus:1997:BEB

[BGT97] V. Bentkus, F. Götze, and A. Tikhomoirov. Berry–Esseen bounds for statistics of weakly dependent samples. Bernoulli, 3(3):329–349, September 1997. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1177334459.

Bojdecki:2007:OTF

[BGT07] Tomasz Bojdecki, Luis G. Gorostiza, and Anna Talarczyk. Occupation time fluctuations of an infinite-variance branching system in large dimensions. Bernoulli, 13(1):20–39, February 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1175287718.

Barbour:2015:SFN

[BGX15] A. D. Barbour, H. L. Gan, and A. Xia. Stein factors for negative binomial approximation in Wasserstein distance. Bernoulli, 21(2):1002–1013, May 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1429624968.

Bobkov:1996:VLF

[BH96] Sergei G. Bobkov and Christian Houdré. Variance of Lipschitz functions and an isoperimetric problem for a class of product measures. Bernoulli, 2(3):249–255, September 1996. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1178291721.

Bobkov:2000:WDF

[BH00] Sergey G. Bobkov and Christian Houdré. Weak dimensionfree concentration of measure. Bernoulli, 6(4):621–632, August 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1081449597.

Breton:2010:ARY

[BH10] Jean-Christophe Breton and Christian Houdré. Asymptotics for random Young diagrams when the word length and alphabet size simultaneously grow to infinity. Bernoulli, 16(2):471– 492, May 2010. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1274821080.

Barthe:2017:PUE

[BH17] Franck Barthe and Benoît Huou. Perimeters, uniform enlargement and high dimensions. Bernoulli, 23(2):1056–1081, May 2017. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1486177392.

Bhattacharya:2023:LDE

[Bha23]  Ayan Bhattacharya.  Large deviations of extremes in

branching random walk with regularly varying displacements. Bernoulli, 29(2):1007–1034, May 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-2/Large-deviations-of-extremes-in-branching-
random-walk-with-regularly/10.3150/22-BEJ1488.full.

Ben-Hamou:2017:CII

[BHBO17] Anna Ben-Hamou, Stéphane Boucheron, and Mesrob I. Ohannessian. Concentration inequalities in the infinite urn scheme for occupancy counts and the missing mass, with applications. Bernoulli, 23(1):249–287, February 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1475001355.

Berkes:2003:GPS

[BHK03] István Berkes, Lajos Horváth, and Piotr Kokoszka. GARCH processes: structure and estimation. Bernoulli, 9(2):201–227, April 2003. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1068128975.

Brigo:1999:ANF

[BHL99] Damiano Brigo, Bernard Hanzon, and François Le Gland. Approximate nonlinear filtering by projection on exponential manifolds of densities. Bernoulli, 5(3):495–534, June 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1172617201.

Barral:2018:MBR

[BHM18] Julien Barral, Yueyun Hu, and Thomas Madaule. The minimum of a branching random walk outside the boundary case. Bernoulli, 24(2):801–841, May 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1505980879.

Braunsteins:2022:PEB

[BHM22] Peter Braunsteins, Sophie Hautphenne, and Carmen Minuesa. Parameter estimation in branching processes with almost sure extinction. Bernoulli, 28(1):33–63, February 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-28/issue-1/Parameter-estimation-in-
branching-processes-with-almost-sure-extinction/10.
3150/21-BEJ1332.full.

Bai:2015:CES

[BHPZ15] Zhidong Bai, Jiang Hu, Guangming Pan, and Wang Zhou. Convergence of the empirical spectral distribution function of beta matrices. Bernoulli, 21(3):1538–1574, August 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1432732029.

Berkes:2011:ATC

[BHS11] István Berkes, Lajos Horváth, and Johannes Schauer. Asymptotics of trimmed CUSUM statistics. Bernoulli, 17(4):1344– 1367, November 2011. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL http://projecteuclid.org/
euclid.bj/1320417507.

Bucher:2017:NTD

[BHVD17] Axel Bücher, Michael Hoffmann, Mathias Vetter, and Holger Dette. Nonparametric tests for detecting breaks in the jump behaviour of a time-continuous process. Bernoulli, 23(2):1335– 1364, May 2017. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1486177401.

Butucea:2013:DSS

[BI13] Cristina Butucea and Yuri I. Ingster. Detection of a sparse submatrix of a high-dimensional noisy matrix. Bernoulli, 19(5B): 2652–2688, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1386078616.

Bianconcini:2014:APA

[Bia14] Silvia Bianconcini. Asymptotic properties of adaptive maximum likelihood estimators in latent variable models. Bernoulli, 20(3):1507–1531, August 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1402488948.

Belomestny:2022:EVM

[BIPZ22] Denis Belomestny, Leonid Iosipoi, Quentin Paris, and Nikita Zhivotovskiy. Empirical variance minimization with applications in variance reduction and optimal control. Bernoulli, 28(2):1382–1407, May 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-28/issue-
2/Empirical-variance-minimization-with-applications-
in-variance-reduction-and-optimal/10.3150/21-BEJ1392.
full.

Birge:2004:MSG

[Bir04] Lucien Birgé. Model selection for Gaussian regression with random design. Bernoulli, 10(6):1039–1051, December 2004. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1106314849.

Banjevic:2002:RMF

[BIZ02] Dragan Banjevic, Hemant Ishwaran, and Mahmoud Zarepour. A recursive method for functionals of Poisson processes. Bernoulli, 8(3):295–311, April 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1078779872.

Bardina:2000:WAB

[BJ00] Xavier Bardina and Maria Jolis. Weak approximation of the Brownian sheet from a Poisson process in the plane. Bernoulli, 6(4):653–665, August 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1081449599.

Bonnefont:2022:NEE

[BJ22] Michel Bonnefont and Aldéric Joulin. A note on eigenvalues estimates for one-dimensional diffusion operators. Bernoulli, 28(1):64–86, February 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-28/issue-
1/A-note-on-eigenvalues-estimates-for-one-dimensional-
diffusion-operators/10.3150/21-BEJ1333.full.

Bucher:2024:SHT

[BJ24] Axel Bücher and Tobias Jennessen. Statistics for heteroscedastic time series extremes. Bernoulli, 30(1):46–71, February 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-30/issue-1/Statistics-
for-heteroscedastic-time-series-extremes/10.3150/22-
BEJ1560.full.

Bossy:2021:WCR

[BJM21] Mireille Bossy, Jean-François Jabir, and Kerlyns Martínez. On the weak convergence rate of an exponential Euler scheme for SDEs governed by coefficients with superlinear growth. Bernoulli, 27(1):312–347, February 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-1/On-the-weak-convergence-rate-of-an-exponential-
Euler-scheme/10.3150/20-BEJ1241.full.

Bentkus:2007:LDN

[BJSZ07] Vidmantas Bentkus, Bing-Yi Jing, Qi-Man Shao, and Wang Zhou. Limiting distributions of the non-central t-statistic and their applications to the power of t-tests under nonnormality. Bernoulli, 13(2):346–364, May 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1179498752.

Blower:2002:NKD

[BK02] Gordon Blower and Julia E. Kelsall. Nonlinear kernel density estimation for binned data: convergence in entropy. Bernoulli, 8(4):423–449, August 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1078681378.

Buchmann:2006:FIE

[BK06] Boris Buchmann and Claudia Klüppelberg. Fractional integral equations and state space transforms. Bernoulli, 12(3):431– 456, June 2006. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1151525129.

Bucher:2016:DMB

[BK16] Axel Bücher and Ivan Kojadinovic. A dependent multiplier bootstrap for the sequential empirical copula process under strong mixing. Bernoulli, 22(2):927–968, May 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1447077765.

Bolin:2023:EMA

[BK23] David Bolin and Kristin Kirchner. Equivalence of measures and asymptotically optimal linear prediction for Gaussian random fields with fractional-order covariance operators. Bernoulli, 29(2):1476–1504, May 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-2/Equivalence-of-measures-and-asymptotically-
optimal-linear-prediction-for-Gaussian/10.3150/22-BEJ1507.
full.

Bardet:2023:ECM

[BKK23] Jean-Marc Bardet, Kamila Kare, and William Kengne. Efficient and consistent model selection procedures for time series. Bernoulli, 29(4):2652–2690, November 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-4/Efficient-and-consistent-model-selection-
procedures-for-time-series/10.3150/22-BEJ1556.full.

Baldi:2009:SNC

[BKMP09] P. Baldi, G. Kerkyacharian, D. Marinucci, and D. Picard. Subsampling needlet coefficients on the sphere. Bernoulli, 15(2):438–463, May 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1241444897.

Balkema:1999:LLE

[BKR99] August A. Balkema, Claudia Klüppelberg, and Sidney I. Resnick. Limit laws for exponential families. Bernoulli, 5 (6):951–968, December 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1143122297.

Burdzy:2019:IED

[BKT19] Krzysztof Burdzy, Bartosz Kołodziejek, and Tvrtko Tadić. Inverse exponential decay: Stochastic fixed point equation and ARMA models. Bernoulli, 25(4B):3939–3977, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL https://projecteuclid.org/euclid.bj/1569398790.

Brunel:2021:ECS

[BKY21] Victor-Emmanuel Brunel, Jason M. Klusowski, and Dana Yang. Estimation of convex supports from noisy measurements. Bernoulli, 27(2):772–793, May 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-2/Estimation-of-convex-supports-from-noisy-
measurements/10.3150/20-BEJ1229.full.

Bickel:2004:STF

[BL04] Peter J. Bickel and Elizaveta Levina. Some theory for Fisher’s linear discriminant function, ‘naive Bayes’, and some alternatives when there are many more variables than observations. Bernoulli, 10(6):989–1010, December 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1106314847.

Bouaziz:2010:CDE

[BL10] Olivier Bouaziz and Olivier Lopez. Conditional density estimation in a censored single-index regression model. Bernoulli, 16(2):514–542, May 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1274821082.

Beltran:2013:QNP

[BL13] Johel Beltran and Jean-François Le Gall. Quadrangulations with no pendant vertices. Bernoulli, 19(4):1150–1175, September 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1377612847.

Blanchet:2014:TVA

[BL14] Jose Blanchet and Jingchen Liu. Total variation approximations and conditional limit theorems for multivariate regularly varying random walks conditioned on ruin. Bernoulli, 20(2):416–456, May 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1393593993.

Bednorz:2015:ICT

[BŁ15a] Witold Marek Bednorz and Rafał Marcin Łochowski. Integrability and concentration of the truncated variation for the sample paths of fractional Brownian motions, diffusions and Lévy processes. Bernoulli, 21(1):437–464, February 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1426597077.

Bierme:2015:MCS

[BL15b] Hermine Biermé and Céline Lacaux. Modulus of continuity of some conditionally sub-Gaussian fields, application to stable random fields. Bernoulli, 21(3):1719–1759, August 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1432732035.

Biscio:2016:QRD

[BL16] Christophe Ange Napoléon Biscio and Frédéric Lavancier. Quantifying repulsiveness of determinantal point processes. Bernoulli, 22(4):2001–2028, November 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1462297673.

Brecheteau:2020:PBD

[BL20] Claire Brécheteau and Clément Levrard. A k-points-based distance for robust geometric inference. Bernoulli, 26(4):3017– 3050, November 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-26/issue-4/A-k-points-
based-distance-for-robust-geometric-inference/10.3150/
20-BEJ1214.full.

Berger:2022:CWD

[BL22] David Berger and Alexander Lindner. A Cramér–Wold device for infinite divisibility of Z^{d}-valued distributions. Bernoulli, 28(2):1276–1283, May 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-28/issue-
2/A-Cram%c3%a9rWold-device-for-infinite-divisibility-
of-Zd-valued-distributions/10.3150/21-BEJ1386.full.

Bednorz:2023:MTH

[BL23] Witold M. Bednorz and RafaL M. Lochowski. Moments and tails of hitting times of Bessel processes and convolutions of elementary mixtures of exponential distributions. Bernoulli, 29(3):2219–2246, August 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-3/Moments-and-tails-of-hitting-times-of-Bessel-
processes-and/10.3150/22-BEJ1539.full.

Boissard:2015:DTE

[BLL15] Emmanuel Boissard, Thibaut Le Gouic, and Jean-Michel Loubes. Distribution’s template estimate with Wasserstein metrics. Bernoulli, 21(2):740–759, May 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1429624959.

Bar-Lev:2012:STB

[BLLS12] Shaul K. Bar-Lev, Andreas Löpker, and Wolfgang Stadje. On the small-time behavior of subordinators. Bernoulli, 18(3):823– 835, August 2012. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1340887004.

Buchmann:2019:WSM

[BLM19] Boris Buchmann, Kevin W. Lu, and Dilip B. Madan. Weak subordination of multivariate Lévy processes and variance generalised gamma convolutions. Bernoulli, 25(1):742–770, February 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1544605262.

Blower:2004:TMY

[Blo04] Gordon Blower. Transportation of measure, Young diagrams and random matrices. Bernoulli, 10(5):755–782, October 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1099579155.

Beiglbock:2019:DAM

[BLO19] Mathias Beiglböck, Tongseok Lim, and Jan Obłój. Dual attainment for the martingale transport problem. Bernoulli, 25 (3):1640–1658, August 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1560326422.

Bai:2019:CLT

[BLP19] Zhidong Bai, Huiqin Li, and Guangming Pan. Central Limit Theorem for linear spectral statistics of large dimensional separable sample covariance matrices. Bernoulli, 25(3):1838–1869, August 2019. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1560326430.

Birge:1998:MCE

[BM98] Lucien Birgé and Pascal Massart. Minimum contrast estimators on sieves: exponential bounds and rates of convergence. Bernoulli, 4(3):329–375, September 1998. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1174324984.

Butucea:2005:MEN

[BM05] Cristina Butucea and Catherine Matias. Minimax estimation of the noise level and of the deconvolution density in a semiparametric convolution model. Bernoulli, 11(2):309–340, April 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1116340297.

Banerjee:2007:EOS

[BM07] Moulinath Banerjee and Ian W. McKeague. Estimating optimal step-function approximations to instantaneous hazard rates. Bernoulli, 13(1):279–299, February 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1175287733.

Bender:2008:SCC

[BM08] Christian Bender and Tina Marquardt. Stochastic calculus for convoluted Lévy processes. Bernoulli, 14(2):499–518, May 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1208872115.

Bouchard:2009:SAB

[BM09] Bruno Bouchard and Stéphane Menozzi. Strong approximations of BSDEs in a domain. Bernoulli, 15(4):1117–1147, November 2009. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1262962228.

Buchmann:2012:LEG

[BM12] Boris Buchmann and Gernot Müller. Limit experiments of GARCH. Bernoulli, 18(1):64–99, February 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1327068618.

Blocker:2013:PPP

[BM13] Alexander W. Blocker and Xiao-Li Meng. The potential and perils of preprocessing: Building new foundations. Bernoulli, 19 (4):1176–1211, September 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1377612848.

Bai:2015:SDE

[BM15a] Lihua Bai and Jin Ma. Stochastic differential equations driven by fractional Brownian motion and Poisson point process. Bernoulli, 21(1):303–334, February 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1426597071.

Bardenet:2015:CIS

[BM15b] Rémi Bardenet and Odalric-Ambrym Maillard. Concentration inequalities for sampling without replacement. Bernoulli, 21 (3):1361–1385, August 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1432732023.

Bayraktar:2016:SLT

[BM16] Erhan Bayraktar and Alexander Munk. An α-stable limit theorem under sublinear expectation. Bernoulli, 22(4):2548–2578, November 2016. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1462297689.

Bhattacharya:2018:IIM

[BM18] Bhaswar B. Bhattacharya and Sumit Mukherjee. Inference in Ising models. Bernoulli, 24(1):493–525, February 2018. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1501142453.

Behme:2012:DEI

[BMMS12] Anita Behme, Makoto Maejima, Muneya Matsui, and Noriyoshi Sakuma. Distributions of exponential integrals of independent increment processes related to generalized gamma convolutions. Bernoulli, 18(4):1172–1187, November 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1352727806.

Berg:2008:DFI

[BMP08] Christian Berg, Jorge Mateu, and Emilio Porcu. The Dagum family of isotropic correlation functions. Bernoulli, 14(4): 1134–1149, November 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1225980574.

Bladt:2021:CSS

[BMS21] Mogens Bladt, Marcin Mider, and Michael Sørensen. Corrigendum to “Simple simulation of diffusion bridges with application to likelihood inference for diffusions”. Bernoulli, 27(1):218– 220, February 2021. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-27/issue-1/Corrigendum-to-
Simple-simulation-of-diffusion-bridges-with-application-
to/10.3150/20-BEJ1234.full. See [BS14].

Betz:2021:PAL

[BMSZ21] Volker Betz, Julian Mühlbauer, Helge Schäfer, and Dirk Zeindler. Precise asymptotics of longest cycles in random permutations without macroscopic cycles. Bernoulli, 27(3):1529– 1555, August 2021. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-27/issue-3/Precise-asymptotics-
of-longest-cycles-in-random-permutations-without-macroscopic/
10.3150/20-BEJ1282.full.

Bobrowski:2017:TCK

[BMT17] Omer Bobrowski, Sayan Mukherjee, and Jonathan E. Taylor. Topological consistency via kernel estimation. Bernoulli, 23 (1):288–328, February 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1475001356.

Butucea:2005:EAE

[BN05] Cristina Butucea and Michael H. Neumann. Exact asymptotics for estimating the marginal density of discretely observed diffusion processes. Bernoulli, 11(3):411–444, June 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1120591183.

Belomestny:2017:MPS

[BN17] Denis Belomestny and Tigran Nagapetyan. Multilevel path simulation for weak approximation schemes with application to Lévy-driven SDEs. Bernoulli, 23(2):927–950, May 2017. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1486177388.

Belitser:2020:NSH

[BN20] Eduard Belitser and Nurzhan Nurushev. Needles and straw in a haystack: Robust confidence for possibly sparse sequences. Bernoulli, 26(1):191–225, February 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1574758826.

Barndorff-Nielsen:2011:QOU

[BNBO11] Ole E. Barndorff-Nielsen and Andreas Basse-O’Connor. Quasi Ornstein–Uhlenbeck processes. Bernoulli, 17(3):916–941, August 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1310042850.

Barndorff-Nielsen:2013:MES

[BNBV13] Ole E. Barndorff-Nielsen, Fred Espen Benth, and Almut E. D. Veraart. Modelling energy spot prices by volatility modulated Lévy-driven Volterra processes. Bernoulli, 19(3):803–845, August 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1372251144.

Barndorff-Nielsen:1996:PA

[BNC96] O. E. Barndorff-Nielsen and David R. Cox. Prediction and asymptotics. Bernoulli, 2(4):319–340, December 1996. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1178291834.

Barndorff-Nielsen:2011:MVB

[BNCP11] Ole E. Barndorff-Nielsen, José Manuel Corcuera, and Mark Podolskij. Multipower variation for Brownian semistationary processes. Bernoulli, 17(4):1159–1194, November 2011. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1320417500.

Barndorff-Nielsen:2008:PML

[BNH08] Ole E. Barndorff-Nielsen and Friedrich Hubalek. Probability measures, Lévy measures and analyticity in time. Bernoulli, 14(3):764–790, August 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1219669629.

Barndorff-Nielsen:2006:SCM

[BNMS06] Ole E. Barndorff-Nielsen, Makoto Maejima, and Ken-Iti Sato. Some classes of multivariate infinitely divisible distributions admitting stochastic integral representations. Bernoulli, 12(1):1– 33, February 2006. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1141136646.

Barndorff-Nielsen:2003:RPV

[BNS03] Ole E. Barndorff-Nielsen and Neil Shephard. Realized power variation and stochastic volatility models. Bernoulli, 9(2):243– 265, April 2003. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1068128977.

Barndorff-Nielsen:2002:SDL

[BNT02] Ole E. Barndorff-Nielsen and Steen Thorbjørnsen. Selfdecomposability and Lévy processes in free probability. Bernoulli, 8(3):323–366, April 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1078779874.

Barndorff-Nielsen:1998:LDC

[BNW98] Ole E. Barndorff-Nielsen and Andrew T. A. Wood. On large deviations and choice of ancillary for p∗ and r∗. Bernoulli, 4(1):35–63, March 1998. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1175865489.

Bossy:2018:SCS

[BO18] Mireille Bossy and Héctor Olivero. Strong convergence of the symmetrized Milstein scheme for some CEV-like SDEs. Bernoulli, 24(3):1995–2042, August 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1517540466.

Bonnier:2020:SCO

[BO20] Patric Bonnier and Harald Oberhauser. Signature cumulants, ordered partitions, and independence of stochastic processes. Bernoulli, 26(4):2727–2757, November 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
26/issue-4/Signature-cumulants-ordered-partitions-and-
independence-of-stochastic-processes/10.3150/20-BEJ1205.
full.

Basse-OConnor:2018:LTL

[BOHP18] Andreas Basse-O’Connor, Claudio Heinrich, and Mark Podolskij. On limit theory for Lévy semi-stationary processes. Bernoulli, 24(4A):3117–3146, November 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1522051235.

Borowiecka-Olszewska:2015:LPS

[BOJGMR15] M. Borowiecka-Olszewska, B. H. Jasiulis-Gołdyn, J. K. Misiewicz, and J. Rosiński. Lévy processes and stochastic integrals in the sense of generalized convolutions. Bernoulli, 21 (4):2513–2551, November 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1438777603.

Barp:2022:RSK

[BOPG22] Alessandro Barp, Chris. J. Oates, Emilio Porcu, and Mark Girolami. A Riemann–Stein kernel method. Bernoulli, 28(4):2181–2208, November 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-28/issue-
4/A-RiemannStein-kernel-method/10.3150/21-BEJ1415.full.

Borkovec:2001:ABS

[Bor01] Milan Borkovec. Asymptotic behaviour of the sample autocovariance and autocorrelation function of the AR(1) process with ARCH(1) errors. Bernoulli, 7(6):847–872, December 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1078951126.

Borda:2021:ERW

[Bor21] Bence Borda. Equidistribution of random walks on compact groups II. The Wasserstein metric. Bernoulli, 27(4):2598– 2623, November 2021. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-27/issue-4/Equidistribution-
of-random-walks-on-compact-groups-II-The-Wasserstein/
10.3150/21-BEJ1324.full.

Boutsikas:2006:CPP

[Bou06] Michael V. Boutsikas. Compound Poisson process approximation for locally dependent real-valued random variables via a new coupling inequality. Bernoulli, 12(3):501–514, June 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1151525133.

Bourguin:2015:PCF

[Bou15] Solesne Bourguin. Poisson convergence on the free Poisson algebra. Bernoulli, 21(4):2139–2156, November 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1438777589.

Butler:1998:ADV

[BP98] Ronald W. Butler and Marc S. Paolella. Approximate distributions for the various serial correlograms. Bernoulli, 4(4): 497–518, December 1998. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1173883818.

Bally:2003:QAS

[BP03a] Vlad Bally and Gilles Pagès. A quantization algorithm for solving multidimensional discrete-time optimal stopping problems. Bernoulli, 9(6):1003–1049, December 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1072215199.

Bellanger:2003:CPL

[BP03b] Lise Bellanger and Gonzalo Perera. Compound Poisson limit theorems for high-level exceedances of some non-stationary processes. Bernoulli, 9(3):497–515, June 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1065444815.

Baldi:2006:ECS

[BP06] Paolo Baldi and Barbara Pacchiarotti. Explicit computation of second-order moments of importance sampling estimators for fractional Brownian motion. Bernoulli, 12(4):663–688, August 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1155735931.

Butler:2008:USA

[BP08] Ronald W. Butler and Marc S. Paolella. Uniform saddlepoint approximations for ratios of quadratic forms. Bernoulli, 14 (1):140–154, February 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1202492788.

Bender:2010:AGF

[BP10] Christian Bender and Peter Parczewski. Approximating a geometric fractional Brownian motion and related processes via discrete Wick calculus. Bernoulli, 16(2):389–417, May 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1274821076.

Balabdaoui:2011:DMD

[BP11] Fadoua Balabdaoui and Jim Pitman. The distribution of the maximal difference between a Brownian bridge and its concave majorant. Bernoulli, 17(1):466–483, February 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1297173851.

Barrera:2016:QLT

[BP16] David Barrera and Magda Peligrad. Quenched limit theorems for Fourier transforms and periodogram. Bernoulli, 22 (1):275–301, February 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1443620850.

Bertail:2019:RCM

[BP19a] Patrice Bertail and François Portier. Rademacher complexity for Markov chains: Applications to kernel smoothing and Metropolis–Hastings. Bernoulli, 25(4B):3912–3938, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1569398789.

Bickel:2019:SCC

[BP19b] David R. Bickel and Alexandre G. Patriota. Self-consistent confidence sets and tests of composite hypotheses applicable to restricted parameters. Bernoulli, 25(1):47–74, February 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL https://projecteuclid.org/euclid.bj/1544605238.

Basrak:2021:CPA

[BP21] Bojan Basrak and Hrvoje Planinić. Compound Poisson approximation for regularly varying fields with application to sequence alignment. Bernoulli, 27(2):1371–1408, May 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-2/Compound-Poisson-approximation-for-regularly-
varying-fields-with-application-to/10.3150/20-BEJ1278.
full.

Bachoc:2022:AEP

[BPB^{+}22] François Bachoc, Emilio Porcu, Moreno Bevilacqua, Reinhard Furrer, and Tarik Faouzi. Asymptotically equivalent prediction in multivariate geostatistics. Bernoulli, 28(4):2518–2545, November 2022. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-28/issue-4/Asymptotically-
equivalent-prediction-in-multivariate-geostatistics/
10.3150/21-BEJ1427.full.

Bui:2023:IPO

[BPD23] Mai Ngoc Bui, Yvo Pokern, and Petros Dellaportas. Inference for partially observed Riemannian Ornstein–Uhlenbeck diffusions of covariance matrices. Bernoulli, 29(4):2961–2986, November 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-29/issue-4/Inference-
for-partially-observed-Riemannian-OrnsteinUhlenbeck-
diffusions-of-covariance-matrices/10.3150/22-BEJ1570.
full.

Beskos:2006:RES

[BPR06] Alexandros Beskos, Omiros Papaspiliopoulos, and Gareth O. Roberts. Retrospective exact simulation of diffusion sample paths with applications. Bernoulli, 12(6):1077–1098, December 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1165269151.

Beskos:2012:SSB

[BPR12] Alexandros Beskos, Stefano Peluchetti, and Gareth Roberts. ε-strong simulation of the Brownian path. Bernoulli, 18(4): 1223–1248, November 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1352727808.

Beskos:2013:OTH

[BPR^{+}13] Alexandros Beskos, Natesh Pillai, Gareth Roberts, Jesus-Maria Sanz-Serna, and Andrew Stuart. Optimal tuning of the hybrid Monte Carlo algorithm. Bernoulli, 19(5A):1501–1534, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1383661192.

Belomestny:2019:LFE

[BPW19] Denis Belomestny, Vladimir Panov, and Jeannette H. C. Woerner. Low-frequency estimation of continuous-time moving average Lévy processes. Bernoulli, 25(2):902–931, May 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL https://projecteuclid.org/euclid.bj/1551862839.

Bao:2015:LLR

[BPZ15] Zhigang Bao, Guangming Pan, and Wang Zhou. The logarithmic law of random determinant. Bernoulli, 21(3):1600–1628, August 2015. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1432732031.

Bickel:1996:IHM

[BR96] Peter J. Bickel and Ya’acov Ritov. Inference in hidden Markov models i: Local asymptotic normality in the stationary case. Bernoulli, 2(3):199–228, September 1996. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1178291719.

Balaji:1997:RTD

[BR97] Srinivasan Balaji and Sundareswaran Ramasubramanian. Recurrence and transience of diffusions in a half-space. Bernoulli, 3(1):97–119, February 1997. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1178291934.

Bardina:2007:IFE

[BR07] Xavier Bardina and Carles Rovira. On Itô’s formula for elliptic diffusion processes. Bernoulli, 13(3):820–830, August 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1186503488.

Besalu:2012:SDE

[BR12] Mireia Besalú and Carles Rovira. Stochastic delay equations with non-negativity constraints driven by fractional Brownian motion. Bernoulli, 18(1):24–45, February 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1327068616.

Bandini:2018:SWD

[BR18a] Elena Bandini and Francesco Russo. Special weak Dirichlet processes and BSDEs driven by a random measure. Bernoulli, 24(4A):2569–2609, November 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1522051218.

Blancas:2018:BPR

[BR18b] Airam Blancas and Víctor Rivero. On branching process with rare neutral mutation. Bernoulli, 24(2):1576–1612, May 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL https://projecteuclid.org/euclid.bj/1505980904.

Bradley:1999:GVC

[Bra99] Richard C. Bradley. On the growth of variances in a central limit theorem for strongly mixing sequences. Bernoulli, 5(1):67– 80, February 1999. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1173707095.

Bradley:2002:PSD

[Bra02] Richard C. Bradley. On positive spectral density functions. Bernoulli, 8(2):175–193, April 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1078866866.

Bravo:2014:BLP

[Bra14] Gerónimo Uribe Bravo. Bridges of Lévy processes conditioned to stay positive. Bernoulli, 20(1):190–206, February 2014. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1390407285.

Bras:2022:CRG

[Bra22] Pierre Bras. Convergence rates of Gibbs measures with degenerate minimum. Bernoulli, 28(4):2431–2458, November 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-28/issue-4/Convergence-rates-of-Gibbs-
measures-with-degenerate-minimum/10.3150/21-BEJ1424.
full.

Bou-Rabee:2023:MTG

[BRE23] Nawaf Bou-Rabee and Andreas Eberle. Mixing time guarantees for unadjusted Hamiltonian Monte Carlo. Bernoulli, 29(1): 75–104, February 2023. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-29/issue-1/Mixing-time-guarantees-
for-unadjusted-Hamiltonian-Monte-Carlo/10.3150/21-BEJ1450.
full.

Berenguer-Rico:2019:CAF

[BRJN19] Vanessa Berenguer-Rico, Søren Johansen, and Bent Nielsen. Corrigendum: Analysis of the forward search using some new results for martingales and empirical processes. Bernoulli, 25(4A):3201, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1568362057. See [JN16].

Barbour:2019:EBL

[BRR19] A. D. Barbour, Adrian Röllin, and Nathan Ross. Error bounds in local limit theorems using Stein’s method. Bernoulli, 25(2): 1076–1104, May 2019. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1551862844.

Briand:2007:ODB

[BRS07] Philippe Briand, Jean-Pier Relepeltier, and Jaime San Martín. One-dimensional backward stochastic differential equations whose coefficient is monotonic in y and non-Lipschitz in z. Bernoulli, 13(1):80–91, February 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1175287721.

Bouchet:2016:SEC

[BRS16] Élodie Bouchet, Alejandro F. Ramírez, and Christophe Sabot. Sharp ellipticity conditions for ballistic behavior of random walks in random environment. Bernoulli, 22(2):969–994, May 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1447077766.

Brunel:2019:UBR

[Bru19] Victor-Emmanuel Brunel. Uniform behaviors of random polytopes under the Hausdorff metric. Bernoulli, 25(3):1770–1793, August 2019. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1560326427.

Brunel:2020:DIR

[Bru20] Victor-Emmanuel Brunel. Deviation inequalities for random polytopes in arbitrary convex bodies. Bernoulli, 26(4):2488– 2502, November 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-26/issue-4/Deviation-
inequalities-for-random-polytopes-in-arbitrary-convex-
bodies/10.3150/19-BEJ1164.full.

Breslow:2000:SPE

[BRW00] Norman E. Breslow, James M. Robins, and Jon A. Wellner. On the semi-parametric efficiency of logistic regression under casecontrol sampling. Bernoulli, 6(3):447–455, June 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1081616700.

Bibby:1995:MEF

[BS95] Bo Martin Bibby and Michael Sørensen. Martingale estimation functions for discretely observed diffusion processes. Bernoulli, 1(1–2):17–39, March 1995. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1186078360.

Beghdadi-Sakrani:2002:UCC

[BS02] Samia Beghdadi-Sakrani. The uniqueness class of continuous local martingales. Bernoulli, 8(2):207–217, April 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1078866868.

Beran:2009:APM

[BS09] Jan Beran and Martin Schützner. On approximate pseudomaximum likelihood estimation for LARCH-processes. Bernoulli, 15(4):1057–1081, November 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1262962226.

Bardet:2011:MRR

[BS11] Jean-Marc Bardet and Donatas Surgailis. Measuring the roughness of random paths by increment ratios. Bernoulli, 17(2):749– 780, May 2011. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1302009246.

Beran:2012:AOW

[BS12] Jan Beran and Yevgen Shumeyko. On asymptotically optimal wavelet estimation of trend functions under long-range dependence. Bernoulli, 18(1):137–176, February 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1327068621.

Bladt:2014:SSD

[BS14] Mogens Bladt and Michael Sørensen. Simple simulation of diffusion bridges with application to likelihood inference for diffusions. Bernoulli, 20(2):645–675, May 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:/

/projecteuclid.org/euclid.bj/1393594001. See corrigen-

dum [BMS21].

Behme:2015:CIM

[BS15a] Anita Behme and Alexander Schnurr. A criterion for invariant measures of Itô processes based on the symbol. Bernoulli, 21 (3):1697–1718, August 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1432732034.

Beiglbock:2015:PVB

[BS15b] Mathias Beiglböck and Pietro Siorpaes. Pathwise versions of the Burkholder–Davis–Gundy inequality. Bernoulli, 21(1):360–373, February 2015. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1426597073.

Bosch:2015:IDI

[BS15c] Pierre Bosch and Thomas Simon. On the infinite divisibility of inverse beta distributions. Bernoulli, 21(4):2552–2568, November 2015. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1438777604.

Benth:2016:ITI

[BS16] Fred Espen Benth and André Süss. Integration theory for infinite dimensional volatility modulated Volterra processes. Bernoulli, 22(3):1383–1430, August 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1458132986.

Bruggeman:2018:MCS

[BS18a] Cameron Bruggeman and Andrey Sarantsev. Multiple collisions in systems of competing Brownian particles. Bernoulli, 24(1): 156–201, February 2018. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1501142439.

Bucher:2018:MLE

[BS18b] Axel Bücher and Johan Segers. Maximum likelihood estimation for the Fréchet distribution based on block maxima extracted from a time series. Bernoulli, 24(2):1427–1462, May 2018. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1505980900.

Balan:2019:SOL

[BS19] Raluca M. Balan and Jian Song. Second order Lyapunov exponents for parabolic and hyperbolic Anderson models. Bernoulli, 25(4A):3069–3089, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1568362052.

Barber:2021:LCL

[BS21] Rina Foygel Barber and Richard J. Samworth. Local continuity of log-concave projection, with applications to estimation under model misspecification. Bernoulli, 27(4):2437–2472, November 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-27/issue-4/Local-continuity-of-log-

concave-projection-with-applications-to-estimation/
10.3150/20-BEJ1316.full.

Behme:2022:MMG

[BS22] Anita Behme and Apostolos Sideris. Markov-modulated generalized Ornstein–Uhlenbeck processes and an application in risk theory. Bernoulli, 28(2):1309–1339, May 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-2/Markov-modulated-generalized-Ornstein--Uhlenbeck-
processes-and-an-application-in/10.3150/21-BEJ1389.
full.

Byun:2023:RNM

[BS23] Sung-Soo Byun and Seong-Mi Seo. Random normal matrices in the almost-circular regime. Bernoulli, 29(2):1615–1637, May 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-29/issue-2/Random-normal-matrices-
in-the-almost-circular-regime/10.3150/22-BEJ1514.full.

Bhaumik:2022:TSB

[BSG22] Prithwish Bhaumik, Wenli Shi, and Subhashis Ghosal. Twostep Bayesian methods for generalized regression driven by partial differential equations. Bernoulli, 28(3):1625–1647, August 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-28/issue-3/Two-step-Bayesian-methods-
for-generalized-regression-driven-by-partial/10.3150/
21-BEJ1363.full.

Bibby:2005:DTM

[BSS05] Bo Martin Bibby, Ib Michael Skovgaard, and Michael Sørensen. Diffusion-type models with given marginal distribution and autocorrelation function. Bernoulli, 11(2):191–220, April 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1116340291.

Bulinski:2012:CLT

[BST12] Alexander Bulinski, Evgeny Spodarev, and Florian Timmermann. Central limit theorems for the excursion set volumes of weakly dependent random fields. Bernoulli, 18(1):100–118, February 2012. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1327068619.

Bolin:2024:GWM

[BSW24] David Bolin, Alexandre B. Simas, and Jonas Wallin. Gaussian Whittle–Matérn fields on metric graphs. Bernoulli, 30(2): 1611–1639, May 2024. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-30/issue-2/Gaussian-WhittleMat%
c3%a9rn-fields-on-metric-graphs/10.3150/23-BEJ1647.
full.

Bai:2017:IDP

[BT17a] Shuyang Bai and Murad S. Taqqu. The impact of the diagonals of polynomial forms on limit theorems with long memory. Bernoulli, 23(1):710–742, February 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1475001372.

Bandyopadhyay:2017:PUS

[BT17b] Antar Bandyopadhyay and Debleena Thacker. Pólya urn schemes with infinitely many colors. Bernoulli, 23(4B):3243– 3267, November 2017. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1495505092.

Bai:2020:LTL

[BT20] Shuyang Bai and Murad S. Taqqu. Limit theorems for longmemory flows on Wiener chaos. Bernoulli, 26(2):1473–1503, May 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1580461587.

Belomestny:2019:SCM

[BTT19] Denis Belomestny, Mathias Trabs, and Alexandre B. Tsybakov. Sparse covariance matrix estimation in high-dimensional deconvolution. Bernoulli, 25(3):1901–1938, August 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1560326432.

Barthelme:2023:DPP

[BTUA23] Simon Barthelmé, Nicolas Tremblay, Konstantin Usevich, and Pierre-Olivier Amblard. Determinantal point processes in the flat limit. Bernoulli, 29(2):957–983, May 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-2/Determinantal-point-processes-in-the-flat-
limit/10.3150/22-BEJ1486.full.

Butucea:2017:STM

[BTV17] C. Butucea, R. Ngueyep Tzoumpe, and P. Vandekerkhove. Semiparametric topographical mixture models with symmetric errors. Bernoulli, 23(2):825–862, May 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1486177385.

Buchsteiner:2018:FIS

[Buc18] Jannis Buchsteiner. The function-indexed sequential empirical process under long-range dependence. Bernoulli, 24(3):2154– 2175, August 2018. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1517540471.

Buhlmann:1997:SBT

[Büh97] Peter Bühlmann. Sieve bootstrap for time series. Bernoulli, 3(2):123–148, June 1997. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1177526726.

Buhlmann:2013:SSH

[Büh13] Peter Bühlmann. Statistical significance in high-dimensional linear models. Bernoulli, 19(4):1212–1242, September 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1377612849.

Bull:2015:ATB

[Bul15] Adam D. Bull. Adaptive-treed bandits. Bernoulli, 21(4): 2289–2307, November 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1438777594.

Butler:2017:AEH

[But17] Ronald W. Butler. Asymptotic expansions and hazard rates for compound and first-passage distributions. Bernoulli, 23 (4B):3508–3536, November 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1495505100.

Butler:2024:ADT

[But24] Ronald W. Butler. Approximate double-transform inversion when time is one of the variables. Bernoulli, 30(4):2878–2903, November 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

Boutsikas:2010:NMO

[BV10] Michael V. Boutsikas and Eutichia Vaggelatou. A new method for obtaining sharp compound Poisson approximation error estimates for sums of locally dependent random variables. Bernoulli, 16(2):301–330, May 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1274821072.

Bansaye:2017:SPC

[BV17] Vincent Bansaye and Vladimir Vatutin. On the survival probability for a class of subcritical branching processes in random environment. Bernoulli, 23(1):58–88, February 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1475001348.

Bhamidi:2013:WDS

[BvdHH13] Shankar Bhamidi, Remco van der Hofstad, and Gerard Hooghiemstra. Weak disorder in the stochastic mean-field model of distance II. Bernoulli, 19(2):363–386, May 2013. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1363192031.

Boufoussi:2007:GBD

[BVM07] Brahim Boufoussi, Jan Van Casteren, and N. Mrhardy. Generalized backward doubly stochastic differential equations and SPDEs with nonlinear Neumann boundary conditions. Bernoulli, 13(2):423–446, May 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1179498755.

Backhoff-Veraguas:2022:AWT

[BVP22] J. Backhoff-Veraguas and G. Pammer. Applications of weak transport theory. Bernoulli, 28(1):370–394, February 2022. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-1/Applications-of-weak-transport-theory/10.
3150/21-BEJ1346.full.

Beirlant:1996:EFE

[BVT96] Jan Beirlant, Petra Vynckier, and Josef L. Teugels. Excess functions and estimation of the extreme-value index. Bernoulli, 2(4):293–318, December 1996. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1178291833.

Bobecka:2008:KTI

[BW08] Konstancja Bobecka and Jacek Wesołowski. Kshirsagar–Tan independence property of beta matrices and related characterizations. Bernoulli, 14(3):749–763, August 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1219669628.

Borovskikh:2010:ADC

[BW10] Yuri V. Borovskikh and N. C. Weber. Asymptotic distributions for a class of generalized L-statistics. Bernoulli, 16(4): 1177–1190, November 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1290092901.

Balabdaoui:2014:CDL

[BW14] Fadoua Balabdaoui and Jon A. Wellner. Chernoff’s density is log-concave. Bernoulli, 20(1):231–244, February 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1390407287.

Butler:2019:LSR

[BW19] Ronald W. Butler and Andrew T. A. Wood. Limiting saddlepoint relative errors in large deviation regions under purely Tauberian conditions. Bernoulli, 25(4B):3379–3399, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1569398769.

Betken:2022:RBC

[BW22] Annika Betken and Martin Wendler. Rank-based changepoint analysis for long-range dependent time series. Bernoulli, 28(4):2209–2233, November 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-28/issue-
4/Rank-based-change-point-analysis-for-long-range-dependent-
time/10.3150/21-BEJ1416.full.

Bai:2023:TPS

[BW23a] Shuyang Bai and Yizao Wang. Tail processes for stableregenerative multiple-stable model. Bernoulli, 29(4):3255–3279, November 2023. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-29/issue-4/Tail-processes-
for-stable-regenerative-multiple-stable-model/10.3150/
22-BEJ1582.full.

Bao:2023:EED

[BW23b] Jianhai Bao and Jian Wang. Exponential ergodicity for damping Hamiltonian dynamics with state-dependent and non-local collisions. Bernoulli, 29(3):2442–2465, August 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-29/issue-3/Exponential-ergodicity-
for-damping-Hamiltonian-dynamics-with-state-dependent-
and/10.3150/22-BEJ1548.full.

Broderick:2018:PCE

[BWJ18] Tamara Broderick, Ashia C. Wilson, and Michael I. Jordan. Posteriors, conjugacy, and exponential families for completely random measures. Bernoulli, 24(4B):3181–3221, November 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1524038752.

Burd:2000:SSI

[BWW00] Gregory A. Burd, Edward C. Waymire, and Ronald D. Winn. A self-similar invariance of critical binary Galton–Watson trees. Bernoulli, 6(1):1–21, February 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1082665377.

Bai:2010:FCS

[BWZ10] Zhidong Bai, Xiaoying Wang, and Wang Zhou. Functional CLT for sample covariance matrices. Bernoulli, 16(4):1086–1113, November 2010. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1290092897.

Barbour:2000:ESC

[BX00] A. D. Barbour and Aihua Xia. Estimating Stein’s constants for compound Poisson approximation. Bernoulli, 6(4):581–590, August 2000. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1081449593.

Barbour:2006:SFP

[BX06] A. D. Barbour and Aihua Xia. On Stein’s factors for Poisson approximation in Wasserstein distance. Bernoulli, 12(6):943–954, December 2006. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1165269145.

Bunea:2015:SCM

[BX15] Florentina Bunea and Luo Xiao. On the sample covariance matrix estimator of reduced effective rank population matrices, with applications to fPCA. Bernoulli, 21(2):1200–1230, May 2015. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1429624975.

Bai:2005:CSE

[BY05] Z. D. Bai and J. Yao. On the convergence of the spectral empirical process of Wigner matrices. Bernoulli, 11(6):1059–1092, December 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1137421640;   http://www.jstor.org/stable/25464779.

Bao:2017:TTS

[BYY17] Jianhai Bao, George Yin, and Chenggui Yuan. Two-time-scale stochastic partial differential equations driven by α-stable noises: Averaging principles. Bernoulli, 23(1):645–669, February 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1475001370.

Beutner:2012:DAD

[BZ12] Eric Beutner and Henryk Zähle. Deriving the asymptotic distribution of U- and V -statistics of dependent data using weighted empirical processes. Bernoulli, 18(3):803–822, August 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1340887003.

Beutner:2014:CMA

[BZ14] Eric Beutner and Henryk Zähle. Continuous mapping approach to the asymptotics of U- and V -statistics. Bernoulli, 20(2):846– 877, May 2014. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1393594008.

Blanchard:2019:CWD

[BZ19] Gilles Blanchard and Oleksandr Zadorozhnyi. Concentration of weakly dependent Banach-valued sums and applications to statistical learning methods. Bernoulli, 25(4B):3421–3458, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1569398772.

Bellec:2022:BLD

[BZ22] Pierre C. Bellec and Cun-Hui Zhang. De-biasing the lasso with degrees-of-freedom adjustment. Bernoulli, 28(2):713– 743, May 2022. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-28/issue-2/De-biasing-the-
lasso-with-degrees-of-freedom-adjustment/10.3150/21-
BEJ1348.full.

Beutner:2023:DRE

[BZ23] Eric Beutner and Henryk Zähle. Donsker results for the empirical process indexed by functions of locally bounded variation and applications to the smoothed empirical process. Bernoulli, 29(1):205–228, February 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-1/Donsker-results-for-the-empirical-process-
indexed-by-functions-of/10.3150/21-BEJ1455.full.

Chotard:2019:VCI

[CA19] Alexandre Chotard and Anne Auger. Verifiable conditions for the irreducibility and aperiodicity of Markov chains by analyzing underlying deterministic models. Bernoulli, 25(1):112– 147, February 2019. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1544605241.

Cherief-Abdellatif:2022:FSP

[CAA22] Badr-Eddine Chérief-Abdellatif and Pierre Alquier. Finite sample properties of parametric MMD estimation: Robustness to misspecification and dependence. Bernoulli, 28(1):181– 213, February 2022. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-28/issue-1/Finite-sample-
properties-of-parametric-MMD-estimation--Robustness-
to/10.3150/21-BEJ1338.full.

Caponera:2024:AIH

[Cap24] Alessia Caponera. Asymptotics for isotropic Hilbert-valued spherical random fields. Bernoulli, 30(3):1723–1745, August 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-30/issue-3/Asymptotics-for-isotropic-
Hilbert-valued-spherical-random-fields/10.3150/23-BEJ1650.
full.

Carter:2006:CGA

[Car06] Andrew V. Carter. A continuous Gaussian approximation to a nonparametric regression in two dimensions. Bernoulli, 12 (1):143–156, February 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1141136654.

Carpentier:2015:TRS

[Car15] Alexandra Carpentier. Testing the regularity of a smooth signal. Bernoulli, 21(1):465–488, February 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1426597078.

Castillo:2007:SPS

[Cas07] I. Castillo. Semi-parametric second-order efficient estimation of the period of a signal. Bernoulli, 13(4):910–932, November 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1194625595.

Castro:2014:ASP

[Cas14] Rui M. Castro. Adaptive sensing performance lower bounds for sparse signal detection and support estimation. Bernoulli, 20 (4):2217–2246, November 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1411134458.

Cator:2011:AOM

[Cat11] Eric Cator. Adaptivity and optimality of the monotone leastsquares estimator. Bernoulli, 17(2):714–735, May 2011. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1302009244.

Cavalier:2001:PLA

[Cav01] Laurent Cavalier. On the problem of local adaptive estimation in tomography. Bernoulli, 7(1):63–78, February 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1080572339.

Chatterji:2022:OLB

[CBL22] Niladri S. Chatterji, Peter L. Bartlett, and Philip M. Long. Oracle lower bounds for stochastic gradient sampling algorithms. Bernoulli, 28(2):1074–1092, May 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-2/Oracle-lower-bounds-for-stochastic-gradient-
sampling-algorithms/10.3150/21-BEJ1377.full.

Chao:2000:SIL

[CC00] Tsung-Ming Chao and Ching-Sung Chou. On some inequalities of local times for Azéma. Bernoulli, 6(3):435–445, June 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1081616699.

Cheng:2004:SDE

[CC04] K. F. Cheng and C. K. Chu. Semiparametric density estimation under a two-sample density ratio model. Bernoulli, 10(4):583– 604, August 2004. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1093265631.

Chowdhury:2019:NDQ

[CC19] Joydeep Chowdhury and Probal Chaudhuri. Nonparametric depth and quantile regression for functional data. Bernoulli, 25(1):395–423, February 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1544605251.

Carpentier:2022:E

[CCC^{+}22] Alexandra Carpentier, Olivier Collier, Laetitia Comminges, Alexandre B. Tsybakov, and Yuhao Wang. Estimation of the. Bernoulli, 28(4):2744–2787, November 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-4/Estimation-of-the-%e2%84%932-norm-and-testing-
in-sparse-linear/10.3150/21-BEJ1436.full.

Cheng:2020:MTL

[CCF^{+}20] Dan Cheng, Valentina Cammarota, Yabebal Fantaye, Domenico Marinucci, and Armin Schwartzman. Multiple testing of local maxima for detection of peaks on the (celestial) sphere. Bernoulli, 26(1):31–60, February 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1574758821.

Calonico:2022:CEO

[CCF22] Sebastian Calonico, Matias D. Cattaneo, and Max H. Farrell. Coverage error optimal confidence intervals for local polynomial regression. Bernoulli, 28(4):2998–3022, November 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-28/issue-4/Coverage-error-optimal-
confidence-intervals-for-local-polynomial-regression/
10.3150/21-BEJ1445.full.

Cheng:2000:SMT

[CCFH00] Ming-Yen Cheng, Edwin Choi, Jianqing Fan, and Peter Hall. Skewing methods for two-parameter locally parametric density estimation. Bernoulli, 6(1):169–182, February 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1082665384.

Cattaneo:2024:BAL

[CCJM24] Matias D. Cattaneo, Rajita Chandak, Michael Jansson, and Xinwei Ma. Boundary adaptive local polynomial conditional density estimators. Bernoulli, 30(4):3193–3223, November 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

Choiruddin:2023:ALD

[CCL23] Achmad Choiruddin, Jean-François Coeurjolly, and Frédérique Letué. Adaptive lasso and Dantzig selector for spatial point processes intensity estimation. Bernoulli, 29(3):1849–1876, August 2023. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-29/issue-3/Adaptive-lasso-
and-Dantzig-selector-for-spatial-point-processes-intensity/
10.3150/22-BEJ1523.full.

Cenac:2021:VLM

[CCN^{+}21] Peggy Cénac, Brigitte Chauvin, Camille Noûs, Frédéric Paccaut, and Nicolas Pouyanne. Variable length memory chains: Characterization of stationary probability measures. Bernoulli, 27(3):2011–2039, August 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-3/Variable-Length-Memory-Chains-Characterization-
of-stationary-probability-measures/10.3150/20-BEJ1299.
full.

Carcamo:2020:DDS

[CCR20] Javier Cárcamo, Antonio Cuevas, and Luis-Alberto Rodríguez. Directional differentiability for supremum-type functionals: Statistical applications. Bernoulli, 26(3):2143–2175, August 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1587974536.

Collier:2020:ENF

[CCT20] O. Collier, L. Comminges, and A. B. Tsybakov. On estimation of nonsmooth functionals of sparse normal means. Bernoulli, 26(3):1989–2020, August 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1587974531.

Chen:2019:PSJ

[CCTY19] Xiaoshan Chen, Zhen-Qing Chen, Ky Tran, and George Yin. Properties of switching jump diffusions: Maximum principles and Harnack inequalities. Bernoulli, 25(2):1045–1075, May 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1551862843.

Cardot:2013:EFE

[CCZ13] Hervé Cardot, Peggy Cénac, and Pierre-André Zitt. Efficient and fast estimation of the geometric median in Hilbert spaces with an averaged stochastic gradient algorithm. Bernoulli, 19 (1):18–43, February 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1358531739.

Cox:1995:HZV

[CD95] J. T. Cox and R. Durrett. Hybrid zones and voter model interfaces. Bernoulli, 1(4):343–370, December 1995. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1193758711.

Caramellino:2001:LIL

[CD01] Lucia Caramellino and Valentina Di Vincenzo. Law of the iterated logarithm for random walks on nilpotent groups. Bernoulli, 7(4):605–628, August 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1079559465.

Coutin:2011:FPT

[CD11] Laure Coutin and Diana Dorobantu. First passage time law for some Lévy processes with compound Poisson: Existence of a density. Bernoulli, 17(4):1127–1135, November 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1320417498.

Canale:2017:PAN

[CD17] Antonio Canale and Pierpaolo De Blasi. Posterior asymptotics of nonparametric location-scale mixtures for multivariate density estimation. Bernoulli, 23(1):379–404, February 2017. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1475001358.

Cator:2021:EBC

[CD21] E. Cator and H. Don. Explicit bounds for critical infection rates and expected extinction times of the contact process on finite random graphs. Bernoulli, 27(3):1556–1582, August 2021. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-3/Explicit-bounds-for-critical-infection-rates-
and-expected-extinction-times/10.3150/20-BEJ1283.full.

Chong:2023:PVF

[CD23a] Carsten Chong and Robert C. Dalang. Power variations in fractional Sobolev spaces for a class of parabolic stochastic PDEs. Bernoulli, 29(3):1792–1820, August 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-3/Power-variations-in-fractional-Sobolev-spaces-
for-a-class-of/10.3150/22-BEJ1521.full.

Cont:2023:QVQ

[CD23b] Rama Cont and Purba Das. Quadratic variation and quadratic roughness. Bernoulli, 29(1):496–522, February 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-29/issue-1/Quadratic-variation-and-
quadratic-roughness/10.3150/22-BEJ1466.full.

Clement:2014:ALB

[CDG14] Emmanuelle Clément, Sylvain Delattre, and Arnaud Gloter. Asymptotic lower bounds in estimating jumps. Bernoulli, 20 (3):1059–1096, August 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1402488934.

Carlstein:1998:MBB

[CDH^{+}98] Edward Carlstein, Kim-Anh Do, Peter Hall, Tim Hesterberg, and Hans R. Künsch. Matched-block bootstrap for dependent data. Bernoulli, 4(3):305–328, September 1998. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1174324983.

Chzhen:2021:MSS

[CDH21] Evgenii Chzhen, Christophe Denis, and Mohamed Hebiri. Minimax semi-supervised set-valued approach to multi-class classification. Bernoulli, 27(4):2389–2412, November 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-4/Minimax-semi-supervised-set-valued-approach-
to-multi-class-classification/10.3150/20-BEJ1313.full.

Cardot:2013:CBH

[CDJ13] Hervé Cardot, David Degras, and Etienne Josserand. Confidence bands for Horvitz–Thompson estimators using sampled noisy functional data. Bernoulli, 19(5A):2067–2097, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1383661214.

Chang:2024:PTM

[CDL24] Yinshan Chang, Hang Du, and Xinyi Li. Percolation threshold for metric graph loop soup. Bernoulli, 30(4):3324–3333, November 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

Cenac:2020:RMP

[CdLOR20] Peggy Cénac, Basile de Loynes, Yoann Offret, and Arnaud Rousselle. Recurrence of multidimensional persistent random walks. Fourier and series criteria. Bernoulli, 26(2):858–892, May 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1580461566.

Cordero-Erausquin:2017:TPW

[CEG17] Dario Cordero-Erausquin and Nathael Gozlan. Transport proofs of weighted Poincaré inequalities for log-concave distributions. Bernoulli, 23(1):134–158, February 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1475001351.

Cekanavicius:2004:LES

[Ček04] Vydas Čekanavičius. On local estimates and the Stein method. Bernoulli, 10(4):665–683, August 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1093265635.

Can:2020:GFT

[CEL20] Sami Umut Can, John H. J. Einmahl, and Roger J. A. Laeven. Goodness-of-fit testing for copulas: a distribution-free approach. Bernoulli, 26(4):3163–3190, November 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
26/issue-4/Goodness-of-fit-testing-for-copulas--A-distribution-
free/10.3150/20-BEJ1219.full.

Chen:2015:EBC

[CF15] Louis H. Y. Chen and Xiao Fang. On the error bound in a combinatorial central limit theorem. Bernoulli, 21(1):335–359, February 2015. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1426597072.

Cattiaux:2016:LTB

[CFKR16] Patrick Cattiaux, Myriam Fradon, Alexei M. Kulik, and Sylvie Roelly. Long time behavior of stochastic hard ball systems. Bernoulli, 22(2):681–710, May 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1447077758.

Cattiaux:1999:LDV

[CG99a] Patrick Cattiaux and Fabrice Gamboa. Large deviations and variational theorems for marginal problems. Bernoulli, 5(1): 81–108, February 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1173707096.

Corcuera:1999:RBC

[CG99b] José M. Corcuera and Federica Giummolè. On the relationship between α connections and the asymptotic properties of predictive distributions. Bernoulli, 5(1):163–176, February 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1173707099.

Corcuera:2006:MP

[CG06] José Manuel Corcuera and Federica Giummolè. Multivariate prediction. Bernoulli, 12(1):157–168, February 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1141136655.

Chen:2021:PRR

[CG21] Qizhao Chen and Christina Goldschmidt. Parking on a random rooted plane tree. Bernoulli, 27(1):93–106, February 2021. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-1/Parking-on-a-random-rooted-plane-tree/10.
3150/20-BEJ1227.full.

Cattiaux:2022:FIP

[CG22a] Patrick Cattiaux and Arnaud Guillin. Functional inequalities for perturbed measures with applications to log-concave measures and to some Bayesian problems. Bernoulli, 28(4):2294– 2321, November 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-28/issue-4/Functional-
inequalities-for-perturbed-measures-with-applications-
to-log-concave/10.3150/21-BEJ1419.full.

Chakraborty:2022:RCM

[CG22b] Moumita Chakraborty and Subhashis Ghosal. Rates and coverage for monotone densities using projection-posterior. Bernoulli, 28(2):1093–1119, May 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-28/issue-
2/Rates-and-coverage-for-monotone-densities-using-projection-
posterior/10.3150/21-BEJ1379.full.

Chang:2023:NSC

[CG23] Jie Chang and Michael Grabchak. Necessary and sufficient conditions for the asymptotic normality of higher order Turing estimators. Bernoulli, 29(4):3369–3395, November 2023. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-29/issue-4/Necessary-and-
sufficient-conditions-for-the-asymptotic-normality-
of-higher/10.3150/23-BEJ1587.full.

Comte:2007:PNM

[CGCR07] Fabienne Comte, Valentine Genon-Catalot, and Yves Rozenholc. Penalized nonparametric mean square estimation of the coefficients of diffusion processes. Bernoulli, 13(2):514–543, May 2007. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1179498759.

Cupidon:2007:DMA

[CGER07] J. Cupidon, D. S. Gilliam, R. Eubank, and F. Ruymgaart. The delta method for analytic functions of random operators with application to functional data. Bernoulli, 13(4):1179–1194, November 2007. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1194625607.

Cleanthous:2020:KWD

[CGK^{+}20] Galatia Cleanthous, Athanasios G. Georgiadis, Gerard Kerkyacharian, Pencho Petrushev, and Dominique Picard. Kernel and wavelet density estimators on manifolds and more general metric spaces. Bernoulli, 26(3):1832–1862, August 2020. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1587974525.

Chen:2012:ESP

[CGL12] Jia Chen, Jiti Gao, and Degui Li. Estimation in semiparametric regression with non-stationary regressors. Bernoulli, 18(2):678–702, May 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1334580729.

Casanellas:2024:IRE

[CGLZ24] Marta Casanellas, Marina Garrote-López, and Piotr Zwiernik. Identifiability in robust estimation of tree structured models. Bernoulli, 30(1):1–21, February 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
30/issue-1/Identifiability-in-robust-estimation-of-
tree-structured-models/10.3150/22-BEJ1477.full.

Cioczek-Georges:1995:SFS

[CGMST95] Renata Cioczek-Georges, Benoît B. Mandelbrot, Gennady Samorodnitsky, and Murad S. Taqqu. Stable fractal sums of pulses: The cylindrical case. Bernoulli, 1(3):201–216, September 1995. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1193667815;   http://www.jstor.org/stable/3318477.

Chatterjee:2018:MEU

[CGS18] Sabyasachi Chatterjee, Adityanand Guntuboyina, and Bodhisattva Sen. On matrix estimation under monotonicity constraints. Bernoulli, 24(2):1072–1100, May 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1505980890.

Chafai:2022:EOD

[CGZJ22] Djalil Chafaï, David García-Zelada, and Paul Jung. At the edge of a one-dimensional jellium. Bernoulli, 28(3):1784– 1809, August 2022. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-28/issue-3/At-the-edge-of-
a-one-dimensional-jellium/10.3150/21-BEJ1397.full.

Cox:2007:UCS

[CH07] A. M. G. Cox and D. G. Hobson. A unifying class of Skorokhod embeddings: connecting the Azéma–Yor and Vallois embeddings. Bernoulli, 13(1):114–130, February 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1175287723.

Cheng:2008:BEB

[CH08] Tsung-Lin Cheng and Hwai-Chung Ho. On Berry–Esseen bounds for non-instantaneous filters of linear processes. Bernoulli, 14(2):301–321, May 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1208872106.

Cloez:2015:EEM

[CH15] Bertrand Cloez and Martin Hairer. Exponential ergodicity for Markov processes with random switching. Bernoulli, 21 (1):505–536, February 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1426597080.

Coupier:2018:CEG

[CH18] David Coupier and Christian Hirsch. Coalescence of Euclidean geodesics on the Poisson–Delaunay triangulation. Bernoulli, 24(4A):2721–2751, November 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1522051223.

Chandler:1997:SME

[Cha97] Richard E. Chandler. A spectral method for estimating parameters in rainfall models. Bernoulli, 3(3):301–322, September 1997. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1177334457.

Chakraborty:1999:MMR

[Cha99a] Biman Chakraborty. On multivariate median regression. Bernoulli, 5(4):683–703, August 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1171899324.

Chamberlin:1999:SSM

[Cha99b] Stephen Raymond Chamberlin. Self-similarity of mathematical likelihood. Bernoulli, 5(4):677–682, August 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1171899323.

Chan:2003:UBI

[Cha03] Hock Peng Chan. Upper bounds and importance sampling of p-values for DNA and protein sequence alignments. Bernoulli, 9(2):183–199, April 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1068128974.

Chauvet:2012:COP

[Cha12] Guillaume Chauvet. On a characterization of ordered pivotal sampling. Bernoulli, 18(4):1320–1340, November 2012. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1352727813.

Chae:2023:ABI

[Cha23] Minwoo Chae. Adaptive Bayesian inference for current status data on a grid. Bernoulli, 29(1):403–427, February 2023. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-1/Adaptive-Bayesian-inference-for-current-
status-data-on-a-grid/10.3150/22-BEJ1462.full.

Cheridito:2001:MFB

[Che01] Patrick Cheridito. Mixed fractional Brownian motion. Bernoulli, 7(6):913–934, December 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1078951129.

Chen:2002:ARS

[Che02] Zehua Chen. Adaptive ranked-set sampling with multiple concomitant variables: an effective way to observational economy. Bernoulli, 8(3):313–322, April 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1078779873.

Chen:2024:ERF

[Che24] Juntong Chen. Estimating a regression function in exponential families by model selection. Bernoulli, 30(2):1669–1693, May 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-30/issue-2/Estimating-a-regression-
function-in-exponential-families-by-model-selection/
10.3150/23-BEJ1649.full.

Campbell:2019:TRM

[CHHB19] Trevor Campbell, Jonathan H. Huggins, Jonathan P. How, and Tamara Broderick. Truncated random measures. Bernoulli, 25(2):1256–1288, May 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1551862850.

Chang:2017:MDA

[CHI17] Chih-Hao Chang, Hsin-Cheng Huang, and Ching-Kang Ing. Mixed domain asymptotics for a stochastic process model with time trend and measurement error. Bernoulli, 23(1):159–190, February 2017. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1475001352.

Chorowski:2018:NVE

[Cho18] Jakub Chorowski. Nonparametric volatility estimation in scalar diffusions: Optimality across observation frequencies. Bernoulli, 24(4A):2934–2990, November 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1522051230.

Cassart:2011:COT

[CHP11] Delphine Cassart, Marc Hallin, and Davy Paindaveine. A class of optimal tests for symmetry based on local Edgeworth approximations. Bernoulli, 17(3):1063–1094, August 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1310042857.

Cheng:2024:SWC

[CHR24] Mengyu Cheng, Zimo Hao, and Michael Röckner. Strong and weak convergence for the averaging principle of DDSDE with singular drift. Bernoulli, 30(2):1586–1610, May 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
30/issue-2/Strong-and-weak-convergence-for-the-averaging-
principle-of-DDSDE/10.3150/23-BEJ1646.full.

Chzhen:2019:LRS

[CHS19] Evgenii Chzhen, Mohamed Hebiri, and Joseph Salmon. On lasso refitting strategies. Bernoulli, 25(4A):3175–3200, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1568362056.

Cheng:1998:ODC

[CHT98] Ming-Yen Cheng, Peter Hall, and D. Michael Titterington. Optimal design for curve estimation by local linear smoothing. Bernoulli, 4(1):3–14, March 1998. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1175865487.

Chang:2005:ATN

[CHWW05] I-Shou Chang, Chao Agnes Hsiung, Mei-Chuan Wang, and Chi-Chung Wen. An asymptotic theory for the nonparametric maximum likelihood estimator in the Cox gene model. Bernoulli, 11(5):863–892, October 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1130077598.

Cazares:2021:RBJ

[CI21] Jorge González Cázares and Jevgenijs Ivanovs. Recovering Brownian and jump parts from high-frequency observations of a Lévy process. Bernoulli, 27(4):2413–2436, November 2021. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-4/Recovering-Brownian-and-jump-parts-from-
high-frequency-observations-of/10.3150/20-BEJ1314.full.

Chafai:2013:ICR

[CJ13] Djalil Chafaï and Aldéric Joulin. Intertwining and commutation relations for birth–death processes. Bernoulli, 19(5A): 1855–1879, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1383661206.

Costa:2022:PTN

[CJ22] Marcelo Costa and Jonathan Jordan. Phase transitions in non-linear urns with interacting types. Bernoulli, 28(4):2546– 2562, November 2022. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-28/issue-4/Phase-transitions-
in-non-linear-urns-with-interacting-types/10.3150/21-
BEJ1428.full.

Chen:2017:RDP

[CJL17] Jinwen Chen, Siqi Jian, and Haitao Li. Representations for the decay parameter of Markov chains. Bernoulli, 23(3):2058–2082, August 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1489737633.

Clemencon:2023:CBE

[CJL^{+}23] Stéphan Clémençon, Hamid Jalalzai, Stéphane Lhaut, Anne Sabourin, and Johan Segers. Concentration bounds for the empirical angular measure with statistical learning applications. Bernoulli, 29(4):2797–2827, November 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-4/Concentration-bounds-for-the-empirical-angular-
measure-with-statistical-learning/10.3150/22-BEJ1562.
full.

Chaudhuri:2021:ULS

[CJP21] Rohit Chaudhuri, Vishesh Jain, and Natesh S. Pillai. Universality and least singular values of random matrix products: a simplified approach. Bernoulli, 27(4):2519–2531, November 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-27/issue-4/Universality-and-least-
singular-values-of-random-matrix-products/10.3150/20-
BEJ1320.full.

Cekanavicius:2000:SPA

[ČK00a] Vydas Čekanavičius and Julius Kruopis. Signed Poisson approximation: a possible alternative to normal and Poisson laws. Bernoulli, 6(4):591–606, August 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1081449594.

Christensen:2000:PDL

[CK00b] Bent Jesper Christensen and Nicholas M. Kiefer. Panel data, local cuts and orthogeodesic models. Bernoulli, 6(4):667–678, August 2000. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1081449600.

Cvitanic:2001:GNP

[CK01] Jaksa Cvitanic and Ioannis Karatzas. Generalized Neyman– Pearson lemma via convex duality. Bernoulli, 7(1):79–97, February 2001. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1080572340.

Cao:2011:SCV

[CK11a] Hongyuan Cao and Michael R. Kosorok. Simultaneous critical values for t-tests in very high dimensions. Bernoulli, 17 (1):347–394, February 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1297173846.

Chambeu:2011:SPS

[CK11b] Sébastien Chambeu and Aline Kurtzmann. Some particular self-interacting diffusions: Ergodic behaviour and almost sure convergence. Bernoulli, 17(4):1248–1267, November 2011. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1320417503.

Chong:2015:ICS

[CK15] Carsten Chong and Claudia Klüppelberg. Integrability conditions for space–time stochastic integrals: Theory and applications. Bernoulli, 21(4):2190–2216, November 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1438777591.

Chakraborty:2019:CES

[CK19] Saptarshi Chakraborty and Kshitij Khare. Consistent estimation of the spectrum of trace class data augmentation algorithms. Bernoulli, 25(4B):3832–3863, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1569398786.

Cammarata:2023:PEP

[CK23] Louis V. Cammarata and Zheng Tracy Ke. Power enhancement and phase transitions for global testing of the mixed membership stochastic block model. Bernoulli, 29(3):1741–1763, August 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-29/issue-3/Power-enhancement-and-phase-
transitions-for-global-testing-of-the/10.3150/22-BEJ1519.
full.

Cazares:2023:JDS

[CKHM23] Jorge Ignacio González Cázares, Arturo Kohatsu-Higa, and Aleksandar Mijatović. Joint density of the stable process and its supremum: Regularity and upper bounds. Bernoulli, 29(4):3443–3469, November 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-29/issue-
4/Joint-density-of-the-stable-process-and-its-supremum/
10.3150/23-BEJ1590.full.

Carpentier:2018:ACS

[CKLN18] Alexandra Carpentier, Olga Klopp, Matthias Löffler, and Richard Nickl. Adaptive confidence sets for matrix completion. Bernoulli, 24(4A):2429–2460, November 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1522051214.

Chigansky:2020:EGB

[CKM20] Pavel Chigansky, Marina Kleptsyna, and Dmytro Marushkevych. On the eigenproblem for Gaussian bridges. Bernoulli, 26(3):1706–1726, August 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1587974521.

Chaudhuri:2002:NEC

[CL02] Probal Chaudhuri and Wei-Yin Loh. Nonparametric estimation of conditional quantiles using quantile regression trees. Bernoulli, 8(5):561–576, October 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1078435218.

Cai:2005:AEL

[CL05] T. Tony Cai and Mark G. Low. Adaptive estimation of linear functionals under different performance measures. Bernoulli, 11(2):341–358, April 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1116340298.

Cator:2012:CLT

[CL12] Eric A. Cator and Hendrik P. Lopuhaä. Central limit theorem and influence function for the MCD estimators at general multivariate distributions. Bernoulli, 18(2):520–551, May 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1334580723.

Chichignoud:2014:RAE

[CL14] Michaël Chichignoud and Johannes Lederer. A robust, adaptive M-estimator for pointwise estimation in heteroscedastic regression. Bernoulli, 20(3):1560–1599, August 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1402488950.

Coeurjolly:2017:PEP

[CL17] Jean-François Coeurjolly and Frédéric Lavancier. Parametric estimation of pairwise Gibbs point processes with infinite range interaction. Bernoulli, 23(2):1299–1334, May 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1486177400.

Chatterjee:2019:ARB

[CL19] Sabyasachi Chatterjee and John Lafferty. Adaptive risk bounds in unimodal regression. Bernoulli, 25(1):1–25, February 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL https://projecteuclid.org/euclid.bj/1544605234.

Chatla:2024:IRS

[CL24] Suneel Babu Chatla and Ruiqi Liu. Inverse regression for spatially distributed functional data. Bernoulli, 30(4):3334–3355, November 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

Cascos:2008:CTP

[CLD08] Ignacio Cascos and Miguel López-Díaz. Consistency of the αtrimming of a probability. Applications to central regions. Bernoulli, 14(2):580–592, May 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1208872119.

Chakar:2017:RAE

[CLLLR17] S. Chakar, E. Lebarbier, C. Lévy-Leduc, and S. Robin. A robust approach for estimating change-points in the mean of an AR(1) process. Bernoulli, 23(2):1408–1447, May 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1486177403.

Chu:2014:SVP

[CLR14] Weijuan Chu, Wenbo V. Li, and Yan-Xia Ren. Small value probabilities for supercritical branching processes with immigration. Bernoulli, 20(1):377–393, February 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1390407293.

Coutinho:2020:MSE

[CLR20] Adriana Coutinho, Rodrigo Lambert, and Jérôme Rousseau. Matching strings in encoded sequences. Bernoulli, 26(3):2021– 2050, August 2020. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1587974532.

Chen:2006:RNP

[CLZ06] Dayue Chen, Juxin Liu, and Fuxi Zhang. The reversible nearest particle system on a finite interval. Bernoulli, 12(1):101–111, February 2006. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1141136651.

Cai:2007:TOB

[CLZ07] T. Tony Cai, Mark G. Low, and Linda H. Zhao. Trade-offs between global and local risks in nonparametric function estimation. Bernoulli, 13(1):1–19, February 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1175287717.

Cao:2018:SNR

[CLZ18] Hongyuan Cao, Weidong Liu, and Zhou Zhou. Simultaneous nonparametric regression analysis of sparse longitudinal data. Bernoulli, 24(4A):3013–3038, November 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1522051232.

Casgrain:2024:STE

[CLZ24] Philippe Casgrain, Martin Larsson, and Johanna Ziegel. Sequential testing for elicitable functionals via supermartingales. Bernoulli, 30(2):1347–1374, May 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
30/issue-2/Sequential-testing-for-elicitable-functionals-
via-supermartingales/10.3150/23-BEJ1634.full.

Csorgo:1999:RDR

[CM99] Sándor Csörgö and Jan Mielniczuk. Random-design regression under long-range dependent errors. Bernoulli, 5(2):209–224, April 1999. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1173147903.

Cont:2011:NTP

[CM11] Rama Cont and Cecilia Mancini. Nonparametric tests for pathwise properties of semimartingales. Bernoulli, 17(2):781–813, May 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1302009247.

Cai:2013:OHT

[CM13a] T. Tony Cai and Zongming Ma. Optimal hypothesis testing for high dimensional covariance matrices. Bernoulli, 19(5B): 2359–2388, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1386078606.

Crambes:2013:APF

[CM13b] Christophe Crambes and André Mas. Asymptotics of prediction in functional linear regression with functional outputs. Bernoulli, 19(5B):2627–2651, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1386078615.

Coeurjolly:2014:VAS

[CM14a] Jean-François Coeurjolly and Jesper Møller. Variational approach for spatial point process intensity estimation. Bernoulli, 20(3):1097–1125, August 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1402488935.

Crisan:2014:PKE

[CM14b] Dan Crisan and Joaquín Míguez. Particle-kernel estimation of the filter density in state-space models. Bernoulli, 20(4): 1879–1929, November 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1411134448.

Crisan:2018:NPF

[CM18] Dan Crisan and Joaquín Míguez. Nested particle filters for online parameter estimation in discrete-time state-space Markov models. Bernoulli, 24(4A):3039–3086, November 2018. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1522051233.

Conti:2020:UPF

[CMMA20] Pier Luigi Conti, Daniela Marella, Fulvia Mecatti, and Federico Andreis. A unified principled framework for resampling based on pseudo-populations: Asymptotic theory. Bernoulli, 26(2): 1044–1069, May 2020. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1580461572.

Comets:2019:RPC

[CMr19] Francis Comets, Gregorio Moreno, and Alejandro F. Ramí rez. Random polymers on the complete graph. Bernoulli, 25(1):683– 711, February 2019. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1544605260.

Can:2010:WCF

[CMS10] Sami Umut Can, Thomas Mikosch, and Gennady Samorodnitsky. Weak convergence of the function-indexed integrated periodogram for infinite variance processes. Bernoulli, 16(4): 995–1015, November 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1290092893.

Chen:2015:OMM

[CN15] Fuqi Chen and Sévérien Nkurunziza. Optimal method in multiple regression with structural changes. Bernoulli, 21(4): 2217–2241, November 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1438777592.

Csorgo:2017:RPM

[CNOH17] Miklós Csörgő, Masoud M. Nasari, and Mohamedou Ould-Haye. Randomized pivots for means of short and long memory linear processes. Bernoulli, 23(4A):2558–2586, November 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL https://projecteuclid.org/euclid.bj/1494316825.

Corcuera:2006:PVS

[CNW06] José Manuel Corcuera, David Nualart, and Jeannette H. C. Woerner. Power variation of some integral fractional processes. Bernoulli, 12(4):713–735, August 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1155735933.

Coeurjolly:2005:IMB

[Coe05] Jean-François Coeurjolly. Identification of multifractional Brownian motion. Bernoulli, 11(6):987–1008, December 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/

1137421637. See erratum [Coe06].

Coeurjolly:2006:EIM

[Coe06] Jean-François Coeurjolly. Erratum: Identification of multifractional Brownian motion. Bernoulli, 12(2):381, April 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/

1145993980. See [Coe05].

Comte:2001:AES

[Com01] Fabienne Comte. Adaptive estimation of the spectrum of a stationary Gaussian sequence. Bernoulli, 7(2):267–298, April 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1080222090.

Corlay:2014:PFQ

[Cor14] Sylvain Corlay. Partial functional quantization and generalized bridges. Bernoulli, 20(2):716–746, May 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1393594004.

Cortines:2016:GSP

[Cor16] Aser Cortines. The genealogy of a solvable population model under selection with dynamics related to directed polymers. Bernoulli, 22(4):2209–2236, November 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1462297680.

Cheng:2001:CIT

[CP01] Shihong Cheng and Liang Peng. Confidence intervals for the tail index. Bernoulli, 7(5):751–760, October 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1079399540.

Crimaldi:2005:CRC

[CP05] Irene Crimaldi and Luca Pratelli. Convergence results for conditional expectations. Bernoulli, 11(4):737–745, August 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1126126767.

Chen:2012:CLE

[CP12] B. B. Chen and G. M. Pan. Convergence of the largest eigenvalue of normalized sample covariance matrices when p and n both tend to infinity with their ratio converging to zero. Bernoulli, 18(4):1405–1420, November 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1352727817.

Chen:2015:CLS

[CP15] Binbin Chen and Guangming Pan. CLT for linear spectral statistics of normalized sample covariance matrices with the dimension much larger than the sample size. Bernoulli, 21 (2):1089–1133, May 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1429624972.

Chaudhuri:2016:NGS

[CP16] Ritwik Chaudhuri and Vladas Pipiras. Non-Gaussian semistable laws arising in sampling of finite point processes. Bernoulli, 22(2):1055–1092, May 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1447077769.

Clinet:2018:SID

[CP18] Simon Clinet and Yoann Potiron. Statistical inference for the doubly stochastic self-exciting process. Bernoulli, 24(4B):3469– 3493, November 2018. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1524038760.

Chakraborty:2019:HRA

[CP19] Anirvan Chakraborty and Victor M. Panaretos. Hybrid regularisation and the (in)admissibility of ridge regression in infinite dimensional Hilbert spaces. Bernoulli, 25(3):1939–1976, August 2019. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1560326433.

Chakraborty:2021:FRL

[CP21] Anirvan Chakraborty and Victor M. Panaretos. Functional registration and local variations: Identifiability, rank, and tuning. Bernoulli, 27(2):1103–1130, May 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-2/Functional-registration-and-local-variations-
Identifiability-rank-and-tuning/10.3150/20-BEJ1267.
full.

Coutin:2023:RCN

[CP23] Laure Coutin and Liliana Peralta. Rates of convergence for the number of zeros of random trigonometric polynomials. Bernoulli, 29(3):1983–2007, August 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-3/Rates-of-convergence-for-the-number-of-zeros-
of-random/10.3150/22-BEJ1528.full.

Caballero:2011:EIL

[CPP11] M. E. Caballero, J. C. Pardo, and J. L. Pérez. Explicit identities for Lévy processes associated to symmetric stable processes. Bernoulli, 17(1):34–59, February 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1297173832.

Chaumont:2013:LRR

[CPR13] Loïc Chaumont, Henry Pantí, and Víctor Rivero. The Lamperti representation of real-valued self-similar Markov processes. Bernoulli, 19(5B):2494–2523, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1386078611.

Chen:2013:PEM

[CPY13] Song X. Chen, Liang Peng, and Cindy L. Yu. Parameter estimation and model testing for Markov processes via conditional characteristic functions. Bernoulli, 19(1):228–251, February 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1358531748.

Cohen:2007:GAM

[CR07] Serge Cohen and Jan Rosinski. Gaussian approximation of multivariate Lévy processes with applications to simulation of tempered stable processes. Bernoulli, 13(1):195–210, February 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1175287729.

Chigansky:2011:VPC

[CR11] Pavel Chigansky and Yaacov Ritov. On the Viterbi process with continuous state space. Bernoulli, 17(2):609–627, May 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1302009239.

Chen:2013:ADR

[CR13] Louis H. Y. Chen and Adrian Röllin. Approximating dependent rare events. Bernoulli, 19(4):1243–1267, September 2013.

CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1377612850.

Conforti:2017:BMR

[CR17] Giovanni Conforti and Sylvie Roelly. Bridge mixtures of random walks on an Abelian group. Bernoulli, 23(3):1518–1537, August 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1489737616.

Conforti:2020:AET

[CR20] Giovanni Conforti and Luigia Ripani. Around the entropic Talagrand inequality. Bernoulli, 26(2):1431–1452, May 2020. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1580461585.

Cerny:2021:PJS

[CR21] Ales Cerný and Johannes Ruf. Pure-jump semimartingales. Bernoulli, 27(4):2624–2648, November 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-4/Pure-jump-semimartingales/10.3150/21-BEJ1325.
full.

Cosso:2022:CLV

[CR22] Andrea Cosso and Francesco Russo. Crandall-Lions viscosity solutions for path-dependent PDEs: The case of heat equation. Bernoulli, 28(1):481–503, February 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-1/CrandallLions-viscosity-solutions-for-path-
dependent-PDEs--The-case/10.3150/21-BEJ1353.full.

Crane:2015:LPP

[Cra15a] Harry Crane. Lipschitz partition processes. Bernoulli, 21 (3):1386–1411, August 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1432732024.

Crane:2015:TVN

[Cra15b] Harry Crane. Time-varying network models. Bernoulli, 21 (3):1670–1696, August 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1432732033.

Csaki:2000:FSF

[CRS00] Endre Csáki, Pál Révész, and Zhan Shi. Favourite sites, favourite values and jump sizes for random walk and Brownian motion. Bernoulli, 6(6):951–975, December 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1081194154.

Chen:2021:EBN

[CRT21] Louis H. Y. Chen, Martin Raic, and Lê Văn Thành. On the error bound in the normal approximation for Jack measures. Bernoulli, 27(1):442–468, February 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-1/On-the-error-bound-in-the-normal-approximation-
for-Jack/10.3150/20-BEJ1245.full.

Chee:2024:RDE

[CRW24] Nicholas Chee, Franz Rembart, and Matthias Winkel. A recursive distributional equation for the stable tree. Bernoulli, 30(2): 1029–1054, May 2024. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-30/issue-2/A-recursive-distributional-
equation-for-the-stable-tree/10.3150/23-BEJ1623.full.

Coquet:1999:CDP

[CS99] François Coquet and Leszek Słomiński. On the convergence of Dirichlet processes. Bernoulli, 5(4):615–639, August 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1171899320.

Csorgo:2006:PSS

[CS06] Sandor Csörgö and Gordon Simons. Pooling strategies for St Petersburg gamblers. Bernoulli, 12(6):971–1002, December 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1165269147.

Chen:2007:NAN

[CS07a] Louis H. Y. Chen and Qi-Man Shao. Normal approximation for nonlinear statistics using a concentration inequality approach. Bernoulli, 13(2):581–599, May 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1179498762.

Christmann:2007:CRK

[CS07b] Andreas Christmann and Ingo Steinwart. Consistency and robustness of kernel-based regression in convex risk minimization. Bernoulli, 13(3):799–819, August 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1186503487.

Cheridito:2013:BBN

[CS13] Patrick Cheridito and Mitja Stadje. BSΔEs and BSDEs with non-Lipschitz drivers: Comparison, convergence and robustness. Bernoulli, 19(3):1047–1085, August 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1372251153.

Chopin:2015:PGS

[CS15] Nicolas Chopin and Sumeetpal S. Singh. On particle Gibbs sampling. Bernoulli, 21(3):1855–1883, August 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1432732040.

Caputo:2018:EPN

[CS18a] Pietro Caputo and Alistair Sinclair. Entropy production in nonlinear recombination models. Bernoulli, 24(4B):3246–3282, November 2018. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1524038754.

Cheng:2018:ENH

[CS18b] Dan Cheng and Armin Schwartzman. Expected number and height distribution of critical points of smooth isotropic Gaussian random fields. Bernoulli, 24(4B):3422–3446, November 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1524038758.

Castillo:2020:SSE

[CS20] Ismaël Castillo and Botond Szabó. Spike and slab empirical Bayes sparse credible sets. Bernoulli, 26(1):127–158, February 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1574758824.

Chen:2024:NSE

[CS24] Zao-Li Chen and Gennady Samrodnitsky. A new shape of extremal clusters for certain stationary semi-exponential processes with moderate long range dependence. Bernoulli, 30(2): 1105–1153, May 2024. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-30/issue-2/A-new-shape-of-
extremal-clusters-for-certain-stationary-semi/10.3150/
23-BEJ1626.full.

Cohen:1995:CTN

[CSSC95] Arthur Cohen, Harold B. Sackrowitz, and Ester Samuel-Cahn. Constructing tests for normal order-restricted inference. Bernoulli, 1(4):321–333, December 1995. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1193758709.

Cosma:2007:MWB

[CSvS07] Antonio Cosma, Olivier Scaillet, and Rainer von Sachs. Multivariate wavelet-based shape-preserving estimation for dependent observations. Bernoulli, 13(2):301–329, May 2007. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1179498750.

Csorgo:1999:SAP

[CSY99] Miklós Csörgö, Zhan Shi, and Marc Yor. Some asymptotic properties of the local time of the uniform empirical process. Bernoulli, 5(6):1035–1058, December 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1143122302.

Comets:2003:DPR

[CSY03] Francis Comets, Tokuzo Shiga, and Nobuo Yoshida. Directed polymers in a random environment: path localization and strong disorder. Bernoulli, 9(4):705–723, August 2003. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1066223275.

Campbell:2023:LE

[CSY^{+}23] Trevor Campbell, Saifuddin Syed, Chiao-Yu Yang, Michael I. Jordan, and Tamara Broderick. Local exchangeability. Bernoulli, 29(3):2084–2100, August 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-29/issue-
3/Local-exchangeability/10.3150/22-BEJ1533.full.

Chan:2002:NET

[CT02] K.-S. Chan and H. Tong. A note on the equivalence of two approaches for specifying a Markov process. Bernoulli, 8(1): 117–122, February 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1078951092.

Chan:2012:ISL

[CT12] Kung-Sik Chan and Henghsiu Tsai. Inference of seasonal longmemory aggregate time series. Bernoulli, 18(4):1448–1464, November 2012. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1352727819.

Castro:2019:TNM

[CT19] Rui M. Castro and Ervin Tánczos. Are there needles in a moving haystack? Adaptive sensing for detection of dynamically evolving signals. Bernoulli, 25(2):977–1012, May 2019. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1551862841.

Cui:2022:UPA

[CT22] Tiangang Cui and Xin T. Tong. A unified performance analysis of likelihood-informed subspace methods. Bernoulli, 28(4):2788–2815, November 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-28/issue-
4/A-unified-performance-analysis-of-likelihood-informed-
subspace-methods/10.3150/21-BEJ1437.full.

Cardona-Tobon:2021:YLC

[CTP21] Natalia Cardona-Tobón and Sandra Palau. Yaglom’s limit for critical Galton–Watson processes in varying environment: a probabilistic approach. Bernoulli, 27(3):1643–1665, August 2021. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-27/issue-3/Yagloms-limit-
for-critical-GaltonWatson-processes-in-varying-environment/
10.3150/20-BEJ1286.full.

Cuny:2015:CLM

[Cun15] Christophe Cuny. A compact LIL for martingales in 2-smooth Banach spaces with applications. Bernoulli, 21(1):374–400, February 2015. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1426597074.

Cutler:2000:CPF

[Cut00] Colleen D. Cutler. Computing pointwise fractal dimension by conditioning in multivariate distributions and time series. Bernoulli, 6(3):381–399, June 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1081616697.

Csorgo:1997:ANL

[CV97] Sándor Csörgö and László Viharos. Asymptotic normality of least-squares estimators of tail indices. Bernoulli, 3(3):351–370, September 1997. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1177334460.

Chen:2009:GFT

[CV09] Song Xi Chen and Ingrid Van Keilegom. A goodness-of-fit test for parametric and semi-parametric models in multiresponse regression. Bernoulli, 15(4):955–976, November 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1262962222.

Cekanavicius:2010:CPS

[ČV10] V. Čekanavičius and P. Vellaisamy. Compound Poisson and signed compound Poisson approximations to the Markov binomial law. Bernoulli, 16(4):1114–1136, November 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1290092898.

Colling:2019:EFN

[CV19] Benjamin Colling and Ingrid Van Keilegom. Estimation of fully nonparametric transformation models. Bernoulli, 25(4B):3762– 3795, November 2019. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1569398784.

Carpentier:2021:OST

[CV21] Alexandra Carpentier and Nicolas Verzelen. Optimal sparsity testing in linear regression model. Bernoulli, 27(2):727–750, May 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-27/issue-2/Optimal-sparsity-testing-
in-linear-regression-model/10.3150/20-BEJ1224.full.

Cardon-Weber:2001:CHS

[CW01] Caroline Cardon-Weber. Cahn–Hilliard stochastic equation: existence of the solution and of its density. Bernoulli, 7(5):777– 816, October 2001. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1079399542.

Chae:2019:BCN

[CW19] Minwoo Chae and Stephen G. Walker. Bayesian consistency for a nonparametric stationary Markov model. Bernoulli, 25 (2):877–901, May 2019. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1551862838.

Cao:2022:TEC

[CW22] Hongyuan Cao and Wei Biao Wu. Testing and estimation for clustered signals. Bernoulli, 28(1):525–547, February 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-28/issue-1/Testing-and-estimation-
for-clustered-signals/10.3150/21-BEJ1355.full.

Chen:2011:PPA

[CX11] Louis H. Y. Chen and Aihua Xia. Poisson process approximation for dependent superposition of point processes. Bernoulli, 17(2):530–544, May 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1302009235.

Cheng:2016:EPG

[CX16] Dan Cheng and Yimin Xiao. Excursion probability of Gaussian random fields on sphere. Bernoulli, 22(2):1113–1130, May 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1447077771.

Cheng:2024:KTS

[CX24] Xiuyuan Cheng and Yao Xie. Kernel two-sample tests for manifold data. Bernoulli, 30(4):2572–2597, November 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

Chiu:2005:PTS

[CY05] Sung Nok Chiu and Chuancun Yin. Passage times for a spectrally negative Lévy process with applications to risk theory. Bernoulli, 11(3):511–522, June 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1120591186.

Chen:2021:HWI

[CY21] Xiaohui Chen and Yun Yang. Hanson–Wright inequality in Hilbert spaces with application to K-means clustering for non-Euclidean data. Bernoulli, 27(1):586–614, February 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-27/issue-1/HansonWright-inequality-
in-Hilbert-spaces-with-application-to-K-means/10.3150/
20-BEJ1251.full.

Chan:2013:MEP

[CZ13] Ngai Hang Chan and Rongmao Zhang. Marked empirical processes for non-stationary time series. Bernoulli, 19(5A): 2098–2119, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1383661215.

Cheng:2014:ESE

[CZH14] Guang Cheng, Lan Zhou, and Jianhua Z. Huang. Efficient semiparametric estimation in generalized partially linear additive models for longitudinal/clustered data. Bernoulli, 20 (1):141–163, February 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1390407283.

Dai:2017:NRS

[Dai17] Hongsheng Dai. A new rejection sampling method without using hat function. Bernoulli, 23(4A):2434–2465, November 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL https://projecteuclid.org/euclid.bj/1494316822.

Daniels:1996:AFC

[Dan96] Henry E. Daniels. Approximating the first crossing-time density for a curved boundary. Bernoulli, 2(2):133–143, June 1996. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1193839220.

Daniels:2000:FCT

[Dan00] Henry E. Daniels. The first crossing-time density for Brownian motion with a perturbed linear boundary. Bernoulli, 6(4):571– 580, August 2000. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1081449592.

Dassios:2005:QBM

[Das05] Angelos Dassios. On the quantiles of Brownian motion and their hitting times. Bernoulli, 11(1):29–36, January 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1110228240.

delBarrio:2022:CBG

[dBCAM22] Eustasio del Barrio, Juan A. Cuesta-Albertos, and Carlos Matran. The complex behaviour of Galton rank-order statistic. Bernoulli, 28(4):2123–2150, November 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-4/The-complex-behaviour-of-Galton-rank-order-
statistic/10.3150/21-BEJ1406.full.

denBoer:2017:CRL

[dBM17] Arnoud V. den Boer and Michel Mandjes. Convergence rates of Laplace-transform based estimators. Bernoulli, 23(4A):2533– 2557, November 2017. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1494316824.

Dhar:2014:CMD

[DCC14] Subhra Sankar Dhar, Biman Chakraborty, and Probal Chaudhuri. Comparison of multivariate distributions using quantile– quantile plots and related tests. Bernoulli, 20(3):1484–1506, August 2014. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1402488947.

deCastro:2023:CQO

[dCCGZ23] Luciano de Castro, Bruno N. Costa, Antonio F. Galvao, and Jorge P. Zubelli. Conditional quantiles: an operator-theoretical approach. Bernoulli, 29(3):2392–2416, August 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-3/Conditional-quantiles-An-operator-theoretical-
approach/10.3150/22-BEJ1546.full.

Dacunha-Castelle:1997:EOM

[DCG97] Didier Dacunha-Castelle and Elisabeth Gassiat. The estimation of the order of a mixture model. Bernoulli, 3(3):279–299, September 1997. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1177334456.

deCooman:2009:ELP

[dCQM09] Gert de Cooman, Erik Quaeghebeur, and Enrique Miranda. Exchangeable lower previsions. Bernoulli, 15(3):721–735, August 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1251463278.

Deleaval:2015:PPP

[DD15] Luc Deleaval and Nizar Demni. Probabilistic proof of product formulas for Bessel functions. Bernoulli, 21(4):2419–2429, November 2015. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1438777599.

Denis:2021:RED

[DDBM21] Christophe Denis, Charlotte Dion-Blanc, and Miguel Martinez. A ridge estimator of the drift from discrete repeated observations of the solution of a stochastic differential equation. Bernoulli, 27(4):2675–2713, November 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-4/A-ridge-estimator-of-the-drift-from-discrete-
repeated-observations/10.3150/21-BEJ1327.full.

Draisma:2004:BTE

[DDFD04] Gerrit Draisma, Holger Drees, Ana Ferreira, and Laurens De Haan. Bivariate tail estimation: dependence in asymptotic independence. Bernoulli, 10(2):251–280, April 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1082380219.

Drees:2015:EFP

[DdH15] Holger Drees and Laurens de Haan. Estimating failure probabilities. Bernoulli, 21(2):957–1001, May 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1429624967.

DelMoral:2012:ARS

[DDJ12] Pierre Del Moral, Arnaud Doucet, and Ajay Jasra. On adaptive resampling strategies for sequential Monte Carlo methods. Bernoulli, 18(1):252–278, February 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1327068625.

Dette:2022:QDS

[DDK22] Holger Dette, Gauthier Dierickx, and Tim Kutta. Quantifying deviations from separability in space–time functional processes. Bernoulli, 28(4):2909–2940, November 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-4/Quantifying-deviations-from-separability-
in-space-time-functional-processes/10.3150/21-BEJ1442.
full.

Duchemin:2023:CIS

[DDL23] Quentin Duchemin, Yohann De Castro, and Claire Lacour. Concentration inequality for U-statistics of order two for uniformly ergodic Markov chains. Bernoulli, 29(2):929– 956, May 2023. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-29/issue-2/Concentration-
inequality-for-U-statistics-of-order-two-for-uniformly/
10.3150/22-BEJ1485.full.

Dehling:2014:ACA

[DDT14] Herold Dehling, Olivier Durieu, and Marco Tusche. Approximating class approach for empirical processes of dependent sequences indexed by functions. Bernoulli, 20(3):1372–1403, August 2014. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1402488943.

Dai:2013:MBD

[DDW13] Bin Dai, Shilin Ding, and Grace Wahba. Multivariate Bernoulli distribution. Bernoulli, 19(4):1465–1483, September 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1377612861.

Deblassie:1999:SID

[Deb99] Richard Dante Deblassie. Scale-invariant diffusions: transience and non-polar points. Bernoulli, 5(4):589–614, August 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1171899319.

Dueck:2014:AID

[DEGR14] Johannes Dueck, Dominic Edelmann, Tilmann Gneiting, and Donald Richards. The affinely invariant distance correlation. Bernoulli, 20(4):2305–2330, November 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1411134461.

Dehay:2005:IDF

[Deh05] Dominique Dehay. On invariant distribution function estimation for continuous-time stationary processes. Bernoulli, 11 (5):933–948, October 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1130077600.

Delyon:2017:CRP

[Del17] Bernard Delyon. Convergence rate of the powers of an operator. Applications to stochastic systems. Bernoulli, 23(4A):2129– 2180, November 2017. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1494316815.

Demni:2007:LPG

[Dem07] Nizar Demni. The Laguerre process and generalized Hartman– Watson law. Bernoulli, 13(2):556–580, May 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1179498761.

Diaconis:1998:CBE

[DF98] Persi W. Diaconis and David Freedman. Consistency of Bayes estimates for nonparametric regression: normal theory. Bernoulli, 4(4):411–444, December 1998. CODEN ????

ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1173883814.

Dombry:2019:MLE

[DF19] Clément Dombry and Ana Ferreira. Maximum likelihood estimators based on the block maxima method. Bernoulli, 25 (3):1690–1723, August 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1560326424.

Dolera:2020:RCF

[DF20] Emanuele Dolera and Stefano Favaro. Rates of convergence in de Finetti’s representation theorem, and Hausdorff moment problem. Bernoulli, 26(2):1294–1322, May 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1580461580.

Denisov:2008:LLE

[DFK08] Denis Denisov, Serguei Foss, and Dmitry Korshunov. On lower limits and equivalences for distribution tails of randomly stopped sums. Bernoulli, 14(2):391–404, May 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1208872110.

Denisov:2010:ARS

[DFK10] Denis Denisov, Serguei Foss, and Dmitry Korshunov. Asymptotics of randomly stopped sums in the presence of heavy tails. Bernoulli, 16(4):971–994, November 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1290092892.

Doring:2016:CTL

[DFK16] Hanna Döring, Gabriel Faraud, and Wolfgang König. Connection times in large ad-hoc mobile networks. Bernoulli, 22(4): 2143–2176, November 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1462297678.

Daouia:2010:FEE

[DFS10] Abdelaati Daouia, Jean-Pierre Florens, and Léopold Simar. Frontier estimation and extreme value theory. Bernoulli, 16 (4):1039–1063, November 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1290092895.

Das:2013:WLE

[DG13] Bikramjit Das and Souvik Ghosh. Weak limits for exploratory plots in the analysis of extremes. Bernoulli, 19(1):308–343, February 2013. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1358531752.

Dutta:2011:SIP

[DGC11] Subhajit Dutta, Anil K. Ghosh, and Probal Chaudhuri. Some intriguing properties of Tukey’s half-space depth. Bernoulli, 17 (4):1420–1434, November 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1320417511.

Daouia:2013:KSE

[DGG13] Abdelaati Daouia, Laurent Gardes, and Stéphane Girard. On kernel smoothing for extremal quantile regression. Bernoulli, 19(5B):2557–2589, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1386078613.

Duarte:2019:EIG

[DGLO19] Aline Duarte, Antonio Galves, Eva Löcherbach, and Guilherme Ost. Estimating the interaction graph of stochastic neural dynamics. Bernoulli, 25(1):771–792, February 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1544605263.

Dombry:2009:DAS

[DGP09] C. Dombry and N. Guillotin-Plantard. Discrete approximation of a stable self-similar stationary increments process. Bernoulli, 15(1):195–222, February 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1233669888.

Decrouez:2018:FSP

[DGP18] Geoffrey Decrouez, Michael Grabchak, and Quentin Paris. Finite sample properties of the mean occupancy counts and probabilities. Bernoulli, 24(3):1910–1941, August 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1517540463.

DiNardo:2008:UFS

[DGS08] Elvira Di Nardo, Giuseppe Guarino, and Domenico Senato. A unifying framework for k-statistics, polykays and their multivariate generalizations. Bernoulli, 14(2):440–468, May 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1208872113.

Daouia:2020:TEP

[DGS20] Abdelaati Daouia, Stéphane Girard, and Gilles Stupfler. Tail expectile process and risk assessment. Bernoulli, 26(1):531– 556, February 2020. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1574758837.

DelBarrio:2005:AFE

[DGU05] Eustasio Del Barrio, Evarist Giné, and Frederic Utzet. Asymptotics for L_{2} functionals of the empirical quantile process, with applications to tests of fit based on weighted Wasserstein distances. Bernoulli, 11(1):131–189, January 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1110228245.

Delattre:2002:AEN

[DH02] Sylvain Delattre and Marc Hoffmann. Asymptotic equivalence for a null recurrent diffusion. Bernoulli, 8(2):139–174, April 2002. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1078866865.

Decrouez:2013:NAS

[DH13] Geoffrey Decrouez and Peter Hall. Normal approximation and smoothness for sums of means of lattice-valued random variables. Bernoulli, 19(4):1268–1293, September 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1377612851.

Deaconu:2017:SHT

[DH17] Madalina Deaconu and Samuel Herrmann. Simulation of hitting times for Bessel processes with non-integer dimension. Bernoulli, 23(4B):3744–3771, November 2017. CODEN ????

ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1495505108.

Doumic:2015:SEG

[DHKR15] Marie Doumic, Marc Hoffmann, Nathalie Krell, and Lydia Robert. Statistical estimation of a growth-fragmentation model observed on a genealogical tree. Bernoulli, 21(3):1760–1799, August 2015. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1432732036.

Dette:2015:CQR

[DHKV15] Holger Dette, Marc Hallin, Tobias Kley, and Stanislav Volgushev. Of copulas, quantiles, ranks and spectra: An L_{1}-approach to spectral analysis. Bernoulli, 21(2):781–831, May 2015. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1429624961.

Dalalyan:2017:PPL

[DHL17] Arnak S. Dalalyan, Mohamed Hebiri, and Johannes Lederer. On the prediction performance of the lasso. Bernoulli, 23 (1):552–581, February 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1475001366.

Diaconis:2013:STW

[Dia13] Persi Diaconis. Some things we’ve learned (about Markov chain Monte Carlo). Bernoulli, 19(4):1294–1305, September 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1377612852.

Diaconu:2023:ECM

[Dia23] Simona Diaconu. On the eigenstructure of covariance matrices with divergent spikes. Bernoulli, 29(2):1275–1296, May 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-29/issue-2/On-the-eigenstructure-of-
covariance-matrices-with-divergent-spikes/10.3150/22-
BEJ1498.full.

Dicker:2016:RRA

[Dic16] Lee H. Dicker. Ridge regression and asymptotic minimax estimation over spheres of growing dimension. Bernoulli, 22(1):1– 37, February 2016. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1443620842.

Dinwoodie:1998:DSA

[Din98] Ian H. Dinwoodie. The Diaconis–Sturmfels algorithm and rules of succession. Bernoulli, 4(3):401–410, September 1998. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1174324986.

Ding:2020:HDD

[Din20] Xiucai Ding. High dimensional deformed rectangular matrices with applications in matrix denoising. Bernoulli, 26(1):387– 417, February 2020. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1574758832.

Diop:2012:CSR

[Dio12] Assane Diop. Convergence of some random functionals of discretized semimartingales. Bernoulli, 18(4):1188–1222, November 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1352727807.

Ditzhaus:2019:SDP

[Dit19] Marc Ditzhaus. Signal detection via phi-divergences for general mixtures. Bernoulli, 25(4A):3041–3068, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL https://projecteuclid.org/euclid.bj/1568362051.

Donoho:1996:NCM

[DJ96] David L. Donoho and Iain M. Johnstone. Neo-classical minimax problems, thresholding and adaptive function estimation. Bernoulli, 2(1):39–62, March 1996. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1193758789.

Delattre:1997:CLT

[DJ97] Sylvain Delattre and Jean Jacod. A central limit theorem for normalized functions of the increments of a diffusion process, in the presence of round-off errors. Bernoulli, 3(1):1–28, February 1997. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1178291930.

DeBlasi:2010:BNE

[DJL10] Pierpaolo De Blasi, Lancelot F. James, and John W. Lau. Bayesian nonparametric estimation and consistency of mixed multinomial logit choice models. Bernoulli, 16(3):679–704, August 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1281099880.

Du:2023:EAI

[DJL23] Kai Du, Yifan Jiang, and Jinfeng Li. Empirical approximation to invariant measures for McKean–Vlasov processes: Meanfield interaction vs self-interaction. Bernoulli, 29(3):2492–2518, August 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-29/issue-3/Empirical-approximation-
to-invariant-measures-for-McKeanVlasov-processes--Mean/
10.3150/22-BEJ1550.full.

Dombry:2018:RTA

[DK18] Clément Dombry and Zakhar Kabluchko. Random tessellations associated with max-stable random fields. Bernoulli, 24(1):30– 52, February 2018. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1501142435.

Doney:2016:PTF

[DKM16] Ron Doney, Claudia Klüppelberg, and Ross Maller. Passage time and fluctuation calculations for subexponential Lévy processes. Bernoulli, 22(3):1491–1519, August 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1458132989.

Dembo:2001:RMC

[DKS01] Amir Dembo, Abram Kagan, and Lawrence A. Shepp. Remarks on the maximum correlation coefficient. Bernoulli, 7(2):343– 350, April 2001. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1080222081.

Dereudre:2009:CEE

[DL09a] David Dereudre and Frédéric Lavancier. Campbell equilibrium equation and pseudo-likelihood estimation for non-hereditary Gibbs point processes. Bernoulli, 15(4):1368–1396, November 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1262962240.

Doukhan:2009:EMR

[DL09b] Paul Doukhan and Gabriel Lang. Evaluation for moments of a ratio with application to regression estimation. Bernoulli, 15 (4):1259–1286, November 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1262962235.

Dubarry:2013:NAD

[DL13] Cyrille Dubarry and Sylvain Le Corff. Non-asymptotic deviation inequalities for smoothed additive functionals in nonlinear state-space models. Bernoulli, 19(5B):2222–2249, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1386078601.

Denisov:2016:LTM

[DL16] Denis Denisov and Nikolai Leonenko. Limit theorems for multifractal products of geometric stationary processes. Bernoulli, 22 (4):2579–2608, November 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1462297690.

Dumont:2017:NRH

[DL17] Thierry Dumont and Sylvain Le Corff. Nonparametric regression on hidden Φ-mixing variables: Identifiability and consistency of a pseudo-likelihood based estimation procedure. Bernoulli, 23(2):990–1021, May 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1486177390.

Das:2019:SOC

[DL19] Debraj Das and S. N. Lahiri. Second order correctness of perturbation bootstrap M-estimator of multiple linear regression parameter. Bernoulli, 25(1):654–682, February 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1544605259.

Diel:2020:LDL

[DLL20] Roland Diel, Sylvain Le Corff, and Matthieu Lerasle. Learning the distribution of latent variables in paired comparison models with round-robin scheduling. Bernoulli, 26(4):2670– 2698, November 2020. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-26/issue-4/Learning-the-distribution-
of-latent-variables-in-paired-comparison-models/10.
3150/20-BEJ1203.full.

Durastanti:2014:GSE

[DLM14] Claudio Durastanti, Xiaohong Lan, and Domenico Marinucci. Gaussian semiparametric estimates on the unit sphere. Bernoulli, 20(1):28–77, February 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1390407280.

Doukhan:2022:MPN

[DLN22] Paul Doukhan, Anne Leucht, and Michael H. Neumann. Mixing properties of non-stationary. Bernoulli, 28(1):663–688, February 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-28/issue-1/Mixing-properties-of-non-
stationary-INGARCH11-processes/10.3150/21-BEJ1362.full.

Dette:2009:AOD

[DLPZ09] Holger Dette, Nikolai Leonenko, Andrey Pepelyshev, and Anatoly Zhigljavsky. Asymptotic optimal designs under longrange dependence error structure. Bernoulli, 15(4):1036–1056, November 2009. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.

bj/1262962225. See correction [DLPZ12].

Dette:2012:CAO

[DLPZ12] Holger Dette, Nikolai Leonenko, Andrey Pepelyshev, and Anatoly Zhigljavsky. Correction to “Asymptotic optimal designs under long-range dependence error structure” Bernoulli 15 (2009) 1036–1056. Bernoulli, 18(2):746, May 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1334580732.  See

[DLPZ09].

DelCastillo:2006:SAE

[DLR06] Joan Del Castillo and Anna López-Ratera. Saddlepoint approximation in exponential models with boundary points. Bernoulli, 12(3):491–500, June 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1151525132.

Duerinckx:2014:MLC

[DLS14] Mitia Duerinckx, Christophe Ley, and Yvik Swan. Maximum likelihood characterization of distributions. Bernoulli, 20(2):775–802, May 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1393594006.

Durot:2009:EJD

[DLT09] C. Durot, E. Lebarbier, and A.-S. Tocquet. Estimating the joint distribution of independent categorical variables via model selection. Bernoulli, 15(2):475–507, May 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1241444899.

Drton:2017:MLM

[DLWZ17] Mathias Drton, Shaowei Lin, Luca Weihs, and Piotr Zwiernik. Marginal likelihood and model selection for Gaussian latent tree and forest models. Bernoulli, 23(2):1202–1232, May 2017. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1486177397.

Douc:2001:AML

[DM01] Randal Douc and Catherine Matias. Asymptotics of the maximum likelihood estimator for general hidden Markov models. Bernoulli, 7(3):381–420, June 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1080004757.

Delaigle:2008:DEH

[DM08a] Aurore Delaigle and Alexander Meister. Density estimation with heteroscedastic error. Bernoulli, 14(2):562–579, May 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1208872118.

Dony:2008:UBC

[DM08b] Julia Dony and David M. Mason. Uniform in bandwidth consistency of conditional U-statistics. Bernoulli, 14(4):1108–1133, November 2008. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1225980573.

Davis:2009:ECE

[DM09] Richard A. Davis and Thomas Mikosch. The extremogram: A correlogram for extreme events. Bernoulli, 15(4):977–1009, November 2009. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1262962223.

Davis:2013:E

[DM13] Richard A. Davis and Thomas Mikosch. Editorial. Bernoulli, 19(4):1087, September 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1377612844.

Durmus:2019:HDB

[DM19] Alain Durmus and Éric Moulines. High-dimensional Bayesian inference via the unadjusted Langevin algorithm. Bernoulli, 25(4A):2854–2882, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1568362045.

Duval:2021:SFE

[DM21] Céline Duval and Ester Mariucci. Spectral-free estimation of Lévy densities in high-frequency regime. Bernoulli, 27(4):2649– 2674, November 2021. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-27/issue-4/Spectral-free-
estimation-of-L%c3%a9vy-densities-in-high-frequency-
regime/10.3150/21-BEJ1326.full.

Derchu:2024:MRB

[DM24] Joffrey Derchu and Thibaut Mastrolia. On Z-mean reflected BSDEs. Bernoulli, 30(2):1502–1524, May 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-

30/issue-2/On-Z-mean-reflected-BSDEs/10.3150/23-BEJ1642.
full.

Dehling:2020:DCD

[DMM^{+}20] Herold Dehling, Muneya Matsui, Thomas Mikosch, Gennady Samorodnitsky, and Laleh Tafakori. Distance covariance for discretized stochastic processes. Bernoulli, 26(4):2758–2789, November 2020. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-26/issue-4/Distance-covariance-
for-discretized-stochastic-processes/10.3150/20-BEJ1206.
full.

Davis:2018:ADC

[DMMW18] Richard A. Davis, Muneya Matsui, Thomas Mikosch, and Phyllis Wan. Applications of distance correlation to time series. Bernoulli, 24(4A):3087–3116, November 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1522051234.

Dedecker:2011:IPL

[DMP11] Jérôme Dedecker, Florence Merlevède, and Magda Peligrad. Invariance principles for linear processes with application to isotonic regression. Bernoulli, 17(1):88–113, February 2011. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1297173834.

Djehiche:2021:FIR

[DMR21] Boualem Djehiche, Othmane Mazhar, and Cristian R. Rojas. Finite impulse response models: a non-asymptotic analysis of the least squares estimator. Bernoulli, 27(2):976–1000, May 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-27/issue-2/Finite-impulse-response-
models--A-non-asymptotic-analysis-of/10.3150/20-BEJ1262.
full.

Douc:2007:CCR

[DMS07] Randal Douc, Eric Moulines, and Philippe Soulier. Computable convergence rates for sub-geometric ergodic Markov chains. Bernoulli, 13(3):831–848, August 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1186503489.

Duarte:2021:DSM

[DMS21] Eliana Duarte, Orlando Marigliano, and Bernd Sturmfels. Discrete statistical models with rational maximum likelihood estimator. Bernoulli, 27(1):135–154, February 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-1/Discrete-statistical-models-with-rational-
maximum-likelihood-estimator/10.3150/20-BEJ1231.full.

Durastanti:2024:FBN

[DMT24] Claudio Durastanti, Domenico Marinucci, and Anna Paola Todino. Flexible-bandwidth needlets. Bernoulli, 30(1):22– 45, February 2024. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-30/issue-1/Flexible-bandwidth-
needlets/10.3150/22-BEJ1513.full.

Donati-Martin:2000:SMV

[DMY00] Catherine Donati-Martin and Marc Yor. Some measure-valued Markov processes attached to occupation times of Brownian motion. Bernoulli, 6(1):63–72, February 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1082665380.

Davydov:2011:SRM

[DMZ11] Youri Davydov, Ilya Molchanov, and Sergei Zuyev. Stability for random measures, point processes and discrete semigroups. Bernoulli, 17(3):1015–1043, August 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1310042854.

Darses:2008:AET

[DN08] Sébastien Darses and Ivan Nourdin. Asymptotic expansions at any time for scalar fractional SDEs with Hurst index H > 1/2. Bernoulli, 14(3):822–837, August 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1219669631.

Deya:2014:IPH

[DN14] Aurélien Deya and Ivan Nourdin. Invariance principles for homogeneous sums of free random variables. Bernoulli, 20(2):586– 603, May 2014. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1393593998.

Drees:2021:ASB

[DN21] Holger Drees and Sebastian Neblung. Asymptotics for sliding blocks estimators of rare events. Bernoulli, 27(2):1239–1269, May 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-27/issue-2/Asymptotics-for-sliding-
blocks-estimators-of-rare-events/10.3150/20-BEJ1272.
full.

Darses:2010:LTN

[DNN10] Sébastien Darses, Ivan Nourdin, and David Nualart. Limit theorems for nonlinear functionals of Volterra processes via white noise analysis. Bernoulli, 16(4):1262–1293, November 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1290092906.

Dette:2006:SNE

[DNP06] Holger Dette, Natalie Neumeyer, and Kay F. Pilz. A simple nonparametric estimator of a strictly monotone regression function. Bernoulli, 12(3):469–490, June 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1151525131.

Davis:2020:SDE

[DNR20] Richard A. Davis, Mikkel Slot Nielsen, and Victor Rohde. Stochastic differential equations with a fractionally filtered delay: a semimartingale model for long-range dependent processes. Bernoulli, 26(2):799–827, May 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1580461564.

Drees:2019:EHT

[DNS19] Holger Drees, Natalie Neumeyer, and Leonie Selk. Estimation and hypotheses testing in boundary regression models. Bernoulli, 25(1):424–463, February 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1544605252.

Dahlhaus:1999:NWE

[DNV99] Rainer Dahlhaus, Michael H. Neumann, and Rainer Von Sachs. Nonlinear wavelet estimation of time-varying autoregressive processes. Bernoulli, 5(5):873–906, October 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1171290403.

Dobra:2003:MBD

[Dob03] Adrian Dobra. Markov bases for decomposable graphical models. Bernoulli, 9(6):1093–1108, December 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1072215202.

Dobler:2023:REP

[Dob23] Dennis Dobler. Randomized empirical processes by algebraic groups, and tests for weak null hypotheses. Bernoulli, 29(2): 1109–1136, May 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-29/issue-2/Randomized-
empirical-processes-by-algebraic-groups-and-tests-for-
weak/10.3150/22-BEJ1492.full.

Dombry:2015:ECM

[Dom15] Clément Dombry. Existence and consistency of the maximum likelihood estimators for the extreme value index within the block maxima framework. Bernoulli, 21(1):420–436, February 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1426597076.

Dorea:1997:SDM

[Dor97] Chang Chung Yu Dorea. Stationary distribution of Markov chains in R^{d} with application to global random optimization. Bernoulli, 3(4):415–427, December 1997. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1175882216.

Dorofeev:1998:UFP

[Dor98] Eugene A. Dorofeev. Upper functions for plane Brownian windings. Bernoulli, 4(4):461–475, December 1998. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1173883816.

Dahlhaus:2009:ESP

[DP09] Rainer Dahlhaus and Wolfgang Polonik. Empirical spectral processes for locally stationary time series. Bernoulli, 15(1):1– 39, February 2009. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1233669881.

Didier:2011:IRP

[DP11] Gustavo Didier and Vladas Pipiras. Integral representations and properties of operator fractional Brownian motions. Bernoulli, 17(1):1–33, February 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1297173831.

Delyon:2016:IAK

[DP16] Bernard Delyon and François Portier. Integral approximation by kernel smoothing. Bernoulli, 22(4):2177–2208, November 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1462297679.

Dobler:2018:GSE

[DP18] Christian Döbler and Giovanni Peccati. The Gamma Stein equation and noncentral de Jong theorems. Bernoulli, 24 (4B):3384–3421, November 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1524038757.

Daouia:2024:OWP

[DPS24] Abdelaati Daouia, Simone A. Padoan, and Gilles Stupfler. Optimal weighted pooling for inference about the tail index and extreme quantiles. Bernoulli, 30(2):1287–1312, May 2024. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
30/issue-2/Optimal-weighted-pooling-for-inference-about-
the-tail-index-and/10.3150/23-BEJ1632.full.

Dette:2010:ODD

[DPSW10] Holger Dette, Andrey Pepelyshev, Piter Shpilev, and Weng Kee Wong. Optimal designs for discriminating between doseresponse models in toxicology studies. Bernoulli, 16(4):1164– 1176, November 2010. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL http://projecteuclid.org/
euclid.bj/1290092900.

DeMasi:2015:ERC

[DPTV15a] Anna De Masi, Errico Presutti, Dimitrios Tsagkarogiannis, and Maria Eulalia Vares. Exponential rate of convergence in current reservoirs. Bernoulli, 21(3):1844–1854, August 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1432732039.

DeMasi:2015:ETR

[DPTV15b] Anna De Masi, Errico Presutti, Dimitrios Tsagkarogiannis, and Maria Eulalia Vares. Extinction time for a random walk in a random environment. Bernoulli, 21(3):1824–1843, August 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1432732038.

Dette:2009:ODD

[DPW09] Holger Dette, Andrey Pepelyshev, and Weng Kee Wong. Optimal designs for dose-finding experiments in toxicity studies. Bernoulli, 15(1):124–145, February 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1233669885.

Ding:2023:LMS

[DPX23] Yujia Ding, Qidi Peng, and Yimin Xiao. Linear multifractional stable sheets in the broad sense: Existence and joint continuity of local times. Bernoulli, 29(1):785–814, February 2023. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-1/Linear-multifractional-stable-sheets-in-
the-broad-sense--Existence/10.3150/22-BEJ1479.full.

Dahlhaus:2007:ROA

[DR07] Rainer Dahlhaus and Suhasini Subba Rao. A recursive online algorithm for the estimation of time-varying ARCH parameters. Bernoulli, 13(2):389–422, May 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1179498754.

Dumbgen:2009:MLE

[DR09] Lutz Dümbgen and Kaspar Rufibach. Maximum likelihood estimation of a log-concave density and its distribution function:

Basic properties and uniform consistency. Bernoulli, 15(1):40– 68, February 2009. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1233669882.

Das:2011:CEC

[DR11] Bikramjit Das and Sidney I. Resnick. Conditioning on an extreme component: Model consistency with regular variation on cones. Bernoulli, 17(1):226–252, February 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1297173841.

Delattre:2016:EDF

[DR16] Sylvain Delattre and Etienne Roquain. On empirical distribution function of high-dimensional Gaussian vector components with an application to multiple testing. Bernoulli, 22 (1):302–324, February 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1443620851.

Dolera:2019:URG

[DR19] Emanuele Dolera and Eugenio Regazzini. Uniform rates of the Glivenko-Cantelli convergence and their use in approximating Bayesian inferences. Bernoulli, 25(4A):2982–3015, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1568362049.

Deo:2023:ACS

[DR23] Neil Deo and Thibault Randrianarisoa. On adaptive confidence sets for the Wasserstein distances. Bernoulli, 29(3):2119–2141, August 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-29/issue-3/On-adaptive-confidence-
sets-for-the-Wasserstein-distances/10.3150/22-BEJ1535.
full.

Dalalyan:2020:SLC

[DRD20] Arnak S. Dalalyan and Lionel Riou-Durand. On sampling from a log-concave density using kinetic Langevin diffusions. Bernoulli, 26(3):1956–1988, August 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1587974530.

Drees:2003:EQE

[Dre03] Holger Drees. Extreme quantile estimation for dependent data, with applications to finance. Bernoulli, 9(4):617–657, August 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1066223272.

Donnet:2018:PCR

[DRRS18] Sophie Donnet, Vincent Rivoirard, Judith Rousseau, and Catia Scricciolo. Posterior concentration rates for empirical Bayes procedures with applications to Dirichlet process mixtures. Bernoulli, 24(1):231–256, February 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1501142441.

Drton:2009:DCG

[Drt09] Mathias Drton. Discrete chain graph models. Bernoulli, 15 (3):736–753, August 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1251463279.

Dattner:2016:AQE

[DRT16] Itai Dattner, Markus Reiß, and Mathias Trabs. Adaptive quantile estimation in deconvolution with unknown error distribution. Bernoulli, 22(1):143–192, February 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1443620846.

Dahlhaus:2019:TGT

[DRW19] Rainer Dahlhaus, Stefan Richter, and Wei Biao Wu. Towards a general theory for nonlinear locally stationary processes. Bernoulli, 25(2):1013–1044, May 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1551862842.

Drton:2020:NCD

[DRW20] Mathias Drton, Elina Robeva, and Luca Weihs. Nested covariance determinants and restricted trek separation in Gaussian graphical models. Bernoulli, 26(4):2503–2540, November 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-26/issue-4/Nested-covariance-determinants-

and-restricted-trek-separation-in-Gaussian-graphical/
10.3150/19-BEJ1179.full.

Delbaen:1996:VOM

[DS96] Freddy Delbaen and Walter Schachermayer. The varianceoptimal martingale measure for continuous processes. Bernoulli, 2(1):81–105, March 1996. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1193758791.

DeGunst:2003:MSH

[DS03] Mathisca C. M. De Gunst and Barry Schouten. Model selection for hidden Markov models of ion channel data by reversible jump Markov chain Monte Carlo. Bernoulli, 9(3):373– 393, June 2003. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1065444810.

Decreusefond:2005:FBM

[DS05] Laurent Decreusefond and Nicholas Savy. Filtered Brownian motions as weak limit of filtered Poisson processes. Bernoulli, 11(2):283–292, April 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1116340295.

Deblassie:2006:BMS

[DS06] Dante Deblassie and Robert Smits. Brownian motion in selfsimilar domains. Bernoulli, 12(1):113–132, February 2006. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1141136652.

Deya:2019:IRN

[DS19] Aurélien Deya and René Schott. Integration with respect to the non-commutative fractional Brownian motion. Bernoulli, 25(3):2137–2162, August 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1560326440.

daSilva:2023:FMV

[dSJMT23] Poly H. da Silva, Arash Jamshidpey, Peter McCullagh, and Simon Tavaré. Fisher’s measure of variability in repeated samples. Bernoulli, 29(2):1166–1194, May 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-2/Fishers-measure-of-variability-in-repeated-
samples/10.3150/22-BEJ1494.full.

Dalang:2010:CHP

[DSS10] Robert C. Dalang and Marta Sanz-Solé. Criteria for hitting probabilities with applications to systems of stochastic wave equations. Bernoulli, 16(4):1343–1368, November 2010. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1290092909.

Dumbgen:2021:BDE

[DSW21] Lutz Dümbgen, Richard J. Samworth, and Jon A. Wellner. Bounding distributional errors via density ratios. Bernoulli, 27 (2):818–852, May 2021. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-27/issue-2/Bounding-distributional-
errors-via-density-ratios/10.3150/20-BEJ1256.full.

Dalalyan:2012:MAS

[DT12] Arnak S. Dalalyan and Alexandre B. Tsybakov. Mirror averaging with sparsity priors. Bernoulli, 18(3):914–944, August 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1340887008.

Debaly:2023:MTS

[DT23a] Zinsou-Max Debaly and Lionel Truquet. Multivariate time series models for mixed data. Bernoulli, 29(1):669–695, February 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-29/issue-1/Multivariate-time-series-
models-for-mixed-data/10.3150/22-BEJ1474.full.

Du:2023:HTE

[DT23b] Xinjie Du and Minh Tang. Hypothesis testing for equality of latent positions in random graphs. Bernoulli, 29(4):3221– 3254, November 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-29/issue-4/Hypothesis-
testing-for-equality-of-latent-positions-in-random-
graphs/10.3150/22-BEJ1581.full.

deUna-Alvarez:2021:EPI

[dUÁV21] Jacobo de Uña-Álvarez and Ingrid Van Keilegom. Efron– Petrosian integrals for doubly truncated data with covariates: An asymptotic analysis. Bernoulli, 27(1):249–273, February 2021. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-27/issue-1/EfronPetrosian-
integrals-for-doubly-truncated-data-with-covariates-
-An/10.3150/20-BEJ1236.full.

Dufresne:1996:SEX

[Duf96] Daniel Dufresne. On the stochastic equation L(X) = L[B(X + C)] and a property of gamma distributions. Bernoulli, 2(3): 287–291, September 1996. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1178291724.

Dumbgen:2003:OCB

[Düm03] Lutz Dümbgen. Optimal confidence bands for shape-restricted curves. Bernoulli, 9(3):423–449, June 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1065444812.

Dawid:1999:PPP

[DV99] A. Philip Dawid and Vladimir G. Vovk. Prequential probability: principles and properties. Bernoulli, 5(1):125–162, February 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1173707098.

Dette:2001:TAK

[DV01] Holger Dette and Carsten Von Lieres Und Wilkau. Testing additivity by kernel-based methods — what is a reasonable test? Bernoulli, 7(4):669–697, August 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1079559469.

Hernandez-del-Valle:2013:HTB

[dV13] Gerardo Hernandez del Valle. On hitting times, Bessel bridges and Schrödinger’s equation. Bernoulli, 19(5A):1559–1575, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1383661194.

DelMoral:2018:EMP

[DV18] Pierre Del Moral and Denis Villemonais. Exponential mixing properties for time inhomogeneous diffusion processes with killing. Bernoulli, 24(2):1010–1032, May 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1505980887.

Drost:2009:ASN

[DvdAW09] Feike C. Drost, Ramon van den Akker, and Bas J. M. Werker. The asymptotic structure of nearly unstable nonnegative integer-valued AR(1) models. Bernoulli, 15(2):297– 324, May 2009. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1241444892.

Delgado-Vences:2014:ASW

[DVSS14] Francisco J. Delgado-Vences and Marta Sanz-Solé. Approximation of a stochastic wave equation in dimension three, with application to a support theorem in Hölder norm. Bernoulli, 20 (4):2169–2216, November 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1411134457.

Delgado-Vences:2016:ASW

[DVSS16] Francisco J. Delgado-Vences and Marta Sanz-Solé. Approximation of a stochastic wave equation in dimension three, with application to a support theorem in Hölder norm: The nonstationary case. Bernoulli, 22(3):1572–1597, August 2016. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1458132992.

Durieu:2019:RPF

[DW19] Olivier Durieu and Yizao Wang. From random partitions to fractional Brownian sheets. Bernoulli, 25(2):1412–1450, May 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1551862855.

Doring:2020:SPC

[DW20] Leif Döring and Philip Weissmann. Stable processes conditioned to hit an interval continuously from the outside. Bernoulli, 26(2):980–1015, May 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1580461570.

Dhar:2023:CTV

[DW23] Subhra Sankar Dhar and Weichi Wu. Comparing time varying regression quantiles under shift invariance. Bernoulli, 29(2): 1527–1554, May 2023. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-29/issue-2/Comparing-time-
varying-regression-quantiles-under-shift-invariance/
10.3150/22-BEJ1509.full.

Drton:2016:WTS

[DX16] Mathias Drton and Han Xiao. Wald tests of singular hypotheses. Bernoulli, 22(1):38–59, February 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1443620843.

Dieker:2014:ACT

[DY14] A. B. Dieker and B. Yakir. On asymptotic constants in the theory of extremes for Gaussian processes. Bernoulli, 20(3): 1600–1619, August 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1402488951.

DelMoral:2003:NLV

[DZ03] P. Del Moral and T. Zajic. A note on the Laplace–Varadhan integral lemma. Bernoulli, 9(1):49–65, February 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1068129010.

Draisma:2017:AGG

[DZ17] Jan Draisma and Piotr Zwiernik. Automorphism groups of Gaussian Bayesian networks. Bernoulli, 23(2):1102–1129, May 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1486177394.

Dumitriu:2024:ESV

[DZ24] Ioana Dumitriu and Yizhe Zhu. Extreme singular values of inhomogeneous sparse random rectangular matrices. Bernoulli, 30(4):2904–2931, November 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

Eftekhari:2023:DCO

[EBR23] Hamid Eftekhari, Moulinath Banerjee, and Ya’acov Ritov. Design of c-optimal experiments for high-dimensional linear models. Bernoulli, 29(1):652–668, February 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-1/Design-of-c-optimal-experiments-for-high-
dimensional-linear-models/10.3150/22-BEJ1472.full.

Eaton:2004:DBA

[EF04] Morris L. Eaton and David A. Freedman. Dutch book against some ‘objective’ priors. Bernoulli, 10(5):861–872, October 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1099579159.

Efromovich:1998:GPA

[Efr98] Sam Efromovich. On global and pointwise adaptive estimation. Bernoulli, 4(2):273–282, June 1998. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1174937296.

Efromovich:2000:CAE

[Efr00] Sam Efromovich. Can adaptive estimators for Fourier series be of interest to wavelets? Bernoulli, 6(4):699–708, August 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1081449602.

Einmahl:2008:SEU

[EFVG08] John H. J. Einmahl, Amélie Fils-Villetard, and Armelle Guillou. Statistics of extremes under random censoring. Bernoulli, 14(1):207–227, February 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1202492791.

Ehm:2003:SCA

[EGG03] Werner Ehm, Marc G. Genton, and Tilmann Gneiting. Stationary covariances associated with exponentially convex functions. Bernoulli, 9(4):607–615, August 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1066223271.

ElBantli:2002:EIQ

[EH02] Faouzi El Bantli and Marc Hallin. Estimation of the innovation quantile density function of an AR(p) process based on autoregression quantiles. Bernoulli, 8(2):255–274, April 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1078866870.

Eltzner:2023:DMG

[EHHS23] Benjamin Eltzner, Pernille E. H. Hansen, Stephan F. Huckemann, and Stefan Sommer. Diffusion means in geometric spaces. Bernoulli, 29(4):3141–3170, November 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-29/issue-4/Diffusion-
means-in-geometric-spaces/10.3150/22-BEJ1578.full.

Evans:2005:MEM

[EHS05] Steven N. Evans, Ben B. Hansen, and Philip B. Stark. Minimax expected measure confidence sets for restricted location parameters. Bernoulli, 11(4):571–590, August 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1126126761.

Estrade:2010:BTS

[EI10] Anne Estrade and Jacques Istas. Ball throwing on spheres. Bernoulli, 16(4):953–970, November 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1290092891.

Esmaeeli:2018:AOA

[EI18] Neda Esmaeeli and Peter Imkeller. American options with asymmetric information and reflected BSDE. Bernoulli, 24(2): 1394–1426, May 2018. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1505980899.

Eisenbaum:2000:LTS

[Eis00] Nathalie Eisenbaum. On local times of a symmetric stable process as a doubly indexed process. Bernoulli, 6(5):871–886, October 2000. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1081282693.

Eberlein:1995:HDF

[EK95] Ernst Eberlein and Ulrich Keller. Hyperbolic distributions in finance. Bernoulli, 1(3):281–299, September 1995. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1193667819.

Einmahl:2011:CLT

[EK11] John H. J. Einmahl and Estáte V. Khmaladze. Central limit theorems for local empirical processes near boundaries of sets. Bernoulli, 17(2):545–561, May 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1302009236.

Eichinger:2018:MPE

[EK18] Birte Eichinger and Claudia Kirch. A MOSUM procedure for the estimation of multiple random change points. Bernoulli, 24(1):526–564, February 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1501142454.

Einmahl:2008:MME

[EKS08] John H. J. Einmahl, Andrea Krajina, and Johan Segers. A method of moments estimator of tail dependence. Bernoulli, 14 (4):1003–1026, November 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1225980569.

Engelke:2015:MIN

[EKS15] Sebastian Engelke, Zakhar Kabluchko, and Martin Schlather. Maxima of independent, non-identically distributed Gaussian vectors. Bernoulli, 21(1):38–61, February 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1426597063.

Erny:2022:MFL

[ELL22] Xavier Erny, Eva Löcherbach, and Dasha Loukianova. Mean field limits for interacting Hawkes processes in a diffusive regime. Bernoulli, 28(1):125–149, February 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-1/Mean-field-limits-for-interacting-Hawkes-
processes-in-a-diffusive/10.3150/21-BEJ1335.full.

Nmeir:2022:NAG

[ELP22] Rancy El Nmeir, Harald Luschgy, and Gilles Pagès. New approach to greedy vector quantization. Bernoulli, 28(1):424– 452, February 2022. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-28/issue-1/New-approach-to-
greedy-vector-quantization/10.3150/21-BEJ1350.full.

ElIdrissi:2024:VES

[ELP24] Yazid Janati El Idrissi, Sylvain Le Corff, and Yohan Petetin. Variance estimation for sequential Monte Carlo algorithms: a backward sampling approach. Bernoulli, 30(2):911–935, May 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-30/issue-2/Variance-estimation-for-
sequential-Monte-Carlo-algorithms--A-backward/10.3150/
23-BEJ1586.full.

Eltzner:2022:GSN

[Elt22] Benjamin Eltzner. Geometrical smeariness - a new phenomenon of Fréchet means. Bernoulli, 28(1):239–254, February 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-28/issue-1/Geometrical-smeariness-
-a-new-phenomenon-of-Fr%c3%a9chet-means/10.3150/21-
BEJ1340.full.

Einmahl:2003:ELB

[EM03] John H. J. Einmahl and Ian W. McKeague. Empirical likelihood based hypothesis testing. Bernoulli, 9(2):267–290, April 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1068128978.

ElBarmi:2013:ELB

[EM13] Hammou El Barmi and Ian W. McKeague. Empirical likelihood-based tests for stochastic ordering. Bernoulli, 19 (1):295–307, February 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1358531751.

Eguchi:2018:STM

[EM18] Shoichi Eguchi and Hiroki Masuda. Schwarz type model comparison for LAQ models. Bernoulli, 24(3):2278–2327, August 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1517540475.

Enikeeva:2020:BDP

[EMPW20] Farida Enikeeva, Axel Munk, Markus Pohlmann, and Frank Werner. Bump detection in the presence of dependency: Does it ease or does it load? Bernoulli, 26(4):3280–3310, November 2020. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-26/issue-4/Bump-detection-
in-the-presence-of-dependency--Does-it/10.3150/20-BEJ1226.
full.

Enikeeva:2018:BDH

[EMW18] Farida Enikeeva, Axel Munk, and Frank Werner. Bump detection in heterogeneous Gaussian regression. Bernoulli, 24(2): 1266–1306, May 2018. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1505980896.

ElMachkouri:2007:ECR

[EO07] M. El Machkouri and L. Ouchti. Exact convergence rates in the central limit theorem for a class of martingales. Bernoulli, 13(4):981–999, November 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1194625599.

Evans:2019:SIS

[ER19] Robin J. Evans and Thomas S. Richardson. Smooth, identifiable supermodels of discrete DAG models with latent variables. Bernoulli, 25(2):848–876, May 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1551862837.

Eichelsbacher:2023:KBD

[ER23] Peter Eichelsbacher and Benedikt Rednoß. Kolmogorov bounds for decomposable random variables and subgraph counting by the Stein-tikhomirov method. Bernoulli, 29(3):1821–1848, August 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-29/issue-3/Kolmogorov-
bounds-for-decomposable-random-variables-and-subgraph-
counting-by/10.3150/22-BEJ1522.full.

Erhardsson:2014:CCR

[Erh14] Torkel Erhardsson. Conditions for convergence of random coefficient AR(1) processes and perpetuities in higher dimensions. Bernoulli, 20(2):990–1005, May 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1393594013.

Evans:2019:LFF

[ERS19] Steven N. Evans, Ronald L. Rivest, and Philip B. Stark. Leading the field: Fortune favors the bold in Thurstonian choice models. Bernoulli, 25(1):26–46, February 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1544605237.

Ernst:2020:FOC

[ERS20] Marie Ernst, Gesine Reinert, and Yvik Swan. First-order covariance inequalities via Stein’s method. Bernoulli, 26(3):2051– 2081, August 2020. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1587974533.

Eaton:1999:CSI

[ES99a] Morris L. Eaton and William D. Sudderth. Consistency and strong inconsistency of group-invariant predictive inferences. Bernoulli, 5(5):833–854, October 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1171290401.

Eisenbaum:1999:UOL

[ES99b] Nathalie Eisenbaum and Zhan Shi. Uniform oscillations of the local time of iterated Brownian motion. Bernoulli, 5(1):49–65, February 1999. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1173707094.

Eyjolfsson:2023:MSE

[ET23] Heidar Eyjolfsson and Dag Tjøstheim. Multivariate selfexciting jump processes with applications to financial data. Bernoulli, 29(3):2167–2191, August 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-29/issue-
3/Multivariate-self-exciting-jump-processes-with-applications-
to-financial-data/10.3150/22-BEJ1537.full.

Erdos:2023:SDE

[EX23] László Erdős and Yuanyuan Xu. Small deviation estimates for the largest eigenvalue of Wigner matrices. Bernoulli, 29(2): 1063–1079, May 2023. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-29/issue-2/Small-deviation-
estimates-for-the-largest-eigenvalue-of-Wigner-matrices/
10.3150/22-BEJ1490.full.

Eichelsbacher:2003:MDM

[EZ03] Peter Eichelsbacher and Tim Zajic. Moderate deviations for mean-field Gibbs measures. Bernoulli, 9(1):67–95, February 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1068129011.

Fang:2014:DNA

[Fan14] Xiao Fang. Discretized normal approximation by Stein’s method. Bernoulli, 20(3):1404–1431, August 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1402488944.

Fan:2021:SPG

[Fan21] Wai-Tong Louis Fan. Stochastic PDEs on graphs as scaling limits of discrete interacting systems. Bernoulli, 27(3):1899– 1941, August 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-27/issue-3/Stochastic-
PDEs-on-graphs-as-scaling-limits-of-discrete-interacting/
10.3150/20-BEJ1296.full.

Fasen:2010:ARS

[Fas10] Vicky Fasen. Asymptotic results for sample autocovariance functions and extremes of integrated generalized Ornstein– Uhlenbeck processes. Bernoulli, 16(1):51–79, February 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1265984704.

Foucart:2014:LEC

[FB14] Clément Foucart and Gerónimo Uribe Bravo. Local extinction in continuous-state branching processes with immigration. Bernoulli, 20(4):1819–1844, November 2014. CODEN ????

ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1411134446.

Fertl:2022:CVE

[FB22a] Lukas Fertl and Efstathia Bura. Conditional variance estimator for sufficient dimension reduction. Bernoulli, 28(3):1862–1891, August 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-28/issue-3/Conditional-variance-estimator-
for-sufficient-dimension-reduction/10.3150/21-BEJ1402.
full.

Fontaine:2022:AMT

[FB22b] Simon Fontaine and Mylène Bédard. An adaptive multipletry Metropolis algorithm. Bernoulli, 28(3):1986–2011, August 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-28/issue-3/An-adaptive-multiple-try-
Metropolis-algorithm/10.3150/21-BEJ1408.full.

Feng:2015:BQR

[FCH15] Yang Feng, Yuguo Chen, and Xuming He. Bayesian quantile regression with approximate likelihood. Bernoulli, 21(2):832– 850, May 2015. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1429624962.

Ferreira:2014:GPP

[FdH14] Ana Ferreira and Laurens de Haan. The generalized Pareto process; with a view towards application and simulation. Bernoulli, 20(4):1717–1737, November 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1411134442.

Ferger:2001:ACP

[Fer01] Dietmar Ferger. Analysis of change-point estimators under the null hypothesis. Bernoulli, 7(3):487–506, June 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1080004762.

Favaro:2023:DSN

[FFP23] Stefano Favaro, Sandra Fortini, and Stefano Peluchetti. Deep stable neural networks: Large-width asymptotics and convergence rates. Bernoulli, 29(3):2574–2597, August 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-3/Deep-stable-neural-networks--Large-width-
asymptotics-and-convergence/10.3150/22-BEJ1553.full.

Fraser:2010:SOA

[FFS10] Ailana M. Fraser, D. A. S. Fraser, and Ana-Maria Staicu. Second order ancillary: A differential view from continuity. Bernoulli, 16(4):1208–1223, November 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1290092903.

Fontes:2022:ABL

[FFZ22] Luiz Renato Fontes, Susana Frómeta, and Leonel Zuaznábar. Asymptotic behavior of a low-temperature non-cascading 2-GREM dynamics at extreme time scales. Bernoulli, 28(4):2716– 2743, November 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-28/issue-4/Asymptotic-
behavior-of-a-low-temperature-non-cascading-2-GREM/
10.3150/21-BEJ1435.full.

Fang:2019:RCC

[FG19] Billy Fang and Adityanand Guntuboyina. On the risk of convex-constrained least squares estimators under misspecification. Bernoulli, 25(3):2206–2244, August 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1560326443.

Fukker:2016:ABG

[FGK16] Gábor Fukker, László Györfi, and Péter Kevei. Asymptotic behavior of the generalized St. Petersburg sum conditioned on its maximum. Bernoulli, 22(2):1026–1054, May 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1447077768.

Fan:2019:SNC

[FGLS19] Xiequan Fan, Ion Grama, Quansheng Liu, and Qi-Man Shao. Self-normalized Cramér type moderate deviations for martingales. Bernoulli, 25(4A):2793–2823, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1568362043.

Fontes:2021:CPU

[FGS21] Luiz Renato Fontes, Pablo Almeida Gomes, and Remy Sanchis. Contact process under heavy-tailed renewals on finite graphs. Bernoulli, 27(3):1745–1763, August 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-3/Contact-process-under-heavy-tailed-renewals-
on-finite-graphs/10.3150/20-BEJ1290.full.

Franceschini:2022:SIP

[FGS22] Chiara Franceschini, Patrícia Gonçalves, and Federico Sau. Symmetric inclusion process with slow boundary: Hydrodynamics and hydrostatics. Bernoulli, 28(2):1340–1381, May 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-28/issue-2/Symmetric-
inclusion-process-with-slow-boundary-Hydrodynamics-
and-hydrostatics/10.3150/21-BEJ1390.full.

Favaro:2012:CMV

[FGW12] Stefano Favaro, Alessandra Guglielmi, and Stephen G. Walker. A class of measure-valued Markov chains and Bayesian nonparametrics. Bernoulli, 18(3):1002–1030, August 2012. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1340887011.

Fan:2005:PLI

[FH05] Jianqing Fan and Tao Huang. Profile likelihood inferences on semiparametric varying-coefficient partially linear models. Bernoulli, 11(6):1031–1057, December 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1137421639.

Fuh:2007:EHM

[FH07] Cheng-Der Fuh and Inchi Hu. Estimation in hidden Markov models via efficient importance sampling. Bernoulli, 13(2):492– 513, May 2007. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1179498758.

Fan:2020:CAS

[FHJK20] Jie Yen Fan, Kais Hamza, Peter Jagers, and Fima Klebaner. Convergence of the age structure of general schemes of population processes. Bernoulli, 26(2):893–926, May 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1580461567.

Fan:2015:MSS

[FHK15] Jie Yen Fan, Kais Hamza, and Fima Klebaner. Mimicking selfsimilar processes. Bernoulli, 21(3):1341–1360, August 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1432732022.

Filipovic:2000:EPF

[Fil00] Damir Filipovic. Exponential-polynomial families and the term structure of interest rates. Bernoulli, 6(6):1081–1107, December 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1081194161.

Fischer:2014:FLD

[Fis14] Markus Fischer. On the form of the large deviation rate function for the empirical measures of weakly interacting systems. Bernoulli, 20(4):1765–1801, November 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1411134444.

Fink:2011:FLD

[FK11] Holger Fink and Claudia Klüppelberg. Fractional Lévy-driven Ornstein–Uhlenbeck processes and stochastic differential equations. Bernoulli, 17(1):484–506, February 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1297173852.

Foondun:2012:ATR

[FK12] Mohammud Foondun and Davar Khoshnevisan. An asymptotic theory for randomly forced discrete nonlinear heat equations. Bernoulli, 18(3):1042–1060, August 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1340887013.

Fasen:2017:ICM

[FK17] Vicky Fasen and Sebastian Kimmig. Information criteria for multivariate CARMA processes. Bernoulli, 23(4A):2860–2886, November 2017. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1494316835.

Fan:2021:DRC

[FK21] Aihua Fan and Davit Karagulyan. On μ-Dvoretzky random covering of the circle. Bernoulli, 27(2):1270–1290, May 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-27/issue-2/On-%ce%bc-
Dvoretzky-random-covering-of-the-circle/10.3150/20-
BEJ1273.full.

Fushiki:2005:NBP

[FKA05] Tadayoshi Fushiki, Fumiyasu Komaki, and Kazuyuki Aihara. Nonparametric bootstrap prediction. Bernoulli, 11(2):293–307, April 2005. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1116340296.

Franke:2002:BKS

[FKM02] Jürgen Franke, Jens-Peter Kreiss, and Enno Mammen. Bootstrap of kernel smoothing in nonlinear time series. Bernoulli, 8(1):1–37, February 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1078951087.

Figueroa-Lopez:2011:SBC

[FL11] José E. Figueroa-López. Sieve-based confidence intervals and bands for Lévy densities. Bernoulli, 17(2):643–670, May 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1302009241.

Figueroa-Lopez:2014:STE

[FLLO14] José E. Figueroa-López, Yankeng Luo, and Cheng Ouyang. Small-time expansions for local jump-diffusion models with infinite jump activity. Bernoulli, 20(3):1165–1209, August 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1402488937.

Favaro:2012:ABN

[FLP12] Stefano Favaro, Antonio Lijoi, and Igor Prünster. Asymptotics for a Bayesian nonparametric estimator of species variety. Bernoulli, 18(4):1267–1283, November 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1352727810.

Ferrari:2001:BMA

[FLS01] Pablo A. Ferrari, Joel L. Lebowitz, and Eugene Speer. Blocking measures for asymmetric exclusion processes via coupling. Bernoulli, 7(6):935–950, December 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1078951130.

Fang:2020:RCT

[FLS20] Xiao Fang, Li Luo, and Qi-Man Shao. A refined Cramértype moderate deviation for sums of local statistics. Bernoulli, 26(3):2319–2352, August 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1587974543.

Fang:2023:CTM

[FLS23] Xiao Fang, Song-Hao Liu, and Qi-Man Shao. Cramértype moderate deviation for quadratic forms with a fast rate. Bernoulli, 29(3):2466–2491, August 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-3/Cram%c3%a9r-type-moderate-deviation-for-
quadratic-forms-with-a-fast/10.3150/22-BEJ1549.full.

Figueroa-Lopez:2014:STA

[FLT14] José E. Figueroa-López and Peter Tankov. Small-time asymptotics of stopped Lévy bridges and simulation schemes with controlled bias. Bernoulli, 20(3):1126–1164, August 2014. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1402488936.

Foucart:2021:TCS

[FLZ21] Clément Foucart, Pei-Sen Li, and Xiaowen Zhou. Timechanged spectrally positive Lévy processes started from infinity. Bernoulli, 27(2):1291–1318, May 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-2/Time-changed-spectrally-positive-L%c3%a9vy-
processes-started-from-infinity/10.3150/20-BEJ1274.
full.

Fernandez:2000:ABI

[FM00] Begoña Fernández and Sylvie Méléard. Asymptotic behaviour for interacting diffusion processes with space-time random birth. Bernoulli, 6(1):91–111, February 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1082665382.

Fournier:2002:WCA

[FM02] Nicolas Fournier and Sylvie Méléard. A weak criterion of absolute continuity for jump processes: application to the Boltzmann equation. Bernoulli, 8(4):537–558, August 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1078681383.

Feldheim:2018:NCI

[FMNW18] Naomi Feldheim, Arnaud Marsiglietti, Piotr Nayar, and Jing Wang. A note on the convex infimum convolution inequality. Bernoulli, 24(1):257–270, February 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1501142442.

Fort:2014:CLT

[FMPV14] G. Fort, E. Moulines, P. Priouret, and P. Vandekerkhove. A central limit theorem for adaptive and interacting Markov chains. Bernoulli, 20(2):457–485, May 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1393593994.

Flammarion:2019:ORS

[FMR19] Nicolas Flammarion, Cheng Mao, and Philippe Rigollet. Optimal rates of statistical seriation. Bernoulli, 25(1):623–653, February 2019. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1544605258.

Fasano:2012:CDR

[FMSY12] María V. Fasano, Ricardo A. Maronna, Mariela Sued, and Víctor J. Yohai. Continuity and differentiability of regression m functionals. Bernoulli, 18(4):1284–1309, November 2012. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1352727811.

Ferrante:1997:ENM

[FN97] Marco Ferrante and David Nualart. An example of a non-Markovian stochastic two-point boundary value problem. Bernoulli, 3(4):371–386, December 1997. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1175882213.

Foondun:2021:OCS

[FN21] Mohammud Foondun and Eulalia Nualart. The Osgood condition for stochastic partial differential equations. Bernoulli, 27(1):295–311, February 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-27/issue-
1/The-Osgood-condition-for-stochastic-partial-differential-
equations/10.3150/20-BEJ1240.full.

Favaro:2023:NOE

[FN23] Stefano Favaro and Zacharie Naulet. Near-optimal estimation of the unseen under regularly varying tail populations. Bernoulli, 29(4):3423–3442, November 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-4/Near-optimal-estimation-of-the-unseen-under-
regularly-varying-tail/10.3150/23-BEJ1589.full.

Fukasawa:2024:MCT

[FO24] Masaaki Fukasawa and Teppei Ogihara. Malliavin calculus techniques for local asymptotic mixed normality and their application to hypoelliptic diffusions. Bernoulli, 30(2):983– 1006, May 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-30/issue-2/Malliavin-
calculus-techniques-for-local-asymptotic-mixed-normality-
and-their/10.3150/23-BEJ1621.full.

Fournier:2001:STS

[Fou01] Nicolas Fournier. Support theorem for the solution of a whitenoise-driven parabolic stochastic partial differential equation with temporal Poissonian jumps. Bernoulli, 7(1):165–190, February 2001. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1080572344.

Fournier:2010:ACS

[FP10] Nicolas Fournier and Jacques Printems. Absolute continuity for some one-dimensional processes. Bernoulli, 16(2):343–360, May 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1274821074.

Fissler:2017:TMR

[FP17a] Tobias Fissler and Mark Podolskij. Testing the maximal rank of the volatility process for continuous diffusions observed with noise. Bernoulli, 23(4B):3021–3066, November 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1495505084.

Fortini:2017:PCM

[FP17b] Sandra Fortini and Sonia Petrone. Predictive characterization of mixtures of Markov chains. Bernoulli, 23(3):1538–1565, August 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1489737617.

Fox:2017:AGS

[FP17c] Colin Fox and Albert Parker. Accelerated Gibbs sampling of normal distributions using matrix splittings and polynomials. Bernoulli, 23(4B):3711–3743, November 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1495505107.

Follmer:1995:QCE

[FPS95] Hans Föllmer, Philip Protter, and Albert N. Shiryayev. Quadratic covariation and an extension of Itô’s formula. Bernoulli, 1(1–2):149–169, March 1995. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1186078365.

Fang:2019:LTR

[FPSS19] Xiao Fang, Shige Peng, Qi-Man Shao, and Yongsheng Song. Limit theorems with rate of convergence under sublinear expectations. Bernoulli, 25(4A):2564–2596, November 2019. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1568362036.

Franke:2017:CSI

[FPW17] Brice Franke, Françoise Pène, and Martin Wendler. Convergence of U-statistics indexed by a random walk to stochastic integrals of a Lévy sheet. Bernoulli, 23(1):329–378, February 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1475001357.

Ferrante:2006:SDD

[FR06] Marco Ferrante and Carles Rovira. Stochastic delay differential equations driven by fractional Brownian motion with Hurst parameter H > 1/2. Bernoulli, 12(1):85–100, February 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1141136650.

Fryzlewicz:2011:MPA

[FR11] Piotr Fryzlewicz and Suhasini Subba Rao. Mixing properties of ARCH and time-varying ARCH processes. Bernoulli, 17 (1):320–346, February 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1297173845.

Fang:2015:RCM

[FR15] Xiao Fang and Adrian Röllin. Rates of convergence for multivariate normal approximation with applications to dense graphs and doubly indexed permutation statistics. Bernoulli, 21(4): 2157–2189, November 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1438777590.

Franceschi:2019:IES

[FR19] Sandro Franceschi and Kilian Raschel. Integral expression for the stationary distribution of reflected Brownian motion in a wedge. Bernoulli, 25(4B):3673–3713, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1569398781.

Franke:2006:SLB

[Fra06] Rice Franke. The scaling limit behaviour of periodic stable-like processes. Bernoulli, 12(3):551–570, June 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:/

/projecteuclid.org/euclid.bj/1151525136. See correction

[Fra07].

Franke:2007:CSL

[Fra07] Brice Franke. Correction to: The scaling limit behaviour of periodic stable-like processes. Bernoulli, 13(2):600, May 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/

1179498763. See [Fra06].

Frikha:2018:WAS

[Fri18] Noufel Frikha. On the weak approximation of a skew diffusion by an Euler-type scheme. Bernoulli, 24(3):1653–1691, August 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1517540457.

Fritz:2024:CAC

[Fri24] Tobias Fritz. Characterizing the asymptotic and catalytic stochastic orders on topological abelian groups. Bernoulli, 30 (4):2932–2954, November 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

Fay:2007:EMP

[FRS07] Gilles Faÿ, François Roueff, and Philippe Soulier. Estimation of the memory parameter of the infinite-source Poisson process. Bernoulli, 13(2):473–491, May 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1179498757.

Fermanian:2004:WCE

[FRW04] Jean-David Fermanian, Dragan Radulovic, and Marten Wegkamp. Weak convergence of empirical copula processes. Bernoulli, 10(5):847–860, October 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1099579158.

Fermanian:2015:ATV

[FRW15] Jean-David Fermanian, Dragan Radulović, and Marten Wegkamp. Asymptotic total variation tests for copulas. Bernoulli, 21(3):1911–1945, August 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1432732042.

Ferreyra:2000:CSV

[FS00] Guillermo Ferreyra and Padamanbhan Sundar. Comparison of stochastic Volterra equations. Bernoulli, 6(6):1001–1006, December 2000. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1081194156.

Franke:2010:SSP

[FS10] Brice Franke and Tatsuhiko Saigo. A self-similar process arising from a random walk with random environment in random scenery. Bernoulli, 16(3):825–857, August 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1281099886.

Follmer:2013:PAF

[FS13] Hans Föllmer and Alexander Schied. Probabilistic aspects of finance. Bernoulli, 19(4):1306–1326, September 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1377612853.

Fathi:2018:CTI

[FS18] Max Fathi and Yan Shu. Curvature and transport inequalities for Markov chains in discrete spaces. Bernoulli, 24(1):672– 698, February 2018. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1501142459.

Fokianos:2020:MCA

[FSTD20] Konstantinos Fokianos, Bård Støve, Dag Tjøstheim, and Paul Doukhan. Multivariate count autoregression. Bernoulli, 26(1): 471–499, February 2020. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1574758835.

Francis:2017:BEC

[FSW17] Andrew R. Francis, Milan Stehlík, and Henry P. Wynn. “Building” exact confidence nets. Bernoulli, 23(4B):3145–3165, November 2017. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1495505087.

Fukasawa:2019:AEE

[FT19] Masaaki Fukasawa and Tetsuya Takabatake. Asymptotically efficient estimators for self-similar stationary Gaussian noises under high frequency observations. Bernoulli, 25(3):1870–1900, August 2019. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1560326431.

Fushiki:2005:BPB

[Fus05] Tadayoshi Fushiki. Bootstrap prediction and Bayesian prediction under misspecified models. Bernoulli, 11(4):747–758, August 2005. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1126126768.

Fiocco:2003:DCS

[FV03a] Marta Fiocco and Willem R. Van Zwet. Decaying correlations for the supercritical contact process conditioned on survival. Bernoulli, 9(5):763–781, October 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1066418877.

Fiocco:2003:PES

[FV03b] Marta Fiocco and Willem R. Van Zwet. Parameter estimation for the supercritical contact process. Bernoulli, 9(6):1071–1092, December 2003. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1072215201.

Fujiwara:2020:NLD

[FY20] Akio Fujiwara and Koichi Yamagata. Noncommutative Lebesgue decomposition and contiguity with applications in quantum statistics. Bernoulli, 26(3):2105–2142, August 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL https://projecteuclid.org/euclid.bj/1587974535.

Francq:2004:MLE

[FZ04] Christian Francq and Jean-Michel Zakoïan. Maximum likelihood estimation of pure GARCH and ARMA-GARCH processes. Bernoulli, 10(4):605–637, August 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1093265632.

Francq:2022:AED

[FZ22] Christian Francq and Jean-Michel Zakoïan. Adaptiveness of the empirical distribution of residuals in semi-parametric conditional location scale models. Bernoulli, 28(1):548–578, February 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-28/issue-1/Adaptiveness-of-the-empirical-
distribution-of-residuals-in-semi-parametric/10.3150/
21-BEJ1357.full.

Fontana:2023:CPU

[FZV23] Matteo Fontana, Gianluca Zeni, and Simone Vantini. Conformal prediction: a unified review of theory and new challenges. Bernoulli, 29(1):1–23, February 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-1/Conformal-prediction--A-unified-review-of-
theory-and-new/10.3150/21-BEJ1447.full.

Ganesan:2021:MST

[Gan21] Ghurumuruhan Ganesan. Minimum spanning trees of random geometric graphs with location dependent weights. Bernoulli, 27(4):2473–2493, November 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-27/issue-
4/Minimum-spanning-trees-of-random-geometric-graphs-
with-location-dependent/10.3150/20-BEJ1318.full.

Gao:2020:RRM

[Gao20] Chao Gao. Robust regression via mutivariate regression depth. Bernoulli, 26(2):1139–1170, May 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1580461575.

Gass:2023:ASA

[Gas23] Louis Gass. Almost-sure asymptotics for Riemannian random waves. Bernoulli, 29(1):625–651, February 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-1/Almost-sure-asymptotics-for-Riemannian-random-
waves/10.3150/22-BEJ1471.full.

Gates:2006:SSD

[Gat06] John Gates. The shape of a sequence of dual random triangles. Bernoulli, 12(1):55–63, February 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1141136648.

Gaunt:2017:SMP

[Gau17] Robert E. Gaunt. On Stein’s method for products of normal random variables and zero bias couplings. Bernoulli, 23(4B):3311–3345, November 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1495505094.

Gayraud:2001:MHT

[Gay01] Ghislaine Gayraud. Minimax hypothesis testing about the density support. Bernoulli, 7(3):507–525, June 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1080004763.

Gaynanova:2020:PEC

[Gay20] Irina Gaynanova. Prediction and estimation consistency of sparse multi-class penalized optimal scoring. Bernoulli, 26(1): 286–322, February 2020. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1574758829.

Gaynanova:2022:EPE

[Gay22] Irina Gaynanova. Erratum for prediction and estimation consistency of sparse multi-class penalized optimal scoring. Bernoulli, 28(1):601–605, February 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-1/Erratum-for-Prediction-and-estimation-consistency-
of-sparse-multi-class/10.3150/21-BEJ1359.full.

Ghosh:2005:DDD

[GC05] Anil K. Ghosh and Probal Chaudhuri. On data depth and distribution-free discriminant analysis using separating surfaces. Bernoulli, 11(1):1–27, January 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1110228239.

Gerber:2017:CSQ

[GC17] Mathieu Gerber and Nicolas Chopin. Convergence of sequential quasi-Monte Carlo smoothing algorithms. Bernoulli, 23(4B):2951–2987, November 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1495505081.

Genon-Catalot:1998:LTD

[GCJL98] Valentine Genon-Catalot, Thierry Jeantheau, and Catherine Larédo. Limit theorems for discretely observed stochastic volatility models. Bernoulli, 4(3):283–303, September 1998. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1174324982.

Genon-Catalot:1999:PED

[GCJL99] Valentine Genon-Catalot, Thierry Jeantheau, and Catherine Laredo. Parameter estimation for discretely observed stochastic volatility models. Bernoulli, 5(5):855–872, October 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1171290402.

Genon-Catalot:2000:SVM

[GCJL00] Valentine Genon-Catalot, Thierry Jeantheau, and Catherine Larédo. Stochastic volatility models as hidden Markov models and statistical applications. Bernoulli, 6(6):1051–1079, December 2000. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1081194160.

Genon-Catalot:2004:RSP

[GCK04] Valentine Genon-Catalot and Mathieu Kessler. Random scale perturbation of an AR(1) process and its properties as a nonlinear explicit filter. Bernoulli, 10(4):701–720, August 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1093265637.

Geenens:2017:PTN

[GCP17] Gery Geenens, Arthur Charpentier, and Davy Paindaveine. Probit transformation for nonparametric kernel estimation of the copula density. Bernoulli, 23(3):1848–1873, August 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL https://projecteuclid.org/euclid.bj/1489737627.

Giraud:2017:NAA

[GD17] François Giraud and Pierre Del Moral. Nonasymptotic analysis of adaptive and annealed Feynman–Kac particle models. Bernoulli, 23(1):670–709, February 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1475001371.

Gaillac:2022:AEL

[GG22] Christophe Gaillac and Eric Gautier. Adaptive estimation in the linear random coefficients model when regressors have limited variation. Bernoulli, 28(1):504–524, February 2022. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-1/Adaptive-estimation-in-the-linear-random-
coefficients-model-when-regressors/10.3150/21-BEJ1354.
full.

Gamboa:2022:GSA

[GGKL22] Fabrice Gamboa, Pierre Gremaud, Thierry Klein, and Agnès Lagnoux. Global sensitivity analysis: a novel generation of mighty estimators based on rank statistics. Bernoulli, 28(4):2345–2374, November 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-4/Global-sensitivity-analysis--A-novel-generation-
of-mighty-estimators/10.3150/21-BEJ1421.full.

Gadat:2020:OFS

[GGM20] Sébastien Gadat, Sébastien Gerchinovitz, and Clément Marteau. Optimal functional supervised classification with separation condition. Bernoulli, 26(3):1797–1831, August 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1587974524.

Gaigall:2022:EPC

[GGT22] Daniel Gaigall, Julian Gerstenberg, and Thi Thu Ha Trinh. Empirical process of concomitants for partly categorial data and applications in statistics. Bernoulli, 28(2):803–829, May 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-28/issue-2/Empirical-
process-of-concomitants-for-partly-categorial-data-
and-applications/10.3150/21-BEJ1367.full.

Guo:2005:ZSC

[GHL05] Xianping Guo and Onésimo Hernández-Lerma. Zero-sum continuous-time Markov games with unbounded transition and discounted payoff rates. Bernoulli, 11(6):1009–1029, December 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1137421638.

Ghosh:2017:GDS

[GHM^{+}17] Abhik Ghosh, Ian R. Harris, Avijit Maji, Ayanendranath Basu, and Leandro Pardo. A generalized divergence for statistical inference. Bernoulli, 23(4A):2746–2783, November 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1494316831.

Guha:2021:PCP

[GHN21] Aritra Guha, Nhat Ho, and XuanLong Nguyen. On posterior contraction of parameters and interpretability in Bayesian mixture modeling. Bernoulli, 27(4):2159–2188, November 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-27/issue-4/On-posterior-contraction-
of-parameters-and-interpretability-in-Bayesian-mixture/
10.3150/20-BEJ1275.full.

Ghosal:1999:ANP

[Gho99] Subhashis Ghosal. Asymptotic normality of posterior distributions in high-dimensional linear models. Bernoulli, 5(2):315– 331, April 1999. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1173147909.

Gine:1996:EPA

[Gin96a] Evarist Giné. Empirical processes and applications: an overview. Bernoulli, 2(1):1–28, March 1996. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1193758786.

Gine:1996:R

[Gin96b] Evarist Giné. Rejoinder. Bernoulli, 2(1):29–38, March 1996. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1193758788.

Giraud:2008:MLS

[Gir08] Christophe Giraud. Mixing least-squares estimators when the variance is unknown. Bernoulli, 14(4):1089–1107, November 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1225980572.

Giulini:2018:RDF

[Giu18] Ilaria Giulini. Robust dimension-free Gram operator estimates. Bernoulli, 24(4B):3864–3923, November 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1524038772.

Groeneboom:2013:TMH

[GJ13] Piet Groeneboom and Geurt Jongbloed. Testing monotonicity of a hazard: Asymptotic distribution theory. Bernoulli, 19(5A): 1965–1999, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1383661210.

Giordano:2020:SFN

[GJQS20] Luca M. Giordano, Maria Jolis, and Lluís Quer-Sardanyons. SPDEs with fractional noise in space: Continuity in law with respect to the Hurst index. Bernoulli, 26(1):352–386, February 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1574758831.

Gut:2010:BLL

[GJS10] Allan Gut, Fredrik Jonsson, and Ulrich Stadtmüller. Between the LIL and the LSL. Bernoulli, 16(1):1–22, February 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1265984702.

Groisman:2022:NEF

[GJS22] Pablo Groisman, Matthieu Jonckheere, and Facundo Sapienza. Nonhomogeneous Euclidean first-passage percolation and distance learning. Bernoulli, 28(1):255–276, February 2022. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-1/Nonhomogeneous-Euclidean-first-passage-percolation-
and-distance-learning/10.3150/21-BEJ1341.full.

Going-Jaeschke:2003:SSG

[GJY03] Anja Göing-Jaeschke and Marc Yor. A survey and some generalizations of Bessel processes. Bernoulli, 9(2):313–349, April 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1068128980.

Gao:2023:CTM

[GJZ23] Fuqing Gao, Hui Jiang, and Xingqiu Zhao. Cramér type moderate deviations for the Grenander estimator near the boundaries of the support. Bernoulli, 29(4):2854–2878, November 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-29/issue-4/Cram%c3%a9r-type-moderate-
deviations-for-the-Grenander-estimator-near-the/10.
3150/22-BEJ1566.full.

Gushchin:1999:AIL

[GK99] Alexander A. Gushchin and Uwe Küchler. Asymptotic inference for a linear stochastic differential equation with time delay. Bernoulli, 5(6):1059–1098, December 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1143122303.

Gushchin:2001:AAI

[GK01] Alexander A. Gushchin and Uwe Küchler. Addendum to ‘asymptotic inference for a linear stochastic differential equation with time delay’. Bernoulli, 7(4):629–632, August 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1079559466.

Gaigalas:2003:CSR

[GK03] Raimundas Gaigalas and Ingemar Kaj. Convergence of scaled renewal processes and a packet arrival model. Bernoulli, 9 (4):671–703, August 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1066223274.

Gugushvili:2012:CPE

[GK12] Shota Gugushvili and Chris A. J. Klaassen. \sqrt{n}-consistent parameter estimation for systems of ordinary differential equations: bypassing numerical integration via smoothing. Bernoulli, 18(3):1061–1098, August 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1340887014.

Giraitis:2013:ADW

[GK13] Liudas Giraitis and Hira L. Koul. On asymptotic distributions of weighted sums of periodograms. Bernoulli, 19(5B): 2389–2413, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1386078607.

Gissibl:2018:MLM

[GK18] Nadine Gissibl and Claudia Klüppelberg. Max-linear models on directed acyclic graphs. Bernoulli, 24(4A):2693–2720, November 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1522051222.

Genest:2011:GFT

[GKNY11] Christian Genest, Ivan Kojadinovic, Johanna Nešlehová, and Jun Yan. A goodness-of-fit test for bivariate extreme-value copulas. Bernoulli, 17(1):253–275, February 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1297173842.

Gamboa:2012:CLM

[GKP12] Fabrice Gamboa, Thierry Klein, and Clémentine Prieur. Conditional large and moderate deviations for sums of discrete random variables. Combinatoric applications. Bernoulli, 18(4): 1341–1360, November 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1352727814.

Gusakova:2024:SVT

[GKT24] Anna Gusakova, Zakhar Kabluchko, and Christoph Thäle. Sectional Voronoi tessellations: Characterization and highdimensional limits. Bernoulli, 30(2):1482–1501, May 2024. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
30/issue-2/Sectional-Voronoi-tessellations-Characterization-
and-high-dimensional-limits/10.3150/23-BEJ1641.full.

Gill:1995:AVT

[GL95] Richard D. Gill and Boris Y. Levit. Applications of the van Trees inequality: a Bayesian Cramér–Rao bound. Bernoulli, 1(1–2):59–79, March 1995. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1186078362.

Gamboa:2005:WEM

[GL05] Fabrice Gamboa and Jean-Michel Loubes. Wavelet estimation of a multifractal function. Bernoulli, 11(2):221–246, April 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1116340292.

Girardin:2006:ESM

[GL06] Valerie Girardin and Nikolaos Limnios. Entropy for semi-Markov processes with Borel state spaces: asymptotic equirepartition properties and invariance principles. Bernoulli, 12(3):515–533, June 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1151525134.

Goldenshluger:2008:UPS

[GL08] Alexander Goldenshluger and Oleg Lepski. Universal pointwise selection rule in multivariate function estimation. Bernoulli, 14 (4):1150–1190, November 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1225980575.

Gao:2018:STO

[GL18] ZhiQiang Gao and Quansheng Liu. Second and third orders asymptotic expansions for the distribution of particles in a branching random walk with a random environment in time. Bernoulli, 24(1):772–800, February 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1501142462.

Goldenshluger:2022:MENa

[GL22a] Alexander Goldenshluger and Oleg V. Lepski. Minimax estimation of norms of a probability density: I. Lower bounds. Bernoulli, 28(2):1120–1154, May 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-2/Minimax-estimation-of-norms-of-a-probability-
density--I/10.3150/21-BEJ1380.full.

Goldenshluger:2022:MENb

[GL22b] Alexander Goldenshluger and Oleg V. Lepski. Minimax estimation of norms of a probability density: II. Rate-optimal estimation procedures. Bernoulli, 28(2):1155–1178, May 2022. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-2/Minimax-estimation-of-norms-of-a-probability-
density--II/10.3150/21-BEJ1381.full.

Gontscharuk:2016:GFT

[GLF16] Veronika Gontscharuk, Sandra Landwehr, and Helmut Finner. Goodness of fit tests in terms of local levels with special emphasis on higher criticism tests. Bernoulli, 22(3):1331–1363, August 2016. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1458132984.

Geiss:2020:RWA

[GLL20] Christel Geiss, Céline Labart, and Antti Luoto. Random walk approximation of BSDEs with Hölder continuous terminal condition. Bernoulli, 26(1):159–190, February 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1574758825.

Gloaguen:2022:PMS

[GLO22] Pierre Gloaguen, Sylvain Le Corff, and Jimmy Olsson. A pseudo-marginal sequential Monte Carlo online smoothing algorithm. Bernoulli, 28(4):2606–2633, November 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-4/A-pseudo-marginal-sequential-Monte-Carlo-
online-smoothing-algorithm/10.3150/21-BEJ1431.full.

Gouet:2007:ANC

[GLS07] Raúl Gouet, F. Javier López, and Gerardo Sanz. Asymptotic normality for the counting process of weak records and δ-records in discrete models. Bernoulli, 13(3):754–781, August 2007. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1186503485.

Grahovac:2019:UPA

[GLST19] Danijel Grahovac, Nikolai N. Leonenko, Alla Sikorskii, and Murad S. Taqqu. The unusual properties of aggregated superpositions of Ornstein–Uhlenbeck type processes. Bernoulli, 25 (3):2029–2050, August 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1560326436.

Golubev:1996:AEE

[GLT96] Yuri K. Golubev, Boris Y. Levit, and Alexander B. Tsybakov. Asymptotically efficient estimation of analytic functions in Gaussian noise. Bernoulli, 2(2):167–181, June 1996. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1193839222.

Garzon:2023:FIB

[GLT23] Johanna Garzón, Jorge A. León, and Soledad Torres. Forward integration of bounded variation coefficients with respect to Hölder continuous processes. Bernoulli, 29(3):1877–1904, August 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-29/issue-3/Forward-integration-of-
bounded-variation-coefficients-with-respect-to-H%c3%
b6lder/10.3150/22-BEJ1524.full.

Gine:2004:LIL

[GM04] Evarist Giné and David M. Mason. The law of the iterated logarithm for the integrated squared deviation of a kernel density estimator. Bernoulli, 10(4):721–752, August 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1093265638.

Giacomin:2015:WNN

[GM15] Giambattista Giacomin and Mathieu Merle. Weak noise and non-hyperbolic unstable fixed points: Sharp estimates on transit and exit times. Bernoulli, 21(4):2242–2288, November 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1438777593.

Graczyk:2019:SBP

[GM19] Piotr Graczyk and Jacek Malecki. On squared Bessel particle systems. Bernoulli, 25(2):828–847, May 2019. CODEN ????

ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1551862836.

Gonzalez:2005:CCB

[GMD05] Miguel González, Manuel Molina, and Inés Del Puerto. On L^{2}- convergence of controlled branching processes with random control function. Bernoulli, 11(1):37–46, January 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1110228241.

Gonzalez:2006:GGS

[GMD06] Miguel González, Manuel Molina, and Inés Del Puerto. Geometric growth for stochastic difference equations with application to branching populations. Bernoulli, 12(5):931–942, October 2006. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1161614953.

Gonzalez:2005:UGC

[GMM05] Miguel González, Rodrigo Martínez, and Manuel Mota. On the unlimited growth of a class of homogeneous multitype Markov chains. Bernoulli, 11(3):559–570, June 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1120591189.

Greenshtein:2019:NEB

[GMR19] Eitan Greenshtein, Ariel Mantzura, and Ya’acov Ritov. Nonparametric empirical Bayes improvement of shrinkage estimators with applications to time series. Bernoulli, 25(4B):3459– 3478, November 2019. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1569398773.

Guasoni:2021:MDP

[GMZ21] Paolo Guasoni, Eberhard Mayerhofer, and Mingchuan Zhao. Minimal L^{p}-densities with prescribed marginals. Bernoulli, 27(1):576–585, February 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-27/issue-
1/Minimal-mathcalLp-densities-with-prescribed-marginals/
10.3150/20-BEJ1250.full.

Gine:2008:SAE

[GN08] Evarist Giné and Richard Nickl. A simple adaptive estimator of the integrated square of a density. Bernoulli, 14(1):47–61, February 2008. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1202492784.

Gine:2010:AED

[GN10] Evarist Giné and Richard Nickl. Adaptive estimation of a distribution function and its density in sup-norm loss by wavelet and spline projections. Bernoulli, 16(4):1137–1163, November 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1290092899.

Gnedin:2004:BS

[Gne04] Alexander V. Gnedin. The Bernoulli sieve. Bernoulli, 10(1):79– 96, February 2004. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1077544604.

Gneiting:2013:SNS

[Gne13] Tilmann Gneiting. Strictly and non-strictly positive definite functions on spheres. Bernoulli, 19(4):1327–1349, September 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1377612854.

Garino:2021:LTI

[GNNS21] Valentin Garino, Ivan Nourdin, David Nualart, and Majid Salamat. Limit theorems for integral functionals of Hermite-driven processes. Bernoulli, 27(3):1764–1788, August 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-3/Limit-theorems-for-integral-functionals-
of-Hermite-driven-processes/10.3150/20-BEJ1291.full.

Genest:2014:EMC

[GNR14] Christian Genest, Johanna G. Nešlehová, and Bruno Rémillard. On the empirical multilinear copula process for count data. Bernoulli, 20(3):1344–1371, August 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1402488942.

Gamboa:2019:SRL

[GNR19] Fabrice Gamboa, Jan Nagel, and Alain Rouault. Sum rules and large deviations for spectral matrix measures. Bernoulli, 25(1): 712–741, February 2019. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1544605261.

Gotze:2019:LBP

[GNSU19] Friedrich Götze, Alexey Naumov, Vladimir Spokoiny, and Vladimir Ulyanov. Large ball probabilities, Gaussian comparison and anti-concentration. Bernoulli, 25(4A):2538–2563, November 2019. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1568362035.

Gotze:2017:DLS

[GNT17] Friedrich Götze, Alexey Naumov, and Alexander Tikhomirov. Distribution of linear statistics of singular values of the product of random matrices. Bernoulli, 23(4B):3067–3113, November 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1495505085.

Gotze:2018:LSL

[GNTT18] Friedrich Götze, Alexey Naumov, Alexander Tikhomirov, and Dmitry Timushev. On the local semicircular law for Wigner ensembles. Bernoulli, 24(3):2358–2400, August 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1517540477.

Ganesh:2000:LDP

[GO00] Ayalvadi J. Ganesh and Neil O’Connell. A large-deviation principle for Dirichlet posteriors. Bernoulli, 6(6):1021–1034, December 2000. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1081194158.

Gobet:2001:LAM

[Gob01] Emmanuel Gobet. Local asymptotic mixed normality property for elliptic diffusion: a Malliavin calculus approach. Bernoulli, 7(6):899–912, December 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1078951128.

Goldenshluger:1999:PAN

[Gol99] Alexander Goldenshluger. On pointwise adaptive nonparametric deconvolution. Bernoulli, 5(5):907–925, October 1999. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1171290404.

Gorostiza:1996:AFC

[Gor96] Luis G. Gorostiza. Asymptotic fluctuations and critical dimension for a two-level branching system. Bernoulli, 2(2):109–132, June 1996. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1193839219.

Goldenshluger:2003:AIE

[GP03] Alexander Goldenshluger and Sergei V. Pereverzev. On adaptive inverse estimation of linear functionals in Hilbert scales. Bernoulli, 9(5):783–807, October 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1066418878.

Garrett:2004:NUA

[GP04] Elizabeth S. Garrett and Giovanni Parmigiani. A nested unsupervised approach to identifying novel molecular subtypes. Bernoulli, 10(6):951–969, December 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1106314845.

Gordin:2011:FCL

[GP11] Mikhail Gordin and Magda Peligrad. On the functional central limit theorem via martingale approximation. Bernoulli, 17 (1):424–440, February 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1297173848.

Guillotte:2012:BEB

[GP12] Simon Guillotte and François Perron. Bayesian estimation of a bivariate copula using the Jeffreys prior. Bernoulli, 18(2):496– 519, May 2012. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1334580722.

Galtchouk:2015:EPE

[GP15] L. I. Galtchouk and S. M. Pergamenshchikov. Efficient pointwise estimation based on discrete data in ergodic nonparametric diffusions. Bernoulli, 21(4):2569–2594, November 2015. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1438777605.

Guillotte:2016:PPF

[GP16] Simon Guillotte and François Perron. Polynomial Pickands functions. Bernoulli, 22(1):213–241, February 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1443620848.

Garner:2018:LBB

[GP18] William Garner and Dimitris N. Politis. Local block bootstrap for inhomogeneous Poisson marked point processes. Bernoulli, 24(1):592–615, February 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1501142456.

Giurcanu:2023:BIC

[GP23] Mihai Giurcanu and Brett Presnell. Bootstrap inference for a class of non-regular estimators. Bernoulli, 29(3):2192– 2218, August 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-29/issue-3/Bootstrap-
inference-for-a-class-of-non-regular-estimators/10.
3150/22-BEJ1538.full.

Garcia-Portugues:2023:PBC

[GPNECA23] Eduardo García-Portugués, Paula Navarro-Esteban, and Juan A. Cuesta-Albertos. On a projection-based class of uniformity tests on the hypersphere. Bernoulli, 29(1):181–204, February 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-29/issue-1/On-a-projection-based-class-
of-uniformity-tests-on-the/10.3150/21-BEJ1454.full.

Guo:2023:CSH

[GQ23] Shaojun Guo and Xinghao Qiao. On consistency and sparsity for high-dimensional functional time series with application to autoregressions. Bernoulli, 29(1):451–472, February 2023. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-29/issue-1/On-consistency-
and-sparsity-for-high-dimensional-functional-time-series/
10.3150/22-BEJ1464.full.

Greenshtein:2000:SSP

[GR00] Eitan Greenshtein and Ya’acov Ritov. Sampling from a stationary process and detecting a change in the mean of a stationary distribution. Bernoulli, 6(4):679–697, August 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1081449601.

Greenshtein:2004:PHD

[GR04] Eitan Greenshtein and Ya’acov Ritov. Persistence in highdimensional linear predictor selection and the virtue of overparametrization. Bernoulli, 10(6):971–988, December 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1106314846.

Gassiat:2014:APD

[GR14] Elisabeth Gassiat and Judith Rousseau. About the posterior distribution in hidden Markov models with unknown number of states. Bernoulli, 20(4):2039–2075, November 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1411134453.

Gassiat:2016:NFT

[GR16] Elisabeth Gassiat and Judith Rousseau. Nonparametric finite translation hidden Markov models and extensions. Bernoulli, 22(1):193–212, February 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1443620847.

Gaudlitz:2023:ERT

[GR23a] Sascha Gaudlitz and Markus Reiß. Estimation for the reaction term in semi-linear SPDEs under small diffusivity. Bernoulli, 29(4):3033–3058, November 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-4/Estimation-for-the-reaction-term-in-semi-
linear-SPDEs-under/10.3150/22-BEJ1573.full.

Gaunt:2023:BCS

[GR23b] Robert E. Gaunt and Gesine Reinert. Bounds for the chi-square approximation of Friedman’s statistic by Stein’s method. Bernoulli, 29(3):2008–2034, August 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-3/Bounds-for-the-chi-square-approximation-
of-Friedmans-statistic-by/10.3150/22-BEJ1530.full.

Grandits:1999:OMM

[Gra99] Peter Grandits. The p-optimal martingale measure and its asymptotic relation with the minimal-entropy martingale measure. Bernoulli, 5(2):225–247, April 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1173147904.

Grandits:2021:AHP

[Gra21] Peter Grandits. Asymptotics of the hitting probability for a small sphere and a two dimensional Brownian motion with discontinuous anisotropic drift. Bernoulli, 27(2):853– 865, May 2021. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-27/issue-2/Asymptotics-of-
the-hitting-probability-for-a-small-sphere-and/10.3150/
20-BEJ1257.full.

Goldstein:2011:SCS

[GRS11] Larry Goldstein, Yosef Rinott, and Marco Scarsini. Stochastic comparisons of stratified sampling techniques for some Monte Carlo estimators. Bernoulli, 17(2):592–608, May 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1302009238.

Gruet:2000:NHM

[Gru00] Jean-Claude Gruet. A note on hyperbolic von Mises distributions. Bernoulli, 6(6):1007–1020, December 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1081194157.

Garren:2000:ESL

[GS00a] Steven T. Garren and Richard L. Smith. Estimating the second largest eigenvalue of a Markov transition matrix. Bernoulli, 6(2):215–242, April 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1081788027.

Graversen:2000:EPL

[GS00b] Svend Erik Graversen and Albert N. Shiryaev. An extension of P. Lévy’s distributional. Bernoulli, 6(4):615–620, August 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1081449596.

Girardin:2003:SSP

[GS03] Valerie Girardin and Rachid Senoussi. Semigroup stationary processes and spectral representation. Bernoulli, 9(5):857–876, October 2003. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1066418881.

Gut:2008:LSL

[GS08a] Allan Gut and Ulrich Stadtmüller. Laws of the single logarithm for delayed sums of random fields. Bernoulli, 14(1):249–276, February 2008. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1202492793.

Gyongy:2008:RS

[GŠ08b] István Gyöngy and David Šiška. On randomized stopping. Bernoulli, 14(2):352–361, May 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1208872108.

Griffiths:2011:MJL

[GS11] Robert C. Griffiths and Dario Spanò. Multivariate Jacobi and Laguerre polynomials, infinite-dimensional extensions, and their probabilistic connections with multivariate Hahn and Meixner polynomials. Bernoulli, 17(3):1095–1125, August 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1310042858.

Griffiths:2013:OPK

[GS13] Robert C. Griffiths and Dario Spanò. Orthogonal polynomial kernels and canonical correlations for Dirichlet measures. Bernoulli, 19(2):548–598, May 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1363192038.

Geiss:2016:VLD

[GS16] Christel Geiss and Alexander Steinicke. L_{2}-variation of Lévy driven BSDEs with non-smooth terminal conditions. Bernoulli, 22(2):995–1025, May 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1447077767.

Gorgen:2018:ECS

[GS18a] Christiane Görgen and Jim Q. Smith. Equivalence classes of staged trees. Bernoulli, 24(4A):2676–2692, November 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL https://projecteuclid.org/euclid.bj/1522051221.

Gross:2018:MLT

[GS18b] Elizabeth Gross and Seth Sullivant. The maximum likelihood threshold of a graph. Bernoulli, 24(1):386–407, February 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL https://projecteuclid.org/euclid.bj/1501142448.

Gao:2020:FLE

[GS20a] Zheng Gao and Stilian Stoev. Fundamental limits of exact support recovery in high dimensions. Bernoulli, 26(4):2605– 2638, November 2020. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-26/issue-4/Fundamental-limits-
of-exact-support-recovery-in-high-dimensions/10.3150/
20-BEJ1197.full.

Gotze:2020:SOC

[GS20b] Friedrich Götze and Holger Sambale. Second order concentration via logarithmic Sobolev inequalities. Bernoulli, 26(1):93– 126, February 2020. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1574758823.

Grzywny:2020:LPC

[GS20c] Tomasz Grzywny and Karol Szczypkowski. Lévy processes: Concentration function and heat kernel bounds. Bernoulli, 26(4):3191–3223, November 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-26/issue-
4/L%c3%a9vy-processes-Concentration-function-and-heat-
kernel-bounds/10.3150/20-BEJ1220.full.

Gnedin:2021:ARA

[GS21] Alexander Gnedin and Amirlan Seksenbayev. Asymptotics and renewal approximation in the online selection of increasing subsequence. Bernoulli, 27(3):1851–1869, August 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-27/issue-3/Asymptotics-and-renewal-
approximation-in-the-online-selection-of-increasing/
10.3150/20-BEJ1294.full.

Gong:2022:CBB

[GS22] Fuzhou Gong and Xiaoxia Sun. On the characterization of Brownian bridge measure on the pinned path space over a compact Riemannian manifold. Bernoulli, 28(4):2322–2344, November 2022. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-28/issue-4/On-the-characterization-
of-Brownian-bridge-measure-on-the-pinned/10.3150/21-
BEJ1420.full.

Germain:2023:NOD

[GS23] Gilles Germain and Yvik Swan. A note on one-dimensional Poincaré inequalities by Stein-type integration. Bernoulli, 29(2):1714–1740, May 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-29/issue-
2/A-note-on-one-dimensional-Poincar%c3%a9-inequalities-
by-Stein-type/10.3150/22-BEJ1518.full.

Gottschalk:2008:FGR

[GST08] Hanno Gottschalk, Boubaker Smii, and Horst Thaler. The Feynman graph representation of convolution semigroups and its applications to Lévy statistics. Bernoulli, 14(2):322–351, May 2008. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1208872107.

Guglielmi:2001:MCM

[GT01] Alessandra Guglielmi and Richard L. Tweedie. Markov chain Monte Carlo estimation of the law of the mean of a Dirichlet process. Bernoulli, 7(4):573–592, August 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1079559463.

Gotze:2004:RCP

[GT04] Friedrich Götze and Alexander Tikhomirov. Rate of convergence in probability to the Marchenko–Pastur law. Bernoulli, 10(3):503–548, June 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1089206408.

Geiss:2009:WCE

[GT09] Stefan Geiss and Anni Toivola. Weak convergence of error processes in discretizations of stochastic integrals and Besov spaces. Bernoulli, 15(4):925–954, November 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1262962221.

Gobet:2016:ABS

[GT16a] Emmanuel Gobet and Plamen Turkedjiev. Approximation of backward stochastic differential equations using Malliavin weights and least-squares regression. Bernoulli, 22(1):530–562, February 2016. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1443620859.

Gulisashvili:2016:TBS

[GT16b] Archil Gulisashvili and Peter Tankov. Tail behavior of sums and differences of log-normal random variables. Bernoulli, 22 (1):444–493, February 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1443620857.

Grote:2018:CMD

[GT18] Julian Grote and Christoph Thäle. Concentration and moderate deviations for Poisson polytopes and polyhedra. Bernoulli, 24(4A):2811–2841, November 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1522051226.

Gouere:2019:ESS

[GT19] Jean-Baptiste Gouéré and Marie Théret. Equivalence of some subcritical properties in continuum percolation. Bernoulli, 25(4B):3714–3733, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1569398782.

Ghosh:2022:DRS

[GT22] Satyajit Ghosh and Zhiqiang Tan. Doubly robust semiparametric inference using regularized calibrated estimation with high-dimensional data. Bernoulli, 28(3):1675–1703, August 2022. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-28/issue-3/Doubly-robust-
semiparametric-inference-using-regularized-calibrated-
estimation-with-high/10.3150/21-BEJ1378.full.

Gut:2001:MSM

[Gut01] Allan Gut. Mixed shock models. Bernoulli, 7(3):541–555, June 2001. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1080004765.

Gut:2002:MP

[Gut02] Allan Gut. On the moment problem. Bernoulli, 8(3):407–421, April 2002. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1078779876.

Granelli:2019:CLT

[GV19] Andrea Granelli and Almut E. D. Veraart. A central limit theorem for the realised covariation of a bivariate Brownian semistationary process. Bernoulli, 25(3):2245–2278, August 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1560326444.

Greenwood:1998:CFR

[GW98] Priscilla E. Greenwood and Wolfgang Wefelmeyer. Cox’s factoring of regression model likelihoods for continuous-time processes. Bernoulli, 4(1):65–80, March 1998. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1175865490.

Greenwood:1999:RMC

[GW99] Priscilla E. Greenwood and Wolfgang Wefelmeyer. Reversible Markov chains and optimality of symmetrized empirical estimators. Bernoulli, 5(1):109–123, February 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1173707097.

Giuliano:2017:ALL

[GW17a] Rita Giuliano and Michel Weber. Approximate local limit theorems with effective rate and application to random walks in random scenery. Bernoulli, 23(4B):3268–3310, November 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL https://projecteuclid.org/euclid.bj/1495505093.

Greene:2017:EBH

[GW17b] Evan Greene and Jon A. Wellner. Exponential bounds for the hypergeometric distribution. Bernoulli, 23(3):1911–1950, August 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1489737629.

Gao:2013:LRD

[GWY13] Jiti Gao, Qiying Wang, and Jiying Yin. Long-range dependent time series specification. Bernoulli, 19(5A):1714–1749, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1383661200.

Grossmann:2004:LDG

[GY04] Steffen Grossmann and Benjamin Yakir. Large deviations for global maxima of independent superadditive processes with negative drift and an application to optimal sequence alignments. Bernoulli, 10(5):829–845, October 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1099579157.

Goldstein:2009:EMP

[GZ09] Larry Goldstein and Haimeng Zhang. Efficiency of the maximum partial likelihood estimator for nested case control sampling. Bernoulli, 15(2):569–597, May 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1241444903.

Guo:2017:CTU

[GZ17] Xianping Guo and Yi Zhang. Constrained total undiscounted continuous-time Markov decision processes. Bernoulli, 23(3):1694–1736, August 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1489737622.

Gao:2018:LDA

[GZ18] Xuefeng Gao and Lingjiong Zhu. Large deviations and applications for Markovian Hawkes processes with a large initial intensity. Bernoulli, 24(4A):2875–2905, November 2018. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1522051228.

Gao:2021:PDH

[GZ21a] Fuqing Gao and Lingjiong Zhu. Precise deviations for Hawkes processes. Bernoulli, 27(1):221–248, February 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-1/Precise-deviations-for-Hawkes-processes/
10.3150/20-BEJ1235.full.

Gentil:2021:FBI

[GZ21b] Ivan Gentil and Simon Zugmeyer. A family of beckner inequalities under various curvature-dimension conditions. Bernoulli, 27(2):751–771, May 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-2/A-family-of-Beckner-inequalities-under-various-
curvature-dimension-conditions/10.3150/20-BEJ1228.full.

Haff:2013:PEP

[Haf13] Ingrid Hobæk Haff. Parameter estimation for pair-copula constructions. Bernoulli, 19(2):462–491, May 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1363192035.

Hamadene:2003:MBS

[Ham03] Saïd Hamadène. Multidimensional backward stochastic differential equations with uniformly continuous coefficients. Bernoulli, 9(3):517–534, June 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1065444816.

Hansen:2003:GED

[Han03] Niels Richard Hansen. Geometric ergodicity of discrete-time approximations to multivariate diffusions. Bernoulli, 9(4):725– 743, August 2003. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1066223276.

Handa:2009:TPP

[Han09] Kenji Handa. The two-parameter Poisson–Dirichlet point process. Bernoulli, 15(4):1082–1116, November 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1262962227.

Han:2022:MUP

[Han22] Qiyang Han. Multiplier U-processes: Sharp bounds and applications. Bernoulli, 28(1):87–124, February 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-1/Multiplier-U-processes-Sharp-bounds-and-
applications/10.3150/21-BEJ1334.full.

Harper:2017:PCD

[Har17] Adam J. Harper. Pickands’ constant H_{α} does not equal 1/Γ(1/α), for small α. Bernoulli, 23(1):582–602, February 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1475001367.

Hashorva:2013:MME

[Has13] Enkelejd Hashorva. Minima and maxima of elliptical arrays and spherical processes. Bernoulli, 19(3):886–904, August 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1372251146.

Hara:2010:MMI

[HAT10] Hisayuki Hara, Satoshi Aoki, and Akimichi Takemura. Minimal and minimal invariant Markov bases of decomposable models for contingency tables. Bernoulli, 16(1):208–233, February 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1265984709.

Hayashi:2022:IGA

[Hay22] Masahito Hayashi. Information geometry approach to parameter estimation in hidden Markov model. Bernoulli, 28(1):307– 342, February 2022. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-28/issue-1/Information-geometry-
approach-to-parameter-estimation-in-hidden-Markov-model/
10.3150/21-BEJ1344.full.

Hall:2004:NMI

[HB04] Peter Hall and Efstathia Bura. Nonparametric methods of inference for finite-state, inhomogeneous Markov processes. Bernoulli, 10(5):919–938, October 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1099579162.

Heine:2023:MBP

[HB23] Kari Heine and Daniel Burrows. Multilevel bootstrap particle filter. Bernoulli, 29(1):551–579, February 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-1/Multilevel-bootstrap-particle-filter/10.
3150/22-BEJ1468.full.

Huang:2022:HNP

[HDS22] Qiao Huang, Jinqiao Duan, and Renming Song. Homogenization of nonlocal partial differential equations related to stochastic differential equations with Lévy noise. Bernoulli, 28(3):1648–1674, August 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-28/issue-
3/Homogenization-of-nonlocal-partial-differential-equations-
related-to-stochastic-differential/10.3150/21-BEJ1365.
full.

Heino:2018:UMD

[Hei18] Joonas Heino. Uniform measure density condition and game regularity for tug-of-war games. Bernoulli, 24(1):408–432, February 2018. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1501142449.

Huang:1999:NEQ

[HF99] Li-Shan Huang and Jianqing Fan. Nonparametric estimation of quadratic regression functionals. Bernoulli, 5(5):927–949, October 1999. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1171290405.

Hooghiemstra:1997:DAS

[HG97] Gerard Hooghiemstra and Priscilla E. Greenwood. The domain of attraction of the α-sun operator for type II and type III distributions. Bernoulli, 3(4):479–489, December 1997. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1175882220.

He:2024:PSE

[HHH24] Shiyuan He, Jianhua Z. Huang, and Kejun He. Penalized spline estimation of principal components for sparse functional data: Rates of convergence. Bernoulli, 30(4):2795–2820, November 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

Hien:2017:CLE

[Hie17] Le Thi Thu Hien. On the continuity of Lyapunov exponents of random walk in random potential. Bernoulli, 23(1):522–538, February 2017. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1475001364.

Hill:2015:REI

[Hil15] Jonathan B. Hill. Robust estimation and inference for heavy tailed GARCH. Bernoulli, 21(3):1629–1669, August 2015. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1432732032.

Hinz:2003:OPP

[Hin03] Juri Hinz. Optimizing a portfolio of power-producing plants. Bernoulli, 9(4):659–669, August 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1066223273.

Hirth:1997:PAD

[Hir97] Ulrich Martin Hirth. A Poisson approximation for the Dirichlet law, the Ewens sampling formula and the Griffiths–Engen– McCloskey law by the Stein–Chen coupling method. Bernoulli, 3(2):225–232, June 1997. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1177526730.

Hirose:2016:DID

[Hir16] Yuichi Hirose. On differentiability of implicitly defined function in semi-parametric profile likelihood estimation. Bernoulli, 22 (1):589–614, February 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1443620861.

Honda:2019:AWG

[HIW19] Toshio Honda, Ching-Kang Ing, and Wei-Ying Wu. Adaptively weighted group lasso for semiparametric quantile regression models. Bernoulli, 25(4B):3311–3338, November 2019. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1569398767.

Hillion:2017:PSO

[HJ17] Erwan Hillion and Oliver Johnson. A proof of the Shepp– Olkin entropy concavity conjecture. Bernoulli, 23(4B):3638– 3649, November 2017. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1495505104.

Haggstrom:2002:CBR

[HJL02] Olle Häggström, Johan Jonasson, and Russell Lyons. Coupling and Bernoullicity in random-cluster and Potts models. Bernoulli, 8(3):275–294, April 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1078779871.

Hall:2014:FSW

[HJM14] Peter Hall, Jiashun Jin, and Hugh Miller. Feature selection when there are many influential features. Bernoulli, 20(3): 1647–1671, August 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1402488953.

Helmers:2004:SAT

[HJQZ04] Roelof Helmers, Bing-Yi Jing, Gengsheng Qin, and Wang Zhou. Saddlepoint approximations to the trimmed mean. Bernoulli, 10(3):465–501, June 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1089206407.

Houdre:2007:LSP

[HK07] Christian Houdré and Reiichiro Kawai. On layered stable processes. Bernoulli, 13(1):252–278, February 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1175287732.

Horvath:2008:SAL

[HK08] Lajos Horváth and Piotr Kokoszka. Sample autocovariances of long-memory time series. Bernoulli, 14(2):405–418, May 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1208872111.

Hoffmann:2011:SAS

[HK11] Marc Hoffmann and Nathalie Krell. Statistical analysis of self-similar conservative fragmentation chains. Bernoulli, 17 (1):395–423, February 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1297173847.

Hormann:2013:CMP

[HK13] Siegfried Hörmann and Piotr Kokoszka. Consistency of the mean and the principal components of spatially distributed functional data. Bernoulli, 19(5A):1535–1558, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1383661193.

Henning:2019:GNG

[HKK19] Florian Henning, Richard C. Kraaij, and Christof Külske. Gibbs–non-Gibbs transitions in the fuzzy Potts model with a Kac-type interaction: Closing the Ising gap. Bernoulli, 25 (3):2051–2074, August 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1560326437.

Horowitz:2006:OEA

[HKM06] Joel Horowitz, Jussi Klemelä, and Enno Mammen. Optimal estimation in additive regression models. Bernoulli, 12(2):271– 298, April 2006. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1145993975.

Huckemann:2016:CSI

[HKM^{+}16] Stephan Huckemann, Kwang-Rae Kim, Axel Munk, Florian Rehfeldt, Max Sommerfeld, Joachim Weickert, and Carina Wollnik. The circular SiZer, inferred persistence of shape parameters and application to early stem cell differentiation. Bernoulli, 22 (4):2113–2142, November 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1462297677.

Hundrieser:2024:UAD

[HKMS24] Shayan Hundrieser, Marcel Klatt, Axel Munk, and Thomas Staudt. A unifying approach to distributional limits for empirical optimal transport. Bernoulli, 30(4):2846–2877, November 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

Hult:2006:KCC

[HL06] Henrik Hult and Filip Lindskog. On Kesten’s counterexample to the Cramér–Wold device for regular variation. Bernoulli, 12 (1):133–142, February 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1141136653.

Hallin:2012:SSD

[HL12a] Marc Hallin and Christophe Ley. Skew-symmetric distributions and Fisher information — a tale of two densities. Bernoulli, 18(3):747–763, August 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1340887000.

Hirose:2012:RLF

[HL12b] Yuichi Hirose and Alan Lee. Reparametrization of the least favorable submodel in semi-parametric multisample models. Bernoulli, 18(2):586–605, May 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1334580725.

Hallin:2014:SSD

[HL14] Marc Hallin and Christophe Ley. Skew-symmetric distributions and Fisher information: The double sin of the skewnormal. Bernoulli, 20(3):1432–1453, August 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1402488945.

Han:2017:SAL

[HL17] Fang Han and Han Liu. Statistical analysis of latent generalized correlation matrix estimation in transelliptical distribution. Bernoulli, 23(1):23–57, February 2017. CODEN ????

ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1475001347.

Hallin:2023:RBT

[HLL23] Marc Hallin, Davide La Vecchia, and Hang Liu. Rankbased testing for semiparametric VAR models: a measure transportation approach. Bernoulli, 29(1):229–273, February 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-29/issue-1/Rank-based-
testing-for-semiparametric-VAR-models--A-measure/10.
3150/21-BEJ1456.full.

Herve:2012:UBE

[HLP12] Loïc Hervé, James Ledoux, and Valentin Patilea. A uniform Berry–Esseen theorem on M-estimators for geometrically ergodic Markov chains. Bernoulli, 18(2):703–734, May 2012. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1334580730.

Hong:2023:ARP

[HLP23] Hyerim Hong, Young Kyung Lee, and Byeong U. Park. Additive regression with parametric help. Bernoulli, 29(4):3059– 3092, November 2023. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-29/issue-4/Additive-regression-
with-parametric-help/10.3150/22-BEJ1575.full.

Hall:2009:TRB

[HLPP09] Peter Hall, Young K. Lee, Byeong U. Park, and Debashis Paul. Tie-respecting bootstrap methods for estimating distributions of sets and functions of eigenvalues. Bernoulli, 15(2):380–401, May 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1241444895.

Hallin:2015:LBM

[HLPŠ15] Marc Hallin, Zudi Lu, Davy Paindaveine, and Miroslav Šiman. Local bilinear multiple-output quantile/depth regression. Bernoulli, 21(3):1435–1466, August 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1432732026.

Heinrich:2014:AGF

[HLS14] Lothar Heinrich, Sebastian Lück, and Volker Schmidt. Asymptotic goodness-of-fit tests for the Palm mark distribution of stationary point processes with correlated marks. Bernoulli, 20 (4):1673–1697, November 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1411134440.

He:2019:CMC

[HLS19] Yuanzhen He, C. Devon Lin, and Fasheng Sun. Construction of marginally coupled designs by subspace theory. Bernoulli, 25(3):2163–2182, August 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1560326441.

Hallin:2001:DES

[HLT01] Marc Hallin, Zudi Lu, and Lanh Tat Tran. Density estimation for spatial linear processes. Bernoulli, 7(4):657–668, August 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1079559468.

Hopfner:2017:SDT

[HLT17] R. Höpfner, E. Löcherbach, and M. Thieullen. Strongly degenerate time inhomogeneous SDEs: Densities and support properties. Application to Hodgkin–Huxley type systems. Bernoulli, 23(4A):2587–2616, November 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1494316826.

Hallin:2009:LLS

[HLY09] Marc Hallin, Zudi Lu, and Keming Yu. Local linear spatial quantile regression. Bernoulli, 15(3):659–686, August 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1251463276.

Hwang:2020:LLT

[HLY20] Jong Yun Hwang, Ji Oon Lee, and Wooseok Yang. Local law and Tracy–Widom limit for sparse stochastic block models. Bernoulli, 26(3):2400–2435, August 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1587974546.

Hu:2023:SRP

[HLZZ23] Xiangbin Hu, Li Liu, Ying Zhang, and Xingqiu Zhao. Semiparametric regression of panel count data with informative terminal event. Bernoulli, 29(4):2828–2853, November 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-4/Semiparametric-regression-of-panel-count-
data-with-informative-terminal-event/10.3150/22-BEJ1565.
full.

Haas:2011:SSS

[HM11] Bénédicte Haas and Grégory Miermont. Self-similar scaling limits of non-increasing Markov chains. Bernoulli, 17(4): 1217–1247, November 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1320417502.

Houdre:2015:SLD

[HM15] Christian Houdré and Jinyong Ma. Simultaneous large deviations for the shape of Young diagrams associated with random words. Bernoulli, 21(3):1494–1537, August 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1432732028.

Heiny:2019:ESC

[HM19] Johannes Heiny and Thomas Mikosch. The eigenstructure of the sample covariance matrices of high-dimensional stochastic volatility models with heavy tails. Bernoulli, 25(4B):3590– 3622, November 2019. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1569398778.

Hauser:2020:MPS

[HM20a] Raphael Andreas Hauser and Heinrich Matzinger. Microscopic path structure of optimally aligned random sequences. Bernoulli, 26(1):1–30, February 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1574758820.

Hirsch:2020:DLD

[HM20b] Christian Hirsch and Christian Mönch. Distances and large deviations in the spatial preferential attachment model. Bernoulli, 26(2):927–947, May 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1580461568.

Holzmann:2020:RON

[HM20c] Hajo Holzmann and Alexander Meister. Rate-optimal nonparametric estimation for random coefficient regression models. Bernoulli, 26(4):2790–2814, November 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
26/issue-4/Rate-optimal-nonparametric-estimation-for-
random-coefficient-regression-models/10.3150/20-BEJ1207.
full.

Huang:2023:VFA

[HM23] Lu-Jing Huang and Yong-Hua Mao. Variational formulas for asymptotic variance of general discrete-time Markov chains. Bernoulli, 29(1):300–322, February 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-1/Variational-formulas-for-asymptotic-variance-
of-general-discrete-time-Markov/10.3150/21-BEJ1458.
full.

Han:2018:SBA

[HMP18a] Kyunghee Han, Hans-Georg Müller, and Byeong U. Park. Smooth backfitting for additive modeling with small errors-in-variables, with an application to additive functional regression for multiple predictor functions. Bernoulli, 24(2):1233–1265, May 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1505980895.

Hauser:2018:UBC

[HMP18b] Raphael Hauser, Heinrich Matzinger, and Ionel Popescu. An upper bound on the convergence rate of a second functional in optimal sequence alignment. Bernoulli, 24(2):971–992, May 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1505980885.

Houdre:2008:CNI

[HMRB08] Christian Houdré, Philippe Marchal, and Patricia Reynaud-Bouret. Concentration for norms of infinitely divisible vectors with independent components. Bernoulli, 14(4):926–948, November 2008. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1225980565.

Hutter:2020:EMM

[HMRR20] Jan-Christian Hütter, Cheng Mao, Philippe Rigollet, and Elina Robeva. Estimation of Monge matrices. Bernoulli, 26(4):3051– 3080, November 2020. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-26/issue-4/Estimation-of-
Monge-matrices/10.3150/20-BEJ1215.full.

Hobert:2004:STM

[HMS04] James P. Hobert, Dobrin Marchev, and Jason Schweinsberg. Stability of the tail Markov chain and the evaluation of improper priors for an exponential rate parameter. Bernoulli, 10(3):549–564, June 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1089206409.

Haeusler:2000:SSR

[HMT00] Erich Haeusler, David M. Mason, and Tatyana S. Turova. A study of serial ranks via random graphs. Bernoulli, 6(3):541– 570, June 2000. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1081616704.

Hazelton:2024:SRM

[HMTvB24] Martin Hazelton, Michael McVeagh, Christopher Tuffley, and Bruce van Brunt. Some rapidly mixing hit-and-run samplers for latent counts in linear inverse problems. Bernoulli, 30(4): 2676–2699, November 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

Hall:2001:BCB

[HMW01] Peter Hall, Gavin Melville, and Alan H. Welsh. Bias correction and bootstrap methods for a spatial sampling scheme. Bernoulli, 7(6):829–846, December 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1078951125.

Huang:2021:AHT

[HMW21] Lu-Jing Huang, Mateusz B. Majka, and Jian Wang. Approximation of heavy-tailed distributions via stable-driven SDEs. Bernoulli, 27(3):2040–2068, August 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-3/Approximation-of-heavy-tailed-distributions-
via-stable-driven-SDEs/10.3150/20-BEJ1300.full.

He:2010:FLR

[HMWY10] Guozhong He, Hans-Georg Müller, Jane-Ling Wang, and Wenjing Yang. Functional linear regression via canonical analysis. Bernoulli, 16(3):705–729, August 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1281099881.

Hinz:2010:FPE

[HN10] Juri Hinz and Alex Novikov. On fair pricing of emission-related derivatives. Bernoulli, 16(4):1240–1261, November 2010. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1290092905.

Ho:2020:REM

[HNR20] Nhat Ho, XuanLong Nguyen, and Ya’acov Ritov. Robust estimation of mixing measures in finite mixture models. Bernoulli, 26(2):828–857, May 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1580461565.

Henderson:2023:CSO

[HNR23] Iain Henderson, Pascal Noble, and Olivier Roustant. Characterization of the second order random fields subject to linear distributional PDE constraints. Bernoulli, 29(4):3396–3422, November 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-29/issue-4/Characterization-of-the-
second-order-random-fields-subject-to-linear/10.3150/
23-BEJ1588.full.

Hu:2015:DCB

[HNTX15] Yaozhong Hu, David Nualart, Samy Tindel, and Fangjun Xu. Density convergence in the Breuer–Major theorem for Gaussian stationary sequences. Bernoulli, 21(4):2336–2350, November 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1438777596.

Hoff:2014:IBG

[HNW14] Peter D. Hoff, Xiaoyue Niu, and Jon A. Wellner. Information bounds for Gaussian copulas. Bernoulli, 20(2):604–622, May 2014. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1393593999.

Hu:2018:LDS

[HNZ18] Yaozhong Hu, David Nualart, and Tusheng Zhang. Large deviations for stochastic heat equation with rough dependence in space. Bernoulli, 24(1):354–385, February 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1501142447.

Hofmann:1999:LED

[Hof99] Marc Hofmann. Lp estimation of the diffusion coefficient. Bernoulli, 5(3):447–481, June 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1172617199.

Hoff:2023:BOP

[Hof23] Peter Hoff. Bayes-optimal prediction with frequentist coverage control. Bernoulli, 29(2):901–928, May 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-2/Bayes-optimal-prediction-with-frequentist-
coverage-control/10.3150/22-BEJ1484.full.

Hooker:2016:CER

[Hoo16] Giles Hooker. Consistency, efficiency and robustness of conditional disparity methods. Bernoulli, 22(2):857–900, May 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1447077763.

Hormann:2008:AGS

[Hör08] Siegfried Hörmann. Augmented GARCH sequences: Dependence structure and asymptotics. Bernoulli, 14(2):543–561, May 2008. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1208872117.

Hossjer:2005:SDS

[Hös05] Ola Hössjer. Spectral decomposition of score functions in linkage analysis. Bernoulli, 11(6):1093–1113, December 2005. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1137421641.

Hall:1995:WME

[HP95] Peter Hall and Prakash Patil. On wavelet methods for estimating smooth functions. Bernoulli, 1(1–2):41–58, March 1995. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1186078361.

Heikkinen:1999:BSE

[HP99] Juha Heikkinen and Antti Penttinen. Bayesian smoothing in the estimation of the pair potential function of Gibbs point processes. Bernoulli, 5(6):1119–1136, December 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1143122305.

Hall:2001:CVC

[HP01] Peter Hall and Spiridon Penev. Cross-validation for choosing resolution level for nonlinear wavelet curve estimators. Bernoulli, 7(2):317–341, April 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1080222088.

Hallin:2002:OPB

[HP02a] Marc Hallin and Davy Paindaveine. Optimal procedures based on interdirections and pseudo-Mahalanobis ranks for testing multivariate elliptic white noise against ARMA dependence. Bernoulli, 8(6):787–815, December 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1076364806.

Houdre:2002:CDI

[HP02b] Christian Houdré and Nicolas Privault. Concentration and deviation inequalities in infinite dimensions via covariance representations. Bernoulli, 8(6):697–720, December 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1076364802.

Hansen:2006:NIM

[HP06] Martin B. Hansen and Susan M. Pitts. Nonparametric inference from the M/G/1 workload. Bernoulli, 12(4):737–759, August 2006. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1155735934.

Hintze:2014:SNA

[HP14] Robert Hintze and Ilya Pavlyukevich. Small noise asymptotics and first passage times of integrated Ornstein–Uhlenbeck processes driven by α-stable Lévy processes. Bernoulli, 20 (1):265–281, February 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1390407289.

Hallin:2013:ORB

[HPV13] Marc Hallin, Davy Paindaveine, and Thomas Verdebout. Optimal rank-based tests for common principal components. Bernoulli, 19(5B):2524–2556, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1386078612.

Han:2018:UMM

[HPY18] Xiao Han, Guangming Pan, and Qing Yang. A unified matrix model including both CCA and f matrices in multivariate analysis: The largest eigenvalue and its applications. Bernoulli, 24(4B):3447–3468, November 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1524038759.

Horvath:2013:TSF

[HR13] Lajos Horváth and Ron Reeder. A test of significance in functional quadratic regression. Bernoulli, 19(5A):2120–2151, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1383661216.

Huang:2015:FPC

[HR15] Qing Huang and Rosemary Renaut. Functional partial canonical correlation. Bernoulli, 21(2):1047–1066, May 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1429624970.

Heaukulani:2016:CSB

[HR16] Creighton Heaukulani and Daniel M. Roy. The combinatorial structure of beta negative binomial processes. Bernoulli, 22(4): 2301–2324, November 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1462297682.

Hart:2022:PPN

[HR22a] Sergiu Hart and Yosef Rinott. Posterior probabilities: Nonmonotonicity, asymptotic rates, log-concavity, and Turán’s inequality. Bernoulli, 28(2):1461–1490, May 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-2/Posterior-probabilities--Nonmonotonicity-
asymptotic-rates-log-concavity-and-Tur%c3%a1ns/10.3150/
21-BEJ1398.full.

Howard:2022:SEQ

[HR22b] Steven R. Howard and Aaditya Ramdas. Sequential estimation of quantiles with applications to A/B testing and best-arm identification. Bernoulli, 28(3):1704–1728, August 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-3/Sequential-estimation-of-quantiles-with-
applications-to-A-B-testing/10.3150/21-BEJ1388.full.

Hansen:2015:LPI

[HRBR15] Niels Richard Hansen, Patricia Reynaud-Bouret, and Vincent Rivoirard. Lasso and probabilistic inequalities for multivariate point processes. Bernoulli, 21(1):83–143, February 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1426597065.

Hoffmann:2013:ELL

[HRY13] M. Hoffmann, M. Rosenbaum, and N. Yoshida. Estimation of the lead-lag parameter from non-synchronous data. Bernoulli, 19(2):426–461, May 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1363192034.

Hu:1997:ELB

[HS97] Yueyun Hu and Zhan Shi. Extreme lengths in Brownian and Bessel excursions. Bernoulli, 3(4):387–402, December 1997. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1175882214.

Haeusler:2007:ACI

[HS07] Erich Haeusler and Johan Segers. Assessing confidence intervals for the tail index by Edgeworth expansions for the Hill estimator. Bernoulli, 13(1):175–194, February 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1175287728.

Hult:2008:TPI

[HS08] Henrik Hult and Gennady Samorodnitsky. Tail probabilities for infinite series of regularly varying random vectors. Bernoulli, 14(3):838–864, August 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1219669632.

He:2010:APM

[HS10] Heping He and Thomas A. Severini. Asymptotic properties of maximum likelihood estimators in models with multiple change points. Bernoulli, 16(3):759–779, August 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1281099883.

Heinrich:2007:LTF

[HSS07] Lothar Heinrich, Hendrik Schmidt, and Volker Schmidt. Limit theorems for functionals on the facets of stationary random tessellations. Bernoulli, 13(3):868–891, August 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1186503491.

Haario:2001:AMA

[HST01] Heikki Haario, Eero Saksman, and Johanna Tamminen. An adaptive Metropolis algorithm. Bernoulli, 7(2):223–242, April 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1080222083.

Hall:2000:DDF

[HT00a] Peter Hall and Nader Tajvidi. Distribution and dependencefunction estimation for bivariate extreme-value distributions. Bernoulli, 6(5):835–844, October 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1081282691.

Hoshino:2000:RFS

[HT00b] Nobuaki Hoshino and Akimichi Takemura. On reduction of finite-sample variance by extended Latin hypercube sampling. Bernoulli, 6(6):1035–1050, December 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1081194159.

Huesmann:2019:BBF

[HT19] Martin Huesmann and Dario Trevisan. A Benamou-Brenier formulation of martingale optimal transport. Bernoulli, 25 (4A):2729–2757, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1568362041.

Huang:2021:PIN

[HT21] De Huang and Joel A. Tropp. From Poincaré inequalities to nonlinear matrix concentration. Bernoulli, 27(3):1724–1744, August 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-27/issue-3/From-Poincar%c3%a9-inequalities-
to-nonlinear-matrix-concentration/10.3150/20-BEJ1289.
full.

Hobert:2007:WEM

[HTL07] James P. Hobert, Aixin Tan, and Ruitao Liu. When is Eaton’s Markov chain irreducible? Bernoulli, 13(3):641–652, August 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1186503480.

Hall:2006:LLA

[HTM06] Peter Hall, Nader Tajvidi, and P. E. Malin. Locating lines among scattered points. Bernoulli, 12(5):821–839, October 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1161614948.

Han:2023:HDG

[HTW23] Yuefeng Han, Ruey S. Tsay, and Wei Biao Wu. High dimensional generalized linear models for temporal dependent data. Bernoulli, 29(1):105–131, February 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-1/High-dimensional-generalized-linear-models-
for-temporal-dependent-data/10.3150/21-BEJ1451.full.

Hilario:2022:NPT

[HU22] Marcelo Hilário and Daniel Ungaretti. A note on the phase transition for independent alignment percolation. Bernoulli, 28(2): 1432–1447, May 2022. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-28/issue-2/A-note-on-the-
phase-transition-for-independent-alignment-percolation/
10.3150/21-BEJ1395.full.

Huber:2012:SBD

[Hub12] Mark Huber. Spatial birth–death swap chains. Bernoulli, 18 (3):1031–1041, August 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1340887012.

Hall:2009:NRW

[HV09] Peter Hall and Ingrid Van Keilegom. Nonparametric “regression” when errors are positioned at end-points. Bernoulli, 15 (3):614–633, August 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1251463274.

Haggstrom:1999:CRM

[HVM99] Olle Häggström, Marie-Colette N. M. Van Lieshout, and Jesper Møller. Characterization results and Markov chain Monte Carlo algorithms including exact simulation for some spatial point processes. Bernoulli, 5(4):641–658, August 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1171899321.

Hestir:1995:SDS

[HW95] Kevin Hestir and Stanley C. Williams. Supports of doubly stochastic measures. Bernoulli, 1(3):217–243, September 1995.

CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1193667816.

Hall:1999:EEP

[HW99] Peter Hall and Julian Z. Wang. Estimating the end-point of a probability distribution using minimum-distance methods. Bernoulli, 5(1):177–189, February 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1173707100.

Hallin:2003:SPE

[HW03] Marc Hallin and Bas J. M. Werker. Semi-parametric efficiency, distribution-freeness and invariance. Bernoulli, 9(1):137–165, February 2003. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1068129013.

Hall:2010:SAL

[HW10] Peter Hall and Qiying Wang. Strong approximations of level exceedences related to multiple hypothesis testing. Bernoulli, 16(2):418–434, May 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1274821077.

Hitczenko:2011:RDP

[HW11] Paweł Hitczenko and Jacek Wesołowski. Renorming divergent perpetuities. Bernoulli, 17(3):880–894, August 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1310042848.

He:2019:GHP

[HW19a] Hui He and Matthias Winkel. Gromov–Hausdorff–Prokhorov convergence of vertex cut-trees of n-leaf Galton–Watson trees. Bernoulli, 25(3):2301–2329, August 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1560326446.

Hu:2019:SDK

[HW19b] Shulan Hu and Xinyu Wang. Subexponential decay in kinetic Fokker–Planck equation: Weak hypocoercivity. Bernoulli, 25(1):174–188, February 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1544605243.

Huang:2024:REE

[HW24] Xing Huang and Feng-Yu Wang. Regularities and exponential ergodicity in entropy for SDEs driven by distribution dependent noise. Bernoulli, 30(4):3303–3323, November 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

Hobson:1999:TEC

[HWW99] David G. Hobson, David Williams, and Andrew T. A. Wood. Taylor expansions of curve-crossing probabilities. Bernoulli, 5 (5):779–795, October 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1171290399.

Hu:2011:TBP

[HWW11] Yonggang Hu, Yong Wang, and Yi Wu. Tensor-based projection depth. Bernoulli, 17(4):1386–1399, November 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1320417509.

Hu:2024:MAS

[HWXZ24] Yaozhong Hu, Xiong Wang, Panqiu Xia, and Jiayu Zheng. Moment asymptotics for super-Brownian motions. Bernoulli, 30 (4):3119–3136, November 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

Hu:2023:SRH

[HWZZ23] Jiang Hu, Shaochen Wang, Yangchun Zhang, and Wang Zhou. Sampling without replacement from a high-dimensional finite population. Bernoulli, 29(4):3198–3220, November 2023. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-4/Sampling-without-replacement-from-a-high-
dimensional-finite-population/10.3150/22-BEJ1580.full.

Han:2018:GCI

[HXZ18] Fang Han, Sheng Xu, and Wen-Xin Zhou. On Gaussian comparison inequality and its application to spectral analysis of large random matrices. Bernoulli, 24(3):1787–1833, August 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL https://projecteuclid.org/euclid.bj/1517540460.

Hayashi:2005:CEN

[HY05] Takaki Hayashi and Nakahiro Yoshida. On covariance estimation of non-synchronously observed diffusion processes. Bernoulli, 11(2):359–379, April 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1116340299.

Hu:2008:ANC

[HY08] Jianhua Hu and Guohua Yan. Asymptotic normality and consistency of a two-stage generalized least squares estimator in the growth curve model. Bernoulli, 14(3):623–636, August 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1219669623.

Hirsch:2013:MTP

[HY13] Francis Hirsch and Marc Yor. On the Mellin transforms of the perpetuity and the remainder variables associated to a subordinator. Bernoulli, 19(4):1350–1377, September 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1377612855.

Hu:2011:EAG

[HYY11] Jianhua Hu, Guohua Yan, and Jinhong You. Estimation for an additive growth curve model with orthogonal design matrices. Bernoulli, 17(4):1400–1419, November 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1320417510.

Hu:2004:ANU

[HZ04] Feifang Hu and Li-Xin Zhang. Asymptotic normality of urn models for clinical trials with delayed response. Bernoulli, 10(3):447–463, June 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1089206406.

Hou:2023:SEN

[HZBH23] Zhiqiang Hou, Xiaozhuo Zhang, Zhidong Bai, and Jiang Hu. Spiked eigenvalues of noncentral Fisher matrix with applications. Bernoulli, 29(4):3171–3197, November 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-4/Spiked-eigenvalues-of-noncentral-Fisher-
matrix-with-applications/10.3150/22-BEJ1579.full.

Ing:2016:EIA

[ICG16] Ching-Kang Ing, Hai-Tang Chiou, and Meihui Guo. Estimation of inverse autocovariance matrices for long memory processes. Bernoulli, 22(3):1301–1330, August 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1458132983.

Irurozki:2019:MGM

[ICL19] Ekhine Irurozki, Borja Calvo, and Jose A. Lozano. Mallows and generalized Mallows model for matchings. Bernoulli, 25(2): 1160–1188, May 2019. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1551862847.

Ichiba:2022:DCT

[IK22] Tomoyuki Ichiba and Ioannis Karatzas. Degenerate competing three-particle systems. Bernoulli, 28(3):2067–2094, August 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-28/issue-3/Degenerate-
competing-three-particle-systems/10.3150/21-BEJ1411.
full.

Ichiba:2013:DRB

[IKP13] Tomoyuki Ichiba, Ioannis Karatzas, and Vilmos Prokaj. Diffusions with rank-based characteristics and values in the nonnegative quadrant. Bernoulli, 19(5B):2455–2493, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1386078610.

Ispany:2014:ABC

[IKP14] Márton Ispány, Kristóf Körmendi, and Gyula Pap. Asymptotic behavior of CLS estimators for 2-type doubly symmetric critical Galton–Watson processes with immigration. Bernoulli, 20(4): 2247–2277, November 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1411134459.

Inoue:2018:BIF

[IKP18] Akihiko Inoue, Yukio Kasahara, and Mohsen Pourahmadi. Baxter’s inequality for finite predictor coefficients of multivariate long-memory stationary processes. Bernoulli, 24(2):1202–1232, May 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1505980894.

Ing:2009:TOM

[ILY09] Ching-Kang Ing, Jin-Lung Lin, and Shu-Hui Yu. Toward optimal multistep forecasts in non-stationary autoregressions. Bernoulli, 15(2):402–437, May 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1241444896.

Iksanov:2009:NCB

[IMM09] Alex Iksanov, Alex Marynych, and Martin Möhle. On the number of collisions in beta(2, b )-coalescents. Bernoulli, 15(3):829– 845, August 2009. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1251463283.

Iksanov:2017:ARPa

[IMM17a] Alexander Iksanov, Alexander Marynych, and Matthias Meiners. Asymptotics of random processes with immigration I: Scaling limits. Bernoulli, 23(2):1233–1278, May 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1486177398.

Iksanov:2017:ARPb

[IMM17b] Alexander Iksanov, Alexander Marynych, and Matthias Meiners. Asymptotics of random processes with immigration II: Convergence to stationarity. Bernoulli, 23(2):1279–1298, May 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1486177399.

Iouditsky:2001:AEF

[IMS01] Anatoli Iouditsky, Eric Moulines, and Philippe Soulier. Adaptive estimation of the fractional differencing coefficient. Bernoulli, 7(5):699–731, October 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1079399538.

Iksanov:2023:FLT

[IPP23] Alexander Iksanov, Andrey Pilipenko, and Ben Povar. Functional limit theorems for random walks perturbed by positive alpha-stable jumps. Bernoulli, 29(2):1638–1662, May 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-29/issue-2/Functional-
limit-theorems-for-random-walks-perturbed-by-positive-
alpha/10.3150/22-BEJ1515.full.

Issoglio:2020:FKR

[IR20] Elena Issoglio and Francesco Russo. A Feynman–Kac result via Markov BSDEs with generalised drivers. Bernoulli, 26(1):728– 766, February 2020. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1574758844.

Iacobelli:2022:TBR

[IRVZ22] Giulio Iacobelli, Rodrigo Ribeiro, Glauco Valle, and Leonel Zuaznábar. Tree builder random walk: Recurrence, transience and ballisticity. Bernoulli, 28(1):150–180, February 2022. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-1/Tree-builder-random-walk-Recurrence-transience-
and-ballisticity/10.3150/21-BEJ1337.full. See erratum

[IRVZ24].

Iacobelli:2024:ETB

[IRVZ24] Giulio Iacobelli, Rodrigo Ribeiro, Glauco Valle, and Leonel Zuaznábar. Erratum: Tree builder random walk: Recurrence, transience and ballisticity. Bernoulli, 30(4):3378–3387, November 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). See [IRVZ22].

Istas:1998:DOW

[IS98] Jacques Istas and Henrik Stryhn. Discretely observing a white noise change-point model in the presence of blur. Bernoulli, 4(2):185–201, June 1998. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1174937294.

Ichiba:2019:SDC

[IS19] Tomoyuki Ichiba and Andrey Sarantsev. Stationary distributions and convergence for Walsh diffusions. Bernoulli, 25 (4A):2439–2478, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1568362032.

Jaber:2021:WEU

[Jab21] Eduardo Abi Jaber. Weak existence and uniqueness for affine stochastic Volterra equations with. Bernoulli, 27(3):1583– 1615, August 2021. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-27/issue-3/Weak-existence-
and-uniqueness-for-affine-stochastic-Volterra-equations-
with/10.3150/20-BEJ1284.full.

Jacobsen:1996:LOO

[Jac96] Martin Jacobsen. Laplace and the origin of the Ornstein– Uhlenbeck process. Bernoulli, 2(3):271–286, September 1996. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1178291723.

Jacobsen:2002:OSO

[Jac02] Martin Jacobsen. Optimality and small Δ-optimality of martingale estimating functions. Bernoulli, 8(5):643–668, October 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1078435222.

Jacod:2006:PID

[Jac06] Jean Jacod. Parametric inference for discretely observed nonergodic diffusions. Bernoulli, 12(3):383–401, June 2006. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1151525127.

Jagers:1995:BPP

[Jag95] Peter Jagers. Branching processes as population dynamics. Bernoulli, 1(1–2):191–200, March 1995. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1186078367.

James:2010:DMI

[Jam10] Lancelot F. James. Dirichlet mean identities and laws of a class of subordinators. Bernoulli, 16(2):361–388, May 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1274821075.

Janson:2007:MAN

[Jan07] Svante Janson. Monotonicity, asymptotic normality and vertex degrees in random graphs. Bernoulli, 13(4):952–965, November 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1194625597.

Jiang:2021:GFM

[JB21a] Dandan Jiang and Zhidong Bai. Generalized four moment theorem and an application to CLT for spiked eigenvalues of high-dimensional covariance matrices. Bernoulli, 27(1):274– 294, February 2021. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-27/issue-1/Generalized-four-
moment-theorem-and-an-application-to-CLT-for/10.3150/
20-BEJ1237.full.

Jiang:2021:PGF

[JB21b] Dandan Jiang and Zhidong Bai. Partial generalized four moment theorem revisited. Bernoulli, 27(4):2337–2352, November 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-27/issue-4/Partial-generalized-four-
moment-theorem-revisited/10.3150/20-BEJ1310.full.

Jiang:2020:DLD

[JCL20] Binyan Jiang, Ziqi Chen, and Chenlei Leng. Dynamic linear discriminant analysis in high dimensional space. Bernoulli, 26(2): 1234–1268, May 2020. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1580461578.

Janssen:2016:SVM

[JD16] Anja Janssen and Holger Drees. A stochastic volatility model with flexible extremal dependence structure. Bernoulli, 22 (3):1448–1490, August 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1458132988.

Jauch:2020:ROM

[JHD20] Michael Jauch, Peter D. Hoff, and David B. Dunson. Random orthogonal matrices and the Cayley transform. Bernoulli, 26(2):1560–1586, May 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1580461590.

Jiang:2013:LTB

[Jia13] Tiefeng Jiang. Limit theorems for beta-Jacobi ensembles. Bernoulli, 19(3):1028–1046, August 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1372251152.

Jirak:2018:RCH

[Jir18] Moritz Jirak. Rate of convergence for Hilbert space valued processes. Bernoulli, 24(1):202–230, February 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1501142440.

Jirak:2023:BEB

[Jir23] Moritz Jirak. A Berry–Esseen bound with (almost) sharp dependence conditions. Bernoulli, 29(2):1219–1245, May 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-29/issue-2/A-Berry--Esseen-bound-with-
almost-sharp-dependence-conditions/10.3150/22-BEJ1496.
full.

Jun:2022:TWL

[JJLW22] Wen Jun, Xie Jiahui, Yu Long, and Zhou Wang. Tracy–Widom limit for the largest eigenvalue of high-dimensional covariance matrices in elliptical distributions. Bernoulli, 28(4):2941–2967, November 2022. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-28/issue-4/Tracy-Widom-limit-
for-the-largest-eigenvalue-of-high-dimensional/10.3150/
21-BEJ1443.full.

Jiang:2024:ELR

[JJW24] Jiancheng Jiang, Xuejun Jiang, and Haofeng Wang. Empirical likelihood ratio tests for non-nested model selection based on predictive losses. Bernoulli, 30(2):1458–1481, May 2024. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
30/issue-2/Empirical-likelihood-ratio-tests-for-non-
nested-model-selection-based/10.3150/23-BEJ1640.full.

Jureckova:2012:NMR

[JK12] Jana Jurečková and Jan Kalina. Nonparametric multivariate rank tests and their unbiasedness. Bernoulli, 18(1):229–251, February 2012. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1327068624.

Jahnel:2017:STG

[JK17] Benedikt Jahnel and Christof Külske. Sharp thresholds for Gibbs-non-Gibbs transitions in the fuzzy Potts model with a Kac-type interaction. Bernoulli, 23(4A):2808–2827, November 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1494316833.

Jentsch:2021:BHE

[JK21] Carsten Jentsch and Rafał Kulik. Bootstrapping Hill estimator and tail array sums for regularly varying time series. Bernoulli, 27(2):1409–1439, May 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-2/Bootstrapping-Hill-estimator-and-tail-array-
sums-for-regularly-varying/10.3150/20-BEJ1279.full.

Jahnel:2023:GNG

[JK23] Benedikt Jahnel and Christof Külske. Gibbsianness and non-Gibbsianness for Bernoulli lattice fields under removal of isolated sites. Bernoulli, 29(4):3013–3032, November 2023. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-4/Gibbsianness-and-non-Gibbsianness-for-Bernoulli-
lattice-fields-under-removal/10.3150/22-BEJ1572.full.

Jureckova:2016:BEU

[JKNP16] Jana Jurečková, Hira L. Koul, Radim Navrátil, and Jan Picek. Behavior of R-estimators under measurement errors. Bernoulli, 22(2):1093–1112, May 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1447077770.

Juditsky:2004:MDE

[JLL04] Anatoli Juditsky and Sophie Lambert-Lacroix. On minimax density estimation on R. Bernoulli, 10(2):187–220, April 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1082380217.

Jourdain:2014:OST

[JLM14] Benjamin Jourdain, Tony Lelièvre, and Błażej Miasojedow. Optimal scaling for the transient phase of Metropolis Hastings algorithms: The longtime behavior. Bernoulli, 20(4):1930–1978, November 2014. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1411134449.

Jeon:2022:LPH

[JLMP22] Jeong Min Jeon, Young Kyung Lee, Enno Mammen, and Byeong U. Park. Locally polynomial Hilbertian additive regression. Bernoulli, 28(3):2034–2066, August 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-3/Locally-polynomial-Hilbertian-additive-regression/
10.3150/21-BEJ1410.full.

James:2010:PDC

[JLP10] Lancelot F. James, Antonio Lijoi, and Igor Prünster. On the posterior distribution of classes of random means. Bernoulli, 16(1):155–180, February 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1265984707.

Jacod:2008:EBD

[JLT08] Jean Jacod, Antoine Lejay, and Denis Talay. Estimation of the Brownian dimension of a continuous Itô process. Bernoulli, 14(2):469–498, May 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1208872114.

Jorgensen:1997:TTI

[JM97] Bent Jørgensen and José Raúl Martínez. Tauber theory for infinitely divisible variance functions. Bernoulli, 3(2):213–224, June 1997. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1177526729.

Jien:2009:SDE

[JM09] Yu-Juan Jien and Jin Ma. Stochastic differential equations driven by fractional Brownian motions. Bernoulli, 15(3):846– 870, August 2009. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1251463284.

Jovanovski:2017:CRH

[JM17] Oliver Jovanovski and Neal Madras. Convergence rates for a hierarchical Gibbs sampler. Bernoulli, 23(1):603–625, February 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1475001368.

Jourdain:2022:MWI

[JM22] Benjamin Jourdain and William Margheriti. Martingale Wasserstein inequality for probability measures in the convex order. Bernoulli, 28(2):830–858, May 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-2/Martingale-Wasserstein-inequality-for-probability-
measures-in-the-convex-order/10.3150/21-BEJ1368.full.

Jacob:2017:GCP

[JMM17] Christine Jacob, Manuel Molina, and Manuel Mota. A general class of population-dependent two-sex processes with random mating. Bernoulli, 23(3):1737–1758, August 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1489737623.

Janssen:2018:ESC

[JMRX18] Anja Janssen, Thomas Mikosch, Mohsen Rezapour, and Xiaolei Xie. The eigenvalues of the sample covariance matrix of a multivariate heavy-tailed stochastic volatility model. Bernoulli, 24(2):1351–1393, May 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1505980898.

Jammalamadaka:2020:STU

[JMV20] Sreenivasa Rao Jammalamadaka, Simos Meintanis, and Thomas Verdebout. On Sobolev tests of uniformity on the circle with an extension to the sphere. Bernoulli, 26(3):2226–2252, August 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1587974539.

Jourdain:2005:PAI

[JMW05] Benjamin Jourdain, Sylvie Méléard, and Wojbor A. Woyczynski. Probabilistic approximation and inviscid limits for onedimensional fractional conservation laws. Bernoulli, 11(4):689– 714, August 2005. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1126126765.

Joshi:1995:JDM

[JN95] S. N. Joshi and H. N. Nagaraja. Joint distribution of maxima of concomitants of subsets of order statistics. Bernoulli, 1 (3):245–255, September 1995. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1193667817.

Jensen:2000:IMP

[JN00] Eva B. Vedel Jensen and Linda Stougaard Nielsen. Inhomogeneous Markov point processes by transformation. Bernoulli, 6(5):761–782, October 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1081282688.

Juditsky:2015:DHO

[JN15] Anatoli Juditsky and Arkadi Nemirovski. On detecting harmonic oscillations. Bernoulli, 21(2):1134–1165, May 2015. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1429624973.

Johansen:2016:AFS

[JN16] Søren Johansen and Bent Nielsen. Analysis of the forward search using some new results for martingales and empirical processes. Bernoulli, 22(2):1131–1183, May 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:/

/projecteuclid.org/euclid.bj/1447077772. See corrigen-

dum [BRJN19].

Juditsky:2020:ELQ

[JN20] Anatoli Juditsky and Arkadi Nemirovski. Estimating linear and quadratic forms via indirect observations. Bernoulli, 26(4):2639–2669, November 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-26/issue-

4/Estimating-linear-and-quadratic-forms-via-indirect-
observations/10.3150/20-BEJ1200.full.

Jonsson:2011:HTS

[Jon11] Fredrik Jonsson. On the heavy-tailedness of Student’s tstatistic. Bernoulli, 17(1):276–289, February 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1297173843.

Jordan:2013:SCS

[Jor13] Michael I. Jordan. On statistics, computation and scalability. Bernoulli, 19(4):1378–1390, September 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1377612856.

Jordan:2023:SPC

[Jor23] Serres Jordan. Stability of the Poincaré constant. Bernoulli, 29(2):1297–1320, May 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-29/issue-
2/Stability-of-the-Poincar%c3%a9-constant/10.3150/22-
BEJ1499.full.

Joulin:2007:PTD

[Jou07] Aldéric Joulin. Poisson-type deviation inequalities for curved continuous-time Markov chains. Bernoulli, 13(3):782–798, August 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1186503486.

Joulin:2009:NPT

[Jou09] Aldéric Joulin. A new Poisson-type deviation inequality for Markov jump processes with positive Wasserstein curvature. Bernoulli, 15(2):532–549, May 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1241444901.

Jentsch:2015:TES

[JP15] Carsten Jentsch and Markus Pauly. Testing equality of spectral densities using randomization techniques. Bernoulli, 21(2):697– 739, May 2015. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1429624958.

Jurczak:2017:SAH

[JR17] Kamil Jurczak and Angelika Rohde. Spectral analysis of highdimensional sample covariance matrices with missing observations. Bernoulli, 23(4A):2466–2532, November 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1494316823.

Jakubowski:2021:DPD

[JRSK21] Adam Jakubowski, Igor Rodionov, and Natalia Soja-Kukieła. Directional phantom distribution functions for stationary random fields. Bernoulli, 27(2):1028–1056, May 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-2/Directional-phantom-distribution-functions-
for-stationary-random-fields/10.3150/20-BEJ1264.full.

Johnson:2005:CLT

[JS05] Oliver Johnson and Richard Samworth. Central limit theorem and convergence to stable laws in Mallows distance. Bernoulli, 11(5):829–845, October 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1130077596.

Johnson:2006:CLT

[JS06] Oliver Johnson and Richard Samworth. Central limit theorem and convergence to stable laws in Mallows distance. Bernoulli, 12(1):191, February 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1141136657.

Jagers:2008:GBP

[JS08] Peter Jagers and Serik Sagitov. General branching processes in discrete time as random trees. Bernoulli, 14(4):949–962, November 2008. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1225980566.

Jedidi:2013:FEG

[JS13a] Wissem Jedidi and Thomas Simon. Further examples of GGC and HCM densities. Bernoulli, 19(5A):1818–1838, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1383661204.

Johannes:2013:ACD

[JS13b] Jan Johannes and Maik Schwarz. Adaptive circular deconvolution by model selection under unknown error distribution. Bernoulli, 19(5A):1576–1611, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1383661195.

Jakobsen:2017:EED

[JS17] Nina Munkholt Jakobsen and Michael Sørensen. Efficient estimation for diffusions sampled at high frequency over a fixed time interval. Bernoulli, 23(3):1874–1910, August 2017. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1489737628.

Jarner:2003:NCG

[JT03] Søren F. Jarner and Richard L. Tweedie. Necessary conditions for geometric and polynomial ergodicity of random-walk-type. Bernoulli, 9(4):559–578, August 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1066223269.

Janson:2022:CTD

[JT22] Svante Janson and Debleena Thacker. Continuous-time digital search tree and a border aggregation model. Bernoulli, 28(4):2563–2577, November 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-28/issue-
4/Continuous-time-digital-search-tree-and-a-border-
aggregation-model/10.3150/21-BEJ1429.full.

Jureckova:1999:RRS

[Jur99] Jana Jureckova. Regression rank-scores tests against heavytailed alternatives. Bernoulli, 5(4):659–676, August 1999. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1171899322.

Jonker:2001:SPM

[JV01] Marianne A. Jonker and Aad W. Van Der Vaart. A semiparametric model for censored and passively registered data. Bernoulli, 7(1):1–31, February 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1080572337.

Jongbloed:2005:NIL

[JVV05] G. Jongbloed, F. H. Van Der Meulen, and A. W. Van Der Vaart. Nonparametric inference for Lévy-driven Ornstein–Uhlenbeck processes. Bernoulli, 11(5):759–791, October 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1130077593.

Jankowski:2009:NEC

[JW09] Hanna K. Jankowski and Jon A. Wellner. Nonparametric estimation of a convex bathtub-shaped hazard function. Bernoulli, 15(4):1010–1035, November 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1262962224.

Jakubowski:2013:HBP

[JW13] Jacek Jakubowski and Maciej Wiśniewolski. On hyperbolic Bessel processes and beyond. Bernoulli, 19(5B):2437–2454, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1386078609.

Jentsch:2019:BIM

[JW19] Carsten Jentsch and Christian H. Weiß. Bootstrapping INAR models. Bernoulli, 25(3):2359–2408, August 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1560326448.

Jakubowski:2021:CFL

[JW21] Jacek Jakubowski and Maciej Wiśniewolski. A convolution formula for the local time of an Itô diffusion reflecting at 0 and a generalized stroock-Williams equation. Bernoulli, 27(3):1870–1898, August 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-3/A-convolution-formula-for-the-local-time-
of-an-It%c3%b4/10.3150/20-BEJ1295.full.

Jiang:2022:DIC

[JWY22] Hui Jiang, Yilong Wan, and Guangyu Yang. Deviation inequalities and Cramér-type moderate deviations for the explosive autoregressive process. Bernoulli, 28(4):2634–2662, November 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-28/issue-4/Deviation-inequalities-
and-Cram%c3%a9r-type-moderate-deviations-for-the-explosive/
10.3150/21-BEJ1432.full.

Jiang:2019:RDN

[JZ19] Wenhua Jiang and Cun-Hui Zhang. Rate of divergence of the nonparametric likelihood ratio test for Gaussian mixtures. Bernoulli, 25(4B):3400–3420, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1569398770.

Kabluchko:2011:FLT

[Kab11] Zakhar Kabluchko. Functional limit theorems for sums of independent geometric Lévy processes. Bernoulli, 17(3):942–968, August 2011. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1310042851.

Kamatani:2018:ESM

[Kam18] Kengo Kamatani. Efficient strategy for the Markov chain Monte Carlo in high-dimension with heavy-tailed target probability distribution. Bernoulli, 24(4B):3711–3750, November 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL https://projecteuclid.org/euclid.bj/1524038768.

Karvonen:2023:SSS

[Kar23] Toni Karvonen. Small sample spaces for Gaussian processes. Bernoulli, 29(2):875–900, May 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-2/Small-sample-spaces-for-Gaussian-processes/
10.3150/22-BEJ1483.full.

Kargin:2024:JDA

[Kar24] Vladislav Kargin. On the joint distribution of the area and the number of peaks for Bernoulli excursions. Bernoulli, 30(4): 2700–2720, November 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

Kato:2009:DPU

[Kat09] Shogo Kato. A distribution for a pair of unit vectors generated by Brownian motion. Bernoulli, 15(3):898–921, August 2009.

CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1251463286.

Kellner:2019:OST

[KC19] Jérémie Kellner and Alain Celisse. A one-sample test for normality with kernel methods. Bernoulli, 25(3):1816–1837, August 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1560326429.

Kelsall:1995:KER

[KD95] Julia E. Kelsall and Peter J. Diggle. Kernel estimation of relative risk. Bernoulli, 1(1–2):3–16, March 1995. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1186078359.

Keener:2013:IBA

[Kee13] Robert Keener. Improving Brownian approximations for boundary crossing problems. Bernoulli, 19(1):137–153, February 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1358531744.

Kendall:1995:FBS

[Ken95] David G. Kendall. The foundation of the Bernoulli Society. Bernoulli, 1(1–2):1–2, March 1995. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1186078358.

Kendall:2015:CLT

[Ken15] Wilfrid S. Kendall. Coupling, local times, immersions. Bernoulli, 21(2):1014–1046, May 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1429624969.

Kanamori:2014:AID

[KF14] Takafumi Kanamori and Hironori Fujisawa. Affine invariant divergences associated with proper composite scoring rules and their applications. Bernoulli, 20(4):2278–2304, November 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1411134460.

Kasper:2021:WCC

[KFT21] Thimo M. Kasper, Sebastian Fuchs, and Wolfgang Trutschnig. On weak conditional convergence of bivariate Archimedean and Extreme Value copulas, and consequences to nonparametric estimation. Bernoulli, 27(4):2217–2240, November 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-4/On-weak-conditional-convergence-of-bivariate-
Archimedean-and-Extreme-Value/10.3150/20-BEJ1306.full.

Koltchinskii:2000:RMA

[KG00] Vladimir Koltchinskii and Evarist Giné. Random matrix approximation of spectra of integral operators. Bernoulli, 6(1): 113–167, February 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1082665383.

Kohatsu-Higa:1997:SDE

[KHLN97] Arturo Kohatsu-Higa, Jorge A. León, and David Nualart. Stochastic differential equations with random coefficients. Bernoulli, 3(2):233–245, June 1997. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1177526731.

Khmaladze:2016:UTE

[Khm16] Estate Khmaladze. Unitary transformations, empirical processes and distribution free testing. Bernoulli, 22(1):563–588, February 2016. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1443620860.

Kim:2014:MBE

[Kim14] Arlene K. H. Kim. Minimax bounds for estimation of normal mixtures. Bernoulli, 20(4):1802–1818, November 2014. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1411134445.

Kim:2021:CLN

[Kim21] Ilmun Kim. Comparing a large number of multivariate distributions. Bernoulli, 27(1):419–441, February 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-1/Comparing-a-large-number-of-multivariate-
distributions/10.3150/20-BEJ1244.full.

Kato:2013:EFC

[KJ13] Shogo Kato and M. C. Jones. An extended family of circular distributions related to wrapped Cauchy distributions via Brownian motion. Bernoulli, 19(1):154–171, February 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1358531745.

Kokoszka:2020:FDT

[KJ20] Piotr Kokoszka and Neda Mohammadi Jouzdani. Frequency domain theory for functional time series: Variance decomposition and an invariance principle. Bernoulli, 26(3):2383–2399, August 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1587974545.

Kanika:2016:MIE

[KK16] Kanika and Somesh Kumar. Methods for improving estimators of truncated circular parameters. Bernoulli, 22(4):2521–2547, November 2016. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1462297688.

Kim:2017:SMC

[KK17] Sojung Kim and Kyoung-Kuk Kim. Saddlepoint methods for conditional expectations with applications to risk management. Bernoulli, 23(3):1481–1517, August 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1489737615.

Khasminskii:2018:PEH

[KK18] Rafail Z. Khasminskii and Yury A. Kutoyants. On parameter estimation of hidden telegraph process. Bernoulli, 24(3):2064– 2090, August 2018. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1517540468.

Kohler:2021:PDN

[KK21] Michael Kohler and Adam Krzyzak. Over-parametrized deep neural networks minimizing the empirical risk do not generalize well. Bernoulli, 27(4):2564–2597, November 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-27/issue-4/Over-parametrized-deep-
neural-networks-minimizing-the-empirical-risk-do/10.
3150/21-BEJ1323.full.

Klimsiak:2023:HTL

[KK23a] Tomasz Klimsiak and Tomasz Komorowski. Hopf type lemmas for subsolutions of integro-differential equations. Bernoulli, 29(2):1435–1463, May 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-29/issue-
2/Hopf-type-lemmas-for-subsolutions-of-integro-differential-
equations/10.3150/22-BEJ1505.full.

Kohler:2023:RCD

[KK23b] Michael Kohler and Adam Krzyzak. On the rate of convergence of a deep recurrent neural network estimate in a regression problem with dependent data. Bernoulli, 29(2):1663– 1685, May 2023. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-29/issue-2/On-the-rate-of-
convergence-of-a-deep-recurrent-neural/10.3150/22-BEJ1516.
full.

Kluppelberg:2007:ETD

[KKP07] Claudia Klüppelberg, Gabriel Kuhn, and Liang Peng. Estimating the tail dependence function of an elliptical distribution. Bernoulli, 13(1):229–251, February 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1175287731.

Krampe:2021:BBI

[KKP21] Jonas Krampe, Jens-Peter Kreiss, and Efstathios Paparoditis. Bootstrap based inference for sparse high-dimensional time series models. Bernoulli, 27(3):1441–1466, August 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-3/Bootstrap-based-inference-for-sparse-high-
dimensional-time-series-models/10.3150/20-BEJ1239.full.

Kulczycki:2024:DRM

[KKR24] Tadeusz Kulczycki, Oleksii Kulyk, and Michał Ryznar. Drift reduction method for SDEs driven by heterogeneous singular Lévy noise. Bernoulli, 30(4):3089–3118, November 2024. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

Kohler:2002:ASR

[KKS02] Michael Kohler, Adam Krzyzak, and Dominik Schäfer. Application of structural risk minimization to multivariate smoothing spline regression estimates. Bernoulli, 8(4):475–489, August 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1078681380.

Kim:2017:LIL

[KKW17] Panki Kim, Takashi Kumagai, and Jian Wang. Laws of the iterated logarithm for symmetric jump processes. Bernoulli, 23(4A):2330–2379, November 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1494316820.

Kokoszka:2000:CPE

[KL00] Piotr Kokoszka and Remigijus Leipus. Change-point estimation in ARCH models. Bernoulli, 6(3):513–539, June 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1081616703.

Kuhn:2002:OSR

[KL02] Thomas Kühn and Werner Linde. Optimal series representation of fractional Brownian sheets. Bernoulli, 8(5):669–696, October 2002. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1078435223.

Kluppelberg:2005:EVT

[KL05] Claudia Klüppelberg and Alexander Lindner. Extreme value theory for moving average processes with light-tailed innovations. Bernoulli, 11(3):381–410, June 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1120591182.

Kim:2009:TTI

[KL09] Moosup Kim and Sangyeol Lee. Test for tail index change in stationary time series with Pareto-type marginal distribution. Bernoulli, 15(2):325–356, May 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1241444893.

Kolassa:2010:MSA

[KL10] John Kolassa and Jixin Li. Multivariate saddlepoint approximations in tail probability and conditional inference. Bernoulli, 16 (4):1191–1207, November 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1290092902.

Kulik:2013:EMB

[KL13] A. M. Kulik and N. N. Leonenko. Ergodicity and mixing bounds for the Fisher–Snedecor diffusion. Bernoulli, 19(5B):2294–2329, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1386078604.

Koltchinskii:2017:CIM

[KL17] Vladimir Koltchinskii and Karim Lounici. Concentration inequalities and moment bounds for sample covariance operators. Bernoulli, 23(1):110–133, February 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1475001350.

Kouamo:2013:CLT

[KLLM13] O. Kouamo, C. Lévy-Leduc, and E. Moulines. Central limit theorem for the robust log-regression wavelet estimation of the memory parameter in the Gaussian semi-parametric context. Bernoulli, 19(1):172–204, February 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1358531746.

Klok:2006:AER

[Klo06] Marten J. Klok. Asymptotic error rates in third-generation wireless systems. Bernoulli, 12(2):193–219, April 2006. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1145993972.

Klopp:2014:NLR

[Klo14] Olga Klopp. Noisy low-rank matrix completion with general sampling distribution. Bernoulli, 20(1):282–303, February 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1390407290.

Kerkyacharian:2012:RNT

[KLP12] Gérard Kerkyacharian, Erwan Le Pennec, and Dominique Picard. Radon needlet thresholding. Bernoulli, 18(2):391–433, May 2012. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1334580718.

Kwak:2021:EES

[KLP21] Jinwoong Kwak, Ji Oon Lee, and Jaewhi Park. Extremal eigenvalues of sample covariance matrices with general population. Bernoulli, 27(4):2740–2765, November 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-27/issue-4/Extremal-eigenvalues-of-
sample-covariance-matrices-with-general-population/
10.3150/21-BEJ1329.full.

Kozek:1998:USL

[KLS98] Andrzej S. Kozek, Julian R. Leslie, and Eugene F. Schuster. On a universal strong law of large numbers for conditional expectations. Bernoulli, 4(2):143–165, June 1998. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1174937292.

Klopp:2019:SME

[KLTZ19] Olga Klopp, Yu Lu, Alexandre B. Tsybakov, and Harrison H. Zhou. Structured matrix estimation and completion. Bernoulli, 25(4B):3883–3911, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1569398788.

Kim:2018:LVM

[KLW18] Donggyu Kim, Yi Liu, and Yazhen Wang. Large volatility matrix estimation with factor-based diffusion model for highfrequency financial data. Bernoulli, 24(4B):3657–3682, November 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1524038766.

Kluppelberg:1995:EPS

[KM95] Claudia Klüppelberg and Thomas Mikosch. Explosive Poisson shot noise processes with applications to risk reserves. Bernoulli, 1(1–2):125–147, March 1995. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1186078364.

Koshevoy:1998:LZR

[KM98] Gleb Koshevoy and Karl Mosler. Lift zonoids, random convex hulls and the variability of random vectors. Bernoulli, 4 (3):377–399, September 1998. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1174324985.

Konakov:2005:ETE

[KM05] Valentin Konakov and Enno Mammen. Edgeworth-type expansions for transition densities of Markov chains converging to diffusions. Bernoulli, 11(4):591–641, August 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1126126762.

Kokonendji:2006:CPG

[KM06] Célestin C. Kokonendji and Afif Masmoudi. A characterization of Poisson–Gaussian families by generalized variance. Bernoulli, 12(2):371–379, April 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1145993979.

Kevei:2011:NMB

[KM11] Péter Kevei and David M. Mason. A note on a maximal Bernstein inequality. Bernoulli, 17(3):1054–1062, August 2011. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1310042856.

Kalbasi:2020:PDL

[KM20a] Kamran Kalbasi and Thomas Mountford. On the probability distribution of the local times of diagonally operator-self-similar Gaussian fields with stationary increments. Bernoulli, 26(2): 1504–1534, May 2020. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1580461588.

Kato:2020:SPC

[KM20b] Shogo Kato and Peter McCullagh. Some properties of a Cauchy family on the sphere derived from the Möbius transformations. Bernoulli, 26(4):3224–3248, November 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
26/issue-4/Some-properties-of-a-Cauchy-family-on-the-
sphere-derived/10.3150/20-BEJ1222.full.

Kersting:2022:DGW

[KM22] Götz Kersting and Carmen Minuesa. Defective Galton–Watson processes in a varying environment. Bernoulli, 28(2):1408– 1431, May 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-28/issue-2/Defective-
Galton--Watson-processes-in-a-varying-environment/10.
3150/21-BEJ1393.full.

Kampf:2023:TLB

[KM23] Jürgen Kampf and Alexander Meister. Testing for linearity in boundary regression models with application to maximal life expectancies. Bernoulli, 29(3):1764–1791, August 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-3/Testing-for-linearity-in-boundary-regression-
models-with-application-to/10.3150/22-BEJ1520.full.

Kurisu:2024:LPT

[KM24] Daisuke Kurisu and Yasumasa Matsuda. Local polynomial trend regression for spatial data on rd. Bernoulli, 30(4):2770– 2794, November 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

Kuchibhotla:2021:HDC

[KMB21] Arun Kumar Kuchibhotla, Somabha Mukherjee, and Debapratim Banerjee. High-dimensional CLT: Improvements, nonuniform extensions and large deviations. Bernoulli, 27(1):192– 217, February 2021. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-27/issue-1/High-dimensional-

CLT--Improvements-non-uniform-extensions-and-large/
10.3150/20-BEJ1233.full.

Kendall:2024:OMC

[KMM24] Wilfrid S. Kendall, Mateusz B. Majka, and Aleksandar Mijatović. Optimal Markovian coupling for finite activity Lévy processes. Bernoulli, 30(4):2821–2845, November 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

Kagan:1999:SBR

[KMS^{+}99] Abram Kagan, Colin L. Mallows, Larry A. Shepp, Robert J. Vanderbei, and Yehuda Vardi. Symmetrization of binary random variables. Bernoulli, 5(6):1013–1020, December 1999. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1143122300.

Kluppelberg:2003:RVM

[KMS03] Claudia Klüppelberg, Thomas Mikosch, and Anette Schärf. Regular variation in the mean and stable limits for Poisson shot noise. Bernoulli, 9(3):467–496, June 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1065444814.

Kubokawa:2017:PDE

[KMS17] Tatsuya Kubokawa, Éric Marchand, and William E. Strawderman. On predictive density estimation for location families under integrated absolute error loss. Bernoulli, 23(4B):3197– 3212, November 2017. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1495505090.

Konakov:2014:SCM

[KMW14] Valentin Konakov, Enno Mammen, and Jeannette Woerner. Statistical convergence of Markov experiments to diffusion limits. Bernoulli, 20(2):623–644, May 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1393594000.

Korostelev:1999:AMC

[KN99] Alexander Korostelev and Michael Nussbaum. The asymptotic minimax constant for sup-norm loss in nonparametric density estimation. Bernoulli, 5(6):1099–1118, December 1999. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1143122304.

Kallabis:2006:EIU

[KN06] Raoul S. Kallabis and Michael H. Neumann. An exponential inequality under weak dependence. Bernoulli, 12(2):333–350, April 2006. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1145993977.

Khoshnevisan:2008:LSS

[KN08] Davar Khoshnevisan and Eulalia Nualart. Level sets of the stochastic wave equation driven by a symmetric Lévy noise. Bernoulli, 14(4):899–925, November 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1225980564.

Koike:2016:QCE

[Koi16] Yuta Koike. Quadratic covariation estimation of an irregularly observed semimartingale with jumps and noise. Bernoulli, 22 (3):1894–1936, August 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1458133002.

Koloydenko:2008:SMM

[Kol08] Alexey Koloydenko. Symmetric measures via moments. Bernoulli, 14(2):362–390, May 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1208872109.

Koltchinskii:2009:DSS

[Kol09] Vladimir Koltchinskii. The Dantzig selector and sparsity oracle inequalities. Bernoulli, 15(3):799–828, August 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1251463282.

Koster:2002:MCG

[Kos02] Jan T. A. Koster. Marginalizing and conditioning in graphical models. Bernoulli, 8(6):817–840, December 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1076364807.

Kerkyacharian:2002:MM

[KP02] Gérard Kerkyacharian and Dominique Picard. Minimax or maxisets? Bernoulli, 8(2):219–253, April 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1078866869.

Kerkyacharian:2004:RRD

[KP04] Gérard Kerkyacharian and Dominique Picard. Regression in random design and warped wavelets. Bernoulli, 10(6):1053– 1105, December 2004. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL http://projecteuclid.org/
euclid.bj/1106314850.

Kyprianou:2006:QSD

[KP06] Andreas E. Kyprianou and Zbigniew Palmowski. Quasistationary distributions for Lévy processes. Bernoulli, 12 (4):571–581, August 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1155735927.

Kuchibhotla:2020:EES

[KP20] Arun K. Kuchibhotla and Rohit K. Patra. Efficient estimation in single index models through smoothing splines. Bernoulli, 26(2):1587–1618, May 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1580461591.

Konen:2022:MQS

[KP22a] Dimitri Konen and Davy Paindaveine. Multivariate ρquantiles: A spatial approach. Bernoulli, 28(3):1912–1934, August 2022. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-28/issue-3/Multivariate-%
cf%81-quantiles-A-spatial-approach/10.3150/21-BEJ1404.
full.

Kremp:2022:MSD

[KP22b] Helena Kremp and Nicolas Perkowski. Multidimensional SDE with distributional drift and Lévy noise. Bernoulli, 28(3):1757– 1783, August 2022. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-28/issue-3/Multidimensional-

SDE-with-distributional-drift-and-L%c3%a9vy-noise/10.
3150/21-BEJ1394.full.

Kock:2023:CPN

[KP23] Anders Bredahl Kock and David Preinerstorfer. Consistency of p -norm based tests in high dimensions: Characterization, monotonicity, domination. Bernoulli, 29(3):2544– 2573, August 2023. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-29/issue-3/Consistency-of-
p-norm-based-tests-in-high-dimensions/10.3150/22-BEJ1552.
full.

Kerkyacharian:1996:LAD

[KPT96] Gérard Kerkyacharian, Dominique Picard, and Karine Tribouley. Lp adaptive density estimation. Bernoulli, 2(3): 229–247, September 1996. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1178291720.

Kabluchko:2021:NLR

[KPT21] Zakhar Kabluchko, Joscha Prochno, and Christoph Thäle. A new look at random projections of the cube and general product measures. Bernoulli, 27(3):2117–2138, August 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-3/A-new-look-at-random-projections-of-the-
cube-and/10.3150/20-BEJ1303.full.

Krampe:2024:ICO

[KR24] Jonas Krampe and Suhasini Subba Rao. Inverse covariance operators of multivariate nonstationary time series. Bernoulli, 30(2):1177–1196, May 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
30/issue-2/Inverse-covariance-operators-of-multivariate-
nonstationary-time-series/10.3150/23-BEJ1628.full.

Kraaij:2021:FLD

[Kra21] Richard C. Kraaij. Flux large deviations of weakly interacting jump processes via well-posedness of an associated Hamilton–Jacobi equation. Bernoulli, 27(3):1496–1528, August 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-27/issue-3/Flux-large-
deviations-of-weakly-interacting-jump-processes-via-
well/10.3150/20-BEJ1281.full.

Krebs:2021:LIL

[Kre21a] Johannes Krebs. On the law of the iterated logarithm and strong invariance principles in stochastic geometry. Bernoulli, 27(3):1695–1723, August 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-27/issue-
3/On-the-law-of-the-iterated-logarithm-and-strong-invariance/
10.3150/20-BEJ1288.full.

Kreiss:2021:CBM

[Kre21b] Alexander Kreiß. Correlation bounds, mixing and m - dependence under random time-varying network distances with an application to Cox-processes. Bernoulli, 27(3):1666–1694, August 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-27/issue-3/Correlation-bounds-mixing-
and-m-dependence-under-random-time-varying/10.3150/
20-BEJ1287.full.

Krebs:2024:BRS

[Kre24] Johannes Krebs. On the Bahadur representation of sample quantiles for score functionals. Bernoulli, 30(4):3356–3377, November 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

Keeler:2018:WDW

[KRX18] H. Paul Keeler, Nathan Ross, and Aihua Xia. When do wireless network signals appear Poisson? Bernoulli, 24(3):1973–1994, August 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1517540465.

Koul:1997:EEN

[KS97] Hira L. Koul and Anton Schick. Efficient estimation in nonlinear autoregressive time-series models. Bernoulli, 3(3):247–277, September 1997. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1177334455.

Kessler:1999:EEB

[KS99a] Mathieu Kessler and Michael Sørensen. Estimating equations based on eigenfunctions for a discretely observed diffusion process. Bernoulli, 5(2):299–314, April 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1173147908.

Kuchler:1999:NLT

[KS99b] Uwe Küchler and Michael Sørensen. A note on limit theorems for multivariate martingales. Bernoulli, 5(3):483–493, June 1999. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1172617200.

Kyprianou:2001:MCF

[KS01] Andreas E. Kyprianou and A. Rahimzadeh Sani. Martingale convergence and the functional equation in the multi-type branching random walk. Bernoulli, 7(4):593–604, August 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1079559464.

Kubokawa:2013:DPC

[KS13a] Tatsuya Kubokawa and William E. Strawderman. Dominance properties of constrained Bayes and empirical Bayes estimators. Bernoulli, 19(5B):2200–2221, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1386078600.

Kuchler:2013:SID

[KS13b] Uwe Küchler and Michael Sørensen. Statistical inference for discrete-time samples from affine stochastic delay differential equations. Bernoulli, 19(2):409–425, May 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1363192033.

Kabluchko:2016:SIR

[KS16] Zakhar Kabluchko and Stilian Stoev. Stochastic integral representations and classification of sum- and max-infinitely divisible processes. Bernoulli, 22(1):107–142, February 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1443620845.

Knapik:2018:GAP

[KS18] Bartek Knapik and Jean-Bernard Salomond. A general approach to posterior contraction in nonparametric inverse problems. Bernoulli, 24(3):2091–2121, August 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1517540469.

Kolb:2020:CFP

[KS20] Martin Kolb and Mladen Savov. A characterization of the finiteness of perpetual integrals of Lévy processes. Bernoulli, 26(2): 1453–1472, May 2020. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1580461586.

Krell:2021:NEJ

[KS21] Nathalie Krell and Émeline Schmisser. Nonparametric estimation of jump rates for a specific class of piecewise deterministic Markov processes. Bernoulli, 27(4):2362–2388, November 2021. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-27/issue-4/Nonparametric-
estimation-of-jump-rates-for-a-specific-class-of/10.
3150/20-BEJ1312.full.

Koskela:2019:CBN

[KSJ19] Jere Koskela, Dario Spanò, and Paul A. Jenkins. Consistency of Bayesian nonparametric inference for discretely observed jump diffusions. Bernoulli, 25(3):2183–2205, August 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1560326442.

Klebanov:2021:LCE

[KSS21] Ilja Klebanov, Björn Sprungk, and T. J. Sullivan. The linear conditional expectation in Hilbert space. Bernoulli, 27(4):2267– 2299, November 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-27/issue-4/The-linear-
conditional-expectation-in-Hilbert-space/10.3150/20-
BEJ1308.full.

Kessler:2001:IML

[KSW01] Mathieu Kessler, Anton Schick, and Wolfgang Wefelmeyer. The information in the marginal law of a Markov chain. Bernoulli, 7(2):243–266, April 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1080222086.

Kokoszka:2019:PCA

[KSX19] Piotr Kokoszka, Stilian Stoev, and Qian Xiong. Principal components analysis of regularly varying functions. Bernoulli, 25(4B):3864–3882, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1569398787.

Kashimura:2015:SIU

[KT15] Takuya Kashimura and Akimichi Takemura. Standard imsets for undirected and chain graphical models. Bernoulli, 21 (3):1467–1493, August 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1432732027.

Kusuoka:2018:CCT

[KT18] Seiichiro Kusuoka and Ciprian A. Tudor. Characterization of the convergence in total variation and extension of the Fourth Moment Theorem to invariant measures of diffusions. Bernoulli, 24(2):1463–1496, May 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1505980901.

Kao:2018:TIC

[KTU18] Chihwa Kao, Lorenzo Trapani, and Giovanni Urga. Testing for instability in covariance structures. Bernoulli, 24(1):740– 771, February 2018. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1501142461.

Kuhn:2019:PIR

[Küh19] Franziska Kühn. Perpetual integrals via random time changes. Bernoulli, 25(3):1755–1769, August 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1560326426.

Kuhn:2022:UFS

[Küh22] Franziska Kühn. Upper functions for sample paths of Lévy(-type) processes. Bernoulli, 28(4):2874–2908, November 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-28/issue-4/Upper-functions-for-sample-
paths-of-L%c3%a9vy-type-processes/10.3150/21-BEJ1441.
full.

Kujala:2016:AOM

[Kuj16] Janne V. Kujala. Asymptotic optimality of myopic informationbased strategies for Bayesian adaptive estimation. Bernoulli, 22(1):615–651, February 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1443620862.

Kulik:2007:BKT

[Kul07] Rafał Kulik. Bahadur–Kiefer theory for sample quantiles of weakly dependent linear processes. Bernoulli, 13(4):1071–1090, November 2007. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1194625603.

Kunsch:2013:PF

[Kün13] Hans R. Künsch. Particle filters. Bernoulli, 19(4):1391–1403, September 2013. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1377612857.

Kuroki:2016:EBD

[Kur16] Manabu Kuroki. Equivalence between direct and indirect effects with different sets of intermediate variables and covariates. Bernoulli, 22(1):421–443, February 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1443620856.

Kurisu:2022:NRL

[Kur22] Daisuke Kurisu. Nonparametric regression for locally stationary random fields under stochastic sampling design. Bernoulli, 28(2):1250–1275, May 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-28/issue-
2/Nonparametric-regression-for-locally-stationary-random-
fields-under-stochastic-sampling/10.3150/21-BEJ1385.
full.

Kutoyants:1997:EED

[Kut97] Yury A. Kutoyants. Efficiency of the empirical distribution for ergodic diffusion. Bernoulli, 3(4):445–456, December 1997. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1175882218.

Kutoyants:2015:AGF

[Kut15] Yury A. Kutoyants. On ADF goodness-of-fit tests for perturbed dynamical systems. Bernoulli, 21(4):2430–2456, November 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1438777600.

Kutoyants:2021:MSE

[Kut21] Yury A. Kutoyants. On multi-step estimation of delay for SDE. Bernoulli, 27(3):2069–2090, August 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-3/On-multi-step-estimation-of-delay-for-SDE/
10.3150/20-BEJ1301.full.

Koudou:2012:IPM

[KV12] A. E. Koudou and P. Vallois. Independence properties of the Matsumoto–Yor type. Bernoulli, 18(1):119–136, February 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1327068620.

Keles:2004:AOM

[KVD04] Sündüz Keles, Mark Van Der Laan, and Sandrine Dudoit. Asymptotically optimal model selection method with right censored outcomes. Bernoulli, 10(6):1011–1037, December 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1106314848.

Kley:2016:QSP

[KVDH16] Tobias Kley, Stanislav Volgushev, Holger Dette, and Marc Hallin. Quantile spectral processes: Asymptotic analysis and inference. Bernoulli, 22(3):1770–1807, August 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1458132999.

Kabluchko:2019:MAA

[KVZ19] Zakhar Kabluchko, Vladislav Vysotsky, and Dmitry Zaporozhets. A multidimensional analogue of the arcsine law for the number of positive terms in a random walk. Bernoulli, 25(1):521–548, February 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1544605255.

Klaassen:1997:EEB

[KW97] Chris A. J. Klaassen and Jon A. Wellner. Efficient estimation in the bivariate normal copula model: normal margins are least favourable. Bernoulli, 3(1):55–77, February 1997. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1178291932.

Kohler:2013:DBO

[KW13] Michael Kohler and Harro Walk. On data-based optimal stopping under stationarity and ergodicity. Bernoulli, 19(3):931– 953, August 2013. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1372251148.

Klusowski:2020:ENC

[KW20] Jason M. Klusowski and Yihong Wu. Estimating the number of connected components in a graph via subgraph sampling. Bernoulli, 26(3):1635–1664, August 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1587974519.

Kutoyants:2007:MEE

[KY07] Yury A. Kutoyants and Nakahiro Yoshida. Moment estimation for ergodic diffusion processes. Bernoulli, 13(4):933–951, November 2007. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1194625596.

Kim:2022:LTT

[KY22] Kunwoo Kim and Jaeyun Yi. Limit theorems for timedependent averages of nonlinear stochastic heat equations. Bernoulli, 28(1):214–238, February 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-28/issue-

1/Limit-theorems-for-time-dependent-averages-of-nonlinear-
stochastic-heat/10.3150/21-BEJ1339.full.

Kovchegov:2021:IAP

[KZ21a] Yevgeniy Kovchegov and Ilya Zaliapin. Invariance and attraction properties of Galton–Watson trees. Bernoulli, 27(3):1789– 1823, August 2021. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-27/issue-3/Invariance-and-
attraction-properties-of-GaltonWatson-trees/10.3150/
20-BEJ1292.full.

Kupper:2021:LDB

[KZ21b] Michael Kupper and José Miguel Zapata. Large deviations built on max-stability. Bernoulli, 27(2):1001–1027, May 2021. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-2/Large-deviations-built-on-max-stability/
10.3150/20-BEJ1263.full.

Lageraas:2010:CMD

[Lag10] Andreas N. Lagerås. Copulas for Markovian dependence. Bernoulli, 16(2):331–342, May 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1274821073.

Li:2020:HDG

[LAP20] Haoran Li, Alexander Aue, and Debashis Paul. Highdimensional general linear hypothesis tests via non-linear spectral shrinkage. Bernoulli, 26(4):2541–2571, November 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-26/issue-4/High-dimensional-general-
linear-hypothesis-tests-via-non-linear-spectral/10.
3150/19-BEJ1186.full.

Last:2014:PAP

[Las14] Günter Last. Perturbation analysis of Poisson processes. Bernoulli, 20(2):486–513, May 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1393593995.

Laurent:1997:EIF

[Lau97] Béatrice Laurent. Estimation of integral functionals of a density and its derivatives. Bernoulli, 3(2):181–211, June 1997. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1177526728.

Laurent:2004:SSP

[Lau04] Denis Laurent. Solutions of stochastic partial differential equations considered as Dirichlet processes. Bernoulli, 10(5):783– 827, October 2004. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1099579156.

Lau:2013:CCR

[Lau13] John W. Lau. A conjugate class of random probability measures based on tilting and with its posterior analysis. Bernoulli, 19(5B):2590–2626, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1386078614.

Livingstone:2019:GEH

[LBBG19] Samuel Livingstone, Michael Betancourt, Simon Byrne, and Mark Girolami. On the geometric ergodicity of Hamiltonian Monte Carlo. Bernoulli, 25(4A):3109–3138, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL https://projecteuclid.org/euclid.bj/1568362054.

Liu:2019:LOE

[LBDM19] Zhipeng Liu, Jose H. Blanchet, A. B. Dieker, and Thomas Mikosch. On logarithmically optimal exact simulation of maxstable and related random fields on a compact set. Bernoulli, 25(4A):2949–2981, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1568362048.

Lau:2022:TCR

[LC22] John W. Lau and Edward Cripps. Thinned completely random measures with applications in competing risks models. Bernoulli, 28(1):638–662, February 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-1/Thinned-completely-random-measures-with-

applications-in-competing-risks-models/10.3150/21-BEJ1361.
full.

Loeffen:2013:PRP

[LCP13] Ronnie Loeffen, Irmina Czarna, and Zbigniew Palmowski. Parisian ruin probability for spectrally negative Lévy processes. Bernoulli, 19(2):599–609, May 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1363192039.

Li:2011:AFB

[LD11] Yuqiang Li and Hongshuai Dai. Approximations of fractional Brownian motion. Bernoulli, 17(4):1195–1216, November 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1320417501.

Li:2023:RLA

[LE23] Mufan (Bill) Li and Murat A. Erdogdu. Riemannian Langevin algorithm for solving semidefinite programs. Bernoulli, 29 (4):3093–3113, November 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-29/issue-
4/Riemannian-Langevin-algorithm-for-solving-semidefinite-
programs/10.3150/22-BEJ1576.full.

Lecue:2007:ORA

[Lec07] Guillaume Lecué. Optimal rates of aggregation in classification under low noise assumption. Bernoulli, 13(4):1000–1022, November 2007. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1194625600.

Lecue:2013:ERM

[Lec13] Guillaume Lecué. Empirical risk minimization is optimal for the convex aggregation problem. Bernoulli, 19(5B):2153–2166, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1386078598.

Leeb:2008:ESM

[Lee08] Hannes Leeb. Evaluation and selection of models for out-of-sample prediction when the sample size is small relative to the complexity of the data-generating process. Bernoulli, 14 (3):661–690, August 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1219669625.

Lee:2022:HMR

[Lee22] Cheuk Yin Lee. The Hausdorff measure of the range and level sets of Gaussian random fields with sectorial local nondeterminism. Bernoulli, 28(1):277–306, February 2022. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-28/issue-1/The-Hausdorff-
measure-of-the-range-and-level-sets-of/10.3150/21-BEJ1342.
full.

Lehericy:2019:COE

[Leh19] Luc Lehéricy. Consistent order estimation for nonparametric hidden Markov models. Bernoulli, 25(1):464–498, February 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1544605253.

Lei:2006:LDK

[Lei06] Liangzhen Lei. Large deviations of the kernel density estimator in l1(rd) for reversible Markov processes. Bernoulli, 12(1):65– 83, February 2006. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1141136649.

Lei:2019:SLR

[Lei19] Tao Lei. Scaling limit of random forests with prescribed degree sequences. Bernoulli, 25(4A):2409–2438, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL https://projecteuclid.org/euclid.bj/1568362031.

Lei:2020:CCE

[Lei20] Jing Lei. Convergence and concentration of empirical measures under Wasserstein distance in unbounded functional spaces. Bernoulli, 26(1):767–798, February 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1574758845.

Lopes:2023:BON

[LEM23] Miles E. Lopes, N. Benjamin Erichson, and Michael W. Mahoney. Bootstrapping the operator norm in high dimensions: Error estimation for covariance matrices and sketching. Bernoulli, 29(1):428–450, February 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-1/Bootstrapping-the-operator-norm-in-high-
dimensions--Error-estimation/10.3150/22-BEJ1463.full.

Lenstra:2005:CRR

[Len05] Andries Lenstra. Cramér–Rao revisited. Bernoulli, 11(2):263– 282, April 2005. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1116340294.

Lenart:2011:ADS

[Len11] Łukasz Lenart. Asymptotic distributions and subsampling in spectral analysis for almost periodically correlated time series. Bernoulli, 17(1):290–319, February 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1297173844.

Lepski:2016:UFN

[Lep16] Oleg Lepski. Upper functions for L_{p}-norms of Gaussian random fields. Bernoulli, 22(2):732–773, May 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1447077760.

Lepski:2018:NAE

[Lep18] O. V. Lepski. A new approach to estimator selection. Bernoulli, 24(4A):2776–2810, November 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1522051225.

Leucht:2012:DSU

[Leu12] Anne Leucht. Degenerate U- and V -statistics under weak dependence: Asymptotic theory and bootstrap consistency. Bernoulli, 18(2):552–585, May 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1334580724.

Levrard:2018:SOI

[Lev18] Clément Levrard. Sparse oracle inequalities for variable selection via regularized quantization. Bernoulli, 24(1):271– 296, February 2018. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1501142443.

Li:2008:AJD

[LGS08] Bo Li, Marc G. Genton, and Michael Sherman. On the asymptotic joint distribution of sample space–time covariance estimators. Bernoulli, 14(1):228–248, February 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1202492792.

Lemor:2006:RCE

[LGW06] Jean-Philippe Lemor, Emmanuel Gobet, and Xavier Warin. Rate of convergence of an empirical regression method for solving generalized backward stochastic differential equations. Bernoulli, 12(5):889–916, October 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1161614951.

Lin:2001:RZD

[LH01] Gwo Dong Lin and Chin-Yuan Hu. The Riemann zeta distribution. Bernoulli, 7(5):817–828, October 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1079399543.

Limic:2019:EMC

[Lim19] Vlada Limic. The eternal multiplicative coalescent encoding via excursions of Lévy-type processes. Bernoulli, 25(4A):2479– 2507, November 2019. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1568362033.

Lindberg:2009:POW

[Lin09] Carl Lindberg. Portfolio optimization when expected stock returns are determined by exposure to risk. Bernoulli, 15(2):464– 474, May 2009. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1241444898.

Lin:2019:HMB

[Lin19] Shen Lin. Harmonic measure for biased random walk in a supercritical Galton–Watson tree. Bernoulli, 25(4B):3652–3672, November 2019. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1569398780.

Larsen:2003:PIW

[LJ03] P. V. Larsen and P. E. Jupp. Parametrization-invariant Wald tests. Bernoulli, 9(1):167–182, February 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1068129014.

Lember:2008:AVT

[LK08] Jüri Lember and Alexey Koloydenko. The adjusted Viterbi training for hidden Markov models. Bernoulli, 14(1):180–206, February 2008. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1202492790.

Lavielle:2000:MCP

[LL00] Marc Lavielle and Carenne Ludeña. The multiple change-points problem for the spectral distribution. Bernoulli, 6(5):845–869, October 2000. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1081282692.

Le:2024:DAS

[LLBF24] Huiling Le, Alexander Lewis, Karthik Bharath, and Christopher Fallaize. A diffusion approach to Stein’s method on Riemannian manifolds. Bernoulli, 30(2):1079–1104, May 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-30/issue-2/A-diffusion-approach-to-
Steins-method-on-Riemannian-manifolds/10.3150/23-BEJ1625.
full.

Li:2019:PCT

[LLD19] Cheng Li, Lizhen Lin, and David B. Dunson. On posterior consistency of tail index for Bayesian kernel mixture models. Bernoulli, 25(3):1999–2028, August 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1560326435.

Lodhia:2022:MMN

[LLL22] Asad Lodhia, Keith Levin, and Elizaveta Levina. Matrix means and a novel high-dimensional shrinkage phenomenon. Bernoulli, 28(4):2578–2605, November 2022. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-4/Matrix-means-and-a-novel-high-dimensional-
shrinkage-phenomenon/10.3150/21-BEJ1430.full.

Liu:2011:NST

[LLS11] Weidong Liu, Shiqing Ling, and Qi-Man Shao. On nonstationary threshold autoregressive models. Bernoulli, 17(3): 969–986, August 2011. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL http://projecteuclid.org/
euclid.bj/1310042852.

Li:2012:MAM

[LLT12] Dong Li, Shiqing Ling, and Howell Tong. On moving-average models with feedback. Bernoulli, 18(2):735–745, May 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1334580731.

Lu:2007:ESN

[LLTY07] Zudi Lu, Arvid Lundervold, Dag Tjøstheim, and Qiwei Yao. Exploring spatial nonlinearity using additive approximation. Bernoulli, 13(2):447–472, May 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1179498756.

Landriault:2017:MAD

[LLZ17] David Landriault, Bin Li, and Hongzhong Zhang. On magnitude, asymptotics and duration of drawdowns for Lévy models. Bernoulli, 23(1):432–458, February 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1475001360.

Li:2007:VER

[LM07] Yingying Li and Per A. Mykland. Are volatility estimators robust with respect to modeling assumptions? Bernoulli, 13 (3):601–622, August 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1186503478.

Lecue:2010:SLB

[LM10] Guillaume Lecué and Shahar Mendelson. Sharper lower bounds on the performance of the empirical risk minimization algorithm. Bernoulli, 16(3):605–613, August 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1281099877.

Lecue:2013:OAE

[LM13] Guillaume Lecué and Shahar Mendelson. On the optimality of the aggregate with exponential weights for low temperatures. Bernoulli, 19(2):646–675, May 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1363192042.

Loustau:2015:MFR

[LM15] Sébastien Loustau and Clément Marteau. Minimax fast rates for discriminant analysis with errors in variables. Bernoulli, 21 (1):176–208, February 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1426597067.

Lecue:2016:PER

[LM16] Guillaume Lecué and Shahar Mendelson. Performance of empirical risk minimization in linear aggregation. Bernoulli, 22 (3):1520–1534, August 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1458132990.

Lugosi:2019:RSR

[LM19] Gábor Lugosi and Shahar Mendelson. Regularization, sparse recovery, and median-of-means tournaments. Bernoulli, 25 (3):2075–2106, August 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1560326438.

Lindskog:2020:ELT

[LM20] Filip Lindskog and Abhishek Pal Majumder. Exact long time behavior of some regime switching stochastic processes. Bernoulli, 26(4):2572–2604, November 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
26/issue-4/Exact-long-time-behavior-of-some-regime-
switching-stochastic-processes/10.3150/20-BEJ1196.full.

Lee:2023:UCS

[LMC23] Juho Lee, Xenia Miscouridou, and François Caron. A unified construction for series representations and finite approximations of completely random measures. Bernoulli, 29(3):2142– 2166, August 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-29/issue-3/A-unified-
construction-for-series-representations-and-finite-
approximations-of/10.3150/22-BEJ1536.full.

Liese:2019:SGA

[LMK19] Friedrich Liese, Alexander Meister, and Johanna Kappus. Strong Gaussian approximation of the mixture Rasch model. Bernoulli, 25(2):1326–1354, May 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1551862852.

Laurent:2016:NAD

[LMMR16] Béatrice Laurent, Clément Marteau, and Cathy Maugis-Rabusseau. Non-asymptotic detection of two-component mixtures with unknown means. Bernoulli, 22(1):242–274, February 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1443620849.

Latuszynski:2013:NBE

[ŁMN13] Krzysztof Łatuszyński, Błażej Miasojedow, and Wojciech Niemiro. Nonasymptotic bounds on the estimation error of MCMC algorithms. Bernoulli, 19(5A):2033–2066, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1383661213.

Linton:2011:NRF

[LMNV11] Oliver Linton, Enno Mammen, Jens Perch Nielsen, and Ingrid Van Keilegom. Nonparametric regression with filtered data. Bernoulli, 17(1):60–87, February 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1297173833.

Lee:2012:PTE

[LMP12] Young K. Lee, Enno Mammen, and Byeong U. Park. Projection-type estimation for varying coefficient regression models. Bernoulli, 18(1):177–205, February 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1327068622.

Lember:2014:OAL

[LMV14] Jüri Lember, Heinrich Matzinger, and Anna Vollmer. Optimal alignments of longest common subsequences and their path properties. Bernoulli, 20(3):1292–1343, August 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1402488941.

Lanteri:2022:CRS

[LMV22] Alessandro Lanteri, Mauro Maggioni, and Stefano Vigogna. Conditional regression for single-index models. Bernoulli, 28(4):3051–3078, November 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-28/issue-
4/Conditional-regression-for-single-index-models/10.
3150/22-BEJ1482.full.

Lugosi:2020:CSN

[LMZ20] Gábor Lugosi, Shahar Mendelson, and Nikita Zhivotovskiy. Concentration of the spectral norm of Erdős–Rényi random graphs. Bernoulli, 26(3):2253–2274, August 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1587974540.

Li:2024:PFD

[LMZ24] Matthew T. C. Li, Youssef Marzouk, and Olivier Zahm. Principal feature detection via φ-Sobolev inequalities. Bernoulli, 30 (4):2979–3003, November 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

Leonard:2002:EST

[LN02] Christian Léonard and Jamal Najim. An extension of Sanov’s theorem: application to the Gibbs conditioning principle. Bernoulli, 8(6):721–743, December 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1076364803.

Lacour:2014:GFT

[LN14] Claire Lacour and Thanh Mai Pham Ngoc. Goodness-of-fit test for noisy directional data. Bernoulli, 20(4):2131–2168, November 2014. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1411134456.

Lee:2018:OSI

[LN18] Clement Lee and Peter Neal. Optimal scaling of the independence sampler: Theory and practice. Bernoulli, 24(3):1636– 1652, August 2018. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1517540456.

Lijoi:2014:BID

[LNP14] Antonio Lijoi, Bernardo Nipoti, and Igor Prünster. Bayesian inference with dependent normalized completely random measures. Bernoulli, 20(3):1260–1291, August 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1402488940.

Lungu:2001:OHI

[LØ01] Edward Lungu and Bernt Øksendal. Optimal harvesting from interacting populations in a stochastic environment. Bernoulli, 7(3):527–539, June 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1080004764.

Leonenko:2014:SMS

[LO14] Nikolai Leonenko and Andriy Olenko. Sojourn measures of Student and Fisher–Snedecor random fields. Bernoulli, 20 (3):1454–1483, August 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1402488946.

Lockhart:2012:CLL

[Loc12] Richard A. Lockhart. Conditional limit laws for goodness-of-fit tests. Bernoulli, 18(3):857–882, August 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1340887006.

Lounici:2014:HDC

[Lou14] Karim Lounici. High-dimensional covariance matrix estimation with missing observations. Bernoulli, 20(3):1029–1058, August 2014. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1402488933.

Lemdani:1999:LRT

[LP99] Mohamed Lemdani and Odile Pons. Likelihood ratio tests in contamination models. Bernoulli, 5(4):705–719, August 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1171899325.

Leon:2000:AOU

[LP00] José R. León and Gonzalo Perera. Approximation of the Ornstein–Uhlenbeck local time by harmonic oscillators. Bernoulli, 6(2):357–379, April 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1081788033.

Lamberton:2002:RCI

[LP02] Damien Lamberton and Gilles Pagès. Recursive computation of the invariant distribution of a diffusion. Bernoulli, 8(3):367– 405, April 2002. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1078779875.

Luschgy:2007:HRP

[LP07] Harald Luschgy and Gilles Pagès. High-resolution product quantization for Gaussian processes under sup-norm distortion. Bernoulli, 13(3):653–671, August 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1186503481.

Lemaire:2017:MRR

[LP17] Vincent Lemaire and Gilles Pagès. Multilevel Richardson– Romberg extrapolation. Bernoulli, 23(4A):2643–2692, November 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1494316828.

Lejay:2018:SEO

[LP18a] Antoine Lejay and Paolo Pigato. Statistical estimation of the Oscillating Brownian Motion. Bernoulli, 24(4B):3568–3602, November 2018. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1524038763.

Leng:2018:CES

[LP18b] Chenlei Leng and Guangming Pan. Covariance estimation via sparse Kronecker structures. Bernoulli, 24(4B):3833–3863, November 2018. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1524038771.

Letac:2018:DCC

[LP18c] Gérard Letac and Mauro Piccioni. Dirichlet curves, convex order and Cauchy distribution. Bernoulli, 24(1):1–29, February 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1501142434.

Liu:2024:NAS

[LP24] Qingwei Liu and Nicolas Privault. Normal approximation of subgraph counts in the random-connection model. Bernoulli, 30(4):3224–3250, November 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

Leonenko:2018:CCT

[LPSS18] N. N. Leonenko, I. Papić, A. Sikorskii, and N. Suvak. Correlated continuous time random walks and fractional Pearson diffusions. Bernoulli, 24(4B):3603–3627, November 2018. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1524038764.

Lopez:2013:SIR

[LPV13] Olivier Lopez, Valentin Patilea, and Ingrid Van Keilegom. Single index regression models in the presence of censoring depending on the covariates. Bernoulli, 19(3):721–747, August 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1372251141.

Lijoi:2004:EDC

[LPW04] Antonio Lijoi, Igor Prünster, and Stephen G. Walker. Extending Doob’s consistency theorem to nonparametric densities. Bernoulli, 10(4):651–663, August 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1093265634.

Lindgren:1995:HRC

[LR95] Georg Lindgren and Igor Rychlik. How reliable are contour curves? confidence sets for level contours. Bernoulli, 1(4): 301–319, December 1995. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1193758708.

Lahiri:2016:CLT

[LR16] S. N. Lahiri and Peter M. Robinson. Central limit theorems for long range dependent spatial linear processes. Bernoulli, 22 (1):345–375, February 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1443620853.

Law:2021:ICU

[LR21] Michael Law and Ya’acov Ritov. Inference without compatibility: Using exponential weighting for inference on a parameter of a linear model. Bernoulli, 27(3):1467–1495, August 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-27/issue-3/Inference-
without-compatibility--Using-exponential-weighting-
for-inference-on/10.3150/20-BEJ1280.full.

Liu:2023:PVL

[LR23] Suqi Liu and Miklós Z. Rácz. A probabilistic view of latent space graphs and phase transitions. Bernoulli, 29(3):2417–2441, August 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-29/issue-3/A-probabilistic-view-of-
latent-space-graphs-and-phase-transitions/10.3150/22-
BEJ1547.full.

Lo:2002:IMI

[LRD02] Violet S. F. Lo, Gareth O. Roberts, and Henry E. Daniels. Inverse method of images. Bernoulli, 8(1):53–80, February 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1078951089.

Leonenko:2017:NCL

[LRMT17] Nikolai Leonenko, M. Dolores Ruiz-Medina, and Murad S. Taqqu. Non-central limit theorems for random fields subordinated to gamma-correlated random fields. Bernoulli, 23 (4B):3469–3507, November 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1495505099.

Li:2023:NFI

[LRS23] Degui Li, Peter M. Robinson, and Han Lin Shang. Nonstationary fractionally integrated functional time series. Bernoulli, 29(2):1505–1526, May 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-29/issue-
2/Nonstationary-fractionally-integrated-functional-
time-series/10.3150/22-BEJ1508.full.

Letac:1995:RCF

[LS95] Gérard Letac and Vanamamalai Seshadri. A random continued fraction in R^{d+1} with an inverse Gaussian distribution. Bernoulli, 1(4):381–393, December 1995. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1193758713.

Lepski:1999:MNH

[LS99] Oleg V. Lepski and Vladimir G. Spokoiny. Minimax nonparametric hypothesis testing: the case of an inhomogeneous alternative. Bernoulli, 5(2):333–358, April 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1173147910.

Lanzinger:2000:WSD

[LS00] Hartmut Lanzinger and Ulrich Stadtmüller. Weighted sums for i.i.d. random variables with relatively thin tails. Bernoulli, 6(1):45–61, February 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1082665379.

Lepeltier:2002:ENE

[LS02a] Jean-Pierre Lepeltier and Jaime San Martín. On the existence or non-existence of solutions for certain backward stochastic differential equations. Bernoulli, 8(1):123–137, February 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1078951093.

Lifshits:2002:FET

[LS02b] Mikhail Lifshits and Zhan Shi. The first exit time of Brownian motion from a parabolic domain. Bernoulli, 8(6):745–765, December 2002. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1076364804.

Lanzinger:2003:BKL

[LS03] Hartmut Lanzinger and Ulrich Stadtmüller. Baum–Katz laws for certain weighted sums of independent and identically distributed random variables. Bernoulli, 9(6):985–1002, December 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1072215198.

Laukajtys:2013:PMS

[ŁS13] Weronika Łaukajtys and Leszek Słomiński. Penalization methods for the Skorokhod problem and reflecting SDEs with jumps. Bernoulli, 19(5A):1750–1775, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1383661201.

Ludena:2014:ESF

[LS14] Carenne Ludeña and Philippe Soulier. Estimating the scaling function of multifractal measures and multifractal random walks using ratios. Bernoulli, 20(1):334–376, February 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1390407292.

Lacaux:2016:TCE

[LS16a] Céline Lacaux and Gennady Samorodnitsky. Time-changed extremal process as a random sup measure. Bernoulli, 22(4): 1979–2000, November 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1462297672.

Lanconelli:2016:NLL

[LS16b] Alberto Lanconelli and Aurel I. Stan. A note on a local limit theorem for Wiener space valued random variables. Bernoulli, 22(4):2101–2112, November 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1462297676.

Lambert:2019:RBC

[LS19a] Amaury Lambert and Emmanuel Schertzer. Recovering the Brownian coalescent point process from the Kingman coalescent by conditional sampling. Bernoulli, 25(1):148–173, February 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1544605242.

Lee:2019:SLT

[LS19b] Ji Hyung Lee and Kyungchul Song. Stable limit theorems for empirical processes under conditional neighborhood dependence. Bernoulli, 25(2):1189–1224, May 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1551862848.

Leung:2019:APR

[LS19c] Dennis Leung and Qiman Shao. Asymptotic power of Rao’s score test for independence in high dimensions. Bernoulli, 25(1):241–263, February 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1544605246.

Lopusanschi:2020:AAR

[LS20] Olga Lopusanschi and Damien Simon. Area anomaly in the rough path Brownian scaling limit of hidden Markov walks. Bernoulli, 26(4):3111–3138, November 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
26/issue-4/Area-anomaly-in-the-rough-path-Brownian-
scaling-limit-of/10.3150/20-BEJ1217.full.

Lazic:2021:SGE

[LS21] Petra Lazić and Nikola Sandrić. On sub-geometric ergodicity of diffusion processes. Bernoulli, 27(1):348–380, February 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-27/issue-1/On-sub-geometric-ergodicity-
of-diffusion-processes/10.3150/20-BEJ1242.full.

Loh:2023:EPS

[LS23] Wei-Liem Loh and Saifei Sun. Estimating the parameters of some common Gaussian random fields with nugget under fixed-domain asymptotics. Bernoulli, 29(3):2519–2543, August 2023. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-29/issue-3/Estimating-the-
parameters-of-some-common-Gaussian-random-fields-with/
10.3150/22-BEJ1551.full.

Leung:2024:NBE

[LS24] Dennis Leung and Qi-Man Shao. Nonuniform Berry–Esseen bounds for studentized U-statistics. Bernoulli, 30(4):3276– 3302, November 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

Liu:2013:SNC

[LSW13] Weidong Liu, Qi-Man Shao, and Qiying Wang. Self-normalized Cramér type moderate deviations for the maximum of sums. Bernoulli, 19(3):1006–1027, August 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1372251151.

Liang:2020:UAC

[LSW20] Mingjie Liang, René L. Schilling, and Jian Wang. A unified approach to coupling SDEs driven by Lévy noise and some applications. Bernoulli, 26(1):664–693, February 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1574758842.

Liu:2022:NIR

[LSY^{+}22] Li Liu, Wen Su, Guosheng Yin, Xingqiu Zhao, and Ying Zhang. Nonparametric inference for reversed mean models with panel count data. Bernoulli, 28(4):2968–2997, November 2022. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-28/issue-4/Nonparametric-
inference-for-reversed-mean-models-with-panel-count-
data/10.3150/21-BEJ1444.full.

Levy:2000:RRP

[LT00] Joshua B. Levy and Murad S. Taqqu. Renewal reward processes with heavy-tailed inter-renewal times and heavy-tailed rewards. Bernoulli, 6(1):23–44, February 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1082665378.

Leng:2013:QRE

[LT13] Chenlei Leng and Xingwei Tong. A quantile regression estimator for censored data. Bernoulli, 19(1):344–361, February 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1358531753.

Li:2015:IIF

[LT15] Jinzhu Li and Qihe Tang. Interplay of insurance and financial risks in a discrete-time model with strongly regular variation. Bernoulli, 21(3):1800–1823, August 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1432732037.

Lerasle:2016:SOI

[LT16] Matthieu Lerasle and Daniel Y. Takahashi. Sharp oracle inequalities and slope heuristic for specification probabilities estimation in discrete random fields. Bernoulli, 22(1):325–344, February 2016. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1443620852.

Lu:2014:ESQ

[LTC14] Zudi Lu, Qingguo Tang, and Longsheng Cheng. Estimating spatial quantile regression with functional coefficients: A robust semiparametric framework. Bernoulli, 20(1):164–189, February 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1390407284.

Ling:2007:EIT

[LTL07] Shiqing Ling, Howell Tong, and Dong Li. Ergodicity and invertibility of threshold moving-average models. Bernoulli, 13 (1):161–168, February 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1175287726.

Lu:2022:CLT

[LTX22] Jianya Lu, Yuzhen Tan, and Lihu Xu. Central limit theorem and self-normalized Cramér-type moderate deviation for Euler–Maruyama scheme. Bernoulli, 28(2):937– 964, May 2022. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-28/issue-2/Central-limit-
theorem-and-self-normalized-Cram%c3%a9r-type-moderate-
deviation/10.3150/21-BEJ1372.full.

Luati:2011:AQC

[Lua11] Alessandra Luati. An approximate quantum Cramér–Rao bound based on skew information. Bernoulli, 17(2):628–642, May 2011. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1302009240.

Ludena:2000:PEG

[Lud00] Carenne Ludeña. Parametric estimation for Gaussian longrange dependent processes based on the log-periodogram. Bernoulli, 6(4):709–728, August 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1081449603.

LeGuevel:2012:FKL

[LV12] R. Le Guével and J. Lévy Véhel. A Ferguson–klass–LePage series representation of multistable multifractional motions and related processes. Bernoulli, 18(4):1099–1127, November 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1352727803.

Lowe:2015:CAM

[LV15] Matthias Löwe and Franck Vermet. Capacity of an associative memory model on random graph architectures. Bernoulli, 21 (3):1884–1910, August 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1432732041.

Leskela:2017:CCO

[LV17] Lasse Leskelä and Matti Vihola. Conditional convex orders and measurable martingale couplings. Bernoulli, 23(4A):2784– 2807, November 2017. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1494316832.

Lederer:2014:NCI

[LvdG14] Johannes Lederer and Sara van de Geer. New concentration inequalities for suprema of empirical processes. Bernoulli, 20 (4):2020–2038, November 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1411134452.

Lee:2004:IDN

[LW04] Mei-Ling Ting Lee and George A. Whitmore. Intensitydependent normalization in microarray analysis: a note of concern. Bernoulli, 10(6):943–949, December 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1106314844.

Lepski:2017:LBC

[LW17] O. V. Lepski and T. Willer. Lower bounds in the convolution structure density model. Bernoulli, 23(2):884–926, May 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1486177387.

Ledoit:2018:OEL

[LW18] Olivier Ledoit and Michael Wolf. Optimal estimation of a large-dimensional covariance matrix under Stein’s loss. Bernoulli, 24(4B):3791–3832, November 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1524038770.

Ledoit:2022:QSL

[LW22] Olivier Ledoit and Michael Wolf. Quadratic shrinkage for large covariance matrices. Bernoulli, 28(3):1519–1547, August 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-28/issue-3/Quadratic-shrinkage-for-
large-covariance-matrices/10.3150/20-BEJ1315.full.

Liu:2023:LER

[LW23] Mingchang Liu and Hao Wu. Loop-erased random walk branch of uniform spanning tree in topological polygons. Bernoulli, 29(2):1555–1577, May 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-29/issue-
2/Loop-erased-random-walk-branch-of-uniform-spanning-
tree-in/10.3150/22-BEJ1510.full.

Lam-Weil:2022:LMR

[LWCS22] Joseph Lam-Weil, Alexandra Carpentier, and Bharath K. Sriperumbudur. Local minimax rates for closeness testing of discrete distributions. Bernoulli, 28(2):1179–1197, May 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-28/issue-2/Local-minimax-rates-for-
closeness-testing-of-discrete-distributions/10.3150/
21-BEJ1382.full.

Lam-Weil:2022:MOG

[LWLL22] Joseph Lam-Weil, Béatrice Laurent, and Jean-Michel Loubes. Minimax optimal goodness-of-fit testing for densities and multinomials under a local differential privacy constraint. Bernoulli, 28(1):579–600, February 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-28/issue-
1/Minimax-optimal-goodness-of-fit-testing-for-densities-
and-multinomials/10.3150/21-BEJ1358.full.

Li:2015:EMC

[LWX15] Yuqiang Li, Wensheng Wang, and Yimin Xiao. Exact moduli of continuity for operator-scaling Gaussian random fields. Bernoulli, 21(2):930–956, May 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1429624966.

Li:2024:DCT

[LWY24] Weiming Li, Qinwen Wang, and Jianfeng Yao. Distance correlation test for high-dimensional independence. Bernoulli, 30 (4):3165–3192, November 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

Li:2022:EHD

[LWYZ22] Weiming Li, Qinwen Wang, Jianfeng Yao, and Wang Zhou. On eigenvalues of a high-dimensional spatial-sign covariance matrix. Bernoulli, 28(1):606–637, February 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-1/On-eigenvalues-of-a-high-dimensional-spatial-
sign-covariance-matrix/10.3150/21-BEJ1360.full.

Li:2021:FGM

[LX21] Yiting Li and Yuanyuan Xu. On fluctuations of global and mesoscopic linear statistics of generalized Wigner matrices. Bernoulli, 27(2):1057–1076, May 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-2/On-fluctuations-of-global-and-mesoscopic-
linear-statistics-of-generalized/10.3150/20-BEJ1265.
full.

Lee:2023:CTL

[LX23] Cheuk Yin Lee and Yimin Xiao. Chung-type law of the iterated logarithm and exact moduli of continuity for a class of anisotropic Gaussian random fields. Bernoulli, 29(1):523– 550, February 2023. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-29/issue-1/Chung-type-law-
of-the-iterated-logarithm-and-exact-moduli/10.3150/
22-BEJ1467.full.

Liu:2013:TSL

[LXY13] Jingchen Liu, Gongjun Xu, and Zhiliang Ying. Theory of self-learning Q-matrix. Bernoulli, 19(5A):1790–1817, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1383661203.

Lederer:2019:OIH

[LYG19] Johannes Lederer, Lu Yu, and Irina Gaynanova. Oracle inequalities for high-dimensional prediction. Bernoulli, 25(2): 1225–1255, May 2019. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1551862849.

Li:2021:EIO

[LYT21] Xiao-Yang Li, Zhi-Sheng Ye, and Cheng Yong Tang. Estimating the inter-occurrence time distribution from superposed renewal processes. Bernoulli, 27(4):2804–2826, November 2021. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-27/issue-4/Estimating-the-
inter-occurrence-time-distribution-from-superposed-
renewal-processes/10.3150/21-BEJ1331.full.

Ma:2010:SEE

[Ma10] Yanyuan Ma. A semiparametric efficient estimator in casecontrol studies. Bernoulli, 16(2):585–603, May 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1274821085.

Ma:2012:ATW

[Ma12] Zongming Ma. Accuracy of the Tracy–Widom limits for the extreme eigenvalues in white Wishart matrices. Bernoulli, 18 (1):322–359, February 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1327068628.

Machado:2023:COD

[Mac23] Mariela Pentón Machado. Convergence of the one-dimensional contact process with two types of particles and priority. Bernoulli, 29(2):1343–1367, May 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-2/Convergence-of-the-one-dimensional-contact-
process-with-two-types/10.3150/22-BEJ1501.full.

Mai:2014:EML

[Mai14] Hilmar Mai. Efficient maximum likelihood estimation for Lévydriven Ornstein–Uhlenbeck processes. Bernoulli, 20(2):919– 957, May 2014. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1393594010.

Malouche:2009:DFC

[Mal09] Dhafer Malouche. Determining full conditional independence by low-order conditioning. Bernoulli, 15(4):1179–1189, November 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1262962231.

Maller:2016:CLP

[Mal16] Ross A. Maller. Conditions for a Lévy process to stay positive near 0, in probability. Bernoulli, 22(4):1963–1978, November 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1462297671.

Mallein:2017:BRW

[Mal17] Bastien Mallein. Branching random walk with selection at critical rate. Bernoulli, 23(3):1784–1821, August 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1489737625.

Manstavicius:2007:HBD

[Man07] Martynas Manstavičius. Hausdorff–Besicovitch dimension of graphs and p-variation of some Lévy processes. Bernoulli, 13 (1):40–53, February 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1175287719.

Markussen:2003:LID

[Mar03] Bo Markussen. Likelihood inference for a discretely observed stochastic partial differential equation. Bernoulli, 9(5):745–762, October 2003. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1066418876.

Marquardt:2006:FLP

[Mar06] Tina Marquardt. Fractional Lévy processes with an application to long memory moving average processes. Bernoulli, 12(6): 1099–1126, December 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1165269152.

Markussen:2013:FDA

[Mar13] Bo Markussen. Functional data analysis in an operator-based mixed-model framework. Bernoulli, 19(1):1–17, February 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1358531738.

Marchina:2018:CIS

[Mar18] Antoine Marchina. Concentration inequalities for separately convex functions. Bernoulli, 24(4A):2906–2933, November 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1522051229.

Marguet:2019:USS

[Mar19] Aline Marguet. Uniform sampling in a structured branching population. Bernoulli, 25(4A):2649–2695, November 2019. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1568362039.

Masse:2004:ATD

[Mas04a] Jean-Claude Massé. Asymptotics for the Tukey depth process, with an application to a multivariate trimmed mean. Bernoulli, 10(3):397–419, June 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1089206404.

Masuda:2004:MOU

[Mas04b] Hiroki Masuda. On multidimensional Ornstein–Uhlenbeck processes driven by a general Lévy process. Bernoulli, 10(1): 97–120, February 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1077544605.

Mattner:2000:MSS

[Mat00] Lutz Mattner. Minimal sufficient statistics in location-scale parameter models. Bernoulli, 6(6):1121–1134, December 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1081194163.

Matsuda:2024:ICA

[Mat24] Takeru Matsuda. Inadmissibility of the corrected Akaike information criterion. Bernoulli, 30(2):1416–1440, May 2024. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
30/issue-2/Inadmissibility-of-the-corrected-Akaike-
information-criterion/10.3150/23-BEJ1638.full.

Maurer:2012:TC

[Mau12] Andreas Maurer. Thermodynamics and concentration. Bernoulli, 18(2):434–454, May 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1334580719.

Maurer:2019:BTI

[Mau19] Andreas Maurer. A Bernstein-type inequality for functions of bounded interaction. Bernoulli, 25(2):1451–1471, May 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL https://projecteuclid.org/euclid.bj/1551862856.

Mccullagh:2000:REA

[Mcc00] Peter Mccullagh. Resampling and exchangeable arrays. Bernoulli, 6(2):285–301, April 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1081788029.

McCullagh:2008:MLP

[McC08] Peter McCullagh. Marginal likelihood for parallel series. Bernoulli, 14(3):593–603, August 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1219669621.

Ma:2021:TAN

[MCC^{+}21] Yi-An Ma, Niladri S. Chatterji, Xiang Cheng, Nicolas Flammarion, Peter L. Bartlett, and Michael I. Jordan. Is there an analog of Nesterov acceleration for gradient-based MCMC? Bernoulli, 27(3):1942–1992, August 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-3/Is-there-an-analog-of-Nesterov-acceleration-
for-gradient-based/10.3150/20-BEJ1297.full.

McGill:2009:ADF

[McG09] Paul McGill. Asymptotics for diffusion first-passage laws. Bernoulli, 15(1):267–278, February 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1233669891.

Miguel:2023:LFE

[MCJ23] Sergio Brenner Miguel, Fabienne Comte, and Jan Johannes. Linear functional estimation under multiplicative measurement error. Bernoulli, 29(3):2247–2271, August 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-3/Linear-functional-estimation-under-multiplicative-
measurement-error/10.3150/22-BEJ1540.full.

Marquez-Carreras:1999:EDW

[MCSS99] David Márquez-Carreras and Marta Sanz-Solé. Expansion of the density: a Wiener-chaos approach. Bernoulli, 5(2):257– 274, April 1999. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1173147906.

Meister:2016:OCN

[Mei16] Alexander Meister. Optimal classification and nonparametric regression for functional data. Bernoulli, 22(3):1729–1744, August 2016. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1458132997.

Mellouk:2000:LDP

[Mel00] Mohamed Mellouk. A large-deviation principle for random evolution equations. Bernoulli, 6(6):977–999, December 2000. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1081194155.

Mou:2022:IBD

[MFWB22] Wenlong Mou, Nicolas Flammarion, Martin J. Wainwright, and Peter L. Bartlett. Improved bounds for discretization of Langevin diffusions: Near-optimal rates without convexity. Bernoulli, 28(3):1577–1601, August 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-3/Improved-bounds-for-discretization-of-Langevin-
diffusions--Near-optimal/10.3150/21-BEJ1343.full.

Mitra:2003:PCF

[MGS03] Sinjini Mitra, Tilmann Gneiting, and Zoltán Sasvári. Polynomial covariance functions on intervals. Bernoulli, 9(2):229– 241, April 2003. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1068128976.

Mendo:2010:EPO

[MH10] Luis Mendo and José M. Hernando. Estimation of a probability with optimum guaranteed confidence in inverse binomial sampling. Bernoulli, 16(2):493–513, May 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1274821081.

Mogensen:2022:GMS

[MH22] Søren Wengel Mogensen and Niels Richard Hansen. Graphical modeling of stochastic processes driven by correlated noise. Bernoulli, 28(4):3023–3050, November 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-4/Graphical-modeling-of-stochastic-processes-
driven-by-correlated-noise/10.3150/21-BEJ1446.full.

Minsker:2015:GMR

[Min15] Stanislav Minsker. Geometric median and robust estimation in Banach spaces. Bernoulli, 21(4):2308–2335, November 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1438777595.

Meyer:2017:BIS

[MJK17] Marco Meyer, Carsten Jentsch, and Jens-Peter Kreiss. Baxter’s inequality and sieve bootstrap for random fields. Bernoulli, 23(4B):2988–3020, November 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1495505083.

Moka:2020:PSG

[MK20] Sarat B. Moka and Dirk P. Kroese. Perfect sampling for Gibbs point processes using partial rejection sampling. Bernoulli, 26(3):2082–2104, August 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1587974534.

Ma:2010:VSM

[ML10] Yanyuan Ma and Runze Li. Variable selection in measurement error models. Bernoulli, 16(1):274–300, February 2010. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1265984712.

Marchetti:2011:CGM

[ML11] Giovanni M. Marchetti and Monia Lupparelli. Chain graph models of multivariate regression type for categorical data. Bernoulli, 17(3):827–844, August 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1310042846.

Ma:2020:SPC

[ML20] Zhuang Ma and Xiaodong Li. Subspace perspective on canonical correlation analysis: Dimension reduction and minimax rates. Bernoulli, 26(1):432–470, February 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1574758834.

Mladenovic:2008:LDP

[Mla08] Pavle Mladenović. Limit distributions for the problem of collecting pairs. Bernoulli, 14(2):419–439, May 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1208872112.

Mariadassou:2015:CGP

[MM15] Mahendra Mariadassou and Catherine Matias. Convergence of the groups posterior distribution in latent or stochastic block models. Bernoulli, 21(1):537–573, February 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1426597081.

Marchand:2020:ISA

[MM20] David Corlin Marchand and Ioan Manolescu. Influence of the seed in affine preferential attachment trees. Bernoulli, 26(3):1665–1705, August 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1587974520.

Marynych:2021:SRI

[MM21] Alexander Marynych and Ilya Molchanov. Sieving random iterative function systems. Bernoulli, 27(1):34–65, February 2021. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-27/issue-1/Sieving-random-
iterative-function-systems/10.3150/20-BEJ1221.full.

Mammen:2023:NEL

[MM23] Enno Mammen and Marilena Müller. Nonparametric estimation of locally stationary Hawkes processes. Bernoulli, 29(3):2062–2083, August 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-29/issue-
3/Nonparametric-estimation-of-locally-stationary-Hawkes-
processes/10.3150/22-BEJ1532.full.

Mao:2000:SDD

[MMP00] Xuerong Mao, Alexander Matasov, and Aleksey B. Piunovskiy. Stochastic differential delay equations with Markovian switching. Bernoulli, 6(1):73–90, February 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1082665381.

Molina:2006:CBB

[MMR06] Manuel Molina, Manuel Mota, and Alfonso Ramos. On L^{α}- convergence (1 ≤ α ≤ 2) for a bisexual branching process with population-size dependent mating. Bernoulli, 12(3):457– 468, June 2006. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1151525130.

Madiman:2023:BSR

[MMR23] Mokshay Madiman, James Melbourne, and Cyril Roberto. Bernoulli sums and Rényi entropy inequalities. Bernoulli, 29(2): 1578–1599, May 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-29/issue-2/Bernoulli-
sums-and-R%c3%a9nyi-entropy-inequalities/10.3150/22-
BEJ1511.full.

Maller:2008:GMC

[MMS08] Ross A. Maller, Gernot Müller, and Alex Szimayer. GARCH modelling in continuous time for irregularly spaced time series data. Bernoulli, 14(2):519–542, May 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1208872116.

MacPhee:2010:PTM

[MMVW10] Iain M. MacPhee, Mikhail V. Menshikov, Stanislav Volkov, and Andrew R. Wade. Passage-time moments and hybrid zones for the exclusion-voter model. Bernoulli, 16(4):1312–1342, November 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1290092908.

Martin:2017:EBP

[MMW17] Ryan Martin, Raymond Mess, and Stephen G. Walker. Empirical Bayes posterior concentration in sparse high-dimensional linear models. Bernoulli, 23(3):1822–1847, August 2017. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1489737626.

Mikosch:2000:SIE

[MN00] Thomas Mikosch and Rimas Norvaiša. Stochastic integral equations without probability. Bernoulli, 6(3):401–434, June 2000.

CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1081616698.

Michel:2022:LTT

[MN22] Bertrand Michel and Anthony Nouy. Learning with tree tensor networks: Complexity estimates and model selection. Bernoulli, 28(2):910–936, May 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-2/Learning-with-tree-tensor-networks--Complexity-
estimates-and-model/10.3150/21-BEJ1371.full.

Maida:2022:SDF

[MNN^{+}22] Mylène Maïda, Tien Dat Nguyen, Thanh Mai Pham Ngoc, Vincent Rivoirard, and Viet Chi Tran. Statistical deconvolution of the free Fokker–Planck equation at fixed time. Bernoulli, 28(2):771–802, May 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-2/Statistical-deconvolution-of-the-free-Fokker-
-Planck-equation-at-fixed/10.3150/21-BEJ1366.full.

Moller:2018:DPP

[MNPR18] Jesper Møller, Morten Nielsen, Emilio Porcu, and Ege Rubak. Determinantal point process models on the sphere. Bernoulli, 24(2):1171–1201, May 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1505980893.

Morters:2008:SVP

[MO08] Peter Mörters and Marcel Ortgiese. Small value probabilities via the branching tree heuristic. Bernoulli, 14(1):277–299, February 2008. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1202492794.

Mohle:1999:CDA

[Möh99] Martin Möhle. The concept of duality and applications to Markov processes arising in neutral population genetics models. Bernoulli, 5(5):761–777, October 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1171290398.

Mohle:2006:SDC

[Möh06] M. Möhle. On sampling distributions for coalescent processes with simultaneous multiple collisions. Bernoulli, 12(1):35–53, February 2006. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1141136647.

Mohle:2018:HPG

[Möh18] M. Möhle. Hitting probabilities for the Greenwood model and relations to near constancy oscillation. Bernoulli, 24(1):316– 332, February 2018. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1501142445.

Mondal:2015:ADI

[Mon15] Debashis Mondal. Applying Dynkin’s isomorphism: An alternative approach to understand the Markov property of the de Wijs process. Bernoulli, 21(3):1289–1303, August 2015. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1432732020.

Mazur:2020:ELF

[MOP20] Stepan Mazur, Dmitry Otryakhin, and Mark Podolskij. Estimation of the linear fractional stable motion. Bernoulli, 26(1): 226–252, February 2020. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1574758827.

Morien:1999:HBR

[Mor99] Pierre-Luc Morien. The Hölder and the Besov regularity of the density for the solution of a parabolic stochastic partial differential equation. Bernoulli, 5(2):275–298, April 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1173147907.

Mourrat:2013:RCM

[Mou13] Jean-Christophe Mourrat. On the rate of convergence in the martingale central limit theorem. Bernoulli, 19(2):633–645, May 2013. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1363192041.

Mendelson:2006:SVM

[MP06] Shahar Mendelson and Alain Pajor. On singular values of matrices with independent rows. Bernoulli, 12(5):761–773, October 2006. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1161614945.

Marinucci:2010:GRH

[MP10] Domenico Marinucci and Giovanni Peccati. Group representations and high-resolution central limit theorems for subordinated spherical random fields. Bernoulli, 16(3):798–824, August 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1281099885.

Meyer:2023:FDB

[MP23] Marco Meyer and Efstathios Paparoditis. A frequency domain bootstrap for general multivariate stationary processes. Bernoulli, 29(3):2367–2391, August 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-3/A-frequency-domain-bootstrap-for-general-
multivariate-stationary-processes/10.3150/22-BEJ1545.
full.

Maejima:2012:CMI

[MPAiS12] Makoto Maejima, Víctor Pérez-Abreu, and Ken iti Sato. A class of multivariate infinitely divisible distributions related to arcsine density. Bernoulli, 18(2):476–495, May 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1334580721.

Maticiuc:2010:VSS

[MPRZ10] Lucian Maticiuc, Etienne Pardoux, Aurel Răşcanu, and Adrian Zălinescu. Viscosity solutions for systems of parabolic variational inequalities. Bernoulli, 16(1):258–273, February 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1265984711.

Ma:1998:AIC

[MPS98] Jin Ma, Philip Protter, and Jaime San Martíin. Anticipating integrals for a class of martingales. Bernoulli, 4(1):81–114, March 1998. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1175865491.

Marin:2019:CAI

[MPS19a] Jean-Michel Marin, Pierre Pudlo, and Mohammed Sedki. Consistency of adaptive importance sampling and recycling schemes. Bernoulli, 25(3):1977–1998, August 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1560326434.

McKeague:2019:SMA

[MPS19b] Ian W. McKeague, Erol A. Peköz, and Yvik Swan. Stein’s method and approximating the quantum harmonic oscillator. Bernoulli, 25(1):89–111, February 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1544605240.

Merlevede:2019:FCM

[MPU19] Florence Merlevède, Magda Peligrad, and Sergey Utev. Functional CLT for martingale-like nonstationary dependent structures. Bernoulli, 25(4B):3203–3233, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1569398764.

Menshikov:2007:REC

[MPV07] Mikhail Menshikov, Dimitri Petritis, and Stanislav Volkov. Random environment on coloured trees. Bernoulli, 13(4): 966–980, November 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1194625598.

Musta:2022:PAE

[MPV22] Eni Musta, Valentin Patilea, and Ingrid Van Keilegom. A presmoothing approach for estimation in the semiparametric Cox mixture cure model. Bernoulli, 28(4):2689–2715, November 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-28/issue-4/A-presmoothing-approach-
for-estimation-in-the-semiparametric-Cox-mixture/10.
3150/21-BEJ1434.full.

McCullagh:2008:GFT

[MPW08] Peter McCullagh, Jim Pitman, and Matthias Winkel. Gibbs fragmentation trees. Bernoulli, 14(4):988–1002, November 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1225980568.

Mikosch:2010:LDM

[MR10] Thomas Mikosch and Alfredas Račkauskas. The limit distribution of the maximum increment of a random walk with regularly varying jump size distribution. Bernoulli, 16(4):1016–1038, November 2010. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1290092894.

Mikosch:2013:SVM

[MR13] Thomas Mikosch and Mohsen Rezapour. Stochastic volatility models with possible extremal clustering. Bernoulli, 19(5A): 1688–1713, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1383661199.

Maticiuc:2015:BSV

[MR15] Lucian Maticiuc and Aurel Răşcanu. Backward stochastic variational inequalities on random interval. Bernoulli, 21(2):1166– 1199, May 2015. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1429624974.

Mena:2016:DDE

[MR16] Ramsés H. Mena and Matteo Ruggiero. Dynamic density estimation with diffusive Dirichlet mixtures. Bernoulli, 22(2):901– 926, May 2016. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1447077764.

Massa:2017:ARG

[MR17] M. Sofia Massa and Eva Riccomagno. Algebraic representations of Gaussian Markov combinations. Bernoulli, 23(1):626–644, February 2017. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1475001369.

Mallein:2019:TSI

[MR19] Bastien Mallein and Sanjay Ramassamy. Two-sided infinitebin models and analyticity for Barak–Erdős graphs. Bernoulli, 25(4B):3479–3495, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1569398774.

Mikosch:2021:PLD

[MR21] Thomas Mikosch and Igor Rodionov. Precise large deviations for dependent subexponential variables. Bernoulli, 27(2):1319– 1347, May 2021. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-27/issue-2/Precise-large-
deviations-for-dependent-subexponential-variables/10.
3150/20-BEJ1276.full.

Mariucci:2020:BNA

[MRS20] Ester Mariucci, Kolyan Ray, and Botond Szabó. A Bayesian nonparametric approach to log-concave density estimation. Bernoulli, 26(2):1070–1097, May 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1580461573.

Maller:2022:SSL

[MRS22] Ross Maller, Sidney Resnick, and Soudabeh Shemehsavar. Splitting the sample at the largest uncensored observation. Bernoulli, 28(4):2234–2259, November 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-4/Splitting-the-sample-at-the-largest-uncensored-
observation/10.3150/21-BEJ1417.full.

Moulines:2007:NIP

[MRST07] E. Moulines, F. Roueff, A. Souloumiac, and T. Trigano. Nonparametric inference of photon energy distribution from indirect measurement. Bernoulli, 13(2):365–388, May 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1179498753.

Madras:2010:QBM

[MS10] Neal Madras and Deniz Sezer. Quantitative bounds for Markov chain convergence: Wasserstein and total variation distances. Bernoulli, 16(3):882–908, August 2010. CODEN ????

ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1281099888.

Marchand:2012:UMR

[MS12] Éric Marchand and William E. Strawderman. A unified minimax result for restricted parameter spaces. Bernoulli, 18(2): 635–643, May 2012. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL http://projecteuclid.org/
euclid.bj/1334580727.

Muller:2017:ETR

[MS17] Ursula U. Müller and Anton Schick. Efficiency transfer for regression models with responses missing at random. Bernoulli, 23(4A):2693–2719, November 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1494316829.

Mukherjee:2018:OAI

[MS18] Rajarshi Mukherjee and Subhabrata Sen. Optimal adaptive inference in random design binary regression. Bernoulli, 24(1): 699–739, February 2018. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1501142460.

Mallein:2023:NSC

[MS23a] Bastien Mallein and Quan Shi. A necessary and sufficient condition for the convergence of the derivative martingale in a branching Lévy process. Bernoulli, 29(1):597–624, February 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-29/issue-1/A-necessary-and-sufficient-
condition-for-the-convergence-of-the/10.3150/22-BEJ1470.
full.

Maruyama:2023:AEM

[MS23b] Yuzo Maruyama and William E. Strawderman. On admissible estimation of a mean vector when the scale is unknown. Bernoulli, 29(1):153–180, February 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-1/On-admissible-estimation-of-a-mean-vector-
when-the-scale/10.3150/21-BEJ1453.full.

Mies:2023:SGA

[MS23c] Fabian Mies and Ansgar Steland. Sequential Gaussian approximation for nonstationary time series in high dimensions. Bernoulli, 29(4):3114–3140, November 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-4/Sequential-Gaussian-approximation-for-nonstationary-
time-series-in-high-dimensions/10.3150/22-BEJ1577.full.

Minchev:2023:ADE

[MS23d] Martin Minchev and Mladen Savov. Asymptotics for densities of exponential functionals of subordinators. Bernoulli, 29(4):3307–3333, November 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-29/issue-
4/Asymptotics-for-densities-of-exponential-functionals-
of-subordinators/10.3150/23-BEJ1584.full.

Millet:2000:AST

[MSS00] Annie Millet and Marta Sanz-Solé. Approximation and support theorem for a wave equation in two space dimensions. Bernoulli, 6(5):887–915, October 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1081282694.

Milstein:2004:TDE

[MSS04] Grigori N. Milstein, John G. M. Schoenmakers, and Vladimir Spokoiny. Transition density estimation for stochastic differential equations via forward-reverse representations. Bernoulli, 10(2):281–312, April 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1082380220.

Molchanov:2014:IPR

[MSS14] Ilya Molchanov, Michael Schmutz, and Kaspar Stucki. Invariance properties of random vectors and stochastic processes based on the zonoid concept. Bernoulli, 20(3):1210–1233, August 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1402488938.

Mai:2016:EES

[MSS16] Jan-Frederik Mai, Steffen Schenk, and Matthias Scherer. Exchangeable exogenous shock models. Bernoulli, 22(2):1278– 1299, May 2016. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1447077775.

Ma:2013:SVS

[MSW13] Shujie Ma, Qiongxia Song, and Li Wang. Simultaneous variable selection and estimation in semiparametric modeling of longitudinal/clustered data. Bernoulli, 19(1):252–274, February 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1358531749.

Ma:2011:TIP

[MSWW11] Yutao Ma, Shi Shen, Xinyu Wang, and Liming Wu. Transportation inequalities: From Poisson to Gibbs measures. Bernoulli, 17(1):155–169, February 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1297173837.

Mukherjee:2020:DSE

[Muk20] Sumit Mukherjee. Degeneracy in sparse ERGMs with functions of degrees as sufficient statistics. Bernoulli, 26(2):1016–1043, May 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1580461571.

Mustapha:2007:GRE

[Mus07] Sami Mustapha. Gambler’s ruin estimates for random walks with symmetric spatially inhomogeneous increments. Bernoulli, 13(1):131–147, February 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1175287724.

Murphy:1999:OIS

[MV99] Susan A. Murphy and Aad W. Van Der Vaart. Observed information in semi-parametric models. Bernoulli, 5(3):381–412, June 1999. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1172617197.

Marriott:2004:GGP

[MV04] Paul Marriott and Paul Vos. On the global geometry of parametric models and information recovery. Bernoulli, 10(4):639– 649, August 2004. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1093265633.

Martin:2018:TSJ

[MV18a] Ole Martin and Mathias Vetter. Testing for simultaneous jumps in case of asynchronous observations. Bernoulli, 24(4B):3522– 3567, November 2018. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1524038762.

Mijatovic:2018:PEM

[MV18b] Aleksandar Mijatović and Jure Vogrinc. On the Poisson equation for Metropolis–Hastings chains. Bernoulli, 24(3):2401– 2428, August 2018. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1517540478.

Mammen:2019:EMR

[MVY19] Enno Mammen, Ingrid Van Keilegom, and Kyusang Yu. Expansion for moments of regression quantiles with applications to nonparametric testing. Bernoulli, 25(2):793–827, May 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL https://projecteuclid.org/euclid.bj/1551862835.

Massam:2004:MYP

[MW04] Hélène Massam and Jacek Wesołowski. The Matsumoto-yor property on trees. Bernoulli, 10(4):685–700, August 2004. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1093265636.

Morvai:2005:CP

[MW05] Gusztáv Morvai and Benjamin Weiss. On classifying processes. Bernoulli, 11(3):523–532, June 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1120591187.

Minsker:2020:RMU

[MW20] Stanislav Minsker and Xiaohan Wei. Robust modifications of U-statistics and applications to covariance estimation problems. Bernoulli, 26(1):694–727, February 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1574758843.

Meerschaert:2008:LTM

[MWX08] Mark Meerschaert, Dongsheng Wu, and Yimin Xiao. Local times of multifractional Brownian sheets. Bernoulli, 14(3):865– 898, August 2008. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1219669633.

Mukherjee:2010:LGM

[MWZ10] Sayan Mukherjee, Qiang Wu, and Ding-Xuan Zhou. Learning gradients on manifolds. Bernoulli, 16(1):181–207, February 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1265984708.

Mukherjee:2023:STS

[MX23] Sumit Mukherjee and Yuanzhe Xu. Statistics of the two star ERGM. Bernoulli, 29(1):24–51, February 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-1/Statistics-of-the-two-star-ERGM/10.3150/
21-BEJ1448.full.

Madan:2002:MMM

[MY02] Dilip B. Madan and Marc Yor. Making Markov martingales meet marginals: with explicit constructions. Bernoulli, 8 (4):509–536, August 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1078681382.

Ma:2023:SCC

[MY23] Zongming Ma and Fan Yang. Sample canonical correlation coefficients of high-dimensional random vectors with finite rank correlations. Bernoulli, 29(3):1905–1932, August 2023. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-3/Sample-canonical-correlation-coefficients-
of-high-dimensional-random-vectors-with/10.3150/22-
BEJ1525.full.

Mykland:2001:LCB

[Myk01] Per Aslak Mykland. Likelihood computations without Bartlett identities. Bernoulli, 7(3):473–485, June 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1080004761.

Matsuda:2006:SMD

[MYT06] Yasumasa Matsuda, Yoshihiro Yajima, and Howell Tong. Selecting models with different spectral density matrix structures by the cross-validated log likelihood criterion. Bernoulli, 12(2):221–249, April 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1145993973.

Mikosch:2014:FAE

[MZ14] Thomas Mikosch and Yuwei Zhao. A Fourier analysis of extreme events. Bernoulli, 20(2):803–845, May 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1393594007.

Muratov:2017:NDP

[MZ17] Anton Muratov and Sergei Zuyev. Neighbour-dependent point shifts and random exchange models: Invariance and attractors. Bernoulli, 23(1):539–551, February 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1475001365.

Mayer:2020:FWL

[MZZ20] Ulrike Mayer, Henryk Zähle, and Zhou Zhou. Functional weak limit theorem for a local empirical process of non-stationary time series and its application. Bernoulli, 26(3):1891–1911, August 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1587974527.

Nakamura:2015:CRZ

[Nak15] Takashi Nakamura. A complete Riemann zeta distribution and the Riemann hypothesis. Bernoulli, 21(1):604–617, February 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1426597083.

Naulet:2022:ABD

[Nau22] Zacharie Naulet. Adaptive Bayesian density estimation in sup-norm. Bernoulli, 28(2):1284–1308, May 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-2/Adaptive-Bayesian-density-estimation-in-
sup-norm/10.3150/21-BEJ1387.full.

Nicolussi:2017:TIC

[NC17] Federica Nicolussi and Roberto Colombi. Type II chain graph models for categorical data: A smooth subclass. Bernoulli, 23(2):863–883, May 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1486177386.

Nicolussi:2021:CSI

[NC21] Federica Nicolussi and Manuela Cazzaro. Context-specific independencies in stratified chain regression graphical models. Bernoulli, 27(3):2091–2116, August 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-3/Context-specific-independencies-in-stratified-
chain-regression-graphical-models/10.3150/20-BEJ1302.
full.

Ning:2024:BMA

[NC24] Bo Y.-C. Ning and Ismaël Castillo. Bayesian multiscale analysis of the Cox model. Bernoulli, 30(2):1525–1554, May 2024. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
30/issue-2/Bayesian-multiscale-analysis-of-the-Cox-
model/10.3150/23-BEJ1643.full.

Ndaoud:2023:HAM

[Nda23] Mohamed Ndaoud. Harmonic analysis meets stationarity: a general framework for series expansions of special Gaussian processes. Bernoulli, 29(3):2295–2317, August 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-3/Harmonic-analysis-meets-stationarity--A-
general-framework-for-series/10.3150/22-BEJ1542.full.

Neumann:2011:ARE

[Neu11] Michael H. Neumann. Absolute regularity and ergodicity of Poisson count processes. Bernoulli, 17(4):1268–1284, November 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1320417504.

Newton:2016:IDS

[New16] Nigel J. Newton. Infinite-dimensional statistical manifolds based on a balanced chart. Bernoulli, 22(2):711–731, May 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1447077759.

Nguyen:2015:PCP

[Ngu15] XuanLong Nguyen. Posterior contraction of the population polytope in finite admixture models. Bernoulli, 21(1):618–646, February 2015. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1426597084.

Nguyen:2016:BSH

[Ngu16] XuanLong Nguyen. Borrowing strengh in hierarchical Bayes: Posterior concentration of the Dirichlet base measure. Bernoulli, 22(3):1535–1571, August 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1458132991.

Nagy:2021:FIF

[NHV^{+}21] Stanislav Nagy, Sami Helander, Germain Van Bever, Lauri Viitasaari, and Pauliina Ilmonen. Flexible integrated functional depths. Bernoulli, 27(1):673–701, February 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-1/Flexible-integrated-functional-depths/10.
3150/20-BEJ1254.full.

Nicolaescu:2018:CPM

[Nic18] Liviu I. Nicolaescu. Critical points of multidimensional random Fourier series: Central limits. Bernoulli, 24(2):1128–1170, May 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1505980892.

Nielsen:2000:SEA

[Nie00] Søren Feodor Nielsen. The stochastic EM algorithm: estimation and asymptotic results. Bernoulli, 6(3):457–489, June 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1081616701.

Nieuwenhuis:2013:AMS

[Nie13] Gert Nieuwenhuis. Asymptotic mean stationarity and absolute continuity of point process distributions. Bernoulli, 19(5A): 1612–1636, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1383661196.

Ning:2021:SMC

[NIR21] Ning Ning, Edward L. Ionides, and Ya’acov Ritov. Scalable Monte Carlo inference and rescaled local asymptotic normality. Bernoulli, 27(4):2532–2555, November 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-4/Scalable-Monte-Carlo-inference-and-rescaled-
local-asymptotic-normality/10.3150/20-BEJ1321.full.

Ng:2014:MCC

[NJ14] Chi Tim Ng and Harry Joe. Model comparison with composite likelihood information criteria. Bernoulli, 20(4):1738–1764, November 2014. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1411134443.

Ning:2020:BLR

[NJG20] Bo Ning, Seonghyun Jeong, and Subhashis Ghosal. Bayesian linear regression for multivariate responses under group sparsity. Bernoulli, 26(3):2353–2382, August 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1587974544.

Na:2021:HDI

[NK21] Sen Na and Mladen Kolar. High-dimensional index volatility models via Stein’s identity. Bernoulli, 27(2):794–817, May 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-27/issue-2/High-dimensional-index-

volatility-models-via-Steins-identity/10.3150/20-BEJ1238.
full.

Nkurunziza:2021:IPG

[Nku21] Sévérien Nkurunziza. Inference problem in generalized fractional Ornstein–Uhlenbeck processes with change-point. Bernoulli, 27(1):107–134, February 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-27/issue-
1/Inference-problem-in-generalized-fractional-OrnsteinUhlenbeck-
processes-with-change-point/10.3150/20-BEJ1230.full.

Nordman:2014:CRE

[NL14] Daniel J. Nordman and Soumendra N. Lahiri. Convergence rates of empirical block length selectors for block bootstrap. Bernoulli, 20(2):958–978, May 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1393594011.

Ng:2018:CPE

[NLL18] Chi Tim Ng, Woojoo Lee, and Youngjo Lee. Change-point estimators with true identification property. Bernoulli, 24(1):616– 660, February 2018. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1501142457.

Niazadeh:2023:CBF

[NLS23] Rad Niazadeh, Renato Paes Leme, and Jon Schneider. Combinatorial Bernoulli factories. Bernoulli, 29(2):1246–1274, May 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-29/issue-2/Combinatorial-Bernoulli-
factories/10.3150/22-BEJ1497.full.

Najafi:2020:RCB

[NMR20] Amir Najafi, Seyed Abolfazl Motahari, and Hamid R. Rabiee. Reliable clustering of Bernoulli mixture models. Bernoulli, 26(2):1535–1559, May 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1580461589.

Nazarov:2022:SEG

[NN22] Alexander Nazarov and Yakov Nikitin. Spectral equivalence of Gaussian random functions: Operator approach. Bernoulli, 28(2):1448–1460, May 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-28/issue-
2/Spectral-equivalence-of-Gaussian-random-functions-
Operator-approach/10.3150/21-BEJ1396.full.

Nobel:2006:HTF

[Nob06] Andrew B. Nobel. Hypothesis testing for families of ergodic processes. Bernoulli, 12(2):251–269, April 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1145993974.

Norvaisa:2015:WPV

[Nor15] Rimas Norvaiša. Weighted power variation of integrals with respect to a Gaussian process. Bernoulli, 21(2):1260–1288, May 2015. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1429624978.

Novak:2014:LBA

[Nov14] S. Y. Novak. Lower bounds to the accuracy of inference on heavy tails. Bernoulli, 20(2):979–989, May 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1393594012.

Neumann:2008:GFT

[NP08] Michael H. Neumann and Efstathios Paparoditis. Goodnessof-fit tests for Markovian time series models: Central limit theory and bootstrap approximations. Bernoulli, 14(1):14–46, February 2008. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1202492783.

Niu:2021:BGS

[NPM21] Yabo Niu, Debdeep Pati, and Bani K. Mallick. Bayesian graph selection consistency under model misspecification. Bernoulli, 27(1):637–672, February 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-27/issue-
1/Bayesian-graph-selection-consistency-under-model-
misspecification/10.3150/20-BEJ1253.full.

Nourdin:2020:STD

[NPS20] Ivan Nourdin, Giovanni Peccati, and Stéphane Seuret. Sojourn time dimensions of fractional Brownian motion. Bernoulli, 26(3):1619–1634, August 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1587974518.

Nualart:2000:LDS

[NR00] David Nualart and Carles Rovira. Large deviations for stochastic Volterra equations. Bernoulli, 6(2):339–355, April 2000. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1081788032.

Neumann:2009:NEL

[NR09] Michael H. Neumann and Markus Reiß. Nonparametric estimation for Lévy processes from low-frequency observations. Bernoulli, 15(1):223–248, February 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1233669889.

Nardi:2012:LLG

[NR12] Yuval Nardi and Alessandro Rinaldo. The log-linear group-lasso estimator and its asymptotic properties. Bernoulli, 18(3):945– 974, August 2012. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1340887009.

Naulet:2024:AAS

[NRC24] Zacharie Naulet, Judith Rousseau, and François Caron. Asymptotic analysis of statistical estimators related to MultiGraphex processes under misspecification. Bernoulli, 30(4):2644–2675, November 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

NZi:1998:BSC

[NRT98] Modeste N’Zi, Bruno Rémillard, and Radu Theodorescu. Between Strassen and Chung normalizations for Lévy’s area process. Bernoulli, 4(1):115–125, March 1998. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1175865492.

Nualart:2001:BSD

[NS01] David Nualart and Wim Schoutens. Backward stochastic differential equations and Feynman–Kac formula for Lévy processes, with applications in finance. Bernoulli, 7(5):761–776, October 2001. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1079399541.

Nourdin:2007:CNC

[NS07] Ivan Nourdin and Thomas Simon. Correcting Newton–Côtes integrals by Lévy areas. Bernoulli, 13(3):695–711, August 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1186503483.

Nitinawarat:2017:USO

[NV17] Sirin Nitinawarat and Venugopal V. Veeravalli. Universal scheme for optimal search and stop. Bernoulli, 23(3):1759–1783, August 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1489737624.

Norros:1999:EAG

[NVV99] Ilkka Norros, Esko Valkeila, and Jorma Virtamo. An elementary approach to a Girsanov formula and other analytical results on fractional Brownian motions. Bernoulli, 5(4):571–587, August 1999. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1171899318.

Niles-Weed:2022:EWD

[NWR22] Jonathan Niles-Weed and Philippe Rigollet. Estimation of Wasserstein distances in the Spiked Transport Model. Bernoulli, 28(4):2663–2688, November 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-28/issue-
4/Estimation-of-Wasserstein-distances-in-the-Spiked-
Transport-Model/10.3150/21-BEJ1433.full.

Nagaev:2005:NLD

[NZ05] Alexander Nagaev and Alexander Zaigraev. New largedeviation local theorems for sums of independent and identically distributed random vectors when the limit distribution is α-stable. Bernoulli, 11(4):665–687, August 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1126126764.

Ng:2023:NHP

[NZM23] Tin Lok James Ng and Andrew Zammit-Mangion. Nonhomogeneous Poisson process intensity modeling and estimation using measure transport. Bernoulli, 29(1):815–838, February 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-29/issue-1/Non-homogeneous-Poisson-
process-intensity-modeling-and-estimation-using-measure/
10.3150/22-BEJ1480.full.

Owada:2020:CPD

[OB20] Takashi Owada and Omer Bobrowski. Convergence of persistence diagrams for topological crackle. Bernoulli, 26(3):2275– 2310, August 2020. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1587974541.

Obloj:2006:CCL

[Obl06] Jan Obloj. A complete characterization of local martingales which are functions of Brownian motion and its maximum. Bernoulli, 12(6):955–969, December 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1165269146.

Oates:2019:CRC

[OCBG19] Chris J. Oates, Jon Cockayne, François-Xavier Briol, and Mark Girolami. Convergence rates for a class of estimators based on Stein’s method. Bernoulli, 25(2):1141–1159, May 2019. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1551862846.

Olsson:2008:SMC

[OCDM08] Jimmy Olsson, Olivier Cappé, Randal Douc, and Éric Moulines. Sequential Monte Carlo smoothing with application to parameter estimation in nonlinear state space models. Bernoulli, 14 (1):155–179, February 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1202492789.

Olsson:2019:NSO

[OD19] Jimmy Olsson and Randal Douc. Numerically stable online estimation of variance in particle filters. Bernoulli, 25(2):1504– 1535, May 2019. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1551862858.

Ogihara:2015:LAM

[Ogi15] Teppei Ogihara. Local asymptotic mixed normality property for nonsynchronously observed diffusion processes. Bernoulli, 21 (4):2024–2072, November 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1438777585.

Ogihara:2018:PIN

[Ogi18] Teppei Ogihara. Parametric inference for nonsynchronously observed diffusion processes in the presence of market microstructure noise. Bernoulli, 24(4B):3318–3383, November 2018. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1524038756.

Ojeda:2023:IPF

[OH23] Gabriel Berzunza Ojeda and Cecilia Holmgren. Invariance principle for fragmentation processes derived from conditioned stable Galton–Watson trees. Bernoulli, 29(4):2745– 2770, November 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-29/issue-4/Invariance-
principle-for-fragmentation-processes-derived-from-
conditioned-stable-GaltonWatson/10.3150/22-BEJ1559.
full.

Ogata:1998:CEL

[OKM^{+}98] Yosihiko Ogata, Akio Kobayashi, Naoya Mikami, Yasuaki Murata, and Koichi Katsura. Correction of earthquake location estimation in a small-seismic-array system. Bernoulli, 4(2):167– 184, June 1998. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1174937293.

Ohn:2023:OBE

[OL23] Ilsang Ohn and Lizhen Lin. Optimal Bayesian estimation of Gaussian mixtures with growing number of components. Bernoulli, 29(2):1195–1218, May 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-2/Optimal-Bayesian-estimation-of-Gaussian-
mixtures-with-growing-number-of/10.3150/22-BEJ1495.
full.

Olofsson:2009:SBB

[Olo09] Peter Olofsson. Size-biased branching population measures and the multi-type x log x condition. Bernoulli, 15(4):1287–1304, November 2009. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1262962236.

Orsingher:2010:FPB

[OP10] Enzo Orsingher and Federico Polito. Fractional pure birth processes. Bernoulli, 16(3):858–881, August 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1281099887.

Orsingher:2011:FLB

[OP11] Enzo Orsingher and Federico Polito. On a fractional linear birth–death process. Bernoulli, 17(1):114–137, February 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1297173835.

Ottobre:2016:FSH

[OPPS16] Michela Ottobre, Natesh S. Pillai, Frank J. Pinski, and Andrew M. Stuart. A function space HMC algorithm with second order Langevin diffusion limit. Bernoulli, 22(1):60–106, February 2016. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1443620844.

Ohashi:2024:RPS

[OR24] Alberto Ohashi and Francesco Russo. Rough paths and symmetric-Stratonovich integrals driven by singular covariance Gaussian processes. Bernoulli, 30(2):1197–1230, May 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-30/issue-2/Rough-paths-and-symmetric-
Stratonovich-integrals-driven-by-singular-covariance/
10.3150/23-BEJ1629.full.

Orenstein:2022:RIS

[Ore22] Paulo Orenstein. Robust importance sampling with adaptive Winsorization. Bernoulli, 28(4):2862–2873, November 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-28/issue-4/Robust-importance-sampling-
with-adaptive-winsorization/10.3150/21-BEJ1440.full.

Owada:2015:MLM

[OS15] Takashi Owada and Gennady Samorodnitsky. Maxima of long memory stationary symmetric α-stable processes, and selfsimilar processes with stationary max-increments. Bernoulli, 21(3):1575–1599, August 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1432732030.

Osekowski:2007:IDM

[Osȩ07] Adam Osȩkowski. Inequalities for dominated martingales. Bernoulli, 13(1):54–79, February 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1175287720.

Osekowski:2009:SWT

[Osȩ09] Adam Osȩkowski. Sharp weak-type inequalities for differentially subordinated martingales. Bernoulli, 15(3):871–897, August 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1251463285.

Osekowski:2011:SMI

[Osȩ11] Adam Osȩkowski. Sharp maximal inequalities for the moments of martingales and non-negative submartingales. Bernoulli, 17 (4):1327–1343, November 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1320417506.

Osekowski:2020:BCM

[Ose20] Adam Osekowski. On the best constant in the martingale version of Fefferman’s inequality. Bernoulli, 26(3):1912–1926, August 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1587974528.

Oesting:2018:EFS

[OSZ18] Marco Oesting, Martin Schlather, and Chen Zhou. Exact and fast simulation of max-stable processes on a compact set using the normalized spectral representation. Bernoulli, 24(2): 1497–1530, May 2018. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1505980902.

Ogihara:2023:LAN

[OU23] Teppei Ogihara and Yuma Uehara. Local asymptotic normality for ergodic jump-diffusion processes via transition density approximation. Bernoulli, 29(3):2342–2366, August 2023. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-3/Local-asymptotic-normality-for-ergodic-jump-
diffusion-processes-via-transition/10.3150/22-BEJ1544.
full.

Oudshoorn:1998:AME

[Oud98] Catharina G. M. Oudshoorn. Asymptotically minimax estimation of a function with jumps. Bernoulli, 4(1):15–33, March 1998. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1175865488.

Ovcharov:2018:PSR

[Ovc18] Evgeni Y. Ovcharov. Proper scoring rules and Bregman divergence. Bernoulli, 24(1):53–79, February 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1501142436.

Olsson:2017:EPB

[OW17] Jimmy Olsson and Johan Westerborn. Efficient particle-based online smoothing in general hidden Markov models: The PaRIS algorithm. Bernoulli, 23(3):1951–1996, August 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1489737630.

Paindaveine:2022:MAG

[Pai22] Davy Paindaveine. On the measure of anchored Gaussian simplices, with applications to multivariate medians. Bernoulli, 28 (2):965–996, May 2022. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-28/issue-2/On-the-measure-
of-anchored-Gaussian-simplices-with-applications-to/
10.3150/21-BEJ1373.full.

Pakes:1999:RCL

[Pak99] Anthony G. Pakes. Revisiting conditional limit theorems for the mortal simple branching process. Bernoulli, 5(6):969–998, December 1999. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1143122298.

Pannier:2022:PLD

[Pan22] Alexandre Pannier. Pathwise large deviations for white noise chaos expansions. Bernoulli, 28(3):1961–1985, August 2022. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-28/issue-3/Pathwise-large-
deviations-for-white-noise-chaos-expansions/10.3150/
21-BEJ1407.full.

Paparoditis:2009:TTC

[Pap09] Efstathios Paparoditis. Testing temporal constancy of the spectral structure of a time series. Bernoulli, 15(4):1190–1221, November 2009. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1262962232.

Patie:2011:RFE

[Pat11] P. Patie. A refined factorization of the exponential law. Bernoulli, 17(2):814–826, May 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1302009248.

Pierce:2015:BFO

[PB15] Donald A. Pierce and Ruggero Bellio. Beyond first-order asymptotics for Cox regression. Bernoulli, 21(1):401–419, February 2015. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1426597075.

Proksch:2015:CBM

[PBD15] Katharina Proksch, Nicolai Bissantz, and Holger Dette. Confidence bands for multivariate and time dependent inverse regression models. Bernoulli, 21(1):144–175, February 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1426597066.

Poinas:2019:MPC

[PDL19] Arnaud Poinas, Bernard Delyon, and Frédéric Lavancier. Mixing properties and Central Limit Theorem for associated point processes. Bernoulli, 25(3):1724–1754, August 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1560326425.

Penda:2017:TCI

[PEBG17] S. Valère Bitseki Penda, Mikael Escobar-Bach, and Arnaud Guillin. Transportation and concentration inequalities for bifurcating Markov chains. Bernoulli, 23(4B):3213–3242, November 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1495505091.

Peccati:2003:EFT

[Pec03] Giovanni Peccati. Explicit formulae for time-space Brownian chaos. Bernoulli, 9(1):25–48, February 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1068129009.

Peccati:2008:MIR

[Pec08] Giovanni Peccati. Multiple integral representation for functionals of Dirichlet processes. Bernoulli, 14(1):91–124, February 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1202492786.

Pedersen:1995:CAN

[Ped95] Asger Roer Pedersen. Consistency and asymptotic normality of an approximate maximum likelihood estimator for discretely observed diffusion processes. Bernoulli, 1(3):257–279, September 1995. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1193667818.

Penrose:2007:LLN

[Pen07] Mathew D. Penrose. Laws of large numbers in stochastic geometry with statistical applications. Bernoulli, 13(4):1124–1150, November 2007. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1194625605.

Perera:2002:ISC

[Per02] Gonzalo Perera. Irregular sets and central limit theorems. Bernoulli, 8(5):627–642, October 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1078435221.

Pettersson:1997:PSR

[Pet97] Roger Pettersson. Penalization schemes for reflecting stochastic differential equations. Bernoulli, 3(4):403–414, December 1997. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1175882215.

Peyre:2017:FBM

[Pey17] Rémi Peyre. Fractional Brownian motion satisfies two-way crossing. Bernoulli, 23(4B):3571–3597, November 2017. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1495505102.

Porcu:2023:RET

[PFEP23] Emilio Porcu, Samuel F. Feng, Xavier Emery, and Ana P. Peron. Rudin extension theorems on product spaces, turning bands, and random fields on balls cross time. Bernoulli, 29(2): 1464–1475, May 2023. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-29/issue-2/Rudin-extension-
theorems-on-product-spaces-turning-bands-and-random/
10.3150/22-BEJ1506.full.

Page:2021:GLM

[PG21] Stephen Page and Steffen Grünewälder. The Goldenshluger– Lepski method for constrained least-squares estimators over RKHSs. Bernoulli, 27(4):2241–2266, November 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-4/The-GoldenshlugerLepski-method-for-constrained-
least-squares-estimators-over-RKHSs/10.3150/20-BEJ1307.
full.

Portier:2017:EBP

[PGV17] François Portier, Anouar El Ghouch, and Ingrid Van Keilegom. Efficiency and bootstrap in the promotion time cure model. Bernoulli, 23(4B):3437–3468, November 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1495505098.

Penda:2017:AEB

[PHO17] S. Valère Bitseki Penda, Marc Hoffmann, and Adélaïde Olivier. Adaptive estimation for bifurcating Markov chains. Bernoulli, 23(4B):3598–3637, November 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1495505103.

Pitman:1997:PSD

[Pit97] Jim Pitman. Partition structures derived from Brownian motion and stable subordinators. Bernoulli, 3(1):79–96, February 1997. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1178291933.

Pollock:2016:ESS

[PJR16] Murray Pollock, Adam M. Johansen, and Gareth O. Roberts. On the exact and ε-strong simulation of (jump) diffusions. Bernoulli, 22(2):794–856, May 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1447077762.

Paulin:2019:EBS

[PJT19] Daniel Paulin, Ajay Jasra, and Alexandre Thiery. Error bounds for sequential Monte Carlo samplers for multimodal distributions. Bernoulli, 25(1):310–340, February 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1544605248.

Park:2009:BKR

[PLH09] B. U. Park, Y. K. Lee, and S. Ha. L_{2} boosting in kernel regression. Bernoulli, 15(3):599–613, August 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1251463273.

Pons:2007:EAC

[Pon07] Odile Pons. Estimation of absolutely continuous distributions for censored variables in two-sample nonparametric and semiparametric models. Bernoulli, 13(1):92–113, February 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1175287722.

Pages:2005:OQM

[PP05] Gilles Pagès and Huyên Pham. Optimal quantization methods for nonlinear filtering with discrete-time observations. Bernoulli, 11(5):893–932, October 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1130077599.

Pages:2009:ADS

[PP09] Gilles Pagès and Fabien Panloup. Approximation of the distribution of a stationary Markov process with application to option pricing. Bernoulli, 15(1):146–177, February 2009. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1233669886.

Perkowski:2016:PSI

[PP16] Nicolas Perkowski and David J. Prömel. Pathwise stochastic integrals for model free finance. Bernoulli, 22(4):2486–2520, November 2016. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1462297687.

Poon:2022:DFG

[PP22] Clarice Poon and Gabriel Peyré. Degrees of freedom for off-the-grid sparse estimation. Bernoulli, 28(3):2095–2121, August 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-28/issue-3/Degrees-of-
freedom-for-off-the-grid-sparse-estimation/10.3150/
21-BEJ1414.full.

Peng:2013:WED

[PQY13] Liang Peng, Linyi Qian, and Jingping Yang. Weighted estimation of the dependence function for an extreme-value distribution. Bernoulli, 19(2):492–520, May 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1363192036.

Pistone:1999:ESM

[PR99] Giovanni Pistone and Maria Piera Rogantin. The exponential statistical manifold: mean parameters, orthogonality and space transformations. Bernoulli, 5(4):721–760, August 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1171899326.

Palmowski:2002:TEC

[PR02a] Zbigniew Palmowski and Tomasz Rolski. A technique for exponential change of measure for Markov processes. Bernoulli, 8(6):767–785, December 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1076364805.

Philippe:2002:NIP

[PR02b] Anne Philippe and Judith Rousseau. Non-informative priors in the case of Gaussian long-memory processes. Bernoulli, 8 (4):451–473, August 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1078681379.

Peszat:2005:LNA

[PR05] Szymon Peszat and Francesco Russo. Large-noise asymptotics for one-dimensional diffusions. Bernoulli, 11(2):247–262, April 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1116340293.

Patilea:2006:PLE

[PR06] Valentin Patilea and Jean-Marie Rolin. Product-limit estimators of the survival function for two modified forms of currentstatus data. Bernoulli, 12(5):801–819, October 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1161614947.

Prunster:2013:BNA

[PR13] Igor Prünster and Matteo Ruggiero. A Bayesian nonparametric approach to modeling market share dynamics. Bernoulli, 19(1):64–92, February 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1358531741.

Papaspiliopoulos:2014:OFD

[PR14] Omiros Papaspiliopoulos and Matteo Ruggiero. Optimal filtering and the dual process. Bernoulli, 20(4):1999–2019, November 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1411134451.

Pakkanen:2016:FLT

[PR16] Mikko S. Pakkanen and Anthony Réveillac. Functional limit theorems for generalized variations of the fractional Brownian sheet. Bernoulli, 22(3):1671–1708, August 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1458132995.

Phandoidaen:2022:EPT

[PR22] Nathawut Phandoidaen and Stefan Richter. Empirical process theory for locally stationary processes. Bernoulli, 28(1):453– 480, February 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-28/issue-1/Empirical-
process-theory-for-locally-stationary-processes/10.
3150/21-BEJ1351.full.

Puccetti:2023:GCC

[PR23] Giovanni Puccetti and Ludger Rüschendorf. General construction and classes of explicit. Bernoulli, 29(1):839–874, February 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-29/issue-1/General-construction-and-
classes-of-explicit-L1-optimal-couplings/10.3150/22-
BEJ1481.full.

Profeta:2024:MDS

[Pro24] Christophe Profeta. Maximal displacement of spectrally negative branching Lévy processes. Bernoulli, 30(2):961–982, May 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-30/issue-2/Maximal-displacement-of-
spectrally-negative-branching-L%c3%a9vy-processes/10.
3150/23-BEJ1620.full.

Pekoz:2013:TVE

[PRR13] Erol A. Peköz, Adrian Röllin, and Nathan Ross. Total variation error bounds for geometric approximation. Bernoulli, 19(2):610–632, May 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1363192040.

Pekoz:2019:PUI

[PRR19] Erol Peköz, Adrian Röllin, and Nathan Ross. Pólya urns with immigration at random times. Bernoulli, 25(1):189–220, February 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1544605244.

Paindaveine:2020:STW

[PRV20] Davy Paindaveine, Julien Remy, and Thomas Verdebout. Sign tests for weak principal directions. Bernoulli, 26(4):2987–3016, November 2020. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-26/issue-4/Sign-tests-for-
weak-principal-directions/10.3150/20-BEJ1213.full.

Pardo:2013:DEF

[PRvS13] J. C. Pardo, V. Rivero, and K. van Schaik. On the density of exponential functionals of Lévy processes. Bernoulli, 19(5A): 1938–1964, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1383661209.

Puhalskii:1998:LDE

[PS98] Anatolii Puhalskii and Vladimir Spokoiny. On large-deviation efficiency in statistical inference. Bernoulli, 4(2):203–272, June 1998. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1174937295.

Pigorsch:2009:DSD

[PS09] Christian Pigorsch and Robert Stelzer. On the definition, stationary distribution and second order structure of positive semidefinite Ornstein–Uhlenbeck type processes. Bernoulli, 15 (3):754–773, August 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1251463280.

Porcu:2011:SPN

[PS11] Emilio Porcu and René L. Schilling. From Schoenberg to Pick–Nevanlinna: Toward a complete picture of the variogram class. Bernoulli, 17(1):441–455, February 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1297173849. See addendum

[PS13c].

Pages:2012:AMR

[PS12] Gilles Pagès and Abass Sagna. Asymptotics of the maximal radius of an L^{r}-optimal sequence of quantizers. Bernoulli, 18 (1):360–389, February 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1327068629.

Peng:2013:ELA

[PS13a] Hanxiang Peng and Anton Schick. Empirical likelihood approach to goodness of fit testing. Bernoulli, 19(3):954–981, August 2013. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1372251149.

Picard:2013:SLM

[PS13b] Jean Picard and Catherine Savona. Smoothness of the law of manifold-valued Markov processes with jumps. Bernoulli, 19(5A):1880–1919, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1383661207.

Porcu:2013:ASP

[PS13c] Emilio Porcu and René L. Schilling. Addendum to “From Schoenberg to Pick–Nevanlinna: Towards a complete picture of the variogram class”. Bernoulli, 19(5B):2768, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/

1386078620. See [PS11].

Puplinskaite:2016:AAR

[PS16] Donata Puplinskaitė and Donatas Surgailis. Aggregation of autoregressive random fields and anisotropic long-range dependence. Bernoulli, 22(4):2401–2441, November 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1462297685.

Pavlyukevich:2020:SSD

[PS20] Ilya Pavlyukevich and Georgiy Shevchenko. Stratonovich stochastic differential equation with irregular coefficients: Girsanov’s example revisited. Bernoulli, 26(2):1381–1409, May 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1580461583.

Padoan:2022:JIE

[PS22a] Simone A. Padoan and Gilles Stupfler. Joint inference on extreme expectiles for multivariate heavy-tailed distributions. Bernoulli, 28(2):1021–1048, May 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-2/Joint-inference-on-extreme-expectiles-for-
multivariate-heavy-tailed-distributions/10.3150/21-
BEJ1375.full.

Pilipauskaite:2022:LSL

[PS22b] Vytaute Pilipauskaitė and Donatas Surgailis. Local scaling limits of Lévy driven fractional random fields. Bernoulli, 28(4):2833–2861, November 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-28/issue-
4/Local-scaling-limits-of-L%c3%a9vy-driven-fractional-
random-fields/10.3150/21-BEJ1439.full.

Palmowski:2024:MEF

[PSS24] Zbigniew Palmowski, Hristo Sariev, and Mladen Savov. Moments of exponential functionals of Lévy processes on a deterministic horizon - identities and explicit expressions. Bernoulli, 30(4):2547–2571, November 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

Privault:2000:CKF

[PSV00] Nicolas Privault, Josep Lluís Solé, and Josep Vives. Chaotic Kabanov formula for the Azéma martingales. Bernoulli, 6(4):633– 651, August 2000. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1081449598.

Pipiras:2000:LRR

[PT00] Vladas Pipiras and Murad S. Taqqu. The limit of a renewal reward process with heavy-tailed rewards is not a linear fractional stable motion. Bernoulli, 6(4):607–614, August 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1081449595.

Pipiras:2001:CDI

[PT01] Vladas Pipiras and Murad S. Taqqu. Are classes of deterministic integrands for fractional Brownian motion on an interval complete? Bernoulli, 7(6):873–897, December 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1078951127.

Pipiras:2003:CLT

[PT03] Vladas Pipiras and Murad S. Taqqu. Central limit theorems for partial sums of bounded functionals of infinite-variance moving averages. Bernoulli, 9(5):833–855, October 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1066418880.

Peccati:2008:CLT

[PT08] Giovanni Peccati and Murad S. Taqqu. Central limit theorems for double Poisson integrals. Bernoulli, 14(3):791–821, August 2008. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1219669630.

Pitman:2015:SBP

[PT15] Jim Pitman and Ngoc M. Tran. Size-biased permutation of a finite sequence with independent and identically distributed terms. Bernoulli, 21(4):2484–2512, November 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1438777602.

Peng:2018:JEL

[PT18a] Hanxiang Peng and Fei Tan. Jackknife empirical likelihood goodness-of-fit tests for U-statistics based general estimating equations. Bernoulli, 24(1):449–464, February 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1501142451.

Pitman:2018:TFM

[PT18b] Jim Pitman and Wenpin Tang. Tree formulas, mean first passage times and Kemeny’s constant of a Markov chain. Bernoulli, 24(3):1942–1972, August 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1517540464.

Pete:2021:FEI

[PT21] Gábor Pete and Ádám Timár. Finite-energy infinite clusters without anchored expansion. Bernoulli, 27(4):2353–2361, November 2021. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-27/issue-4/Finite-energy-
infinite-clusters-without-anchored-expansion/10.3150/
20-BEJ1311.full.

Pipiras:2007:BCF

[PTA07] Vladas Pipiras, Murad S. Taqqu, and Patrice Abry. Bounds for the covariance of functions of infinite variance stable random variables with applications to central limit theorems and wavelet-based estimation. Bernoulli, 13(4):1091–1123, November 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1194625604.

Pipiras:2004:SFA

[PTL04] Vladas Pipiras, Murad S. Taqqu, and Joshua B. Levy. Slow, fast and arbitrary growth conditions for renewal-reward processes when both the renewals and the rewards are heavytailed. Bernoulli, 10(1):121–163, February 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1077544606.

Puig:2007:CCD

[PV07] Pedro Puig and Jordi Valero. Characterization of count data distributions involving additivity and binomial subsampling. Bernoulli, 13(2):544–555, May 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1179498760.

Podolskij:2009:EVF

[PV09] Mark Podolskij and Mathias Vetter. Estimation of volatility functionals in the simultaneous presence of microstructure noise and jumps. Bernoulli, 15(3):634–658, August 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1251463275.

Paindaveine:2015:NCD

[PV15] Davy Paindaveine and Germain Van Bever. Nonparametrically consistent depth-based classifiers. Bernoulli, 21(1):62–82, February 2015. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1426597064.

Paindaveine:2016:HDS

[PV16]  Davy Paindaveine and Thomas Verdebout.  On high-

dimensional sign tests. Bernoulli, 22(3):1745–1769, August 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1458132998.

Preuss:2013:TSB

[PVD13] Philip Preuß, Mathias Vetter, and Holger Dette. A test for stationarity based on empirical processes. Bernoulli, 19(5B): 2715–2749, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1386078618.

Patil:2004:CPI

[PW04] Prakash N. Patil and Andrew T. A. Wood. Counting process intensity estimation by orthogonal wavelet methods. Bernoulli, 10(1):1–24, February 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1077544601.

Pronzato:2017:EGV

[PWZ17] Luc Pronzato, Henry P. Wynn, and Anatoly A. Zhigljavsky. Extended generalised variances, with applications. Bernoulli, 23(4A):2617–2642, November 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1494316827.

Peng:2010:RBC

[PX10] Shige Peng and Mingyu Xu. Reflected BSDE with a constraint and its applications in an incomplete market. Bernoulli, 16 (3):614–640, August 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1281099878.

Pan:2013:SIL

[PXH13] Xiaoqing Pan, Maochao Xu, and Taizhong Hu. Some inequalities of linear combinations of independent random variables: II. Bernoulli, 19(5A):1776–1789, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1383661202.

Pitman:1999:LTR

[PY99] Jim Pitman and Marc Yor. Laplace transforms related to excursions of a one-dimensional diffusion. Bernoulli, 5(2):249– 255, April 1999. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1173147905.

Putter:2001:ECF

[PY01] Hein Putter and G. Alastair Young. On the effect of covariance function estimation on the accuracy of kriging predictors. Bernoulli, 7(3):421–438, June 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1080004758.

Pitman:2003:HOI

[PY03] Jim Pitman and Marc Yor. Hitting, occupation and inverse local times of one-dimensional diffusions: martingale and excursion approaches. Bernoulli, 9(1):1–24, February 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1068129008.

Pitman:2019:GIP

[PY19] Jim Pitman and Yuri Yakubovich. Gaps and interleaving of point processes in sampling from a residual allocation model. Bernoulli, 25(4B):3623–3651, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1569398779.

Peccati:2014:UGF

[PZ14] Giovanni Peccati and Cengbo Zheng. Universal Gaussian fluctuations on the discrete Poisson chaos. Bernoulli, 20(2):697– 715, May 2014. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1393594003.

Peligrad:2019:CLT

[PZ19] Magda Peligrad and Na Zhang. Central limit theorem for Fourier transform and periodogram of random fields. Bernoulli, 25(1):499–520, February 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1544605254.

Qiao:2021:ACR

[Qia21a] Wanli Qiao. Asymptotic confidence regions for density ridges. Bernoulli, 27(2):946–975, May 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-2/Asymptotic-confidence-regions-for-density-
ridges/10.3150/20-BEJ1261.full.

Qiao:2021:NES

[Qia21b]  Wanli Qiao.  Nonparametric estimation of surface in-

tegrals on level sets. Bernoulli, 27(1):155–191, February 2021. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-27/issue-1/Nonparametric-
estimation-of-surface-integrals-on-level-sets/10.3150/
20-BEJ1232.full.

Qiu:2018:HTT

[Qiu18] Jinniao Qiu. Hörmander-type theorem for Itô processes and related backward SPDEs. Bernoulli, 24(2):956–970, May 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL https://projecteuclid.org/euclid.bj/1505980884.

Qiu:2019:RSD

[Qiu19] Yanqi Qiu. Rigid stationary determinantal processes in non-Archimedean fields. Bernoulli, 25(1):75–88, February 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL https://projecteuclid.org/euclid.bj/1544605239.

Qin:2022:CRT

[QJ22] Qian Qin and Galin L. Jones. Convergence rates of twocomponent MCMC samplers. Bernoulli, 28(2):859–885, May 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-28/issue-2/Convergence-rates-of-two-
component-MCMC-samplers/10.3150/21-BEJ1369.full.

Qiu:2021:USB

[QLC21] Hongxiang Qiu, Alex Luedtke, and Marco Carone. Universal sieve-based strategies for efficient estimation using machine learning tools. Bernoulli, 27(4):2300–2336, November 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-27/issue-4/Universal-
sieve-based-strategies-for-efficient-estimation-using-
machine-learning/10.3150/20-BEJ1309.full.

Qiao:2018:ERL

[QP18] Wanli Qiao and Wolfgang Polonik. Extrema of rescaled locally stationary Gaussian fields on manifolds. Bernoulli, 24(3):1834– 1859, August 2018. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1517540461.

Quer-Sardanyons:2004:SWE

[QSSS04] Lluís Quer-Sardanyons and Marta Sanz-Solé. A stochastic wave equation in dimension 3: smoothness of the law. Bernoulli, 10 (1):165–186, February 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1077544607.

Qin:2005:ELB

[QT05] Gengsheng Qin and Min Tsao. Empirical likelihood based inference for the derivative of the nonparametric regression function. Bernoulli, 11(4):715–735, August 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1126126766.

Rabhi:2017:IUB

[RA17] Yassir Rabhi and Masoud Asgharian. Inference under biased sampling and right censoring for a change point in the hazard function. Bernoulli, 23(4A):2720–2745, November 2017. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1494316830.

Raic:2019:MBE

[Rai19] Martin Raic. A multivariate Berry–Esseen theorem with explicit constants. Bernoulli, 25(4A):2824–2853, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL https://projecteuclid.org/euclid.bj/1568362044.

Randrianarisoa:2022:SAS

[Ran22] Thibault Randrianarisoa. Smoothing and adaptation of shifted Pólya tree ensembles. Bernoulli, 28(4):2492–2517, November 2022. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-28/issue-4/Smoothing-and-
adaptation-of-shifted-P%c3%b3lya-tree-ensembles/10.
3150/21-BEJ1426.full.

Rassoul-Agha:2017:VFD

[RASY17] Firas Rassoul-Agha, Timo Seppäläinen, and Atilla Yilmaz. Variational formulas and disorder regimes of random walks in random potentials. Bernoulli, 23(1):405–431, February 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1475001359.

Rath:2019:FPM

[Rát19] Balázs Ráth. Feller property of the multiplicative coalescent with linear deletion. Bernoulli, 25(1):221–240, February 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL https://projecteuclid.org/euclid.bj/1544605245.

Reynaud-Bouret:2006:PPE

[RB06] Patricia Reynaud-Bouret. Penalized projection estimators of the Aalen multiplicative intensity. Bernoulli, 12(4):633–661, August 2006. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1155735930.

Ravner:2019:EIL

[RBM19] Liron Ravner, Onno Boxma, and Michel Mandjes. Estimating the input of a Lévy-driven queue by Poisson sampling of the workload process. Bernoulli, 25(4B):3734–3761, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1569398783.

Rotnitzky:2000:LBI

[RCBR00] Andrea Rotnitzky, David R. Cox, Matteo Bottai, and James Robins. Likelihood-based inference with singular information matrix. Bernoulli, 6(2):243–284, April 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1081788028.

Rebelles:2015:PAE

[Reb15] Gilles Rebelles. Pointwise adaptive estimation of a multivariate density under independence hypothesis. Bernoulli, 21(4): 1984–2023, November 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1438777584.

Reiss:2005:AEA

[Rei05] Markus Reiss. Adaptive estimation for affine stochastic delay differential equations. Bernoulli, 11(1):67–102, January 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1110228243.

Reid:2013:ALI

[Rei13] Nancy Reid. Aspects of likelihood inference. Bernoulli, 19(4): 1404–1418, September 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1377612858.

Ren:2008:SWE

[Ren08] Jian-Jian Ren. Smoothed weighted empirical likelihood ratio confidence intervals for quantiles. Bernoulli, 14(3):725–748, August 2008. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1219669627.

Revesz:1996:DPC

[Rév96] Pál Révész. Distribution of the particles of a critical branching Wiener process. Bernoulli, 2(1):63–80, March 1996. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1193758790.

Rigollet:2006:ADE

[Rig06] Philippe Rigollet. Adaptive density estimation using the blockwise Stein method. Bernoulli, 12(2):351–370, April 2006. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1145993978.

Rivero:2005:RES

[Riv05] Víctor Rivero. Recurrent extensions of self-similar Markov processes and Cramér’s condition. Bernoulli, 11(3):471–509, June 2005. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1120591185.

Rivoirard:2006:NEW

[Riv06] Vincent Rivoirard. Nonlinear estimation over weak Besov spaces and minimax Bayes method. Bernoulli, 12(4):609–632, August 2006. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1155735929.

Rivero:2007:RES

[Riv07] Víctor Rivero. Recurrent extensions of self-similar Markov processes and Cramér’s condition II. Bernoulli, 13(4):1053–1070, November 2007. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1194625602.

Rudloff:2010:TCH

[RK10] Birgit Rudloff and Ioannis Karatzas. Testing composite hypotheses via convex duality. Bernoulli, 16(4):1224–1239, November 2010. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1290092904.

Ronn-Nielsen:2023:ECC

[RNS23] Anders Rønn-Nielsen and Mads Stehr. Extremal clustering and cluster counting for spatial random fields. Bernoulli, 29(4):2771–2796, November 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-29/issue-
4/Extremal-clustering-and-cluster-counting-for-spatial-
random-fields/10.3150/22-BEJ1561.full.

Robert:2008:EME

[Rob08] Christian Y. Robert. Estimating the multivariate extremal index function. Bernoulli, 14(4):1027–1064, November 2008. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1225980570.

Robinson:2012:IPL

[Rob12] Peter M. Robinson. Inference on power law spatial trends. Bernoulli, 18(2):644–677, May 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1334580728.

Rollin:2005:ASC

[Röl05] Adrian Röllin. Approximation of sums of conditionally independent variables by the translated Poisson distribution. Bernoulli, 11(6):1115–1128, December 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1137421642.

Roos:1999:ASB

[Roo99] Bero Roos. Asymptotics and sharp bounds in the Poisson approximation to the Poisson-binomial distribution. Bernoulli, 5 (6):1021–1034, December 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1143122301.

Roos:2005:HCP

[Roo05] Bero Roos. On Hipp’s compound Poisson approximations via concentration functions. Bernoulli, 11(3):533–557, June 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1120591188.

Roos:2010:CCP

[Roo10] Bero Roos. Closeness of convolutions of probability measures. Bernoulli, 16(1):23–50, February 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1265984703.

Rosenbaum:2009:IVR

[Ros09] Mathieu Rosenbaum. Integrated volatility and round-off error. Bernoulli, 15(3):687–720, August 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1251463277.

Roysland:2011:MAC

[Røy11] Kjetil Røysland. A martingale approach to continuous-time marginal structural models. Bernoulli, 17(3):895–915, August 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1310042849.

Ramosaj:2023:CUV

[RP23] Burim Ramosaj and Markus Pauly. Consistent and unbiased variable selection under independent features using Random Forest permutation importance. Bernoulli, 29(3):2101–2118, August 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-29/issue-3/Consistent-and-unbiased-
variable-selection-under-indepedent-features-using-
Random/10.3150/22-BEJ1534.full.

Roberts:2001:IHP

[RR01] Gareth O. Roberts and Jeffrey S. Rosenthal. Infinite hierarchies and prior distributions. Bernoulli, 7(3):453–471, June 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1080004760.

Rollin:2015:LLT

[RR15] Adrian Röllin and Nathan Ross. Local limit theorems via Landau–Kolmogorov inequalities. Bernoulli, 21(2):851–880, May 2015. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1429624963.

Rosinski:1996:CMS

[RS96] Jan Rosinski and Gennady Samorodnitsky. Classes of mixing stable processes. Bernoulli, 2(4):365–377, December 1996. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1178291836.

Resnick:2001:SSD

[RS01a] Sidney Resnick and Gennady Samorodnitsky. Steady-state distribution of the buffer content for M/G/∞ input fluid queues. Bernoulli, 7(2):191–210, April 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1080222087.

Runggaldier:2001:SCF

[RS01b] Wolfgang J. Runggaldier and Fabio Spizzichino. Sufficient conditions for finite dimensionality of filters in discrete time: a Laplace transform-based approach. Bernoulli, 7(2):211–221, April 2001. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1080222084.

Roberts:2010:LDM

[RS10] Gareth O. Roberts and Laura M. Sangalli. Latent diffusion models for survival analysis. Bernoulli, 16(2):435–458, May 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1274821078.

Reiss:2017:EEF

[RS17] Markus Reiß and Leonie Selk. Efficient estimation of functionals in nonparametric boundary models. Bernoulli, 23(2):1022– 1055, May 2017. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1486177391.

Rasonyi:2018:SPL

[RS18a] Miklós Rásonyi and Hasanjan Sayit. Sticky processes, local and true martingales. Bernoulli, 24(4A):2752–2775, November 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1522051224.

Rudolf:2018:PTM

[RS18b] Daniel Rudolf and Nikolaus Schweizer. Perturbation theory for Markov chains via Wasserstein distance. Bernoulli, 24(4A):2610–2639, November 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1522051219.

Roy:2022:SZO

[RSBG22] Abhishek Roy, Lingqing Shen, Krishnakumar Balasubramanian, and Saeed Ghadimi. Stochastic zeroth-order discretizations of Langevin diffusions for Bayesian inference. Bernoulli, 28(3):1810–1834, August 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-28/issue-
3/Stochastic-zeroth-order-discretizations-of-Langevin-
diffusions-for-Bayesian-inference/10.3150/21-BEJ1400.
full.

Resnick:2003:WAC

[RSGW03] Sidney Resnick, Gennady Samorodnitsky, Anna Gilbert, and Walter Willinger. Wavelet analysis of conservative cascades. Bernoulli, 9(1):97–135, February 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1068129012.

Roueff:2012:FIE

[RSS12] François Roueff, Gennady Samorodnitsky, and Philippe Soulier. Function-indexed empirical processes based on an infinite source Poisson transmission stream. Bernoulli, 18(3):783–802, August 2012. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1340887002.

Rakhlin:2017:EEM

[RST17] Alexander Rakhlin, Karthik Sridharan, and Alexandre B. Tsybakov. Empirical entropy, minimax regret and minimax risk. Bernoulli, 23(2):789–824, May 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1486177384.

Roberts:1996:ECL

[RT96] Gareth O. Roberts and Richard L. Tweedie. Exponential convergence of Langevin distributions and their discrete approximations. Bernoulli, 2(4):341–363, December 1996. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1178291835.

Rootzen:2006:MGP

[RT06] Holger Rootzén and Nader Tajvidi. Multivariate generalized Pareto distributions. Bernoulli, 12(5):917–930, October 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1161614952.

Riedler:2015:STH

[RT15] Martin G. Riedler and Michèle Thieullen. Spatio-temporal hybrid (PDMP) models: Central limit theorem and Langevin approximation for global fluctuations. Application to electrophysiology. Bernoulli, 21(2):647–696, May 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1429624957.

Richard:2020:PTO

[RTT20] Alexandre Richard, Etienne Tanré, and Soledad Torres. Penalisation techniques for one-dimensional reflected rough differential equations. Bernoulli, 26(4):2949–2986, November 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-26/issue-4/Penalisation-techniques-
for-one-dimensional-reflected-rough-differential-equations/
10.3150/20-BEJ1212.full.

Rukhin:2014:RLR

[Ruk14] Andrew L. Rukhin. Restricted likelihood representation and decision-theoretic aspects of meta-analysis. Bernoulli, 20(4): 1979–1998, November 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1411134450.

Rigollet:2009:ORP

[RV09] Philippe Rigollet and Régis Vert. Optimal rates for plug-in estimators of density level sets. Bernoulli, 15(4):1154–1178, November 2009. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1262962230.

Roueff:2019:TFA

[RvS19] François Roueff and Rainer von Sachs. Time-frequency analysis of locally stationary Hawkes processes. Bernoulli, 25(2): 1355–1385, May 2019. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1551862853.

Russo:2001:GCL

[RVW01] Francesco Russo, Pierre Vallois, and Jochen Wolf. A generalized class of lyons-Zheng processes. Bernoulli, 7(2):363–379, April 2001. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1080222085.

Ruschendorf:2002:ETD

[RW02] Ludger Rüschendorf and Jeannette H. C. Woerner. Expansion of transition distributions of Lévy processes in small time. Bernoulli, 8(1):81–96, February 2002. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1078951090.

Reiss:2019:FEH

[RW19] Markus Reiß and Martin Wahl. Functional estimation and hypothesis testing in nonparametric boundary models. Bernoulli, 25(4A):2597–2619, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1568362037.

Rogers:2022:RKR

[RW22] Dane Rogers and Matthias Winkel. A Ray–Knight representation of up-down Chinese restaurants. Bernoulli, 28(1):689– 712, February 2022. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-28/issue-1/A-RayKnight-representation-
of-up-down-Chinese-restaurants/10.3150/21-BEJ1364.full.

Radulovic:2017:WCE

[RWZ17] Dragan Radulović, Marten Wegkamp, and Yue Zhao. Weak convergence of empirical copula processes indexed by functions. Bernoulli, 23(4B):3346–3384, November 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1495505095.

Resnick:2013:CMC

[RZ13] Sidney I. Resnick and David Zeber. Clustering of Markov chain exceedances. Bernoulli, 19(4):1419–1448, September 2013. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1377612859.

Rockner:2021:WPD

[RZ21] Michael Röckner and Xicheng Zhang. Well-posedness of distribution dependent SDEs with singular drifts. Bernoulli, 27(2): 1131–1158, May 2021. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-27/issue-2/Well-posedness-
of-distribution-dependent-SDEs-with-singular-drifts/
10.3150/20-BEJ1268.full.

Rockner:2023:SCT

[RZ23] Michael Röckner and Guohuan Zhao. SDEs with critical time dependent drifts: Weak solutions. Bernoulli, 29(1):757– 784, February 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-29/issue-1/SDEs-with-
critical-time-dependent-drifts-Weak-solutions/10.3150/
22-BEJ1478.full.

Sadeghi:2013:SMG

[Sad13] Kayvan Sadeghi. Stable mixed graphs. Bernoulli, 19(5B): 2330–2358, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1386078605.

Samorodnitsky:1998:LTS

[Sam98] Gennady Samorodnitsky. Lower tails of self-similar stable processes. Bernoulli, 4(1):127–142, March 1998. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1175865493.

Sanguesa:2010:UEB

[San10] Carmen Sangüesa. Uniform error bounds for a continuous approximation of non-negative random variables. Bernoulli, 16(2):561–584, May 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1274821084.

Sandric:2013:RTP

[San13] Nikola Sandrić. Recurrence and transience property for a class of Markov chains. Bernoulli, 19(5B):2167–2199, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1386078599.

Sancetta:2016:GAP

[San16] Alessio Sancetta. Greedy algorithms for prediction. Bernoulli, 22(2):1227–1277, May 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1447077774.

Sart:2016:REP

[Sar16] Mathieu Sart. Robust estimation on a parametric model via testing. Bernoulli, 22(3):1617–1670, August 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1458132994.

Sart:2023:DEU

[Sar23] Mathieu Sart. Density estimation under local differential privacy and Hellinger loss. Bernoulli, 29(3):2318–2341, August 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-29/issue-3/Density-estimation-under-
local-differential-privacy-and-Hellinger-loss/10.3150/
22-BEJ1543.full.

Saussereau:2012:TIS

[Sau12] Bruno Saussereau. Transportation inequalities for stochastic differential equations driven by a fractional Brownian motion. Bernoulli, 18(1):1–23, February 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1327068615.

Saussereau:2014:NIF

[Sau14] Bruno Saussereau. Nonparametric inference for fractional diffusion. Bernoulli, 20(2):878–918, May 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1393594009.

Saumard:2018:OER

[Sau18] Adrien Saumard. On optimality of empirical risk minimization in linear aggregation. Bernoulli, 24(3):2176–2203, August 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL https://projecteuclid.org/euclid.bj/1517540472.

Saumard:2019:WPI

[Sau19] Adrien Saumard. Weighted Poincaré inequalities, concentration inequalities and tail bounds related to Stein kernels in dimension one. Bernoulli, 25(4B):3978–4006, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL https://projecteuclid.org/euclid.bj/1569398791.

Shi:2024:FAN

[SBP24] Zhaoyang Shi, Krishnakumar Balasubramanian, and Wolfgang Polonik. A flexible approach for normal approximation of geometric and topological statistics. Bernoulli, 30(4):3029–3058, November 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

Steinwart:2011:ECQ

[SC11] Ingo Steinwart and Andreas Christmann. Estimating conditional quantiles with the help of the pinball loss. Bernoulli, 17 (1):211–225, February 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1297173840.

Song:2023:OFD

[SC23] Qifan Song and Guang Cheng. Optimal false discovery control of minimax estimators. Bernoulli, 29(3):1959–1982, August 2023. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-29/issue-3/Optimal-false-
discovery-control-of-minimax-estimators/10.3150/22-
BEJ1527.full.

Schmidli:1999:CSS

[Sch99] Hanspeter Schmidli. Compound sums and subexponentiality. Bernoulli, 5(6):999–1012, December 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1143122299.

Schick:2001:SSM

[Sch01a] Anton Schick. Sample splitting with Markov chains. Bernoulli, 7(1):33–61, February 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1080572338.

Schlather:2001:SOC

[Sch01b] Martin Schlather. On the second-order characteristics of marked point processes. Bernoulli, 7(1):99–117, February 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1080572341.

Schuhmacher:2009:SMP

[Sch09] Dominic Schuhmacher. Stein’s method and Poisson process approximation for a class of Wasserstein metrics. Bernoulli, 15(2):550–568, May 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1241444902.

Schlather:2010:SCM

[Sch10] Martin Schlather. Some covariance models based on normal scale mixtures. Bernoulli, 16(3):780–797, August 2010. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1281099884.

Schnurr:2013:GBG

[Sch13] Alexander Schnurr. Generalization of the Blumenthal–Getoor index to the class of homogeneous diffusions with jumps and some applications. Bernoulli, 19(5A):2010–2032, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1383661212.

Schnurr:2020:FCS

[Sch20a] Alexander Schnurr. The fourth characteristic of a semimartingale. Bernoulli, 26(1):642–663, February 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1574758841.

Schweinberger:2020:CSE

[Sch20b] Michael Schweinberger. Consistent structure estimation of exponential-family random graph models with block structure. Bernoulli, 26(2):1205–1233, May 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1580461577.

Schmidt:2024:DCT

[Sch24] Sara Kristin Schmidt. Detecting changes in the trend function of heteroscedastic time series. Bernoulli, 30(4):2598–2622, November 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

Song:2023:SIL

[SCK23] Yanglei Song, Xiaohui Chen, and Kengo Kato. Stratified incomplete local simplex tests for curvature of nonparametric multiple regression. Bernoulli, 29(1):323–349, February 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-29/issue-1/Stratified-
incomplete-local-simplex-tests-for-curvature-of-nonparametric-
multiple/10.3150/22-BEJ1459.full.

Shi:2024:ACC

[SDH24] Hongjian Shi, Mathias Drton, and Fang Han. On Azadkia– Chatterjee’s conditional dependence coefficient. Bernoulli, 30 (2):851–877, May 2024. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-30/issue-2/On-AzadkiaChatterjees-
conditional-dependence-coefficient/10.3150/22-BEJ1529.
full.

Strothmann:2024:RDM

[SDS24] Christopher Strothmann, Holger Dette, and Karl Friedrich Siburg. Rearranged dependence measures. Bernoulli, 30(2): 1055–1078, May 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-30/issue-2/Rearranged-
dependence-measures/10.3150/23-BEJ1624.full.

Segers:2012:AEC

[Seg12] Johan Segers. Asymptotics of empirical copula processes under non-restrictive smoothness assumptions. Bernoulli, 18(3):764– 782, August 2012. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1340887001.

Seneta:2013:THL

[Sen13] Eugene Seneta. A tricentenary history of the law of large numbers. Bernoulli, 19(4):1088–1121, September 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1377612845.

Severini:2004:LFB

[Sev04] Thomas A. Severini. Likelihood functions based on parameterdependent functions. Bernoulli, 10(3):421–446, June 2004. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1089206405.

Sariev:2023:ICR

[SFP23] Hristo Sariev, Sandra Fortini, and Sonia Petrone. Infinitecolor randomly reinforced urns with dominant colors. Bernoulli, 29(1):132–152, February 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-29/issue-
1/Infinite-color-randomly-reinforced-urns-with-dominant-
colors/10.3150/21-BEJ1452.full.

Shen:2016:ABD

[SG16] Weining Shen and Subhashis Ghosal. Adaptive Bayesian density regression for high-dimensional data. Bernoulli, 22(1): 396–420, February 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1443620855.

Sun:2024:RAR

[SGL^{+}24] Dayu Sun, Yuanyuan Guo, Yang Li, Wanzhu Tu, and Jianguo Sun. A robust approach for regression analysis of panel count data with time-varying covariates. Bernoulli, 30(4):3251–3275, November 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

Sinova:2018:ELF

[SGRV18] Beatriz Sinova, Gil González-Rodríguez, and Stefan Van Aelst. M-estimators of location for functional data. Bernoulli, 24(3):2328–2357, August 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1517540476.

Sidorova:2018:SDG

[Sid18] Nadia Sidorova. Small deviations of a Galton–Watson process with immigration. Bernoulli, 24(4B):3494–3521, November 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1524038761.

Siegrist:2007:EGD

[Sie07] Kyle Siegrist. Exponential and gamma distributions on positive semigroups, with applications to Dirichlet distributions. Bernoulli, 13(2):330–345, May 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1179498751.

Siorpaes:2018:APB

[Sio18] Pietro Siorpaes. Applications of pathwise Burkholder–Davis– Gundy inequalities. Bernoulli, 24(4B):3222–3245, November 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1524038753.

Sharrock:2023:TTS

[SK23] Louis Sharrock and Nikolas Kantas. Two-timescale stochastic gradient descent in continuous time with applications to joint online parameter estimation and optimal sensor placement. Bernoulli, 29(2):1137–1165, May 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-2/Two-timescale-stochastic-gradient-descent-
in-continuous-time-with-applications/10.3150/22-BEJ1493.
full.

Skovgaard:1996:ELD

[Sko96] Ib M. Skovgaard. An explicit large-deviation approximation to one-parameter tests. Bernoulli, 2(2):145–165, June 1996. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1193839221.

Sadeghi:2014:MPM

[SL14] Kayvan Sadeghi and Steffen Lauritzen. Markov properties for mixed graphs. Bernoulli, 20(2):676–696, May 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1393594002.

Steinberger:2018:CMH

[SL18] Lukas Steinberger and Hannes Leeb. On conditional moments of high-dimensional random vectors given lower-dimensional projections. Bernoulli, 24(1):565–591, February 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1501142455.

Syring:2023:GPC

[SM23] Nicholas Syring and Ryan Martin. Gibbs posterior concentration rates under sub-exponential type losses. Bernoulli, 29(2): 1080–1108, May 2023. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-29/issue-2/Gibbs-posterior-
concentration-rates-under-sub-exponential-type-losses/
10.3150/22-BEJ1491.full.

Sun:2013:MSL

[SN13] Xing Sun and Andrew B. Nobel. On the maximal size of largeaverage and ANOVA-fit submatrices in a Gaussian random matrix. Bernoulli, 19(1):275–294, February 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1358531750.

Sorensen:2002:EDP

[Sør02] Helle Sørensen. Estimation of diffusion parameters for discretely observed diffusion processes. Bernoulli, 8(4):491–508, August 2002. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1078681381.

Sundararajan:2021:SSA

[SPP21] Raanju R. Sundararajan, Vladas Pipiras, and Mohsen Pourahmadi. Stationary subspace analysis of nonstationary covariance processes: Eigenstructure description and testing. Bernoulli, 27(1):381–418, February 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-1/Stationary-subspace-analysis-of-nonstationary-
covariance-processes--Eigenstructure-description/10.
3150/20-BEJ1243.full.

Sherlock:2009:OSR

[SR09] Chris Sherlock and Gareth Roberts. Optimal scaling of the random walk Metropolis on elliptically symmetric unimodal targets. Bernoulli, 15(3):774–798, August 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1251463281.

Sriperumbudur:2016:OEP

[Sri16] Bharath Sriperumbudur. On the optimal estimation of probability measures in weak and strong topologies. Bernoulli, 22 (3):1839–1893, August 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1458133001.

Sulem:2024:BEN

[SRR24] Déborah Sulem, Vincent Rivoirard, and Judith Rousseau. Bayesian estimation of nonlinear Hawkes processes. Bernoulli, 30(2):1257–1286, May 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-30/issue-
2/Bayesian-estimation-of-nonlinear-Hawkes-processes/
10.3150/23-BEJ1631.full.

Schlemm:2012:MCP

[SS12] Eckhard Schlemm and Robert Stelzer. Multivariate CARMA processes, continuous-time state space models and complete regularity of the innovations of the sampled processes. Bernoulli, 18(1):46–63, February 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1327068617.

Strokorb:2015:EMS

[SS15] Kirstin Strokorb and Martin Schlather. An exceptional maxstable process fully parameterized by its extremal coefficients. Bernoulli, 21(1):276–302, February 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1426597070.

Schachermayer:2018:SCB

[SS18] Walter Schachermayer and Florian Stebegg. The sharp constant for the Burkholder–Davis–Gundy inequality and non-smooth pasting. Bernoulli, 24(4A):2499–2530, November 2018. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1522051216.

Sambale:2020:LSI

[SS20] Holger Sambale and Arthur Sinulis. Logarithmic Sobolev inequalities for finite spin systems and applications. Bernoulli, 26(3):1863–1890, August 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1587974526.

Sabel:2014:AEE

[SSH14] Till Sabel and Johannes Schmidt-Hieber. Asymptotically efficient estimation of a scale parameter in Gaussian time series and closed-form expressions for the Fisher information. Bernoulli, 20(2):747–774, May 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1393594005.

Santacroce:2016:NRM

[SST16] Marina Santacroce, Paola Siri, and Barbara Trivellato. New results on mixture and exponential models by Orlicz spaces. Bernoulli, 22(3):1431–1447, August 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1458132987.

Schilling:2012:CPG

[SSW12] René L. Schilling, Paweł Sztonyk, and Jian Wang. Coupling property and gradient estimates of Lévy processes via the symbol. Bernoulli, 18(4):1128–1149, November 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1352727804.

Song:2020:FSW

[SSX20] Jian Song, Xiaoming Song, and Fangjun Xu. Fractional stochastic wave equation driven by a Gaussian noise rough in space. Bernoulli, 26(4):2699–2726, November 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
26/issue-4/Fractional-stochastic-wave-equation-driven-
by-a-Gaussian-noise-rough/10.3150/20-BEJ1204.full.

Sun:2021:IPB

[SSY21] Rongfeng Sun, Jan M. Swart, and Jinjiong Yu. An invariance principle for biased voter model interfaces. Bernoulli, 27(1): 615–636, February 2021. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-27/issue-1/An-invariance-
principle-for-biased-voter-model-interfaces/10.3150/
20-BEJ1252.full.

Samarov:2004:NIC

[ST04] Alexander Samarov and Alexandre Tsybakov. Nonparametric independent component analysis. Bernoulli, 10(4):565–582, August 2004. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1093265630.

Schreiber:2013:GIS

[ST13] Tomasz Schreiber and Christoph Thäle. Geometry of iteration stable tessellations: Connection with Poisson hyperplanes. Bernoulli, 19(5A):1637–1654, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1383661197.

Shi:2018:DGH

[ST18] Chenlu Shi and Boxin Tang. Designs from good Hadamard matrices. Bernoulli, 24(1):661–671, February 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1501142458.

Shi:2020:CRS

[ST20] Chenlu Shi and Boxin Tang. Construction results for strong orthogonal arrays of strength three. Bernoulli, 26(1):418– 431, February 2020. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1574758833.

Stephanovitch:2024:OWD

[STC^{+}24] Arthur Stéphanovitch, Ugo Tanielian, Benoît Cadre, Nicolas Klutchnikoff, and Gérard Biau. Optimal 1-Wasserstein distance for WGANs. Bernoulli, 30(4):2955–2978, November 2024. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

Stelzer:2010:MP

[Ste10] Robert Stelzer. Multivariate COGARCH(1, 1) processes. Bernoulli, 16(1):80–115, February 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1265984705.

Stein:2013:CST

[Ste13] Michael L. Stein. On a class of space–time intrinsic random functions. Bernoulli, 19(2):387–408, May 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1363192032.

Stoyanov:2000:KCP

[Sto00] Jordan Stoyanov. Krein condition in probabilistic moment problems. Bernoulli, 6(5):939–949, October 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1081282696.

Stufler:2023:UIC

[Stu23] Benedikt Stufler. The uniform infinite cubic planar graph. Bernoulli, 29(4):2902–2926, November 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-29/issue-
4/The-uniform-infinite-cubic-planar-graph/10.3150/22-
BEJ1568.full.

Sorensen:2003:SDA

[SU03] Michael Sørensen and Masayuki Uchida. Small-diffusion asymptotics for discretely sampled stochastic differential equations. Bernoulli, 9(6):1051–1069, December 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1072215200.

Sole:2008:TSH

[SU08] Josep Lluís Solé and Frederic Utzet. Time–space harmonic polynomials relative to a Lévy process. Bernoulli, 14(1):1–13, February 2008. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1202492782.

Surgailis:2004:SLS

[Sur04] Donatas Surgailis. Stable limits of sums of bounded functions of long-memory moving averages with finite variance. Bernoulli, 10(2):327–355, April 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1082380222.

Stanghellini:2013:IDC

[SV13] Elena Stanghellini and Barbara Vantaggi. Identification of discrete concentration graph models with one hidden binary variable. Bernoulli, 19(5A):1920–1937, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1383661208.

Schauer:2017:GPS

[SvdMvZ17] Moritz Schauer, Frank van der Meulen, and Harry van Zanten. Guided proposals for simulating multi-dimensional diffusion bridges. Bernoulli, 23(4A):2917–2950, November 2017.

CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL https://projecteuclid.org/euclid.bj/1494316837.

Steland:2017:LSA

[SvS17] Ansgar Steland and Rainer von Sachs. Large-sample approximations for variance-covariance matrices of high-dimensional time series. Bernoulli, 23(4A):2299–2329, November 2017. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1494316819.

Schick:2004:FCO

[SW04] Anton Schick and Wolfgang Wefelmeyer. Functional convergence and optimality of plug-in estimators for stationary densities of moving average processes. Bernoulli, 10(5):889–917, October 2004. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1099579161.

Shimura:2005:IDG

[SW05] Takaaki Shimura and Toshiro Watanabe. Infinite divisibility and generalized subexponentiality. Bernoulli, 11(3):445–469, June 2005. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1120591184.

Stute:2008:CLT

[SW08] Winfried Stute and Jane-Ling Wang. The central limit theorem under random truncation. Bernoulli, 14(3):604–622, August 2008. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1219669622.

Sen:2012:BCI

[SW12] Bodhisattva Sen and Michael Woodroofe. Bootstrap confidence intervals for isotonic estimators in a stereological problem. Bernoulli, 18(4):1249–1266, November 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1352727809.

Schick:2013:UCC

[SW13] Anton Schick and Wolfgang Wefelmeyer. Uniform convergence of convolution estimators for the response density in nonparametric regression. Bernoulli, 19(5B):2250–2276, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1386078602.

Saumard:2019:ICC

[SW19a] Adrien Saumard and Jon A. Wellner. On the isoperimetric constant, covariance inequalities and L_{p}–Poincaré inequalities in dimension one. Bernoulli, 25(3):1794–1815, August 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1560326428.

Shiozawa:2019:LTH

[SW19b] Yuichi Shiozawa and Jian Wang. Long-time heat kernel estimates and upper rate functions of Brownian motion type for symmetric jump processes. Bernoulli, 25(4B):3796–3831, November 2019. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1569398785.

Shen:2020:OSG

[SW20] Yi Shen and Yizao Wang. Operator-scaling Gaussian random fields via aggregation. Bernoulli, 26(1):500–530, February 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL https://projecteuclid.org/euclid.bj/1574758836.

Shieh:2010:HPD

[SX10] Narn-Rueih Shieh and Yimin Xiao. Hausdorff and packing dimensions of the images of random fields. Bernoulli, 16(4): 926–952, November 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1290092890.

Sun:2022:SWC

[SXX22] Xiaobin Sun, Longjie Xie, and Yingchao Xie. Strong and weak convergence rates for slow-fast stochastic differential equations driven by α-stable process. Bernoulli, 28(1):343– 369, February 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-28/issue-1/Strong-and-
weak-convergence-rates-for-slowfast-stochastic-differential-
equations/10.3150/21-BEJ1345.full.

Sun:2018:UDR

[SXXZ18] Xiaobin Sun, Yimin Xiao, Lihu Xu, and Jianliang Zhai. Uniform dimension results for a family of Markov processes. Bernoulli, 24(4B):3924–3951, November 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1524038773.

Siegmund:2000:TPN

[SY00] David Siegmund and Benjamin Yakir. Tail probabilities for the null distribution of scanning statistics. Bernoulli, 6(2):191– 213, April 2000. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1081788026.

Sei:2024:MID

[SY24] Tomonari Sei and Keisuke Yano. Minimum information dependence modeling. Bernoulli, 30(4):2623–2643, November 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

Synowiecki:2007:CAM

[Syn07] Rafal Synowiecki. Consistency and application of moving block bootstrap for non-stationary time series with periodic and almost periodic structure. Bernoulli, 13(4):1151–1178, November 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1194625606.

Sahu:2003:SRM

[SZ03] Sujit K. Sahu and Anatoly A. Zhigljavsky. Self-regenerative Markov chain Monte Carlo with adaptation. Bernoulli, 9 (3):395–422, June 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1065444811.

Shao:2016:CTM

[SZ16] Qi-Man Shao and Wen-Xin Zhou. Cramér type moderate deviation theorems for self-normalized processes. Bernoulli, 22 (4):2029–2079, November 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1462297674.

Shao:2022:BEB

[SZ22] Qi-Man Shao and Zhuo-Song Zhang. Berry–Esseen bounds for multivariate nonlinear statistics with applications to mestimators and stochastic gradient descent algorithms. Bernoulli, 28(3):1548–1576, August 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-28/issue-
3/BerryEsseen-bounds-for-multivariate-nonlinear-statistics-
with-applications-to-M/10.3150/21-BEJ1336.full.

Takemura:2005:DRM

[TA05] Akimichi Takemura and Satoshi Aoki. Distance-reducing Markov bases for sampling from a discrete sample space. Bernoulli, 11(5):793–813, October 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1130077594.

Talata:2013:DRM

[Tal13] Zsolt Talata. Divergence rates of Markov order estimators and their application to statistical estimation of stationary ergodic processes. Bernoulli, 19(3):846–885, August 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1372251145.

Tang:2006:CEA

[Tan06] Qihe Tang. On convolution equivalence with applications. Bernoulli, 12(3):535–549, June 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1151525135.

Tan:2015:IME

[Tan15] Zhiqiang Tan. Improved minimax estimation of a multivariate normal mean under heteroscedasticity. Bernoulli, 21(1): 574–603, February 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1426597082.

Tang:2017:NTS

[TAS^{+}17] Minh Tang, Avanti Athreya, Daniel L. Sussman, Vince Lyzinski, and Carey E. Priebe. A nonparametric two-sample hypothesis testing problem for random graphs. Bernoulli, 23(3):1599– 1630, August 2017. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1489737619.

Taylor:2013:GLS

[Tay13] Jonathan Taylor. The geometry of least squares in the 21st century. Bernoulli, 19(4):1449–1464, September 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1377612860.

Tsai:2003:NPD

[TC03] Henghsiu Tsai and K. S. Chan. A note on parameter differentiation of matrix exponentials, with applications to continuoustime modelling. Bernoulli, 9(5):895–919, October 2003. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1066418883.

To:2024:MST

[TC24] Hon Kiu To and Kin Wai Chan. Mean stationarity test in time series: a signal variance-based approach. Bernoulli, 30(2): 1231–1256, May 2024. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-30/issue-2/Mean-stationarity-
test-in-time-series--A-signal-variance/10.3150/23-BEJ1630.
full.

Tang:2022:AEE

[TCP22] Minh Tang, Joshua Cape, and Carey E. Priebe. Asymptotically efficient estimators for stochastic blockmodels: The naive MLE, the rank-constrained MLE, and the spectral estimator. Bernoulli, 28(2):1049–1073, May 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-2/Asymptotically-efficient-estimators-for-
stochastic-blockmodels--The-naive-MLE/10.3150/21-BEJ1376.
full.

Torrisi:2023:BPS

[TGL23] Giovanni Luca Torrisi, Michele Garetto, and Emilio Leonardi. Bootstrap percolation on the stochastic block model. Bernoulli, 29(1):696–724, February 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-29/issue-
1/Bootstrap-percolation-on-the-stochastic-block-model/
10.3150/22-BEJ1475.full.

Thorisson:1999:PSD

[Tho99] Hermann Thorisson. Point stationarity in d dimensions and Palm theory. Bernoulli, 5(5):797–831, October 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1171290400.

Thomas:2023:CLT

[Tho23] Andrew M. Thomas. Central limit theorems and asymptotic independence for local U-statistics on diverging halfspaces. Bernoulli, 29(4):3280–3306, November 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-4/Central-limit-theorems-and-asymptotic-independence-
for-local-U-statistics/10.3150/23-BEJ1583.full.

Tindel:1998:QSE

[Tin98] Samy Tindel. Quasilinear stochastic elliptic equations with reflection: the existence of a density. Bernoulli, 4(4):445–459, December 1998. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1173883815.

Takemura:2003:TPT

[TK03] Akimichi Takemura and Satoshi Kuriki. Tail probability via tube formula when the critical radius is zero. Bernoulli, 9 (3):535–558, June 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1065444817.

Tkachov:2020:STW

[Tka20] Pasha Tkachov. On stability of traveling wave solutions for integro-differential equations related to branching Markov processes. Bernoulli, 26(2):1354–1380, May 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1580461582.

Takeuchi:2009:NFA

[TKT09] Kei Takeuchi, Masayuki Kumon, and Akimichi Takemura. A new formulation of asset trading games in continuous time with essential forcing of variation exponent. Bernoulli, 15(4): 1243–1258, November 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1262962234.

Tong:2013:OVE

[TMW13] Tiejun Tong, Yanyuan Ma, and Yuedong Wang. Optimal variance estimation without estimating the mean function. Bernoulli, 19(5A):1839–1854, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1383661205.

Todino:2020:NLS

[Tod20a] Anna Paola Todino. Nodal lengths in shrinking domains for random eigenfunctions on. Bernoulli, 26(4):3081–3110, November 2020. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-26/issue-4/Nodal-lengths-
in-shrinking-domains-for-random-eigenfunctions-on-S2/
10.3150/20-BEJ1216.full.

Todorov:2020:TIF

[Tod20b] Viktor Todorov. Testing and inference for fixed times of discontinuity in semimartingales. Bernoulli, 26(4):2907–2948, November 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-26/issue-4/Testing-and-inference-for-
fixed-times-of-discontinuity-in-semimartingales/10.
3150/20-BEJ1211.full.

Torrisi:2017:PAP

[Tor17] Giovanni Luca Torrisi. Probability approximation of point processes with Papangelou conditional intensity. Bernoulli, 23(4A):2210–2256, November 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1494316817.

Trabs:2014:CSD

[Tra14] Mathias Trabs. Calibration of self-decomposable Lévy models. Bernoulli, 20(1):109–140, February 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1390407282.

Tan:2024:NCE

[TRB24] Kai Tan, Gabriel Romon, and Pierre C. Bellec. Noise covariance estimation in multi-task high-dimensional linear models. Bernoulli, 30(3):1695–1722, August 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
30/issue-3/Noise-covariance-estimation-in-multi-task-
high-dimensional-linear-models/10.3150/23-BEJ1644.full.

Truquet:2019:RCE

[Tru19] Lionel Truquet. Root-n consistent estimation of the marginal density in semiparametric autoregressive time series models. Bernoulli, 25(3):2107–2136, August 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1560326439.

Truquet:2020:CPT

[Tru20a] Lionel Truquet. Coupling and perturbation techniques for categorical time series. Bernoulli, 26(4):3249–3279, November 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-26/issue-4/Coupling-and-perturbation-
techniques-for-categorical-time-series/10.3150/20-BEJ1225.
full.

Truquet:2020:PAM

[Tru20b] Lionel Truquet. A perturbation analysis of Markov chains models with time-varying parameters. Bernoulli, 26(4):2876– 2906, November 2020. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-26/issue-4/A-perturbation-
analysis-of-Markov-chains-models-with-time-varying/
10.3150/20-BEJ1210.full.

Tsai:2009:CTA

[Tsa09] Henghsiu Tsai. On continuous-time autoregressive fractionally integrated moving average processes. Bernoulli, 15(1):178–194, February 2009. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1233669887.

Tanaka:2006:SCM

[TT06] Kentaro Tanaka and Akimichi Takemura. Strong consistency of the maximum likelihood estimator for finite mixtures of location-scale distributions when the scale parameters are exponentially small. Bernoulli, 12(6):1003–1017, December 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1165269148.

Talay:2020:NMV

[TT20a] Denis Talay and Milica Tomasević. A new McKean–Vlasov stochastic interpretation of the parabolic-parabolic Keller–Segel model: The one-dimensional case. Bernoulli, 26(2):1323–1353, May 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1580461581.

Thalmaier:2020:EIE

[TT20b] Anton Thalmaier and James Thompson. Exponential integrability and exit times of diffusions on sub-Riemannian and metric measure spaces. Bernoulli, 26(3):2202–2225, August 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL https://projecteuclid.org/euclid.bj/1587974538.

Tudor:2004:MTS

[Tud04] Ciprian A. Tudor. Martingale-type stochastic calculus for anticipating integral processes. Bernoulli, 10(2):313–325, April 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1082380221.

Turova:2004:BEC

[Tur04] Tatyana S. Turova. Berry–Esseen and central limit theorems for serial rank statistics via graphs. Bernoulli, 10(2):221–250, April 2004. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1082380218.

Troutman:1999:ERE

[TV99] Brent M. Troutman and Aldo V. Vecchia. Estimation of Rényi exponents in random cascades. Bernoulli, 5(2):191–207, April 1999. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1173147902.

Tarigan:2006:CSV

[TV06] Bernadetta Tarigan and Sara A. Van De Geer. Classifiers of support vector machine type with ℓ_{1} complexity regularization. Bernoulli, 12(6):1045–1076, December 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1165269150.

Trashorras:2014:LDB

[TW14] José Trashorras and Olivier Wintenberger. Large deviations for bootstrapped empirical measures. Bernoulli, 20(4):1845–1878, November 2014. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1411134447.

Tudor:2007:SPP

[TX07] Ciprian A. Tudor and Yimin Xiao. Sample path properties of bifractional Brownian motion. Bernoulli, 13(4):1023–1052, November 2007. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1194625601.

Thale:2016:ATS

[TY16] Christoph Thäle and J. E. Yukich. Asymptotic theory for statistics of the Poisson–Voronoi approximation. Bernoulli, 22(4): 2372–2400, November 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1462297684.

Tschiderer:2023:TAR

[TY23] Bertram Tschiderer and Lane Chun Yeung. A trajectorial approach to relative entropy dissipation of McKean–Vlasov diffusions: Gradient flows and HWBI inequalities. Bernoulli, 29(1): 725–756, February 2023. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-29/issue-1/A-trajectorial-
approach-to-relative-entropy-dissipation-of-McKeanVlasov-
diffusions/10.3150/22-BEJ1476.full.

Thulin:2017:ECI

[TZ17] Måns Thulin and Silvelyn Zwanzig. Exact confidence intervals and hypothesis tests for parameters of discrete distributions. Bernoulli, 23(1):479–502, February 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1475001362.

Upadhye:2017:SOD

[UCV17] Neelesh S. Upadhye, Vydas Cekanavicius, and P. Vellaisamy. On Stein operators for discrete approximations. Bernoulli, 23(4A):2828–2859, November 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1494316834.

Uru:2023:CPD

[UDS23] Cagin Uru, Savas Dayanik, and Semih O. Sezer. Compound Poisson disorder problem with uniformly distributed disorder time. Bernoulli, 29(3):2272–2294, August 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-3/Compound-Poisson-disorder-problem-with-uniformly-
distributed-disorder-time/10.3150/22-BEJ1541.full.

Ueda:2021:MCS

[Ued21] Yuki Ueda. Max-convolution semigroups and extreme values in limit theorems for the free multiplicative convolution. Bernoulli, 27(1):502–531, February 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-1/Max-convolution-semigroups-and-extreme-values-
in-limit-theorems-for/10.3150/20-BEJ1247.full.

Utpala:2024:SEH

[US24] Saiteja Utpala and Bharath K. Sriperumbudur. Shrinkage estimation of higher-order Bochner integrals. Bernoulli, 30(4): 2721–2746, November 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

Vergara:2022:GFS

[VAD22] Ricardo Carrizo Vergara, Denis Allard, and Nicolas Desassis. A general framework for SPDE-based stationary random fields. Bernoulli, 28(1):1–32, February 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-1/A-general-framework-for-SPDE-based-stationary-
random-fields/10.3150/20-BEJ1317.full.

VanDerLaan:1995:INM

[Van95] Mark J. Van Der Laan. An identity for the nonparametric maximum likelihood estimator in missing data and biased sampling models. Bernoulli, 1(4):335–341, December 1995. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1193758710.

VanZwet:2004:LLS

[Van04] Erik W. Van Zwet. Laslett’s line segment problem. Bernoulli, 10(3):377–396, June 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1089206403.

VanZanten:2005:RCM

[Van05a] Harry Van Zanten. On the rate of convergence of the maximum likelihood estimator in Brownian semimartingale models. Bernoulli, 11(4):643–664, August 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1126126763.

Vandekerkhove:2005:CAN

[Van05b] Pierre Vandekerkhove. Consistent and asymptotically normal parameter estimates for hidden Markov mixtures of Markov models. Bernoulli, 11(1):103–129, January 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1110228244.

Varvenne:2019:RCE

[Var19] Maylis Varvenne. Rate of convergence to equilibrium for discrete-time stochastic dynamics with memory. Bernoulli, 25(4B):3234–3275, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1569398765.

VanHala:2020:GFD

[VBLN20] Matthew Van Hala, Soutir Bandyopadhyay, Soumendra N. Lahiri, and Daniel J. Nordman. A general frequency domain method for assessing spatial covariance structures. Bernoulli, 26(4):2463–2487, November 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-26/issue-
4/A-general-frequency-domain-method-for-assessing-spatial-
covariance-structures/10.3150/19-BEJ1160.full.

Vetter:2012:MCV

[VD12] Mathias Vetter and Holger Dette. Model checks for the volatility under microstructure noise. Bernoulli, 18(4):1421–1447, November 2012. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1352727818.

Volgushev:2013:NQR

[VD13] Stanislav Volgushev and Holger Dette. Nonparametric quantile regression for twice censored data. Bernoulli, 19(3):748–779, August 2013. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1372251142.

vanDelft:2021:SMS

[vDD21] Anne van Delft and Holger Dette. A similarity measure for second order properties of non-stationary functional time series with applications to clustering and testing. Bernoulli, 27(1):469–501, February 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-1/A-similarity-measure-for-second-order-properties-
of-non-stationary/10.3150/20-BEJ1246.full.

vanDelft:2022:PTR

[vDD22] Anne van Delft and Holger Dette. Pivotal tests for relevant differences in the second order dynamics of functional time series. Bernoulli, 28(4):2260–2293, November 2022. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-28/issue-4/Pivotal-tests-
for-relevant-differences-in-the-second-order-dynamics/
10.3150/21-BEJ1418.full.

vanderMeulen:2013:CNB

[vdMvZ13] Frank van der Meulen and Harry van Zanten. Consistent nonparametric Bayesian inference for discretely observed scalar diffusions. Bernoulli, 19(1):44–63, February 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1358531740.

vanEs:2007:KTN

[vEGS07] Bert van Es, Shota Gugushvili, and Peter Spreij. A kernel type nonparametric density estimator for decompounding. Bernoulli, 13(3):672–694, August 2007. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1186503482.

Vershynin:2020:CIR

[Ver20] Roman Vershynin. Concentration inequalities for random tensors. Bernoulli, 26(4):3139–3162, November 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
26/issue-4/Concentration-inequalities-for-random-tensors/
10.3150/20-BEJ1218.full.

Vetter:2015:EIV

[Vet15] Mathias Vetter. Estimation of integrated volatility of volatility with applications to goodness-of-fit testing. Bernoulli, 21(4): 2393–2418, November 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1438777598.

Vats:2018:SCM

[VFJ18] Dootika Vats, James M. Flegal, and Galin L. Jones. Strong consistency of multivariate spectral variance estimators in Markov chain Monte Carlo. Bernoulli, 24(3):1860–1909, August 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL https://projecteuclid.org/euclid.bj/1517540462.

Verzelen:2018:AEH

[VG18] Nicolas Verzelen and Elisabeth Gassiat. Adaptive estimation of high-dimensional signal-to-noise ratios. Bernoulli, 24 (4B):3683–3710, November 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1524038767.

Viitasaari:2016:IRR

[Vii16] Lauri Viitasaari. Integral representation of random variables with respect to Gaussian processes. Bernoulli, 22(1):376–395, February 2016. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1443620854.

Villa:2014:LAM

[Vil14] Elena Villa. On the local approximation of mean densities of random closed sets. Bernoulli, 20(1):1–27, February 2014. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1390407279.

Villa:2023:MPD

[Vil23] Elena Villa. On the mean perimeter density of inhomogeneous random closed sets. Bernoulli, 29(4):2719–2744, November 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-29/issue-4/On-the-mean-perimeter-density-
of-inhomogeneous-random-closed-sets/10.3150/22-BEJ1558.
full.

VanDenHeuvel:1997:SEN

[VK97] Edwin R. Van Den Heuvel and Chris A. J. Klaassen. Spread, estimators and nuisance parameters. Bernoulli, 3(3):323–328, September 1997. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1177334458.

Vidal-Sanz:2005:PUC

[VS05] Jose M. Vidal-Sanz. Pointwise universal consistency of nonparametric density estimators. Bernoulli, 11(6):971–985, December 2005. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1137421636.

VanEs:2003:NVD

[VSV03] Bert Van Es, Peter Spreij, and Harry Van Zanten. Nonparametric volatility density estimation. Bernoulli, 9(3):451–465, June 2003. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1065444813.

Veillette:2013:PNE

[VT13] Mark S. Veillette and Murad S. Taqqu. Properties and numerical evaluation of the Rosenblatt distribution. Bernoulli, 19(3):982–1005, August 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1372251150.

VanKeilegom:2002:DHE

[VV02] Ingrid Van Keilegom and Noël Veraverbeke. Density and hazard estimation in censored regression models. Bernoulli, 8(5):607– 625, October 2002. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1078435220.

VanDeGeer:2004:HDD

[VV04] Sara A. Van De Geer and Hans C. Van Houwelingen. Highdimensional data: p ≫ n in mathematical statistics and biomedical applications. Bernoulli, 10(6):939–943, December 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1106314843.

VanDerMeulen:2006:CRP

[VVV06] F. H. Van Der Meulen, Aad W. Van Der Vaart, and J. H. Van Zanten. Convergence rates of posterior distributions for Brownian semimartingale models. Bernoulli, 12(5):863–888, October 2006. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1161614950.

Vos:2015:MLE

[VW15] Paul Vos and Qiang Wu. Maximum likelihood estimators uniformly minimize distribution variance among distribution unbiased estimators in exponential families. Bernoulli, 21(4): 2120–2138, November 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1438777588.

Vogel:2017:SQP

[VW17] Daniel Vogel and Martin Wendler. Studentized U-quantile processes under dependence with applications to change-point analysis. Bernoulli, 23(4B):3114–3144, November 2017. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1495505086.

Vakeroudis:2013:IPL

[VY13] Stavros Vakeroudis and Marc Yor. Integrability properties and limit theorems for the exit time from a cone of planar Brownian motion. Bernoulli, 19(5A):2000–2009, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1383661211.

Vaskevicius:2023:SCL

[VZ23] Tomas Vaskevicius and Nikita Zhivotovskiy. Suboptimality of constrained least squares and improvements via non-linear predictors. Bernoulli, 29(1):473–495, February 2023. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-29/issue-1/Suboptimality-
of-constrained-least-squares-and-improvements-via-non-
linear/10.3150/22-BEJ1465.full.

Wakolbinger:1995:LSB

[Wak95] Anton Wakolbinger. Limits of spatial branching populations (with discussion). Bernoulli, 1(1–2):171–189, March 1995. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1186078366.

Walsh:2012:EIC

[Wal12] Alexander Walsh. Extended Itô calculus for symmetric Markov processes. Bernoulli, 18(4):1150–1171, November 2012. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1352727805.

Walsh:2013:SIR

[Wal13] Alexander Walsh. Stochastic integration with respect to additive functionals of zero quadratic variation. Bernoulli, 19(5B): 2414–2436, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1386078608.

Wang:1997:GZO

[Wan97] Hong Wang. Generalized zero-one laws for large-order statistics. Bernoulli, 3(4):429–444, December 1997. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1175882217.

Wang:2011:COU

[Wan11] Feng-Yu Wang. Coupling for Ornstein–Uhlenbeck processes with jumps. Bernoulli, 17(4):1136–1158, November 2011. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1320417499.

Wang:2016:WDS

[Wan16] Jian Wang. L^{p}-Wasserstein distance for stochastic differential equations driven by Lévy processes. Bernoulli, 22(3): 1598–1616, August 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1458132993.

Wang:2017:VCE

[Wan17] Yinghui Wang. Viscosity characterization of the explosion time distribution for diffusions. Bernoulli, 23(1):503–521, February 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1475001363.

Wang:2022:ECA

[Wan22a] Guanyang Wang. Exact convergence analysis of the independent Metropolis–Hastings algorithms. Bernoulli, 28(3):2012– 2033, August 2022. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-28/issue-3/Exact-convergence-
analysis-of-the-independent-Metropolis--Hastings-algorithms/
10.3150/21-BEJ1409.full.

Wang:2022:TPE

[Wan22b] Guanyang Wang. On the theoretical properties of the exchange algorithm. Bernoulli, 28(3):1935–1960, August 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-3/On-the-theoretical-properties-of-the-exchange-
algorithm/10.3150/21-BEJ1405.full.

Wang:2023:EEN

[Wan23a] Feng-Yu Wang. Exponential ergodicity for non-dissipative McKean–Vlasov SDEs. Bernoulli, 29(2):1035–1062, May 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-29/issue-2/Exponential-ergodicity-
for-non-dissipative-McKean--Vlasov-SDEs/10.3150/22-
BEJ1489.full.

Wang:2023:ESE

[Wan23b] Tao Wang. Exponential and strong ergodicity for onedimensional time-changed symmetric stable processes. Bernoulli, 29(1):580–596, February 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-29/issue-
1/Exponential-and-strong-ergodicity-for-one-dimensional-
time-changed-symmetric/10.3150/22-BEJ1469.full.

Wang:2024:TVB

[Wan24] Ruodu Wang. Testing with p∗-values: Between p-values, mid p-values, and e-values. Bernoulli, 30(2):1313–1346, May 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-30/issue-2/Testing-with-p-values--
Between-p-values-mid-p/10.3150/23-BEJ1633.full.

Wang:2017:SAS

[WAP17] Lili Wang, Alexander Aue, and Debashis Paul. Spectral analysis of sample autocovariance matrices of a class of linear time series in moderately high dimensions. Bernoulli, 23(4A):2181– 2209, November 2017. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1494316816.

Weed:2019:SAF

[WB19] Jonathan Weed and Francis Bach. Sharp asymptotic and finitesample rates of convergence of empirical measures in Wasserstein distance. Bernoulli, 25(4A):2620–2648, November 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL https://projecteuclid.org/euclid.bj/1568362038.

Wermuth:1998:AMD

[WC98] Nanny Wermuth and David R. Cox. On association models defined over independence graphs. Bernoulli, 4(4):477–495, December 1998. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1173883817.

Wang:2014:UCR

[WC14] Qiying Wang and Nigel Chan. Uniform convergence rates for a class of martingales with application in non-linear cointegrating regression. Bernoulli, 20(1):207–230, February 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1390407286.

Wang:2018:EEG

[WC18] Li Wang and Guanqun Cao. Efficient estimation for generalized partially linear single-index models. Bernoulli, 24(2): 1101–1127, May 2018. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1505980891.

Wermuth:2006:CC

[WCM06] Nanny Wermuth, D. R. Cox, and Giovanni M. Marchetti. Covariance chains. Bernoulli, 12(5):841–862, October 2006. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1161614949.

Wellner:1996:D

[Wel96] Jon A. Wellner. Discussion. Bernoulli, 2(1):29–38, March 1996. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1193758787.

Werner:1995:BDE

[Wer95] Wendelin Werner. On Brownian disconnection exponents. Bernoulli, 1(4):371–380, December 1995. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1193758712.

Wermuth:2011:PDS

[Wer11] Nanny Wermuth. Probability distributions with summary graph structure. Bernoulli, 17(3):845–879, August 2011. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1310042847.

Wu:2010:VCF

[WFM10] Yichao Wu, Jianqing Fan, and Hans-Georg Müller. Varyingcoefficient functional linear regression. Bernoulli, 16(3):730– 758, August 2010. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1281099882.

Wei:2010:CGS

[WH10] Fengrong Wei and Jian Huang. Consistent group selection in high-dimensional linear regression. Bernoulli, 16(4):1369–1384, November 2010. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1290092910.

Wang:2018:LLS

[WHP18] Xuejun Wang, Xiao Han, and Guangming Pan. The logarithmic law of sample covariance matrices near singularity. Bernoulli, 24(1):80–114, February 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1501142437.

Wiesel:2022:MAW

[Wie22] Johannes C. W. Wiesel. Measuring association with Wasserstein distances. Bernoulli, 28(4):2816–2832, November 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-28/issue-4/Measuring-association-with-
Wasserstein-distances/10.3150/21-BEJ1438.full.

Wolfer:2021:SEE

[WK21] Geoffrey Wolfer and Aryeh Kontorovich. Statistical estimation of ergodic Markov chain kernel over discrete state space. Bernoulli, 27(1):532–553, February 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-1/Statistical-estimation-of-ergodic-Markov-
chain-kernel-over-discrete-state/10.3150/20-BEJ1248.
full.

Whiteley:2016:RIS

[WLH16] Nick Whiteley, Anthony Lee, and Kari Heine. On the role of interaction in sequential Monte Carlo algorithms. Bernoulli, 22 (1):494–529, February 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1443620858.

Wang:2016:CBF

[WM16] Min Wang and Yuzo Maruyama. Consistency of Bayes factor for nonnested model selection when the model dimension grows. Bernoulli, 22(4):2080–2100, November 2016. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1462297675.

Wood:2000:BRE

[Woo00] Andrew T. A. Wood. Bootstrap relative errors and subexponential distributions. Bernoulli, 6(5):809–834, October 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1081282690.

Wang:2022:MFB

[WP22] Yiren Wang and Dimitris N. Politis. Model-free bootstrap for a general class of stationary time series. Bernoulli, 28(2): 744–770, May 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-28/issue-2/Model-free-
bootstrap-for-a-general-class-of-stationary-time/10.
3150/21-BEJ1352.full.

Wang:2023:DBH

[WS23] Runmin Wang and Xiaofeng Shao. Dating the break in high-dimensional data. Bernoulli, 29(4):2879–2901, November 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-29/issue-4/Dating-the-
break-in-high-dimensional-data/10.3150/22-BEJ1567.full.

Wang:2020:MND

[WSX20] Wensheng Wang, Zhonggen Su, and Yimin Xiao. The moduli of non-differentiability for Gaussian random fields with stationary increments. Bernoulli, 26(2):1410–1430, May 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1580461584.

Wu:2012:EEM

[WSZ12] Ping Wu, Winfried Stute, and Li-Xing Zhu. Efficient estimation of moments in linear mixed models. Bernoulli, 18(1):206–228, February 2012. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1327068623.

Wadsworth:2013:NRM

[WT13] J. L. Wadsworth and J. A. Tawn. A new representation for multivariate tail probabilities. Bernoulli, 19(5B):2689–2714, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1386078617.

Wang:2021:TAU

[WT21] Dewei Wang and Chuan-Fa Tang. Testing against uniform stochastic ordering with paired observations. Bernoulli, 27(4):2556–2563, November 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-27/issue-
4/Testing-against-uniform-stochastic-ordering-with-
paired-observations/10.3150/21-BEJ1322.full.

Wu:2003:EPL

[Wu03] Wei Biao Wu. Empirical processes of long-memory sequences. Bernoulli, 9(5):809–831, October 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1066418879.

Wu:2009:CIF

[Wu09] Zhengxiao Wu. A cluster identification framework illustrated by a filtering model for earthquake occurrences. Bernoulli, 15(2):357–379, May 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1241444894.

Wang:2014:CRC

[WW14] Neng-Yi Wang and Liming Wu. Convergence rate and concentration inequalities for Gibbs sampling in high dimension. Bernoulli, 20(4):1698–1716, November 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1411134441.

Wang:2017:ISR

[WXX17] Ran Wang, Jie Xiong, and Lihu Xu. Irreducibility of stochastic real Ginzburg–Landau equation driven by α-stable noises and applications. Bernoulli, 23(2):1179–1201, May 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1486177396.

Wang:2015:VSE

[WXZ15] Tao Wang, Peirong Xu, and Lixing Zhu. Variable selection and estimation for semi-parametric multiple-index models. Bernoulli, 21(1):242–275, February 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1426597069.

Wang:2019:FOC

[WXZ19] Longmin Wang, KaiNan Xiang, and Lang Zou. The first order correction to harmonic measure for random walks of rotationally invariant step distribution. Bernoulli, 25(3):2279–2300, August 2019. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1560326445.

Wegkamp:2011:SVM

[WY11] Marten Wegkamp and Ming Yuan. Support vector machines with a reject option. Bernoulli, 17(4):1368–1385, November 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1320417508.

Wang:2021:OCC

[WYR21] Daren Wang, Yi Yu, and Alessandro Rinaldo. Optimal covariance change point localization in high dimensions. Bernoulli, 27(1):554–575, February 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-1/Optimal-covariance-change-point-localization-
in-high-dimensions/10.3150/20-BEJ1249.full.

Wang:2021:RBM

[WYZ21] Chen Wang, Saisai Yang, and Tusheng Zhang. Reflected Brownian motion with singular drift. Bernoulli, 27(2):866– 898, May 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-27/issue-2/Reflected-
Brownian-motion-with-singular-drift/10.3150/20-BEJ1258.
full.

Wang:2024:CPS

[WYZ24] Yongjin Wang, Chengxin Yan, and Xiaowen Zhou. Comparison principle for stochastic heat equations driven by α-stable white noises. Bernoulli, 30(2):1375–1400, May 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
30/issue-2/Comparison-principle-for-stochastic-heat-
equations-driven-by-%ce%b1-stable/10.3150/23-BEJ1635.
full.

Wegkamp:2016:AEC

[WZ16] Marten Wegkamp and Yue Zhao. Adaptive estimation of the copula correlation matrix for semiparametric elliptical copulas. Bernoulli, 22(2):1184–1226, May 2016. CODEN ????

ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1447077773.

Wu:2018:SQI

[WZ18] Weichi Wu and Zhou Zhou. Simultaneous quantile inference for non-stationary long-memory time series. Bernoulli, 24(4A):2991–3012, November 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1522051231.

Xia:2009:NBE

[Xia09] Yingcun Xia. A note on the backfitting estimation of additive models. Bernoulli, 15(4):1148–1153, November 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1262962229.

Xiao:2020:APP

[Xia20] Luo Xiao. Asymptotic properties of penalized splines for functional data. Bernoulli, 26(4):2847–2875, November 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-26/issue-4/Asymptotic-
properties-of-penalized-splines-for-functional-data/
10.3150/20-BEJ1209.full.

Xu:2010:AMR

[XL10] Xinyi Xu and Feng Liang. Asymptotic minimax risk of predictive density estimation for non-parametric regression. Bernoulli, 16(2):543–560, May 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1274821083.

Xu:2021:ARH

[Xu21] Wei Xu. Asymptotic results for heavy-tailed Lévy processes and their exponential functionals. Bernoulli, 27(4):2766–2803, November 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-27/issue-4/Asymptotic-
results-for-heavy-tailed-L%c3%a9vy-processes-and-their-
exponential/10.3150/21-BEJ1330.full.

Xue:2012:ELS

[XW12] Liugen Xue and Qihua Wang. Empirical likelihood for singleindex varying-coefficient models. Bernoulli, 18(3):836–856, August 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1340887005.

Xiong:2023:SNL

[XY23] Jie Xiong and Xu Yang. SPDEs with non-Lipschitz coefficients and nonhomogeneous boundary conditions. Bernoulli, 29(4):2987–3012, November 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-29/issue-
4/SPDEs-with-non-Lipschitz-coefficients-and-nonhomogeneous-
boundary-conditions/10.3150/22-BEJ1571.full.

Xia:2009:AMB

[XZ09] Aihua Xia and Mei Zhang. On approximation of Markov binomial distributions. Bernoulli, 15(4):1335–1350, November 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1262962238.

Xiong:2018:LDL

[XZ18] Jie Xiong and Jianliang Zhai. Large deviations for locally monotone stochastic partial differential equations driven by Lévy noise. Bernoulli, 24(4A):2842–2874, November 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1522051227.

Yang:2001:NRD

[Yan01] Yuhong Yang. Nonparametric regression with dependent errors. Bernoulli, 7(4):633–655, August 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1079559467.

Yang:2004:ARP

[Yan04] Yuhong Yang. Aggregating regression procedures to improve performance. Bernoulli, 10(1):25–47, February 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1077544602.

Yano:2008:CEP

[Yan08] Kouji Yano. Convergence of excursion point processes and its applications to functional limit theorems of Markov processes on a half-line. Bernoulli, 14(4):963–987, November 2008. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1225980567.

Yang:2024:ABF

[Yan24] Junho Yang. On the asymptotic behavior of a finite section of the optimal causal filter. Bernoulli, 30(4):3137–3164, November 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

Yaroslavtsev:2019:MDW

[Yar19] Ivan S. Yaroslavtsev. Martingale decompositions and weak differential subordination in UMD Banach spaces. Bernoulli, 25 (3):1659–1689, August 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1560326423.

Yao:2006:GML

[YB06] Qiwei Yao and Peter J. Brockwell. Gaussian maximum likelihood estimation for ARMA models II: Spatial processes. Bernoulli, 12(3):403–429, June 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1151525128.

Yeon:2023:BIF

[YDN23] Hyemin Yeon, Xiongtao Dai, and Daniel J. Nordman. Bootstrap inference in functional linear regression models with scalar response. Bernoulli, 29(4):2599–2626, November 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-4/Bootstrap-inference-in-functional-linear-
regression-models-with-scalar-response/10.3150/22-BEJ1554.
full.

Yoo:2019:BMM

[YG19] William Weimin Yoo and Subhashis Ghosal. Bayesian mode and maximum estimation and accelerated rates of contraction. Bernoulli, 25(3):2330–2358, August 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1560326447.

Yin:2022:SSH

[Yin22] Yanqing Yin. Spectral statistics of high dimensional sample covariance matrix with unbounded population spectral norm. Bernoulli, 28(3):1729–1756, August 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
28/issue-3/Spectral-statistics-of-high-dimensional-
sample-covariance-matrix-with-unbounded/10.3150/21-
BEJ1391.full.

Yano:2020:FCE

[YK20] Keisuke Yano and Kengo Kato. On frequentist coverage errors of Bayesian credible sets in moderately high dimensions. Bernoulli, 26(1):616–641, February 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1574758840.

Yano:2021:MPD

[YKK21] Keisuke Yano, Ryoya Kaneko, and Fumiyasu Komaki. Minimax predictive density for sparse count data. Bernoulli, 27(2): 1212–1238, May 2021. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-27/issue-2/Minimax-predictive-
density-for-sparse-count-data/10.3150/20-BEJ1271.full.

Yu:2019:MOE

[YLC19] Zhuqing Yu, Michael Levine, and Guang Cheng. Minimax optimal estimation in partially linear additive models under high dimension. Bernoulli, 25(2):1289–1325, May 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/euclid.bj/1551862851.

Yu:2011:SPR

[YMP11] Kyusang Yu, Enno Mammen, and Byeong U. Park. Semiparametric regression: Efficiency gains from modeling the nonparametric part. Bernoulli, 17(2):736–748, May 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1302009245.

Young:2005:BSR

[You05] James E. Young. Binary sequential representations of random partitions. Bernoulli, 11(5):847–861, October 2005. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1130077597.

Yun:2023:ECG

[YP23] Ho Yun and Byeong U. Park. Exponential concentration for geometric-median-of-means in non-positive curvature spaces. Bernoulli, 29(4):2927–2960, November 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-4/Exponential-concentration-for-geometric-
median-of-means-in-non-positive/10.3150/22-BEJ1569.
full.

Yu:2024:GDP

[YRZ24] Myeonghun Yu, Zhao Ren, and Wen-Xin Zhou. Gaussian differentially private robust mean estimation and inference. Bernoulli, 30(4):3059–3088, November 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

Yuan:2022:SDB

[YS22] Mingao Yuan and Zuofeng Shang. Sharp detection boundaries on testing dense subhypergraph. Bernoulli, 28(4):2459– 2491, November 2022. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-28/issue-4/Sharp-detection-
boundaries-on-testing-dense-subhypergraph/10.3150/21-
BEJ1425.full.

YrJonsdottir:2008:LBG

[ÝSJ08] Kristjana Ýr Jónsdóttir, Jürgen Schmiegel, and Eva B. Vedel Jensen. Lévy-based growth models. Bernoulli, 14(1):62–90, February 2008. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1202492785.

Yu:2010:RLC

[Yu10] Yaming Yu. Relative log-concavity and a pair of triangle inequalities. Bernoulli, 16(2):459–470, May 2010. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1274821079.

Yu:2011:SSI

[Yu11] Yaming Yu. Some stochastic inequalities for weighted sums. Bernoulli, 17(3):1044–1053, August 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1310042855.

Yu:2012:IPS

[Yu12] Yaming Yu. On the inclusion probabilities in some unequal probability sampling plans without replacement. Bernoulli, 18 (1):279–289, February 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1327068626.

Yu:2013:S

[Yu13] Bin Yu. Stability. Bernoulli, 19(4):1484–1500, September 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1377612862.

Yu:2017:SMP

[Yu17] Yaming Yu. Some monotonicity properties of parametric and nonparametric Bayesian bandits. Bernoulli, 23(4B):3685–3710, November 2017. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1495505106.

Yamato:2024:RID

[YY24] Kosuke Yamato and Kouji Yano. Reproduction of initial distributions from the first hitting time distribution for birth-and-death processes. Bernoulli, 30(2):936–960, May 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/journals/
bernoulli/volume-30/issue-2/Reproduction-of-initial-
distributions-from-the-first-hitting-time-distribution/
10.3150/23-BEJ1619.full.

Yor:2001:LDB

[YZ01] Marc Yor and Marguerite Zani. Large deviations for the Bessel clock. Bernoulli, 7(2):351–362, April 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http:

//projecteuclid.org/euclid.bj/1080222082.

Yin:2023:CLT

[YZZ23] Yanqing Yin, Shurong Zheng, and Tingting Zou. Central limit theorem of linear spectral statistics of high-dimensional sample correlation matrices. Bernoulli, 29(2):984–1006, May 2023. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-2/Central-limit-theorem-of-linear-spectral-
statistics-of-high-dimensional/10.3150/22-BEJ1487.full.

Zacharovas:2023:PA

[Zac23] Vytas Zacharovas. Poisson approximation in. Bernoulli, 29(2): 1600–1614, May 2023. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-29/issue-2/Poisson-approximation-
in-%cf%872-distance-by-the-Stein-Chen-approach/10.3150/
22-BEJ1512.full.

Zahle:2015:QRS

[Zäh15] Henryk Zähle. Qualitative robustness of statistical functionals under strong mixing. Bernoulli, 21(3):1412–1434, August 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1432732025.

Zheng:2017:CES

[ZBY17] Shurong Zheng, Zhidong Bai, and Jianfeng Yao. CLT for eigenvalue statistics of large-dimensional general Fisher matrices with applications. Bernoulli, 23(2):1130–1178, May 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1486177395.

Zhang:2018:GAH

[ZC18] Xianyang Zhang and Guang Cheng. Gaussian approximation for high dimensional vector under physical dependence. Bernoulli, 24(4A):2640–2675, November 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1522051220.

Zhang:2023:EWE

[ZC23] Teng Zhang and Sabyasachi Chatterjee. Element-wise estimation error of generalized fused lasso. Bernoulli, 29(4):2691– 2718, November 2023. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-29/issue-4/Element-wise-estimation-
error-of-generalized-Fused-Lasso/10.3150/22-BEJ1557.
full.

Zhang:2023:TCV

[ZDY23] Jiawei Zhang, Jie Ding, and Yuhong Yang. Targeted crossvalidation. Bernoulli, 29(1):377–402, February 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-1/Targeted-cross-validation/10.3150/22-BEJ1461.
full.

Zeineddine:2015:FPV

[Zei15] Raghid Zeineddine. Fluctuations of the power variation of fractional Brownian motion in Brownian time. Bernoulli, 21 (2):760–780, May 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1429624960.

Ziegel:2024:CKH

[ZGD24] Johanna Ziegel, David Ginsbourger, and Lutz Dümbgen. Characteristic kernels on Hilbert spaces, Banach spaces, and on sets of measures. Bernoulli, 30(2):1441–1457, May 2024. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
30/issue-2/Characteristic-kernels-on-Hilbert-spaces-
Banach-spaces-and-on-sets/10.3150/23-BEJ1639.full.

Zhao:2020:ISG

[ZGV20] Yue Zhao, Irène Gijbels, and Ingrid Van Keilegom. Inference for semiparametric Gaussian copula model adjusted for linear regression using residual ranks. Bernoulli, 26(4):2815– 2846, November 2020. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-26/issue-4/Inference-for-
semiparametric-Gaussian-copula-model-adjusted-for-linear-
regression/10.3150/20-BEJ1208.full.

Zwingmann:2020:WCQ

[ZH20] Tobias Zwingmann and Hajo Holzmann. Weak convergence of quantile and expectile processes under general assumptions. Bernoulli, 26(1):323–351, February 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https:/

/projecteuclid.org/euclid.bj/1574758830.

Zhang:1996:SAR

[Zha96] Biao Zhang. Some asymptotic results for kernel density estimation under random censorship. Bernoulli, 2(2):183–198, June 1996. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.
bj/1193839223.

Zhang:2000:QEU

[Zha00] Biao Zhang. Quantile estimation under a two-sample semiparametric model. Bernoulli, 6(3):491–511, June 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1081616702.

Zhang:2006:EES

[Zha06] Lan Zhang. Efficient estimation of stochastic volatility using noisy observations: a multi-scale approach. Bernoulli, 12(6): 1019–1043, December 2006. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1165269149.

Zhang:2013:MSC

[Zha13] Tong Zhang. Multi-stage convex relaxation for feature selection. Bernoulli, 19(5B):2277–2293, November 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1386078603.

Zhang:2018:DAC

[Zha18] Zhiyi Zhang. Domains of attraction on countable alphabets. Bernoulli, 24(2):873–894, May 2018. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1505980881.

Zhao:2022:OUC

[Zha22] Yunpeng Zhao. An optimal uniform concentration inequality for discrete entropies on finite alphabets in the high-dimensional setting. Bernoulli, 28(3):1892–1911, August 2022. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-28/issue-3/An-optimal-
uniform-concentration-inequality-for-discrete-entropies-
on-finite/10.3150/21-BEJ1403.full.

Zhang:2023:CTM

[Zha23] Zhuo-Song Zhang. Cramér-type moderate deviation of normal approximation for unbounded exchangeable pairs. Bernoulli, 29(1):274–299, February 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/journals/bernoulli/volume-29/issue-
1/Cram%c3%a9r-type-moderate-deviation-of-normal-approximation-
for-unbounded-exchangeable/10.3150/21-BEJ1457.full.

Zholud:2014:TAS

[Zho14a] Dmitrii Zholud. Tail approximations for the Student t-, F -, and Welch statistics for non-normal and not necessarily i.i.d. random variables. Bernoulli, 20(4):2102–2130, November 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1411134455.

Zhou:2014:NSN

[Zho14b] Zhou Zhou. Nonparametric specification for non-stationary time series regression. Bernoulli, 20(1):78–108, February 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1390407281.

Zhou:2019:SHW

[Zho19] Shuheng Zhou. Sparse Hanson–Wright inequalities for subgaussian quadratic forms. Bernoulli, 25(3):1603–1639, August 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1560326421.

Zhou:2019:EVA

[ZHZL19] Cheng Zhou, Fang Han, Xin-Sheng Zhang, and Han Liu. An extreme-value approach for testing the equality of large Ustatistic based correlation matrices. Bernoulli, 25(2):1472– 1503, May 2019. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1551862857.

Zorin-Kranich:2020:WLI

[ZK20] Pavel Zorin-Kranich. Weighted Lépingle inequality. Bernoulli, 26(3):2311–2318, August 2020. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1587974542.

Zhang:2003:LPF

[ZL03] Jian Zhang and Anna Liu. Local polynomial fitting based on empirical likelihood. Bernoulli, 9(4):579–605, August 2003. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1066223270.

Zhao:2013:IMS

[ZL13] Zhibiao Zhao and Xiaoye Li. Inference for modulated stationary processes. Bernoulli, 19(1):205–227, February 2013. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1358531747.

Zhao:2016:API

[ZL16] Junlong Zhao and Chenlei Leng. An analysis of penalized interaction models. Bernoulli, 22(3):1937–1961, August 2016. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1458133003.

Zhang:2017:APS

[ZL17] Tonglin Zhang and Ge Lin. Asymptotic properties of spatial scan statistics under the alternative hypothesis. Bernoulli, 23 (1):89–109, February 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1475001349.

Zhao:2017:QRS

[ZLL17] Weihua Zhao, Heng Lian, and Hua Liang. Quantile regression for the single-index coefficient model. Bernoulli, 23(3):1997– 2027, August 2017. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1489737631.

Zastavnyi:2011:CTG

[ZP11] Viktor P. Zastavnyi and Emilio Porcu. Characterization theorems for the Gneiting class of space–time covariances. Bernoulli, 17(1):456–465, February 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1297173850.

Zemel:2019:FMP

[ZP19] Yoav Zemel and Victor M. Panaretos. Fréchet means and Procrustes analysis in Wasserstein space. Bernoulli, 25(2):932–976, May 2019. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL https://projecteuclid.org/euclid.bj/
1551862840.

Zwiernik:2012:TCG

[ZS12] Piotr Zwiernik and Jim Q. Smith. Tree cumulants and the geometry of binary tree models. Bernoulli, 18(1):290–321, February 2012. CODEN ???? ISSN 1350-7265 (print), 1573-9759

(electronic). URL http://projecteuclid.org/euclid.bj/
1327068627.

Zhang:2015:TSI

[ZS15] Xianyang Zhang and Xiaofeng Shao. Two sample inference for the second-order property of temporally dependent functional data. Bernoulli, 21(2):909–929, May 2015. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1429624965.

Zhu:2021:IDB

[ZS21] Changbo Zhu and Xiaofeng Shao. Interpoint distance based two sample tests in high dimension. Bernoulli, 27(2):1189– 1211, May 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://projecteuclid.

org/journals/bernoulli/volume-27/issue-2/Interpoint-
distance-based-two-sample-tests-in-high-dimension/10.
3150/20-BEJ1270.full.

Zhao:2022:ADM

[ZSC22] Qian Zhao, Pragya Sur, and Emmanuel J. Candès. The asymptotic distribution of the MLE in high-dimensional logistic models: Arbitrary covariance. Bernoulli, 28(3):1835– 1861, August 2022. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-28/issue-3/The-asymptotic-
distribution-of-the-MLE-in-high-dimensional-logistic/
10.3150/21-BEJ1401.full.

Zhao:2014:ANR

[ZWL14] Zhibiao Zhao, Ying Wei, and Dennis K. J. Lin. Asymptotics of nonparametric l-1 regression models with dependent data. Bernoulli, 20(3):1532–1559, August 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1402488949.

Zhao:2021:SDR

[ZWS21] Puying Zhao, Lei Wang, and Jun Shao. Sufficient dimension reduction and instrument search for data with nonignorable nonresponse. Bernoulli, 27(2):930–945, May 2021. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
27/issue-2/Sufficient-dimension-reduction-and-instrument-
search-for-data-with-nonignorable/10.3150/20-BEJ1260.
full.

Zhao:2017:SML

[ZWY17] Xingqiu Zhao, Yuanshan Wu, and Guosheng Yin. Sieve maximum likelihood estimation for a general class of accelerated hazards models with bundled parameters. Bernoulli, 23(4B):3385– 3411, November 2017. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
euclid.bj/1495505096.

Zhou:2017:TAE

[ZX17] Yuzhen Zhou and Yimin Xiao. Tail asymptotics for the extremes of bivariate Gaussian random fields. Bernoulli, 23(3):1566–1598, August 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL https://

projecteuclid.org/euclid.bj/1489737618.

Zhang:2024:NEB

[ZX24] Shuxiong Zhang and Jie Xiong. A note on the empty balls of a critical super-Brownian motion. Bernoulli, 30(1):72– 87, February 2024. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-30/issue-1/A-note-on-the-
empty-balls-of-a-critical-super/10.3150/22-BEJ1564.
full.

Zhai:2015:LDD

[ZZ15] Jianliang Zhai and Tusheng Zhang. Large deviations for 2-D stochastic Navier–Stokes equations driven by multiplicative Lévy noises. Bernoulli, 21(4):2351–2392, November 2015. CO-DEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

http://projecteuclid.org/euclid.bj/1438777597.

Zhang:2023:ESS

[ZZ23] Xiaolong Zhang and Xicheng Zhang. Ergodicity of supercritical SDEs driven by α-stable processes and heavy-tailed sampling. Bernoulli, 29(3):1933–1958, August 2023. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL

https://projecteuclid.org/journals/bernoulli/volume-
29/issue-3/Ergodicity-of-supercritical-SDEs-driven-
by-%ce%b1-stable-processes-and/10.3150/22-BEJ1526.full.

Zhang:2022:NAP

[ZZW22] Lyuou Zhang, Wen Zhou, and Haonan Wang. Nonasymptotic properties of spectral decomposition of large Gram-type matrices and applications. Bernoulli, 28(2):1224– 1249, May 2022. CODEN ???? ISSN 1350-7265 (print),

1573-9759 (electronic). URL https://projecteuclid.org/
journals/bernoulli/volume-28/issue-2/Non-asymptotic-
properties-of-spectral-decomposition-of-large-Gram-
type/10.3150/21-BEJ1384.full.

Zhou:2017:TSS

[ZZZ17] Wen-Xin Zhou, Chao Zheng, and Zhen Zhang. Two-sample smooth tests for the equality of distributions. Bernoulli, 23 (2):951–989, May 2017. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1486177389.